### v2

# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

Abstract.
We prove that an apparent horizon can form as a result of Einsteinian evolution in pure vacuum spacetime starting from regular initial data free of apparent horizons due to pure boundary effects. We adapt a Cauchy-double-null framework and use the boundary generalized mean curvature condition for the existence of an interior apparent horizon imposed by the author S-T Yau in [38]. In particular, we prove that the condition of [38] can be met dynamically starting from a configuration that does not verify the same through a focusing mechanism. This is the first part of a two-part sequence, and in the sequel, we will focus on explicitly constructing the Cauchy data.
1. Introduction
In this article, we consider the $3+1$ dimensional pure vacuum Einstein’s equation and investigate the issue of large data semi-global existence and dynamical existence of an apparent horizon. Consider a $3+1$ dimensional globally hyperbolic $C^{\infty}$ connected oriented Lorentzian manifold $(\mathcal{M},g)$. The vacuum Einstein equations correspond to the vanishing of the Ricci curvature of $(\mathcal{M},g)$
(1) $\text{Ric}[g]=0$
and therefore the free gravity is described by the Weyl curvature components. The formation of a black hole is one of the central issues in general relativity and deserves deep attention from a rigorous analytic perspective. We briefly recall the following historical note
1.1. Historical background
Apparent horizons have been an object of significant interest within the classical theory of General Relativity for almost sixty years. After Schwarzschild’s discovery of his eponymous metric in 1915, it took almost twenty years before researchers came to realize the existence, within it, of a region $\mathcal{B}$ with the following surprising yet salient features: First of all, observers situated inside $\mathcal{B}$ cannot send signals to observers situated at an ideal conformal boundary at infinity, called $\mathcal{I}^{+}$. Furthermore, any observer located inside $\mathcal{B}$ lives only for finite proper time [^2] Sbierski [32] moreover showed that the termination of the observer’s proper time manifests in a particularly ferocious way, as they, in fact, get torn apart by infinitely strong tidal forces. (geodesic incompleteness). The characteristics of this region (which later came to be known as a black hole) took most of the researchers of the time aback. The consensus seemed to be that these observed phenomena have to be accidents; pathologies, only present because of the strong (spherical) symmetry inherent in the Schwarzschild solution and that, in general solutions to the Einstein equations [^3] The meaning of this phrase was not rigorous at the time, as the setup for the initial value problem in General Relativity had not yet been discovered., such phenomena would not arise. However, in the 60’s, this belief was spectacularly falsified by Roger Penrose through his celebrated incompleteness theorem [^4] This theorem, in fact, was the main reason why he was awarded the Nobel Prize in Physics back in 2020, ”for the discovery that black hole formation is a robust prediction of the general theory of relativity”.. It was Penrose [29] who introduced the notion of a trapped surface, without which one cannot state his eponymous, celebrated, incompleteness theorem :
Definition 1.
Given a $(3+1)$- dimensional Lorentzian manifold $(\mathcal{M},g)$, a closed spacelike 22-surface $S$ is caled trapped if the following two fundamental forms $\chi$ and $\underline{\chi}$ have everywhere pointwise negative expansions on $S$:
$\chi(X,Y):=g(D_{X}L,Y),\hskip 5.69054pt\underline{\chi}(X,Y):=g(D_{X}\underline{L},Y).$
Here $D$ denotes the Levi-Civita connection of $g$, $L$ and $\underline{L}$ denote a null basis of the 2-dimensional orthogonal complement of $T_{p}S$ in $T_{p}\mathcal{M}$, extended as smooth vector fields and $X,Y$ are arbitrary $S$-tangent vector fields.
In other words, a surface is called trapped if both $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ are pointwise negative everywhere on $S$. These traces signify the infinitesimal changes in area along the null generators normal to $S$, whence one can interpret trapped surfaces as closed, spacelike 22-surfaces that infinitesimally decrease in area "along any possible future direction".
Closely related to the trapped surface is the notion of apparent horizon. The definition of an apparent horizon differs from that of a trapped surface by the fact that the trace $\mbox{tr}\chi$ of the null outgoing second fundamental form $\chi$ vanishes point-wise, while $\mbox{tr}\underline{\chi}$, the trace of the null incoming second fundamental form, is point-wise negative. The formal definition is as follows
Definition 2.
Given a $(3+1)$- dimensional Lorentzian manifold $(\mathcal{M},g)$, a closed spacelike 22-surface $S$ is caled apparent horizon if the fundamental forms $\chi$ and $\underline{\chi}$ have everywhere zero and negative expansions on $S$, respectively i.e.,
(2) $\mbox{tr}\chi=0,~\mbox{tr}\underline{\chi}<0~on~S.$
Formally, an apparent horizon can be interpreted as the outermost boundary of a domain containing closed trapped surfaces in a Cauchy slice.
The incompleteness theorem is now presented.
Theorem 1.1 (Penrose Incompleteness).
Let $(\mathcal{M},g)$ be a spacetime containing a non-compact Cauchy hypersurface. If $(\mathcal{M},g)$ moreover satisfies the null energy codition and contains a closed trapped surface, it is geodesically incomplete.
The existence of an apparent horizon (and trapped surfaces contained in it) is a stable feature in the context of dynamics. Indeed, sufficiently small perturbations of Schwarzschild initial data must also contain such surfaces, by Cauchy stability. As such, incompleteness is not an accident, but rather a recurring theme in the dynamics of the Einstein equations.
At the time, the existence of an apparent horizon in a spacetime was too strong an assumption to begin with. In fact, the only way back then to guarantee its existence was to assume it at the level of initial data, but this itself can be a highly non-trivial question. The first apparent horizon formation result at the level of initial data was given in [37]. Later, [38] provided a stronger result by proving the existence of an apparent horizon at the level of the initial data set because of the boundary effect. In particular, this result did not require the inclusion of matter density (in fact, matter density could be allowed to be negative). This, however, begs the question of whether apparent horizons are dynamical objects, meaning whether they can be formed in evolution starting with data devoid of such surfaces. This problem bears high physical significance and serves as a test of reality of black holes, in the following sense. The mathematical definition of a black hole region would be without physical meaning if it did not accurately capture what physicists perceive as black holes (this is more meaningful now than ever, as scientists recently succeeded in capturing the first-ever image of a black hole). Hence, if "mathematical" black holes describe "physical" black holes, they should mathematically verify certain physical properties, one of which is dynamical formation.
The first results along this direction were obtained by Christodoulou for the Einstein equations coupled to a massless scalar field in spherical symmetry. Through a series of works [16], [14], [15], and [18], Christodoulou managed to not only prove trapped surface formation, but to understand the picture of gravitational collapse in its entirety for the given model and under the given symmetry. The breakthrough in the absence of symmetry came in [17] by the same author, and this study was, in a sense, closely related to the earlier work of S-T Yau [38] and a motivating point of this current study. In this work, Christodoulou introduced a hierarchy of small and large components in the initial data which (almost) persists under the evolution of the Einstein equations. He termed his method the short pulse method. After Christodoulou, the work [22] by Klainerman-Rodnianski reduces the size of Christodoulou’s work from about 600 to approximately 120 pages by using a slightly different hierarchy. Moreover, it reduces the number of derivatives of curvature required to prove semi-global existence from two to one. A few years later, An [3] introduces the signature for decay rates $s_{2}$ on his way to proving an extension of [22] from a finite region to a region close to past null infinity. In 2014, An and Luk [6] proved the first scale-critical trapped surface formation criterion for the vacuum equations in the absence of symmetry. While Christodoulou’s data in [17] were large in $H^{1}(\mathbb{R}^{3})$, An and Luk give data which only have to be large in $H^{\frac{3}{2}}(\mathbb{R}^{3})$, which is a scale-critical norm for the initial data. Taking advantage of the scale criticality in [6], An [2] constructs initial data that give rise not merely to trapped surfaces, but an apparent horizon, a smooth 3-dimensional hypersurface consisting of marginally outer trapped surfaces. In 2019, An [1] produces a 55-page proof of trapped surface formation for the vacuum equations, making use of the signature for decay rates and obtaining an existence result from a region close to past null infinity. In [4], An and the first author extend [1] to the case of the Einstein-Maxwell system. Later, the present authors proved the large data scale-invariant semi-global existence and trapped surface formation result in [4]. Several other studies exist in the context of Einstein-Yang-Mills [11], Einstein-Vlasov system [21, 7, 8, 9, 12], Einstein-Scalar field system, and Einstein-Spinor field system. Recently, [13], in a fundamental study, provided a short proof of the formation of a trapped surface in geodesic foliation.
In this article, we focus on the study by S-T Yau [38] and prove that indeed an apparent horizon can form in a dynamical manner starting from a regular configuration. This is a new idea in the sense that we want to obtain the existence of an apparent horizon as a result of boundary effect in a Cauchy slice that is a product of evolution according to the vacuum Einstein equations 62. Let us make this idea precise. Let $\widehat{\mathcal{M}}_{t}$ be a domain with boundary $S_{t}$ in the Cauchy slice $\mathcal{M}_{t}$. On the Cauchy slice $\mathcal{M}_{t}$, the Einstein constraint equations
(3) $\text{Scal}[h]-|k|^{2}+(\mbox{tr}_{h}k)^{2}=0,~\mbox{div }_{h}(k-\mbox{tr}_{h}kh)=0$
are verified, where $h$ is the induced metric on the slice $\mathcal{M}_{t}$ and $k$ is second fundamental form of this slice. Here $\text{Scal}[h]$ denotes the scalar curvature of the metric $h$. The boundary $S_{t}$ of the domain $\widehat{\mathcal{M}}_{t}\subset\mathcal{M}_{t}$ is co-dimension $2$ in the spacetime $(\mathcal{M},g)$ and hence possesses a time-like and a spacelike second fundamental form. Let $(e_{T},e_{S})$ be the orthonormal pair spanning the tangent bundle of $S_{t}$ where $e_{T}$ is time-like and $e_{S}$ is spacelike. We define the spacelike and timelike second fundamental form of $S_{t}$ as
(4) $s_{AB}:=\langle\nabla_{e_{A}}e_{S},e_{B}\rangle,k_{AB}:=\langle\nabla_{e_{A}}e_{T},e_{B}\rangle,~A=1,2$
where $\{e_{A}\}_{A=1,2}$ are the orthonormal frame tangential to $S_{t}$. If the induced metric on $S_{t}$ is denoted by $\Sigma_{AB}$, then $H=s_{AB}(\Sigma^{-1})^{AB}$ is the spacelike mean curvature of $S_{t}$ while $\kappa:=k_{AB}(\Sigma^{-1})^{AB}$ is the time-like mean curvature of $S_{t}$ also same as the restriction of the trace of the second fundamental form $k_{ij}$ of $\widehat{\mathcal{M}}_{t}$ to $S_{t}$. In addition, we also recall the notion of radius of $\widehat{\mathcal{M}}_{t}$ defined by [38]
Definition 3.
Given a region $\Omega$ and a Jordon curve $\Gamma\subset\partial\Omega$ which bounds an embedded disk in $\Omega$ and a subdomain in $\partial\Omega$, let $R_{\Gamma}$ be the supremum of $r>0$ such that $\Gamma$ does not bound a disk inside the tube of $\Gamma$ with radius $r<R_{\Gamma}$. The radius Rad of $\Omega$ is defined to be the supremum of all such $R_{\Gamma}$.
With these geometric-topological notions, [38] proves a remarkable theorem on the existence of an apparent horizon on a Cauchy slice. More precisely, the theorem is as follows
Theorem 1.2.
[38] Let $M$ be a space-like hypersurface in a spacetime. Let $g_{ij}$ be its induced metric and $k_{ij}$ be its second fundamental form. Assume that the spacelike mean curvature $H$ of $\partial M$ is strictly greater than its time-like mean curvature $\mbox{tr}_{\partial M}k=\kappa$. Let $c:=\min_{\partial M}\bigg({H-|\kappa|}\bigg)$. If $c\geq\frac{3\pi}{2\text{Rad}(M)}$, then $M$ must admit an apparent horizon in its interior.
Note that this is purely a boundary effect in the sense that if the (generalized) boundary mean curvature of a domain in the Cauchy slice is significantly higher, then an apparent horizon must exist inside. This leads to the following natural question: Can one start from a regular configuration free of any apparent horizons and from an apparent horizon in an evolutionary manner in finite time. This statement needs to be made more precise, such as exactly what it means by a regular configuration, etc. We will do this momentarily. First, compare this to the work by [17]. Working in a double null framework, [17] proves that if the incoming gravitational radiation at past null infinity is sufficiently large, then a trapped surface will form in the “interior" as a result of Einstein (pure vacuum) evolution. Since the trapped surface would need to verify $\mbox{tr}\chi<0,\mbox{tr}\underline{\chi}<0$, with enough regularity, one would naturally expect to have $\mbox{tr}\chi=0,\mbox{tr}\chi<0$ (apparent horizon) before formation of a trapped surface. In this work, we indeed adapt a different approach and a different data type that yields the formation of an apparent horizon as a result of evolution.
2. Geometric setup and semi-global development
Let $(\mathcal{M},g)$ be a smooth, time-oriented, four-dimensional Lorentzian manifold. We assume that $(\mathcal{M},g)$ admits a smooth double null foliation in the following sense: there exist smooth optical functions
$(u,\underline{u}):\mathcal{M}\to\mathbb{R}\times\mathbb{R}$
such that for each fixed value of $u$ (respectively $\underline{u}$), the level set
$H_{u}:=\{p\in\mathcal{M}\,:\,u(p)=\mathrm{const}\},\qquad\underline{H}_{\underline{u}}:=\{p\in\mathcal{M}\,:\,\underline{u}(p)=\mathrm{const}\}$
is a smooth null hypersurface, which we refer to as an outgoing (respectively incoming) null hypersurface.
For each pair $(u,\underline{u})$ for which $H_{u}\cap\underline{H}_{\underline{u}}\neq\emptyset$, we define
$S_{u,\underline{u}}:=H_{u}\cap\underline{H}_{\underline{u}}.$
We assume that $S_{u,\underline{u}}$ is a smooth, embedded, spacelike 22–surface diffeomorphic to $\mathbb{S}^{2}$, and we denote by $\not{\gamma}_{u,\underline{u}}$ the Riemannian metric on $S_{u,\underline{u}}$ induced by $g$. When convenient, we abbreviate $\Sigma_{u,\underline{u}}:=(S_{u,\underline{u}},\not{\gamma}_{u,\underline{u}})$.
We prescribe characteristic initial data on two intersecting null hypersurfaces, namely on the incoming null hypersurface
$\underline{H}_{0}:=\{\underline{u}=0\}$
and the outgoing null hypersurface
$H_{u_{\infty}}:=\{u=u_{\infty}\},$
where $u_{\infty}\in\mathbb{R}$ is fixed. The hierarchy and size of the data that we impose on $H_{u_{\infty}}\cup\underline{H}_{0}$ differ in an essential way from those considered in [1, 5, 6, 11, 12]. In particular, we allow large (nonperturbative) incoming radiation along one null direction, while retaining only a degenerate smallness along the transverse direction. This constitutes the first principal new feature of the present work.
From these data, we construct a semi-global causal development, which we denote by
$\mathcal{D}_{a,\epsilon}:=[u_{\infty},-a]\times[0,\epsilon]\times S,$
where $S\simeq\mathbb{S}^{2}$ is a fixed reference two–sphere, $a\gg 1$ is a large parameter, and $\epsilon>0$ is a small parameter. The notation above is to be interpreted as follows: the variables $(u,\underline{u})$ range in the rectangle
$u_{\infty}\leq u\leq-a,\qquad 0\leq\underline{u}\leq\epsilon,$
and the angular variables range in $S$. Hence $\mathcal{D}_{a,\epsilon}$ is the spacetime region covered by the portion of the double null foliation determined by these bounds. A schematic representation of $\mathcal{D}_{a,\epsilon}$ is provided in Figure 1.
We now introduce the associated canonical spacelike foliation. Define the function
$t:=u+\underline{u}.$
For each constant $t\in\mathbb{R}$, we denote by
$\mathcal{M}_{t}:=\{p\in\mathcal{M}\,:\,u(p)+\underline{u}(p)=t\}$
the corresponding level set. We assume that, for the range of $t$ under consideration, $\mathcal{M}_{t}$ is a smooth spacelike Cauchy hypersurface for the relevant portion of $(\mathcal{M},g)$; in particular, $\{\mathcal{M}_{t}\}$ defines a spacelike foliation compatible with the above double null foliation.
Fix $a\gg 1$ as above and consider the slice $\mathcal{M}_{t=-a}$. We define the intersection of this slice with the semi–global development by
$\widehat{\mathcal{M}}_{-a}:=\mathcal{M}_{t=-a}\cap\mathcal{D}_{a,\epsilon}.$
It is convenient to decompose the entire Cauchy slice $\mathcal{M}_{t=-a}$ into three regions: an “interior” region $\widehat{\mathcal{M}}_{\mathrm{int}}$, lying to the past of $\mathcal{D}_{a,\epsilon}$; the interaction region $\widehat{\mathcal{M}}_{-a}$ defined above; and an “exterior” region $\widehat{\mathcal{M}}_{\mathrm{ext}}$, lying to the future/outside of $\mathcal{D}_{a,\epsilon}$. In particular, we write
(5) $\mathcal{M}_{t=-a}\;=\;\widehat{\mathcal{M}}_{\mathrm{int}}\;\cup\;\widehat{\mathcal{M}}_{-a}\;\cup\;\widehat{\mathcal{M}}_{\mathrm{ext}}.$
The decomposition (5) isolates the portion $\widehat{\mathcal{M}}_{-a}$ of the Cauchy hypersurface on which the large characteristic data propagate and interact, separating it from the complementary interior and exterior regions of $\mathcal{M}_{t=-a}$.
Figure 1. The schematics of the current framework: concentration of the generalized mean curvature $H-|\kappa|$ while increasing the radius. The initial data is provided on the null hypersurfaces $u=u_{\infty}$ and $\underline{u}=0$, and the interior Cauchy slice $\mathcal{M}_{1}=\widehat{\mathcal{M}}_{int}$. The Cauchy data on the slice $\widehat{\mathcal{M}}_{t=-a}$ is prescribed by gluing data on $\mathcal{M}_{1}$, the induced data on $\mathcal{M}_{2}$ by the Characteristic development on the slab $\mathcal{D}_{a,\epsilon}:=[u_{\infty},-a]\times[0,\epsilon]\times\mathbb{S}^{2}$, and exterior of a Kerr slice data on $\mathcal{M}_{3}$. The data on $\mathcal{M}_{1}$ is set such that it does not contain an apparent horizon. By the solution of the Characteristic semi-global problem in $\mathcal{D}_{a,\epsilon}$, one does not have any apparent horizon in the entire Cauchy slice $\widehat{\mathcal{M}}_{t=-a}$. In particular, $S_{-a,0}=\underline{H}_{0}\cap H_{-a}$ does not verify the condition of [38], while $S_{-a,\epsilon}:=H_{-a}\cap\underline{H}_{\epsilon}$ does. Therefore, $J^{+}(\mathcal{M}_{1})\cap\widehat{M}_{t=-a+\epsilon}$ has an apparent horizon inside.
3. Gluing of Cauchy data and horizon exclusion
By the semi-global existence result established in the previous section, the development $\mathcal{D}_{a,\epsilon}$ induces canonical Cauchy data on the portion $\widehat{\mathcal{M}}_{-a}:=\mathcal{M}_{t=-a}\cap\mathcal{D}_{a,\epsilon}$. On the complementary regions $\widehat{\mathcal{M}}_{\mathrm{int}}$ and $\widehat{\mathcal{M}}_{\mathrm{ext}}$, one retains the freedom to prescribe data independently, subject only to the constraint and rigidity conditions imposed by the positive mass theorem [33, 34, 35]. In particular, the data on the exterior region $\widehat{\mathcal{M}}_{\mathrm{ext}}$ are far more constrained than those on the interior piece $\widehat{\mathcal{M}}_{\mathrm{int}}$, and a natural choice is to attach a Kerr exterior with prescribed ADM mass and angular momentum via the Corvino–Schoen gluing construction. The objective is then to glue the data on the three components
$\widehat{\mathcal{M}}_{\mathrm{int}},\qquad\widehat{\mathcal{M}}_{-a},\qquad\widehat{\mathcal{M}}_{\mathrm{ext}},$
so as to produce a smooth global Cauchy data set on $\mathcal{M}_{t=-a}$ satisfying the Einstein constraint equations. Note that in the context of dynamical trapped surface formation [23] constructed an initial data set free of trapped surfaces for the Cauchy problem compatible with Christodoulou’s short pulse data type. Later, [10] constructed an initial data set for the dynamical formation of the apparent horizon (though this definition of apparent horizon is slightly different from ours in the sense that in their definition it is a spacelike hypersurface that is foliated by marginally outer trapped surfaces (i.e., $\mbox{tr}\chi=0$) whereas in our context, by an apparent horizon, we simply mean a spacelike topological 22-sphere for which $\mbox{tr}\chi=0,\mbox{tr}\underline{\chi}<0$) and subsequently addressed the Penrose inequality.
The guiding idea is to choose the initial data on the null cones $H_{u_{\infty}}$, $\underline{H}_{0}$, and on the interior Cauchy region $\widehat{\mathcal{M}}_{\mathrm{int}}$ in such a way that the boundary of $\widehat{\mathcal{M}}_{\mathrm{int}}$ fails to satisfy the apparent–horizon condition of [38], namely,
(6) $c_{\partial\widehat{\mathcal{M}}_{\mathrm{int}}}:=\min_{\partial\widehat{\mathcal{M}}_{\mathrm{int}}}\big[H-|\kappa|\big]<\frac{3\pi}{2\,\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})}.$
The strict inequality (6) ensures that $\widehat{\mathcal{M}}_{\mathrm{int}}$ contains no apparent horizon. Assuming that throughout the semi–global region $\mathcal{D}_{a,\epsilon}$ one has
$\mbox{tr}\chi>0,\qquad\mbox{tr}\underline{\chi}<0,$
and that the exterior Kerr portion $\widehat{\mathcal{M}}_{\mathrm{ext}}$ is also free of trapped surfaces, it follows that no apparent horizon occurs anywhere in the spacetime before or up to the slice t=at=-a.
Note that the semi–global domain $\mathcal{D}_{a,\epsilon}$ includes the causal past of the topological sphere $S_{-a,\epsilon}$. By transporting the geometric quantity $c$ along the incoming null direction, each sphere $S_{u,\underline{u}}$ near $(-a,\epsilon)$ can be focused so that $S_{-a,\epsilon}$ attains a large value of $c$. Concurrently, by the standard local existence theory for the Cauchy problem, one obtains a future development of the slice $\mathcal{M}_{t=-a}$, thereby generating a nearby Cauchy hypersurface $\mathcal{M}_{t=-a+\epsilon}$. For $\epsilon>0$ sufficiently small—chosen in a teleological manner depending on the prior data—the slice $\mathcal{M}_{t=-a+\epsilon}$ inherits analogous geometric estimates (e.g. curvature and lapse–shift bounds). Moreover, since $\mbox{tr}\chi>0$ along u=au=-a by the semi–global construction, the intersection of the causal future of $\widehat{\mathcal{M}}_{\mathrm{int}}$ with $\mathcal{M}_{t=-a+\epsilon}$ satisfies
(7) $\mathrm{Rad}\!\left(J^{+}(\widehat{\mathcal{M}}_{\mathrm{int}})\cap\mathcal{M}_{t=-a+\epsilon}\subset\mathcal{M}_{t=-a+\epsilon}\right)\approx\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})+C\epsilon,$
and for a constant $C$ dependent on the initial data and so for $\epsilon$ sufficiently small
(8) $\mathrm{Rad}\!\left(J^{+}(\widehat{\mathcal{M}}_{\mathrm{int}})\cap\mathcal{M}_{t=-a+\epsilon}\subset\mathcal{M}_{t=-a+\epsilon}\right)\approx\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})+\epsilon^{\frac{1}{2}},$
showing that the interior region expands monotonically towards the future for sufficiently small time. This is sufficient for our purpose.
We now analyze the geometric condition on the future Cauchy slice $\mathcal{M}_{t=-a+\epsilon}$ that signals the onset of trapped geometry. Observe that the boundary of the domain
$\widehat{\mathcal{M}}_{\mathrm{int}}\cap\mathcal{M}_{t=-a+\epsilon}$
within $\mathcal{M}_{t=-a+\epsilon}$ coincides with the topological two–sphere $S_{-a,\epsilon}$. Define, analogously to (6), the boundary curvature functional
$c_{S_{-a,\epsilon}}:=\min_{S_{-a,\epsilon}}\big[H-|\kappa|\big],$
where $H$ denotes the mean curvature of $S_{-a,\epsilon}$ in $\mathcal{M}_{t=-a+\epsilon}$ and $\kappa$ is the trace of the second fundamental form of $\mathcal{M}_{t=-a+\epsilon}$ restricted to $S_{-a,\epsilon}$. Suppose that
(9) $c_{S_{-a,\epsilon}}>\frac{3\pi\big(\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})-\epsilon^{\frac{1}{2}}\big)}{2\,\big[\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})\big]^{2}}.$
Then, according to the criterion of [38], the domain $\widehat{\mathcal{M}}_{\mathrm{int}}\cap\mathcal{M}_{t=-a+\epsilon}$ must contain an apparent horizon in its interior.
The key geometric mechanism underlying (9) is as follows. Rather than starting from a configuration dominated by large incoming radiation from past null infinity, we consider the complementary regime in which the outgoing radiation along the incoming null hypersurface $\underline{H}_{0}$ is dominant. Concretely, we assume that the squared shear $|\underline{\hat{\chi}}|^{2}$ is large on $\underline{H}_{0}$, which enhances focusing via the null structure equations. This strong shear, together with nonlinear couplings among the Ricci coefficients and curvature components, amplifies the quantity $H-|\kappa|$ at the boundary sphere $S_{-a,\epsilon}$. Simultaneously, as indicated by (6)–(8), The radius of the interior region increases from $\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})$ to $\mathrm{Rad}\big(J^{+}(\widehat{\mathcal{M}}_{\mathrm{int}})\cap\mathcal{M}_{t=-a+\epsilon}\big)$. The combined effect is that strong outgoing radiation, concentrated within a suitably defined ringed region, distorts the boundary geometry of the interior domain, producing a surface with sufficiently large generalized mean curvature to satisfy the inequality (9). This mechanism provides a new geometric route to apparent-horizon formation.
Let us now try to understand how this mechanism would unfold in a heuristic manner. First, recall that in the double null gauge that we are considering here
(10) $e_{4}=\Omega^{-1}\partial_{\underline{u}},~e_{3}=\Omega^{-1}\bigg(\partial_{u}+b^{A}\partial_{\theta^{A}}\bigg).$
The corresponding connection coefficients that appear in this study are $(\hat{\chi},\mbox{tr}\chi,\underline{\hat{\chi}},\mbox{tr}\underline{\chi},\omega,\underline{\omega},\eta,\underline{\eta})$. In this coordinate, one may explicitly compute the $H-|\kappa|$ for a topological sphere $S_{u,\underline{u}}$ constituting the double null foliation
(11) $H-|\kappa|=H-|\mbox{tr}_{S}K|=\frac{1}{2}\bigg(\mbox{tr}\chi-\mbox{tr}\underline{\chi}\bigg)-\frac{1}{2}\bigg|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\bigg|.$
This leads us to directly control $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ instead of working with separate evolution equations for $H$ and $\kappa$. The main equations that we focus on are the following
(12) $\nabla_{3}\mbox{tr}\underline{\chi}+\frac{1}{2}(\mbox{tr}\underline{\chi})^{2}=-|\underline{\hat{\chi}}|^{2}-2\underline{\omega}\mbox{tr}\underline{\chi},$
(13) $\nabla_{4}\mbox{tr}\chi+\frac{1}{2}(\mbox{tr}\chi)^{2}=-|\hat{\chi}|^{2}-2\omega\mbox{tr}\chi,$
(14) $\nabla_{4}\hat{\underline{\chi}}+\frac{1}{2}\mbox{tr}\chi\hat{\underline{\chi}}=\nabla\hat{\otimes}\underline{\eta}+2\omega\hat{\underline{\chi}}-\frac{1}{2}tr\underline{\chi}\hat{\chi}+\underline{\eta}\hat{\otimes}\underline{\eta}$
Now, the scaling hierarchy of norm that we will use, the lapse $\Omega$, the shift $b$, and the Ricci coefficients will verify the following estimates
[…]
(16) $||\eta,\underline{\eta}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|^{2}},~||\omega||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{1}{|u|},~||\hat{\chi}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{1}{|u|}.$
With this scaling hierarchy, we can integrate $\mbox{tr}\underline{\chi}$ along the incoming direction, i.e., work with the following equation
(17) $\nabla_{3}\mbox{tr}\underline{\chi}+\frac{1}{2}(\mbox{tr}\underline{\chi})^{2}=-|\underline{\hat{\chi}}|^{2}-2\underline{\omega}\mbox{tr}\underline{\chi}.$
Now use the fact that $e_{3}=\Omega^{-1}\bigg(\partial_{u}+b^{A}\partial_{\theta^{A}}\bigg)$ and write
(18) $\nabla_{3}(|u|\mbox{tr}\underline{\chi})=-\bigg(\mbox{tr}\underline{\chi}+\frac{2}{|u|}\bigg)\frac{|u|\mbox{tr}\underline{\chi}}{2}-|u||\underline{\hat{\chi}}|^{2}-2|u|\underline{\omega}\mbox{tr}\underline{\chi}+(1-\Omega^{-1})\mbox{tr}\underline{\chi}$
which, upon integration and using the bounds 15 one obtains for every $\underline{u}\in[0,\epsilon]$ (provided one have the semi-global construction already)
(19) $\mbox{tr}\underline{\chi}(-a,\underline{u})=\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\underline{u})}{a}-\frac{1}{a}\int_{u_{\infty}}^{a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\underline{u})du^{{}^{\prime}}+\frac{C(\mathcal{I}^{0})}{a^{2}},$
where the constants involved depend only on the initial data. In particular, note that the term $\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}du^{{}^{\prime}}$ can make $\mbox{tr}\underline{\chi}(-a,\underline{u})$ large negative. Now we control $\mbox{tr}\chi$ through integrating along $\nabla_{4}$ direction. Recall that the Raichowdhury equation
(20) $\nabla_{4}\mbox{tr}\chi+\frac{1}{2}(\mbox{tr}\chi)^{2}=-|\hat{\chi}|^{2}-2\omega\mbox{tr}\chi.$
But the presence of $|\hat{\chi}|^{2}$ term with a negative sign can reduce $\mbox{tr}\chi$. As it turns out, controlling the size of $\epsilon$ and the new hierarchy where $\hat{\chi}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim|u|^{-1}$, the reduction of $\mbox{tr}\chi$ can be controlled and this potentially dangerous term contributes to a negligible amount. Integration and using the bounds 15 yield
(21) $\mbox{tr}\chi(u,\epsilon)=\mbox{tr}\chi(u,0)-\int_{0}^{\epsilon}\bigg(-\frac{1}{2}(\mbox{tr}\chi)^{2}-|\hat{\chi}|^{2}-2\omega\mbox{tr}\chi\bigg)du^{{}^{\prime}}\approx\mbox{tr}\chi(u,0)-\frac{Ca\epsilon}{|u|^{2}}.$
Now we are at a place to actually compute $H-|\kappa|$ at $(-a,\epsilon)$ and (a,0)(-a,0). Using 11, we obtain
(22) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=\bigg(\mbox{tr}\chi-\mbox{tr}\underline{\chi}\bigg)(-a,\epsilon)-\bigg|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\bigg|(-a,\epsilon)$
(23) $=\bigg(\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a}-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}-\frac{C}{a^{2}}\bigg)$
(24) $-\bigg|\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}+\frac{C}{a^{2}}\bigg|$
and similarly
(25) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,0)=\bigg(\mbox{tr}\chi(-a,0)-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{C}{a^{2}}\bigg)$
(26) $-\bigg|\mbox{tr}\chi(-a,0)+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{C}{a^{2}}\bigg|.$
Now integrating the $\nabla_{4}$ equation one can control $|\underline{\hat{\chi}}|^{2}(u,\epsilon)$ in terms of $|\underline{\hat{\chi}}|^{2}(u,0)$. Recall the following $\nabla_{4}$ equation verified by $\underline{\hat{\chi}}$
(27) $\nabla_{4}\hat{\underline{\chi}}+\frac{1}{2}\mbox{tr}\chi\hat{\underline{\chi}}=\nabla\hat{\otimes}\underline{\eta}+2\omega\hat{\underline{\chi}}-\frac{1}{2}\mbox{tr}\underline{\chi}\hat{\chi}+\underline{\eta}\hat{\otimes}\underline{\eta}$
implying
(28) $\nabla_{4}\bigg[|u||\underline{\hat{\chi}}|^{2}\bigg]=-|u|\mbox{tr}\chi|\underline{\hat{\chi}}|^{2}+2|u|\underline{\hat{\chi}}\nabla\otimes\underline{\eta}+4|u|\omega|\underline{\hat{\chi}}|^{2}-|u|\mbox{tr}\underline{\chi}\hat{\chi}\cdot\underline{\hat{\chi}}+2|u|\underline{\hat{\chi}}\underline{\eta}\otimes\underline{\eta}$
which, after integration and utilizing the estimates 15
(29) $|u||\underline{\hat{\chi}}|^{2}(u,\epsilon)\approx|u||\underline{\hat{\chi}}|^{2}(u,0)+\frac{C\epsilon a^{\frac{3}{2}}}{|u|^{4}}+\frac{C\epsilon a^{\frac{3}{4}}}{|u|^{4}}+\frac{C\epsilon a^{\frac{3}{4}}}{|u|^{3}}+\frac{C\epsilon a^{\frac{7}{4}}}{|u|^{5}}.$
Therefore, $|u||\underline{\hat{\chi}}|^{2}(u,\epsilon)\approx|u||\underline{\hat{\chi}}|^{2}(u,0)$ up to negligible error term. Therefore, we observe the following
(30) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=$ $\bigg(\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{C}{a^{2}}\bigg)$
(31) $-\bigg|\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{C}{a^{2}}\bigg|.$
Subsequently, we would want
(32) $\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{C}{a^{2}}\approx 0$
which is compatible with the data choice. which would lead to the following two expressions
(33) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,0)=-\frac{2|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{2}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{C}{a^{2}}$
and
(34) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=-\frac{2|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{2}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}-\frac{C}{a^{2}}$
since $\mbox{tr}\chi(u_{\infty},0)\approx\mbox{tr}\chi(u_{\infty},\epsilon)$ up to $O(|u_{\infty}|^{-2})$ error terms by the choice of initial data at $u=u_{\infty}$. This provides us with the exact conditions for the non-existence of an apparent horizon in $\widehat{\mathcal{M}}_{int}$ and a dynamical existence of an apparent horizon in $J^{+}(\widehat{\mathcal{M}}_{int})\cap M_{t=-a+\epsilon}$. These are equivalent to the following condition with $\epsilon=O(a^{-\frac{1}{2}})$
(35) $\frac{3\pi(\text{Rad}(\widehat{\mathcal{M}}_{int})-\epsilon^{\frac{1}{2}})}{[\text{Rad}(\widehat{\mathcal{M}}_{int})]^{2}}+\frac{C}{a^{2}}<-\frac{2|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{2}{a}\int_{u_{\infty}}^{a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}<\frac{3\pi}{\text{Rad}(\widehat{\mathcal{M}}_{int})}+\frac{C}{a^{2}}$
Therefore, if the initial outgoing shear and the magnitude of the incoming expansion at the past null inifnity $u=u_{\infty}$ is sufficiently large, then the boundary $\partial\bigg(J^{+}(\widehat{\mathcal{M}}_{int})\cap\mathcal{M}_{t=-a+\epsilon}\bigg)=S_{-a,\epsilon}$ would have sufficiently large generalized mean curvature $H-|\kappa|$ and therefore according to the condition of [38], an apparent horizon will exist in the interior of $J^{+}(\widehat{\mathcal{M}}_{int})\cap\mathcal{M}_{t=-a+\epsilon}$.
Note the scalings of $\underline{\hat{\chi}}$ and $\mbox{tr}\underline{\chi}$ (adapted from [5])
(36) $|\underline{\hat{\chi}}(u,0)|\lesssim a^{\frac{1}{2}}|u|^{-2},~|\mbox{tr}\underline{\chi}(u_{\infty},0)|\lesssim|u_{\infty}|^{-1}$
which seem to be well-suited for the propagation of this scaling norm hierarchy. The remaining task is then to prove the semi-global existence result in the double null framework. In addition, with this scaling, note that the term $\frac{1}{a}\int_{u_{\infty}}^{-a}||u^{{}^{\prime}}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime}}$ and the non-linear contributions in the $\nabla_{3}$ equation for $\mbox{tr}\underline{\chi}$ (equation 12) contributes at the same order as the non-linear error terms ($O(a^{-2})$). Therefore, the initial large-scale-invariant norm of $\underline{\hat{\chi}}$ should be prescribed such that it dominates that of the remaining Ricci coefficients (as in 39). The solution to the subsequent Cauchy problem for t=at=-a to $t=-a+\epsilon$ for sufficiently small $\epsilon$ is a routine process and, as such, incurs significantly less difficulty. Of course, one main task is to actually construct the asymptotically flat Cauchy data on the entire slice $\mathcal{M}_{t=-a}$ from the characteristic initial development and the interior radius condition. The two main points that we want to recall are the following initial data norm
(37) $\begin{split}\Gamma^{0}_{2}:=&\sup_{S_{u,\underline{u}}\subset H_{u_{\infty}},\underline{H}_{0}}\bigg(\sum_{i\leq N}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\mbox{tr}\chi,\hat{\chi},\eta,\underline{\eta},\omega,\underline{\omega})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ &+\sum_{i\leq N}\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\sum_{i\leq N}\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\bigg),\end{split}$
(38) $\Gamma^{0}_{\underline{\hat{\chi}}2}:=\sup_{S_{u,\underline{u}}\subset H_{u_{\infty}},\underline{H}_{0}}\sum_{i\leq N}\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},$
(39) $\Gamma^{0}_{\underline{\hat{\chi}}2}>100\Gamma^{0}_{2}$
(40) $\mathcal{R}^{0}_{2}:=\sup_{S_{u,\underline{u}}\subset H_{u_{\infty}},\underline{H}_{0}}\bigg(\sum_{i\leq N-1}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\sum_{i\leq N-1}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\underline{\alpha},\beta,\underline{\beta},a^{\frac{1}{2}}\rho,a^{\frac{1}{2}}\sigma)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\bigg),$
Finally, for $0\leq i\leq N+4$, we define the norms along the null hypersurfaces:
(41) $\mathcal{R}(u_{\infty},\underline{u}):=\sum_{i\leq N}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+a^{\frac{1}{2}}\sum_{i\leq N}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta,\underline{\beta},\rho,\sigma)}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})},$
(42) $\underline{\mathcal{R}}(u,0):=\sum_{i\leq N}\bigg(\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\underline{\beta},\rho,\sigma,\underline{\alpha})}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\bigg).$
Notice that this data choice is very close to the one developed by [3] and later used by [4, 11, 12] only difference being the shear. In the present framework, we need a large integrated conjugate shear $\underline{\hat{\chi}}$ on the incoming null hypersurface $\underline{H}_{0}$ while we are free to prescribe data on the initial outgoing null hypersurface $H_{u_{\infty}}$ that is compatible with the scaling 37-42. In particular, our data does not have Minkowski initial data on $\underline{H}_{0}$. In this sense, it is different from the previous choices of data considered. On the other hand, the prescription of Cauchy data on the interior part of the Cauchy slice $\mathcal{M}_{-a}$ is rather implicit and relies on the Riemannian geometric arguments of [38]. In the sequel [28], we will address the construction of the Cauchy data explicitly.
Remark 1.
Note a slight scaling difference between [3] and the current study. The $a^{\frac{1}{2}}$ gain in the decay of $\hat{\chi}$ is compensated in $\rho$.
Theorem 3.1 (Main Theorem).
Fix a smooth double null foliation of a four–dimensional Lorentzian manifold $(\mathcal{M},g)$ with optical functions $(u,\underline{u})$ such that the level sets $H_{u}$ and $H_{\underline{u}}$ are outgoing and incoming null hypersurfaces, respectively, and
$S_{u,\underline{u}}:=H_{u}\cap H_{\underline{u}}$
denotes the corresponding two–sphere sections equipped with the induced metric $\gamma$. Let $\nabla,$ $\nabla_{3}$, and $\nabla_{4}$ denote the angular and null derivatives associated with the null frame
$\{e_{1},e_{2},e_{3},e_{4}\}.$
(a) Semi–Global Existence. For every sufficiently large fixed constant $\mathcal{I}>0$, there exists $a_{0}=a_{0}(\mathcal{I})\gg 1$ such that the following holds. Let $a>a_{0}$ and let $N\in\mathbb{N}$ be sufficiently large. Suppose the initial data set is prescribed along the characteristic hypersurfaces $H_{u_{\infty}}$ and $\underline{H}_{0}$, where $u_{\infty}<0$ is large in magnitude, and satisfies the boundedness property of the norms described in 37-42
(43) $\Gamma^{0}_{2}+\Gamma^{0}_{\underline{\hat{\chi}}2}+\mathcal{R}^{0}_{2}+\sup_{\underline{u}}\mathcal{R}(u_{\infty},\underline{u})+\sup_{u}\mathcal{R}(u,0)<\mathcal{I}^{0},~\Gamma^{0}_{\underline{\hat{\chi}}2}>100\Gamma^{0}_{2},$
for sufficiently large $\mathcal{I}^{0}$ such that $\mathcal{I}^{0}a^{-\frac{1}{20}}<1$ and moreover choose the shift $b=0$ and the lapse $\Omega=1$ on the initial null hypersurface $H_{u_{\infty}}$. Then there exists a unique smooth solution $g$ of the vacuum Einstein equations in the region
$\mathcal{D}:=\big\{(u,\underline{u})\mid u_{\infty}\leq u\leq-a,\;\;0\leq\underline{u}\leq\epsilon\big\},$
satisfying the constraint and transport equations, with all geometric quantities $\psi\in\{\mbox{tr}\chi,\hat{\chi},\underline{\hat{\chi}},\mbox{tr}\underline{\chi},\eta,\underline{\eta},\alpha,\underline{\beta},\rho,\sigma\}$ remaining uniformly controlled by constants depending only on the initial data.
(b) Dynamical Existence of an apparent Horizon. Suppose, in addition to the bound 37-42, that the initial data along $\underline{H}_{0}$ satisfies the bound
(44) $\frac{3\pi(\text{Rad}(\widehat{\mathcal{M}}_{int})-\epsilon^{\frac{1}{2}})}{[\text{Rad}(\widehat{\mathcal{M}}_{int})]^{2}}+\frac{C(\Gamma^{0}_{2})}{a^{2}}<-\frac{2|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{2}{a}\int_{u_{\infty}}^{a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}<\frac{3\pi}{\text{Rad}(\widehat{\mathcal{M}}_{int})}+\frac{C(\Gamma^{0}_{2})}{a^{2}},$
for a constant $C$ depending on the scale-invariant initial data norm $\Gamma^{0}_{2}$ controlling the Ricci coefficients $(\hat{\chi},\mbox{tr}\chi,\omega,\underline{\omega},\eta,\underline{\eta})$, uniformly for every direction along $\underline{u}=0$ and the radius $\text{Rad}(\widehat{\mathcal{M}}_{int})$ of the interior Cauchy slice at canonical double null time t=at=-a in the construction is prescribed to construct its Cauchy data, and the null expansion $\mbox{tr}\chi>0$ on $u=u_{\infty}$ compatible with the scaling 37-42, then the interior Cauchy slice $J^{+}(\widehat{\mathcal{M}}_{int})\cap\mathcal{M}_{t=-a+\epsilon}\subset\mathcal{M}_{t=-a+\epsilon}$ at canonical double null time $t=-a+\epsilon$ will contain an apparent horizon in its interior.
3.1. Main Ideas and Novelty
The principal analytical difficulty of the present work lies in the construction of the initial data, whether prescribed on a characteristic or a Cauchy hypersurface. The heuristics discussed in the preceding section suggest that the nonlinearities arising in the transport equations for the null Ricci coefficients contribute, in a delicate sense, only subdominant effects to the leading dynamics. Arbitrary initial data, however, fail to exhibit such cancellation structures, and in fact, by the celebrated small-data result of Christodoulou–Klainerman [17], one already knows that for sufficiently small perturbations of Minkowski data, no apparent horizon or trapped surface can form.
In the large-data regime, the situation is profoundly different: the system is fully nonlinear, no global small parameter exists, and long-time uniform control cannot, in general, be expected in the hyperbolic setting. As a result, one faces the possibility of geometric pathologies, including singularity formation and breakdown of the foliation. A central challenge, therefore, consists in constructing a hierarchy of large but controlled initial data $(g,k)$ that is compatible with a semi-global existence theorem for the Einstein vacuum equations, and whose evolution can be followed up to the onset of an apparent horizon.
From a structural viewpoint, the formation of a trapped surface is governed by the Raychaudhuri equation for the null expansion $\mbox{tr}\chi$ along the outgoing null direction:
[…]
In the traditional approach (cf. [16, 22, 1, 3]), one attempts to produce a negative expansion $\mbox{tr}\chi<0$ within a finite affine parameter time along the null generators. This mechanism is driven by the largeness of the incoming gravitational shear $|\widehat{\chi}|^{2}$, which acts as a source in (45).
By contrast, the present work adopts a complementary perspective. Rather than directly forcing $\mbox{tr}\chi$ to become negative, we study the evolution of the generalized mean curvature quantity
(46) $c:=H-|\kappa|,$
as introduced in Section 1.2. Our objective is to obtain a strictly positive value of $c$ on an appropriate boundary hypersurface, thus triggering the dynamical emergence of an apparent horizon from a previously horizon–free configuration. In the double–null gauge, one has the decomposition
(47) $H-|\kappa|=\tfrac{1}{2}\big(\mbox{tr}\chi-\mbox{tr}\underline{\chi}\big)-\tfrac{1}{2}\big|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\big|.$
The guiding principle is therefore to drive the difference $\mbox{tr}\chi-\mbox{tr}\underline{\chi}$ to become large and positive while simultaneously suppressing the absolute term $|\mbox{tr}\chi+\mbox{tr}\underline{\chi}|$ along the incoming null direction.
This strategy is realized by prescribing a large incoming shear $\underline{\hat{\chi}}$ on the initial null hypersurface $\underline{H}_{0}$ and an appropriately scaled profile for $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ on the distant outgoing hypersurface $u=u_{\infty}$ consistent with the hierarchy of the data. The evolution of $\mbox{tr}\underline{\chi}$ along the incoming direction is then governed by its transport equation, which contains both quadratic and indefinite–sign nonlinearities. The main difficulty arises in ensuring that the positive enforcing term
$a^{-1}\int_{u_{\infty}}^{-a}|u^{\prime}|\,|\underline{\hat{\chi}}|^{2}\,du^{\prime}$
dominates all lower–order contributions. To achieve this dominance, we impose a relative largeness condition
(48) $\Gamma^{0}_{\underline{\hat{\chi}},2}>100\,\Gamma^{0}_{2},$
where $\Gamma^{0}_{\underline{\hat{\chi}},2}$ and $\Gamma^{0}_{2}$ denote the scale–invariant initial data norms controlling $\underline{\hat{\chi}}$ and the remaining Ricci coefficients, respectively. Remarkably, this relative hierarchy is propagated uniformly throughout the entire semi-global development for sufficiently large parameter $a$ used in this study.
An additional constraint arises from the requirement that the radius–mean curvature inequality (9) is satisfied along evolution. The mechanism can be schematically summarized as follows. Along the incoming direction, we enforce $\mbox{tr}\underline{\chi}$ to grow in magnitude near u=au=-a, while $\mbox{tr}\chi$ remains uniformly positive. Owing to the monotonicity property implied by (45), the only viable means of maintaining positivity of $\mbox{tr}\chi$ is to restrict the $\underline{u}$–width of the slab to be sufficiently small. Geometrically, this ensures that the t=at=-a Cauchy slice bounded by $S_{-a,0}$ possesses a smaller radius than its nearby deformation bounded by $S_{-a,\epsilon}$ on $t=-a+\epsilon$, as required for the desired curvature inequality.
The methodology presented here represents a substantial conceptual departure from all previous works on trapped surface formation. The open set of characteristic data constructed herein accommodates a broader class of large data than those treated in the earlier frameworks of Christodoulou and Klainerman–Rodnianski. Moreover, the Cauchy data prescribed on the interior slice $\mathcal{M}_{1}$ is defined implicitly through a Riemannian construction in the spirit of Yau [38], involving the notion of prescribed Yau radius and mean curvature control. This interior data will be rendered explicit and developed further in the companion paper [28].
The analytical framework developed here constitutes a genuinely new approach to the dynamical formation of apparent horizons and trapped surfaces in vacuum general relativity. To the best of our knowledge, no previous result in the literature establishes the emergence of such geometric features through the control of the generalized mean curvature quantity $c=H-|\kappa|$ within a semi–global evolution. In particular, the present construction yields an open set of characteristic initial data exhibiting a substantially broader admissible class than those considered in the works of Christodoulou, Klainerman–Rodnianski, or Luk, thereby enlarging the known landscape of large–data configurations leading to horizon formation.
The corresponding Cauchy data prescribed on the interior slice $\mathcal{M}_{1}$ is, at this stage, defined only implicitly through a Riemannian geometric construction inspired by Yau [38]. Specifically, the data $(g,k)$ on $\mathcal{M}_{1}$ are required to satisfy the boundary mean curvature condition consistent with a prescribed Yau radius, ensuring the geometric compatibility between the interior region and the exterior double–null development. A concrete realization of this construction, including an explicit parametrization of $(g,k)$ and verification of the constraint equations, will be presented in the sequel [28].
A fundamental analytical difficulty arises in propagating the full hierarchy of weighted and scale-invariant data norms from the initial null and Cauchy hypersurfaces throughout the domain of semi-global existence. Once this propagation is established, the Riemannian argument of Yau [38], when applied to the interior slice $\mathcal{M}_{1}$, becomes decisive: the boundary mean curvature condition then determines, in a geometric and quantitative sense, whether an apparent horizon must or must not exist. It is noteworthy that the scaling framework introduced by An [3], after a slight modification consistent with the present gauge choice, integrates seamlessly into this construction and provides the correct normalization for the hierarchy of Ricci coefficients and curvature fluxes.
Another distinctive aspect of the present analysis is that elliptic estimates are not required. By formulating the entire problem at a sufficiently high level of Sobolev regularity within the hyperbolic (double-null) setting, we circumvent any dependence on auxiliary elliptic estimates and maintain full control through transport-type and energy-type arguments (this is well known by now in the context of the characteristic initial value problem). Finally, the structure developed here aligns naturally with the local constructions of Luk [24], suggesting that one can generate more general families of initial data on finite domains satisfying Yau’s boundary condition in a dynamically consistent fashion.
4. Setup
4.1. Construction of the double null gauge
Denote by $L\mathcal{M}$ the frame bundle of $\mathcal{M}$. We construct a double null gauge, meaning a smooth section of this bundle such that, through it, each point $p\in\mathcal{M}$ maps to a renormalized frame $(e_{1},e_{2},e_{3},e_{4})\in L\mathcal{M}$ with $g(e_{3},e_{4})=-2$, $g(e_{A},e_{B})=\epsilon_{AB}$ and $g(e_{3},e_{A})=g(e_{4},e_{A})=0$.
We begin with two null hypersurfaces $H_{u_{\infty}},\underline{H}_{0}$ and their intersection $S_{u_{\infty},0}$, a topological 2-sphere. For any point $q$ on this 22-sphere, the tangent space $T_{q}S_{u_{\infty},u}$ is 2-dimensional and admits a 22-dimensional orthogonal complement $T_{q}^{\perp}{S_{u_{\infty},0}}$, on which we can find two future-directed null vectors $L^{\prime}_{q}$ and $\underline{L}^{\prime}_{q}$, normalized so that
$g(L^{\prime}_{q},\underline{L}^{\prime}_{q})=-2.$
The pair $\begin{Bmatrix}L^{\prime}_{q},\underline{L}^{\prime}_{q}\end{Bmatrix}$ is uniquely determined up to a scaling factor $s>0$
$\begin{Bmatrix}L^{\prime}_{q},\underline{L}^{\prime}_{q}\end{Bmatrix}\mapsto\begin{Bmatrix}sL^{\prime}_{q},s^{-1}\underline{L}^{\prime}_{q}\end{Bmatrix}.$
Starting from $q$ and initially tangent to $L^{\prime}_{q}$, a unique geodesic is sent out. Call this geodesic, $l_{q}$. We extend the vectorfield $L^{\prime}$ along $l_{q}$ by parallel transport: $D_{L^{\prime}}L^{\prime}=0$. It then follows by simple calculation that $l_{q}$ is null, so that $g(L^{\prime},L^{\prime})=0$ along $l_{q}$. Gathering the $\begin{Bmatrix}l_{q}\end{Bmatrix}$ together we get a null hypersurface $H_{u_{\infty}}$. The null hypersurface $\underline{H}_{0}$ is obtained similarly. Note that, by construction, given a point $p$ on $H_{u_{\infty}}$ or $\underline{H}_{0}$, in the corresponding tangent spaces, there is a preferred null vector $L^{\prime}_{p}$ or $\underline{L}^{\prime}_{p}$.
We next choose a lapse function $\Omega$, which we define to be equal to 1 on $S_{u_{\infty},0}$ and then extend as a continuous function along both initial null hypersurfaces [^5] Indeed, there is a gauge freedom in choosing $\Omega$ on the initial hypersurfaces.. Define the vector fields
$L:=\Omega^{2}L^{\prime}\hskip 5.69054pt\text{along}\hskip 5.69054ptH_{u_{\infty}}\hskip 5.69054pt\text{and}\hskip 5.69054pt\underline{L}:=\Omega^{2}\underline{L}^{\prime}\hskip 5.69054pt\text{along}\hskip 5.69054pt\underline{H}_{0}.$
We use these vector fields to define two functions
$\underline{u}\hskip 5.69054pt\text{on}\hskip 5.69054ptH_{u_{\infty}}\hskip 5.69054pt\text{satisfying}\hskip 5.69054ptL\underline{u}=1\hskip 5.69054pt\text{on}\hskip 5.69054ptH_{u_{\infty}}\hskip 5.69054pt\text{and}\hskip 5.69054pt\underline{u}=0\hskip 5.69054pt\text{on}\hskip 5.69054ptS_{u_{\infty},0},$
$u\hskip 5.69054pt\text{on}\hskip 5.69054pt\underline{H}_{0}\hskip 5.69054pt\text{satisfying}\hskip 5.69054pt\underline{L}u=1\hskip 5.69054pt\text{on}\hskip 5.69054pt\underline{H}_{0}\hskip 5.69054pt\text{and}\hskip 5.69054ptu=0\hskip 5.69054pt\text{on}\hskip 5.69054ptS_{u_{\infty},0}.$
We now use these (so-called optical) functions to proceed further with the construction. Let $S_{u_{\infty},\underline{u}^{\prime}}$ be the embedded 22-surface on $H_{u_{\infty}}$ on which $\underline{u}=\underline{u}^{\prime}$ and define $S_{u,0}$ similarly. At each point $p\in S_{u_{\infty},\underline{u}^{\prime}}$ we have constructed a preferred null vector $L^{\prime}_{p}$. It follows that we can uniquely determine an incoming $g$-null vector $\underline{L}^{\prime}_{p}$ satisfying $g(L^{\prime}_{p},\underline{L}^{\prime}_{p})=-2$. Let $\underline{l}_{p}$ be the unique geodesic emanating from $p$ with tangent vector $\underline{L}^{\prime}_{p}$. We extend the definition of $\underline{L}^{\prime}$ along $\underline{l}_{q}$ by parallel transport, so that $D_{\underline{L}^{\prime}}\underline{L}^{\prime}=0$. Gathering all the $\begin{Bmatrix}\underline{l}_{p}\end{Bmatrix}$ on $S_{u_{\infty},\underline{u}^{\prime}}$, we thus obtain the null hypersurface $\underline{H}_{\underline{u}^{\prime}}$. We obtain the null hypersurface $H_{u^{\prime}}$ in an analogous way and define $S_{u,\underline{u}}:=H_{u}\cap\underline{H}_{\underline{u}}$. Having constructed the vector fields $L^{\prime}$ and $\underline{L}^{\prime}$ in all of the spacetime region, we extend the definition of the lapse function $\Omega$ by requiring, at each point $p\in S_{u,\underline{u}}$ that
$g(L^{\prime}_{p},\underline{L}^{\prime}_{p})=-2{\left.\kern-1.2pt\Omega^{-2}\right|_{p}}.$
The incoming null hypersurfaces $\begin{Bmatrix}\underline{H}_{\underline{u}}\end{Bmatrix}_{0\leq\underline{u}\leq 1}$ and outgoing null hypersurfaces $\begin{Bmatrix}H_{u}\end{Bmatrix}_{-a\leq u\leq u_{\infty}}$ along with their pairwise intersections $S_{u,\underline{u}}$ together define a double null foliation on the spacetime. On a given $S_{u,\underline{u}}$, we have $g(\Omega L^{\prime},\Omega\underline{L}^{\prime})=-2$ and hence the vectors
$e_{3}:=\Omega\underline{L}^{\prime},e_{4}:=\Omega L^{\prime}$
define a normalized null pair at each point on the sphere. We make the gauge choice $\Omega\equiv 1$ along both initial hypersurfaces.
4.2. Choice of coordinates and expression of the metric
To define angular coordinates on each $S_{u,\underline{u}}$ in a smooth way, we begin by defining angular coordinates on $S_{u_{\infty},0}$. Since this is a standard 2-sphere in Minkowki space, we can use the stereographic projection coordinates $(\theta^{1},\hskip 1.42262pt\theta^{2})$ on $S_{u_{\infty},0}$. We first extend this coordinate to the whole of $\underline{H}_{0}$ by insisting that $\not{\mathcal{L}}_{\underline{L}}\theta^{A}=0$ on $\underline{H}_{0}$ for $A=1,\hskip 1.42262pt2$ and then to the whole spacetime by insisting that, for all $u$, $\not{\mathcal{L}}_{L}\theta^{A}=0$, where $L$ initially starts normal to some $S_{u,0}$. As such we have established a coordinate system $(u,\underline{u},\theta^{1},\theta^{2})$ in a neighbourhood of the initial sphere. In these coordinates, the vectors $e_{3},\hskip 1.42262pte_{4}$ become
$e_{3}=\Omega^{-1}\left(\frac{\partial}{\partial u}+b^{A}\hskip 1.42262pt\frac{\partial}{\partial\theta^{A}}\right),\hskip 5.69054pte_{4}=\Omega^{-1}\frac{\partial}{\partial\underline{u}}$
and the metric now takes the following form:
[…]
It is a requirement that $b^{A}\equiv 0$ on $\underline{H}_{0}$. The section that maps $p\in\mathcal{M}\mapsto\left({\left.\kern-1.2pt\theta^{1}\right|_{p}},{\left.\kern-1.2pt\theta^{2}\right|_{p}}{\left.\kern-1.2pte_{3}\right|_{p}},{\left.\kern-1.2pte_{4}\right|_{p}}\right)$ is the double null gauge we wanted to construct.
4.3. The vacuum Einstein equations expressed in the double null gauge
In this section we are going to express the Einstein-Yang-Mills system in the double null gauge given above. We begin by decomposing curvature components and Ricci coefficients with respect to the frame $(e_{1},e_{2},e_{3},e_{4})$. Let $A,B$ take values in $\begin{Bmatrix}1,2\end{Bmatrix}$. We define the following (Weyl) null curvature components:
$\alpha_{AB}:=W(e_{A},e_{4},e_{B},e_{4}),\hskip 5.69054pt\underline{\alpha}_{AB}:=W(e_{A},e_{3},e_{B},e_{3}),$
$\beta_{A}:=\frac{1}{2}W(e_{A},e_{4},e_{3},e_{4}),\hskip 5.69054pt\underline{\beta}_{A}:=\frac{1}{2}W(e_{A},e_{3},e_{3},e_{4}),$
$\rho:=\frac{1}{4}W(e_{3},e_{4},e_{3},e_{4}),\hskip 5.69054pt\sigma=\frac{1}{4}\prescript{*}{}{W}(e_{3},e_{4},e_{3},e_{4}).$
For the Ricci coefficients, we decompose as follows:
$\chi_{AB}:=g(D_{A}e_{4},\hskip 1.42262pte_{B}),\hskip 5.69054pt\underline{\chi}_{AB}:=g(D_{A}e_{3},e_{B}),$
$\eta_{A}:=-\frac{1}{2}g(D_{A}e_{3},\hskip 1.42262pte_{4}),\hskip 5.69054pt\underline{\eta}_{A}:=-\frac{1}{2}g(D_{A}e_{4},e_{3}),$
$\omega:=-\frac{1}{4}g(D_{4}e_{3},\hskip 1.42262pte_{4}),\hskip 5.69054pt\underline{\omega}:=-\frac{1}{4}g(D_{3}e_{4},\hskip 1.42262pte_{3}),$
$\zeta_{A}:=\frac{1}{2}g(D_{A}e_{4},\hskip 1.42262pte_{3}).$
Moreover, if $\gamma$ denotes the induced metric on $S_{u,\underline{u}}$, we make the following further decomposition:
$\chi=\hat{\chi}+\frac{1}{2}\mbox{tr}\chi\gamma,\hskip 5.69054pt\underline{\chi}=\underline{\hat{\chi}}+\frac{1}{2}\mbox{tr}\underline{\chi}\hskip 1.42262pt\gamma.$
Before we are ready to present the equations, we introduce a few basic definitions. First of all, denote by $\nabla$ the covariant derivative operators induced by $D$ on $S_{u,\underline{u}}$. Let $\nabla_{3},\nabla_{4}$ denote the projections of the covariant derivatives $D_{3}$ and $D_{4}$ to $S_{u,\underline{u}}$. For two 11-forms $\phi_{A}^{1},\hskip 1.42262pt\phi_{A}^{2}$, we define
$(\phi_{1}\hat{\otimes}\phi_{2})_{AB}:=\phi_{A}^{1}\hskip 1.42262pt\phi_{B}^{2}+\phi_{B}^{1}\hskip 1.42262pt\phi_{A}^{2}-\gamma_{AB}\hskip 1.42262pt(\phi^{1}\cdot\phi^{2}),$
while for symmetric 22-tensors $\phi_{AB}^{1},\hskip 1.42262pt\phi_{AB}^{2}$, we define
$(\phi^{1}\wedge\phi^{2})_{AB}:=\not{\epsilon}^{AB}\hskip 1.42262pt(\gamma^{-1})^{CD}\hskip 1.42262pt\phi_{AB}^{1}\hskip 1.42262pt\phi_{CD}^{2}.$
Here $\not{\epsilon}$ is the volume form associated with the metric $\gamma$. Moreover, by $\phi^{1}\cdot\phi^{2}$ we mean an arbitrary contraction of the tensor product of $\phi^{1}$ and $\phi^{2}$ with respect to the metric $\gamma$. We also define suitable trace, divergence and curl operators. For totally symmetric tensors $\phi$, we define these operators as follows:
$(\mbox{div }\phi)_{A_{1}\dots A_{r}}:=\nabla^{B}\phi_{BA_{1}\dots A_{r}},$
$(\mbox{curl }\phi)_{A_{1}\dots A_{r}}:=\not{\epsilon}^{BC}\nabla_{B}\phi_{CA_{1}\dots A_{r}},$
$(\mbox{tr}\phi)_{A_{1}\dots A_{r-1}}:=(\gamma^{-1})^{BC}\phi_{BCA_{1}\dots A_{r-1}}.$
[…]
$\prescript{*}{}{\phi}_{A}:=\gamma_{AC}\not{\epsilon}^{CB}\hskip 1.42262pt\phi_{B},$
$\prescript{*}{}{\phi}_{AB}:=\gamma_{BD}\hskip 1.42262pt\not{\epsilon}^{DC}\hskip 1.42262pt\phi_{AC}.$
Finally, on a 11-form $\phi$, the operator $\nabla\hat{\otimes}$ is defined as follows:
[…]
The vacuum Einstein equations take the following form in this double null gauge
(50) $\nabla_{4}\mbox{tr}\chi+\frac{1}{2}(\mbox{tr}\chi)^{2}=-|\hat{\chi}|^{2}_{\gamma}-2\omega\mbox{tr}\chi$
(51) $\nabla_{4}\hat{\chi}+\mbox{tr}\chi\hat{\chi}=-2\omega\hat{\chi}-\alpha$
(52) $\nabla_{3}tr\underline{\chi}+\frac{1}{2}(tr\underline{\chi})^{2}=-|\hat{\underline{\chi}}|^{2}_{\gamma}-2\underline{\omega}tr\underline{\chi}$
(53) $\nabla_{3}\hat{\underline{\chi}}+tr\underline{\chi}\hat{\underline{\chi}}=-2\underline{\omega}\hat{\underline{\chi}}-\underline{\alpha}$
(54) $\nabla_{4}\eta_{a}=-\chi\cdot(\eta-\underline{\eta})-\beta$
(55) $\nabla_{3}\underline{\eta}_{a}=-\underline{\chi}\cdot(\underline{\eta}-\eta)+\underline{\beta}$
(56) $\nabla_{4}\underline{\omega}=2\omega\underline{\omega}+\frac{3}{4}|\eta-\underline{\eta}|^{2}-\frac{1}{4}(\eta-\underline{\eta})\cdot(\eta+\underline{\eta})-\frac{1}{8}|\eta+\underline{\eta}|^{2}+\frac{1}{2}\rho$
(57) $\nabla_{3}\omega=2\omega\underline{\omega}+\frac{3}{4}|\eta-\underline{\eta}|^{2}+\frac{1}{4}(\eta-\underline{\eta})\cdot(\eta+\underline{\eta})-\frac{1}{8}|\eta+\underline{\eta}|^{2}+\frac{1}{2}\rho$
(58) $\nabla_{4}tr\underline{\chi}+\frac{1}{2}\mbox{tr}\chi tr\underline{\chi}=2\omega tr\underline{\chi}+2\text{div}\underline{\eta}+2|\underline{\eta}|^{2}_{\gamma}+2\rho-\hat{\chi}\cdot\hat{\underline{\chi}}$
(59) $\nabla_{3}\mbox{tr}\chi+\frac{1}{2}tr\underline{\chi}\mbox{tr}\chi=2\underline{\omega}\mbox{tr}\chi+2\text{div}\eta+2|\eta|^{2}+2\rho-\hat{\chi}\cdot\hat{\underline{\chi}}$
(60) $\nabla_{4}\hat{\underline{\chi}}+\frac{1}{2}\mbox{tr}\chi\hat{\underline{\chi}}=\nabla\hat{\otimes}\underline{\eta}+2\omega\hat{\underline{\chi}}-\frac{1}{2}tr\underline{\chi}\hat{\chi}+\underline{\eta}\hat{\otimes}\underline{\eta}$
(61) $\nabla_{3}\hat{\chi}+\frac{1}{2}tr\underline{\chi}\hat{\chi}=\nabla\hat{\otimes}\eta+2\underline{\omega}\hat{\chi}-\frac{1}{2}\mbox{tr}\chi\hat{\underline{\chi}}+\eta\hat{\otimes}\eta$
(62) $\text{div}\hat{\chi}=\frac{1}{2}\nabla\mbox{tr}\chi-\frac{1}{2}(\eta-\underline{\eta})\cdot(\hat{\chi}-\frac{1}{2}\mbox{tr}\chi\gamma)-\beta$
(63) $\text{div}\hat{\underline{\chi}}=\frac{1}{2}\nabla tr\underline{\chi}-\frac{1}{2}(\underline{\eta}-\eta)\cdot(\hat{\underline{\chi}}-\frac{1}{2}tr\underline{\chi}\gamma)-\underline{\beta}$
(64) $\text{curl}\eta=\hat{\underline{\chi}}\wedge\hat{\chi}+\sigma\epsilon=-\text{curl}\underline{\eta}$
(65) $K-\frac{1}{2}\hat{\chi}\cdot\hat{\underline{\chi}}+\frac{1}{4}\mbox{tr}\chi tr\underline{\chi}=-\rho,$
(66) $\nabla_{4}\zeta=2\nabla\omega+\chi\cdot(\underline{\eta}-\zeta)+2\omega(\zeta+\underline{\eta})-\beta,$
(67) $\nabla_{3}\zeta=-2\nabla\omega-\underline{\chi}\cdot(\zeta+\eta)+2\underline{\omega}(\zeta-\eta)-\underline{\beta}$
The Bianchi equations read in this gauge as follows
(68) $\begin{split}\nabla_{3}\alpha+\frac{1}{2}\mbox{tr}\underline{\chi}\alpha=&\nabla\hat{\otimes}\beta+4\underline{\omega}\alpha-3\left(\hat{\chi}\rho+\prescript{*}{}{\hat{\chi}}\sigma\right)+(\zeta+4\eta)\hat{\otimes}\beta\end{split}$
(69) $\nabla_{4}\beta+2\mbox{tr}\chi\beta=\text{div}\alpha-2\omega\beta+\left(\eta-2\zeta\right)\cdot\alpha,$
(70) $\nabla_{3}\beta+\mbox{tr}\underline{\chi}\beta=\nabla\rho+\prescript{*}{}{\nabla}\sigma+2\underline{\omega}\beta+2\hat{\chi}\cdot\underline{\beta}+3\left(\eta\rho+\prescript{*}{}{\eta}\sigma\right),$
(71) $\nabla_{4}\sigma+\frac{3}{2}\mbox{tr}\chi\sigma=-\text{div}\prescript{*}{}{\beta}+\frac{1}{2}\hskip 1.42262pt\underline{\hat{\chi}}\cdot\prescript{*}{}{\alpha}-(\zeta+2\underline{\eta})\cdot\prescript{*}{}{\beta},$
(72) $\nabla_{3}\sigma+\frac{3}{2}\mbox{tr}\underline{\chi}\sigma=-\text{div}\prescript{*}{}{\underline{\beta}}+\frac{1}{2}\hskip 1.42262pt\hat{\chi}\cdot\prescript{*}{}{\underline{\alpha}}-(\zeta+2\eta)\cdot\prescript{*}{}{\underline{\beta}},$
(73) $\nabla_{4}\rho+\frac{3}{2}\mbox{tr}\chi\rho=\text{\mbox{div }}\beta-\frac{1}{2}\underline{\hat{\chi}}\cdot\alpha+(\zeta+2\underline{\eta})\cdot\beta,$
(74) $\nabla_{3}\rho+\frac{3}{2}\mbox{tr}\underline{\chi}\rho=-\text{div}\underline{\beta}-\frac{1}{2}\hat{\chi}\cdot\underline{\alpha}+(\zeta-2\eta)\cdot\underline{\beta},$
(75) $\nabla_{4}\underline{\beta}+\mbox{tr}\chi\underline{\beta}=-\nabla\rho+\prescript{*}{}{\nabla}\sigma+2\omega\underline{\beta}+2\underline{\hat{\chi}}\cdot\beta-3\left(\underline{\eta}\rho-\prescript{*}{}{\underline{\eta}}\sigma\right),$
(76) $\nabla_{3}\underline{\beta}+2\mbox{tr}\underline{\chi}\underline{\beta}=-\text{div}\underline{\alpha}-2\underline{\omega}\underline{\beta}+\underline{\eta}\cdot\underline{\alpha},$
(77) $\begin{split}\nabla_{4}\underline{\alpha}+\frac{1}{2}\mbox{tr}\chi\hskip 1.42262pt\underline{\alpha}=&-\nabla\hat{\otimes}\underline{\beta}+4\omega\hskip 1.42262pt\underline{\alpha}-3\left(\underline{\hat{\chi}}\rho-\prescript{*}{}{\underline{\hat{\chi}}}\sigma\right)+\left(\zeta-4\underline{\eta}\right)\hat{\otimes}\underline{\beta}.\end{split}$
4.4. Integration
Let $U$ be a coordinate patch on a 22-sphere $S_{u,\underline{u}}$ and let $p_{U}$ be a partition of unity subordinate to $U$. For a function $\phi$, we define its integral on a 22-sphere as well as on the null hypersurfaces $H_{u}$ and $\underline{H}_{\underline{u}}$.
(78) $\int_{S_{u,\underline{u}}}\phi:=\sum_{U}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot p_{U}\cdot\sqrt{\text{det}\gamma}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2},$
(79) $\int_{H_{u}^{(0,\underline{u})}}:=\sum_{U}\int_{0}^{\underline{u}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot 2\hskip 1.42262ptp_{U}\cdot\Omega\cdot\sqrt{\text{det}\gamma}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime},$
(80) $\int_{\underline{H}_{\underline{u}}^{(u_{\infty},u)}}:=\sum_{U}\int_{u_{\infty}}^{u}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot 2\hskip 1.42262ptp_{U}\cdot\Omega\cdot\sqrt{\text{det}g}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2}\hskip 1.42262pt\text{d}u^{\prime},$
For a spacetime region $\mathcal{D}_{u,\underline{u}}:=\begin{Bmatrix}\left(u^{\prime},\underline{u}^{\prime},\theta^{1},\theta^{2}\right)\hskip 1.42262pt\mid\hskip 1.42262ptu_{\infty}\leq u^{\prime}\leq u,0\leq\underline{u}^{\prime}\leq\underline{u}\end{Bmatrix}$, we define the spacetime integral
(81) $\int_{\mathcal{D}_{u,\underline{u}}}\phi:=\sum_{U}\int_{u_{\infty}}^{u}\int_{0}^{\underline{u}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot p_{U}\cdot\Omega^{2}\cdot\sqrt{-\text{det}g}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\hskip 1.42262pt\text{d}u^{\prime}.$
We proceed with the definition of $L^{p}$ norms $(1\leq p<\infty)$ for an arbitrary tensorfield $\phi$:
(82) $\lVert\phi\rVert_{L^{p}(S_{u,\underline{u}})}^{p}:=\int_{S_{u,\underline{u}}}\langle\phi,\phi\rangle_{\gamma}^{\frac{p}{2}}$
(83) $\lVert\phi\rVert_{L^{p}(H_{u}^{(0,\underline{u})})}^{p}:=\int_{H_{u}^{(0,\underline{u})}}\langle\phi,\phi\rangle_{\gamma}^{\frac{p}{2}}$
(84) $\lVert\phi\rVert_{L^{p}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}^{p}:=\int_{\underline{H}_{\underline{u}}^{(u_{\infty},u)}}\langle\phi,\phi\rangle_{\gamma}^{\frac{p}{2}}.$
For the case $p=\infty$, we separately define
(85) $\lVert\phi\rVert_{L^{\infty}(S_{u,\underline{u}})}:=\underset{(\theta^{1},\theta^{2})\in S_{u,\underline{u}}}{\text{sup}}\langle\phi,\phi\rangle_{\gamma}^{\frac{1}{2}}(\theta^{1},\theta^{2}).$
4.5. Signature for decay rates and scale-invariant norms
Perhaps the most challenging aspect of trapped surface formation results, historically, has been the attempt to find initial data that are, in an appropriate sense, large (this is by necessity, as is implied by the monumental work of [19]) but also small enough to allow for an existence result of a spacetime region that gives trapped surfaces the time they would require to form. The first such initial data set, in the absence of symmetries, was given by [17]. Later contributions include [22], [6] and [2]. Moreover, one would have to construct norms which preserve, at least approximately, the hierarchy present in the initial data upon evolution of the Einstein equations. The signature for decay rates, which was first introduced in [3], is the tool we will use in the present paper to build scale-invariant norms. These will be norms that, upon evolution of the initial data, remain bounded above by a uniform constant (with the exception of a few anomalous terms). For another application of this framework, see [4].
To each $\phi\in\begin{Bmatrix}\alpha,\underline{\alpha},\beta^{R},\underline{\beta}^{R},\rho,\sigma,\eta,\underline{\eta},\chi,\underline{\chi},\omega,\underline{\omega},\zeta,\gamma\end{Bmatrix}$ we associate its signature for decay rates $s_{2}(\phi)$:
$s_{2}(\phi)=0\cdot N_{4}(\phi)+\frac{1}{2}N_{A}(\phi)+1\cdot N_{3}(\phi)-1.$
Here $N_{\alpha}(\phi)$ $(\alpha=1,2,3,4)$ denotes the number of times $e_{\alpha}$ appears in the definition of $\phi$. We get the following tables of signatures:
$\phi$ $\alpha$ $\underline{\alpha}$ $\beta$ $\underline{\beta}$ $\rho$ $\sigma$ $\eta$ $\underline{\eta}$ $\chi$ $\underline{\chi}$ $\omega$ $\underline{\omega}$ $\zeta$ $\gamma$
$s_{2}(\phi)$ 0 2 0.5 1.5 1 1 0.5 0.5 0 1 0 1 0.5 0
Several properties of $s_{2}$ follow:
$s_{2}(\nabla_{4}\phi)=s_{2}(\phi),\hskip 5.69054pt$
$s_{2}(\nabla\phi)=s_{2}(\phi)+\frac{1}{2},\hskip 5.69054pt$
$s_{2}(\nabla_{3}\phi)=s_{2}(\phi)+1,\hskip 5.69054pt$
Finally, perhaps the most important property of $s_{2}$ is signature conservation:
(86) $s_{2}(\phi_{1}\cdot\phi_{2})=s_{2}(\phi_{1})+s_{2}(\phi_{2}),\hskip 5.69054pt.$
This allows for the (almost)-preservation of the scale-invariant norms upon evolution, as we shall see.
For any horizontal tensor-field $\phi$ and Yang-Mills component $\mathcal{Y}$, we define the following norms:
(87) $\lVert{\phi}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}:=a^{-s_{2}(\phi)}\lvert u\rvert^{2s_{2}(\phi)+1}\lVert{\phi}\rVert_{L^{\infty}(S_{u,\underline{u}})},$
(88) $\lVert{\phi}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}:=a^{-s_{2}(\phi)}\lvert u\rvert^{2s_{2}(\phi)}\lVert{\phi}\rVert_{L^{2}(S_{u,\underline{u}})},$
(89) $\lVert{\phi}\rVert_{L^{1}_{(sc)}(S_{u,\underline{u}})}:=a^{-s_{2}(\phi)}\lvert u\rvert^{2s_{2}(\phi)-1}\lVert{\phi}\rVert_{L^{1}(S_{u,\underline{u}})},$
Notice the difference in the $u$-weights amongst the definitions.
[…]
(90) $\lVert{\mathrm{\Upsilon}_{1}\cdot\mathrm{\Upsilon}_{2}}\rVert_{L^{1}_{(sc)}(S_{u,\underline{u}})}\leq\frac{1}{\lvert u\rvert}\lVert{\mathrm{\Upsilon}_{1}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lVert{\mathrm{\Upsilon}_{2}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},$
(91) $\lVert{\mathrm{\Upsilon}_{1}\cdot\mathrm{\Upsilon}_{2}}\rVert_{L^{1}_{(sc)}(S_{u,\underline{u}})}\leq\frac{1}{\lvert u\rvert}\lVert{\mathrm{\Upsilon}_{1}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lVert{\mathrm{\Upsilon}_{2}}\rVert_{L^{1}_{(sc)}(S_{u,\underline{u}})},$
(92) $\lVert{\mathrm{\Upsilon}_{1}\cdot\mathrm{\Upsilon}_{2}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\leq\frac{1}{\lvert u\rvert}\lVert{\mathrm{\Upsilon}_{1}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lVert{\mathrm{\Upsilon}_{2}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}.$
Notice that this is possible partly thanks to the signature conservation property (86). In the region of study, the factor $\frac{1}{\lvert u\rvert}$ plays the role of measuring the smallness of the nonlinear terms. The above inequalities are the primary tools that will be used to close the bootstrap argument required for the existence part.
4.6. Norms
Let $N\geq 3$ be a natural number. Let $\psi_{g}\in\begin{Bmatrix}\omega,\mbox{tr}\chi,\hat{\chi},\underline{\omega},\eta,\underline{\eta}\end{Bmatrix}$, $\Psi_{u}\in\begin{Bmatrix}\beta,\rho,\sigma,\underline{\beta}\end{Bmatrix}$ and $\Psi_{\underline{u}}\in\begin{Bmatrix}a^{\frac{1}{2}}\rho,a^{\frac{1}{2}}\sigma,\underline{\beta},\underline{\alpha}\end{Bmatrix}$. Moreover, we will sometimes use $\Psi$ to denote an arbitrary $\Psi_{u}$ or a $\Psi_{\underline{u}}$. Also, define $\widetilde{\mbox{tr}\underline{\chi}}:=\mbox{tr}\underline{\chi}+\frac{2}{\lvert u\rvert}$. For $0\leq i\leq N$, we define
(93) $\Gamma_{i,\infty}(u,\underline{u}):=$ $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\psi_{g}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}$
(94) $+\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})},$
$\mathcal{R}_{i,\infty}(u,\underline{u}):=\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{\underline{u}}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})},$
Furthermore, for $0\leq i\leq N+4$ and $0\leq j\leq N+4$, we define
(95) $\begin{split}\Gamma_{j,2}(u,\underline{u}):=&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ &+\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},\end{split}$
(96) $\mathcal{R}_{i,2}(u,\underline{u}):=\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{\underline{u}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},$
Finally, for $0\leq i\leq N+4$, we define the norms along the null hypersurfaces:
(97) $\mathcal{R}_{i}(u,\underline{u}):=\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})},$
(98) $\underline{\mathcal{R}}_{i}(u,\underline{u}):=\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta^{R}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{\underline{u}}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}$
4.7. Commutation Formulae
Use the definition of the covariant derivatives and project it onto the topological 2-sphere $S_{u,\underline{u}}$ to yield
$[\nabla_{4},\nabla_{B}]\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{n}}=[D_{4},D_{B}]\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{n}}+(\nabla_{B}\log\Omega)\nabla_{4}\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{n}}$
$-\gamma^{CD}\chi_{BD}\nabla_{C}\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{n}}-\sum_{i=1}^{n}\gamma^{CD}\chi_{BD}\underline{\eta}_{A_{i}}\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\hat{A}_{i}C\cdot\cdot A_{n}}$
$+\sum_{i=1}^{n}\gamma^{CD}\chi_{A_{i}B}\underline{\eta}_{D}\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\hat{A}_{i}C\cdot\cdot A_{n}}$
$[D_{4},D_{A}]\mathcal{G}^{P}~_{QA_{1}A_{2}\cdot\cdot\cdot\cdot A_{n}}=-\sum_{i}R(e_{C},e_{A_{i}},e_{4},e_{A})\mathcal{G}^{P}~_{QA_{1}\cdot\cdot\hat{A}_{i},\cdot\cdot A_{n}}+(\nabla_{A}\log\Omega)\nabla_{4}\mathcal{G}^{P}~_{QA_{1}A_{2}\cdot\cdot\cdot\cdot A_{n}}.$
Notice that the last term is redundant since it already appears in the previous expression. We need to take care of the curvature terms.
(99) $[\nabla_{4},\nabla_{A}]\mathcal{G}\sim\beta\mathcal{G}+(\eta+\underline{\eta})\nabla_{4}\mathcal{G}-\chi\nabla\mathcal{G}+\chi\underline{\eta}\mathcal{G}.$
For higher order commutation, we have the following lemma:
Lemma 4.1.
Suppose $\mathcal{G}$ is a section of the product vector bundle $^{k}\otimes T^{*}\mathbb{S}^{2}$, $k\geq 1$, that satisfies $\nabla_{4}\mathcal{G}=\mathcal{F}_{1}$ and $\nabla_{4}\nabla^{I}\mathcal{G}=\mathcal{F}^{I}_{1}$, then $\mathcal{F}^{I}_{1}$ verifies the following schematic expression:
(100) $\begin{split}\mathcal{F}^{I}_{1}\sim&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I-1}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\beta\nabla^{J_{4}}\mathcal{G}\\ &+\sum_{J_{1}+J_{2}+J_{3}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\mathcal{F}_{1}\\ +&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\hat{\nabla}^{J_{3}}\chi\nabla^{J_{4}}\mathcal{G}.\end{split}$
Similarly, for $\nabla_{3}\mathcal{G}=\mathcal{F}_{2}$, and $\nabla_{3}\nabla^{I}\mathcal{G}=\mathcal{F}^{I}_{2}$,
(101) $\begin{split}\mathcal{F}^{I}_{2}\sim&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I-1}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\underline{\beta}\nabla^{J_{4}}\mathcal{G}\\ +&\sum_{J_{1}+J_{2}+J_{3}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\mathcal{F}_{2}\\ +&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\hat{\nabla}^{J_{3}}\underline{\chi}\nabla^{J_{4}}\mathcal{G}.\end{split}$
Proof.
For $I=1$, this identity is clearly satisfied due to the calculations above. Assume it holds for J=I1J=I-1 and show that it holds for $J=I$. We omit the proof and refer to [4]. ∎
Remark 2.
By moving the top derivatives of $\mathcal{G}$ multiplied by $\mbox{tr}\underline{\chi}$ from the right-hand side to the left-hand side, one may also obtain
$\begin{split}\mathcal{F}^{I}_{2}+\frac{I}{2}\mbox{tr}\underline{\chi}\hat{\nabla}^{I}\mathcal{G}\sim&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I-1}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\beta\nabla^{J_{4}}\mathcal{G}\\ +&\sum_{J_{1}+J_{2}+J_{3}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\mathcal{F}_{2}\\ +&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\hat{\nabla}^{J_{3}}\widehat{\underline{\chi}}\nabla^{J_{4}}\mathcal{G}\\ +&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I-1}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}+1}\hat{\nabla}^{J_{3}}\mbox{tr}\underline{\chi}\nabla^{J_{4}}\mathcal{G}.\end{split}$
5. Preliminary estimates
5.1. Preliminary bootstrap assumptions
We will be employing a bootstrap argument to obtain a priori bounds on $\Gamma$ and $\mathcal{R}$. Along the initial hypersurfaces $H_{u_{\infty}}$ and $\underline{H}_{0}$, an analysis of the initial data using transport equations (see, for example, [6]) yields
$\Gamma_{0}+\mathcal{R}_{0}\lesssim\mathcal{I}.$
Our goal is to show that in the entire region
$\mathcal{D}:=\begin{Bmatrix}(u,\underline{u},\theta^{1},\theta^{2})\hskip 2.84526pt\mid\hskip 2.84526ptu_{\infty}\leq u\leq-a,\hskip 1.42262pt0\leq\underline{u}\leq 1\end{Bmatrix}$
there exists a constant $c(\mathcal{I})=\mathcal{I}^{4}+\mathcal{I}^{2}+\mathcal{I}+1$ such that
$\Gamma+\mathcal{R}\lesssim c(\mathcal{I}).$
We assume, as a bootstrap assumption, the following:
(102) $\Gamma\leq\Gamma,\hskip 5.69054pt\mathcal{R}\leq R,$
where $\Gamma$ and $\hskip 1.42262ptR$ are large so that
$\mathcal{I}^{4}+\mathcal{I}^{2}+\mathcal{I}+1\ll\text{min}\begin{Bmatrix}\Gamma,\hskip 1.42262ptR,\hskip 1.42262ptM\end{Bmatrix}$
but also such that
$(\Gamma+R)^{20}\leq a^{\frac{1}{16}}.$
5.2. Estimates on the metric components
For the metric component $\Omega$, the induced metric $\gamma$ of $S_{u,\underline{u}}$ and for the area of $S_{u,\underline{u}}$ the following propositions hold:
Proposition 5.1.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), we have
$\lVert\Omega-1\rVert_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{\Gamma}{\lvert u\rvert}$.
Proposition 5.2.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there exist two constants $c$ and $C$ depending only on the initial data such that the bounds
$c\leq\text{det}\hskip 1.42262pt\gamma\leq C.$
and
$\lvert\gamma_{AB}\rvert+\lvert\gamma^{-1}_{AB}\rvert\leq C$
hold throughout the slab of existence $\mathcal{D}$.
Proposition 5.3.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), fix a point $(u,\theta)$ on the initial hypersurface $\underline{H}_{0}$. Let $\Lambda(u)$ and $\lambda(u)$ be the largest and smallest eigenvalues of $\gamma^{-1}(u,0,\theta)\hskip 1.42262pt\gamma(u,\underline{u},\theta)$ respectively, along the outgoing null geodesics emanating from $(u,\theta)$. There holds
$\lvert\Lambda(u)-1\rvert+\lvert\lambda(u)-1\rvert\lesssim\frac{1}{a^{\frac{1}{2}}}.$
Proposition 5.4.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), for the area of the 22-sphere $S_{u,\underline{u}}$, there holds
$\sup_{\underline{u}}\lvert\text{Area}(S_{u,\underline{u}})-\text{Area}(S_{u,0})\rvert\lesssim\frac{\Gamma^{\frac{1}{2}}}{a^{\frac{1}{2}}}\lvert u\rvert^{2}.$
The proofs of Propositions 5.1 to 5.4 are the same as in [1]
5.3. Estimates for transport equations
We shall be using two fundamental bounds on transport equations throughout this work.
Proposition 5.5.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), the following hold for an arbitrary $\mathcal{G}\in\Gamma(^{N}\otimes T^{*}S):$
(103) $\lVert{\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\lVert\mathcal{G}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime\prime}})}+\int_{\underline{u}^{\prime\prime}}^{\underline{u}}\lVert{\nabla_{3}\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
(104) $\lVert{\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\lVert\mathcal{G}\rVert_{L^{2}_{(sc)}(S_{u^{\prime\prime},\underline{u}})}+\int_{u^{\prime\prime}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla_{4}\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
There are, however, cases that are borderline and require more delicate control than what the above Proposition provides. These have to do with components $X$ satisfying an equation of the form $\nabla_{3}X=-\lambda\mbox{tr}\underline{\chi}X+\dots$, where $\lambda>0$. Keeping in mind that $\mbox{tr}\underline{\chi}$ is the worst Ricci coefficient in terms of peeling, one would hope to be able to get rid of its appearance and thus obtain stronger bounds regarding the peeling properties of $X$. The following weighted transport inequality achieves this.
Proposition 5.6.
Let $\mathcal{G},\mathcal{H}\in\Gamma(^{N}\otimes T^{*}S)$ and assume that the following equation holds:
$\nabla_{3}\mathcal{G}+\lambda_{0}\hskip 1.42262pt\mbox{tr}\underline{\chi}\hskip 1.42262pt\mathcal{G}=\mathcal{H}.$
Then, under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), the following is true:
$\lvert u\rvert^{\lambda_{1}}\lVert{\mathcal{G}}\rVert_{L^{2}(S_{u,\underline{u}})}\lesssim\lvert u_{\infty}\rvert^{\lambda_{1}}\lVert\mathcal{G}\rVert_{L^{2}(S_{u_{\infty},\underline{u}})}+\int_{u_{\infty}}^{u}\lvert u^{\prime}\rvert^{\lambda_{1}}\lVert\mathcal{H}\rVert_{L^{2}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
for $\lambda_{1}=2\lambda_{0}-1$.
Proof.
The variation of area formula for a scalar function $f$ reads:
(105) $\underline{L}\int_{S_{u,\underline{u}}}f=\int_{S_{u,\underline{u}}}\underline{L}f+\Omega\hskip 1.42262pt\mbox{tr}\underline{\chi}\hskip 1.42262ptf=\int_{S_{u,\underline{u}}}\Omega\hskip 1.42262pt\left(e_{3}(f)+\mbox{tr}\underline{\chi}\hskip 1.42262ptf\right).$
Plugging in $f=\lvert u\rvert^{2\lambda_{1}}\lvert\mathcal{G}\rvert_{\gamma}^{2}$, we calculate:
(106) $\begin{split}&\underline{L}\int_{S_{u,\underline{u}}}\lvert u\rvert^{2\lambda_{1}}\lvert\mathcal{G}\rvert_{\gamma}^{2}\\ =&\int_{S_{u,\underline{u}}}\Omega\left(-2\hskip 1.42262pt\lambda_{1}\lvert u\rvert^{2\hskip 1.42262pt\lambda_{1}-1}e_{3}(u)\lvert\mathcal{G}\rvert_{\gamma}^{2}+2\lvert u\rvert^{2\hskip 1.42262pt\lambda_{1}}\langle\mathcal{G},\nabla_{3}\mathcal{G}\rangle_{\gamma}+\mbox{tr}\underline{\chi}\hskip 1.42262pt\lvert u\rvert^{2\lambda_{1}}\hskip 1.42262pt\lvert\mathcal{G}\rvert_{\gamma}^{2}\right)\\ =&\int_{S_{u,\underline{u}}}\Omega\left(2\lvert u\rvert^{2\lambda_{1}}\langle\mathcal{G},\nabla_{3}\mathcal{G}+\lambda_{0}\mbox{tr}\underline{\chi}\mathcal{G}\rangle_{\gamma,\epsilon}\right)+\int_{S_{u,\underline{u}}}\Omega\lvert u\rvert^{2\lambda_{1}}\left(\frac{-2\lambda_{1}e_{3}(u)}{\lvert u\rvert}+(1-2\lambda_{0})\hskip 1.42262pt\mbox{tr}\underline{\chi}\right)\lvert\mathcal{G}\rvert_{\gamma}^{2}.\end{split}$
Notice that
(107) $\begin{split}&\frac{-2\lambda_{1}e_{3}(u)}{\lvert u\rvert}+(1-2\lambda_{0})\hskip 1.42262pt\mbox{tr}\underline{\chi}\\ =&\frac{-2\lambda_{1}(\Omega^{-1}-1)}{\lvert u\rvert}+(1-2\lambda_{0})(\mbox{tr}\underline{\chi}+\frac{2}{\lvert u\rvert})-\frac{2\lambda_{1}+2-4\lambda_{0}}{\lvert u\rvert}\\ \leq&\frac{\Gamma}{\lvert u\rvert^{2}},\end{split}$
where we have used the bootstrap assumption $\lVert{\mbox{tr}\underline{\chi}+\frac{2}{\lvert u\rvert}}\rVert_{L^{\infty}(S_{u,\underline{u}})}\leq\frac{\Gamma}{\lvert u\rvert^{2}}$ and the definition of $\lambda_{1}$. For the first term in the last line of (106) we then use Cauchy-Schwartz and for the second we apply Grönwall’s inequality to get:
(108) $\begin{split}&\lvert u\rvert^{2\lambda_{1}}\lVert{\mathcal{G}}\rVert_{L^{2}(S_{u,\underline{u}})}\\ \lesssim&\mathrm{e}^{\Gamma\lVert u^{-2}\rVert_{L_{u}^{1}}}\left(\lvert u_{\infty}\rvert^{\lambda_{1}}\lVert\mathcal{G}\rVert_{L^{2}(S_{u_{\infty},\underline{u}})}+\int_{u_{\infty}}^{u}\lvert u^{\prime}\rvert^{\lambda_{1}}\lVert\mathcal{H}\rVert_{L^{2}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\right)\\ \lesssim&\lvert u_{\infty}\rvert^{\lambda_{1}}\lVert\mathcal{G}\rVert_{L^{2}(S_{u_{\infty},\underline{u}})}+\int_{u_{\infty}}^{u}\lvert u^{\prime}\rvert^{\lambda_{1}}\lVert\mathcal{H}\rVert_{L^{2}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime},\end{split}$
where we have used the fact that $\mathrm{e}^{\Gamma\lVert u^{-2}\rVert_{L_{u}^{1}}}\lesssim\mathrm{e}^{\epsilon\Gamma/a}\lesssim 1$.
5.4. Sobolev embedding
With the derived estimates for the metric $\gamma$, we can obtain a bound on the isoperimetric constant for a topological 22-sphere $S$:
(109) $I(S):=\sup_{U\subset S,~\partial U\in C^{1}}\frac{\min\left\{Area(U),Area(U^{c})\right\}}{[Perimeter(\partial U)]^{2}}.$
The following proposition yields an upper bound for $I(S)$.
Proposition 5.7.
Under the assumption on the initial data and the bootstrap assumption (2.10), the isoperimetric constant obeys the following bound
(110) $I(S_{u,\underline{u}})\leq\frac{1}{\pi}$
for $u\in[u_{\infty},-a]$ and $\underline{u}\in[0,\epsilon]$.
Proof.
Fix a $u$. For $U_{\underline{u}}$ a subset of $S_{u,\underline{u}}$, denote by $U_{0}\subset S_{u,0}$ the backward image of $U_{\underline{u}}$ under the diffeomorphism generated by the equivariant vector field $L.$ Using Propositions 5.2 and 5.3 and their proof, we can obtain the estimates
$\frac{\text{Perimeter}(\partial U_{\underline{u}})}{\text{Perimeter}(\partial U_{0})}\geq\sqrt{\inf_{S_{u,0}}\lambda(\underline{u})}$
and
$\frac{\text{Area}(U_{\underline{u}})}{\text{Area}(U_{0})}\leq\sup_{S_{u,0}}\frac{\det(\gamma_{\underline{u}})}{\det(\gamma_{0})},\hskip 5.69054pt\frac{\text{Area}(U_{\underline{u}}^{c})}{\text{Area}(U_{0}^{c})}\leq\sup_{S_{u,0}}\frac{\det(\gamma_{\underline{u}})}{\det(\gamma_{0})}.$
The conclusion then follows from the fact that $I(S_{u,0})=\frac{1}{2\pi}$ and the bounds from Propositions 5.2, 5.3 . ∎
Throughout this work, we will be using an $L^{2}-L^{\infty}$ Sobolev estimate. To obtain it, utilizing the basic estimates above, we may proceed to write down the following gauge-invariant Sobolev inequalities for the topological 22- sphere $S$.
Proposition 5.8.
Let $(S,\gamma)$ be a Riemannian 22-manifold with the isoperimetric constant $I(S)$. Then the following Sobolev embedding holds for any $\mathcal{G}\in\Gamma(^{N}\otimes T^{*}S)$
[…]
for any $p\in(2,\infty)$.
Proof.
We know that the standard Sobolev inequality with $p\in(2,\infty)$
[…]
holds for a function $f$. Now we set
(113) $f=\sqrt{\mathcal{G}_{A_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{N}}\mathcal{G}~_{B_{1}B_{2}B_{3}\cdot\cdot\cdot\cdot B_{N}}\gamma^{A_{1}B_{1}}\gamma^{A_{2}B_{2}}\gamma^{A_{3}B_{3}}\cdot\cdot\cdot\cdot\gamma^{A_{N}B_{N}}+\delta}$
with $\delta>0$ to yield
[…]
since $f$ is a function and the covariant derivative is compatible with the metric $\gamma$. Substituting this inequality in (112) and taking the limit $\delta\to 0$ yields the desired gauge-invariant Sobolev inequality. ∎
Proposition 5.9.
Let $(S,\gamma)$ be a Riemannian 22-manifold with the isoperimetric constant $I(S)$. Then the following Sobolev embedding holds for any $\mathcal{G}\in\Gamma(^{N}\otimes T^{*}S)$
[…]
for any $p\in(2,\infty)$.
Proof.
A calculation similar to the previous one and the standard $L^{\infty}-L^{p}$ Sobolev inequality on the Riemannian manifold $(S,\gamma)$ yield the result. ∎
The two inequalities above, together with Propositions 5.1-5.4, allow us to control the $L^{2}$-norm of $\mathcal{G}$ in terms of its $H^{2}$-norm. Following the area estimates, we have $Area(S_{u,\underline{u}})\approx u^{2}$. Therefore, we obtain the following important inequality.
Proposition 5.10.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), the following Sobolev embedding holds for any $\mathcal{G}\in\Gamma(^{N}\otimes T^{*}S_{u,\underline{u}}):$
[…]
which, in scale-invariant norms, reads
[…]
Proof.
Substitute $p=4$ in the previous proposition (114) and estimate the right hand side by means of (111), the estimate on the area given by Proposition 5.4 and on the isoperimetric constant of $S_{u,\underline{u}}$ given by Proposition 110. ∎
5.5. Estimates on the Ricci coefficients
Proposition 5.11.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there holds
$\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\omega}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}\left(1+\hskip 1.42262pt\underline{\mathcal{R}}[\rho]\right).$
Proof.
We begin by recalling that $\omega$ satisfies the schematic equation
$\nabla_{3}\omega=\frac{1}{2}\rho+\psi_{g}\psi_{g}.$
Using the commutation formula 2 and the notation of Section 4.6, we have, for a general $i$:
(117) $\begin{split}\nabla_{3}\nabla^{i}\omega+\frac{i}{2}\mbox{tr}\underline{\chi}\nabla^{i}\omega=&\nabla^{i}\rho+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega\end{split}$
Note that since $\omega$ is not a section of the vector bundle $^{k}\otimes T^{*}S\otimes P_{Ad,\mathfrak{g}}$, the last term on the right-hand side of 2 does not appear. Passing to scale-invariant norms, we get
$\lVert{(a^{\frac{1}{2}}\nabla)^{i}\omega}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}$
$\lesssim$ $\lVert(a^{\frac{1}{2}}\nabla)^{i}\omega\rVert_{L^{2}_{(sc)}(S_{u_{\infty}},0)}+\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\rho}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}=i-1}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}.$
For $0\leq i\leq N+4$, the first term, by virtue of the fact that $\Omega\equiv 1$ initially, vanishes. The second term can be bounded, using Hölder’s inequality, by $\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}\underline{\mathcal{R}}[\rho].$ The third, fourth and fifth terms can be bounded above by
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\leq\frac{a\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert^{2}}\leq\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}.$
The sixth term is controlled as follows:
(118) $\begin{split}&\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{4}}}\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{3}{4}}\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert}\lesssim\frac{a^{\frac{1}{4}}}{\lvert u\rvert^{\frac{1}{2}}}.\end{split}$
The seventh term is controlled as follows:
$\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(119) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\lvert u^{\prime}\rvert^{2}}{a}\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{\Gamma^{3}}{\lvert u\rvert}\lesssim\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}\lesssim 1.$
For the eighth and most borderline term, we estimate:
$\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$=$ $\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\frac{a^{\frac{3}{2}}}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(120) $\lesssim$ $\int_{u_{\infty}}^{u}\int_{u_{\infty}}^{u}\frac{a^{2}}{|u^{{}^{\prime}}|^{2}}\frac{1}{|u^{{}^{\prime}}|^{2}}\frac{|u^{{}^{\prime}}|}{a^{\frac{1}{4}}}\Gamma^{3}du^{{}^{\prime}}\lesssim\frac{a^{\frac{7}{4}}\Gamma^{3}}{|u|^{2}}\lesssim 1$
Proposition 5.12.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there hold
$\sum_{i\leq N+4}\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1,\hskip 5.69054pt\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\alpha]+1.$
Proof.
The proof is exactly similar to that of [11]. ∎
The estimates for $\underline{\omega}$ are, in a sense, dual to those for $\omega$.
Proposition 5.13.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there holds
$\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\leq\mathcal{R}[\rho]+1.$
Proof.
We have, schematically,
$\nabla_{4}\underline{\omega}=\frac{1}{2}\rho+\psi_{g}\psi_{g}.$
As before, the schematic product of the Yang-Mills components is gauge-invariant. Using the commutation formula 4.1, we have, for a general $i$:
(121) $\begin{split}\nabla_{4}\nabla^{i}\underline{\omega}=&\nabla^{i}\rho+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}\\ &++\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}\end{split}$
Passing to scale-invariant norms, we have,
(122) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ \lesssim&\lVert(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}\rVert_{L^{2}_{(sc)}(S_{u,0})}+\int_{0}^{\underline{u}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\rho}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
For $0\leq i\leq N+4$, the first four terms are controlled as in Proposition 5.11 and are bounded above by $\mathcal{R}[\rho]+1$. For the next terms, there holds
(123) $\begin{split}&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{0}^{\underline{u}}a^{\frac{1}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\frac{a^{\frac{1}{2}}\Gamma^{2}}{\lvert u\rvert}\lesssim 1.\end{split}$
Working similarly, there hold
(124) $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{a\hskip 1.42262pt\Gamma^{3}}{\lvert u\rvert^{2}}\lesssim 1.$
The claim follows. ∎
We move on to estimates for $\eta$.
Proposition 5.14.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there holds
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\eta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\beta]+1.$
Proof.
We begin with the schematic structure equation for $\eta$:
$\nabla_{4}\eta=\beta^{R}+(\hat{\chi},\mbox{tr}\chi)\cdot(\eta,\underline{\eta}),$
with the product of Yang-Mills curvature terms being once again gauge-invariant. Using the commutation formula 4.1 for the $\nabla_{4}$-direction, we have
(125) $\begin{split}\nabla_{4}\nabla^{i}\eta=&\nabla^{i}\beta+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\beta^{R}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}(\eta,\underline{\eta})\end{split}$
Estimating in scale-invariant norms, we have
(126) $\begin{split}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\eta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}&\lesssim\int_{0}^{\underline{u}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\beta^{R}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}(\eta,\underline{\eta})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
For $0\leq i\leq N+4$, the first term is bounded by $\mathcal{R}[\beta]$. The second term is bounded by
(127) $\begin{split}\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\beta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\Gamma^{2}}{\lvert u\rvert}\lesssim 1.\end{split}$
Notice that, since $i_{3}\leq i-1\leq N+3$, we can bound $i_{3}$ derivatives of $\beta$ using the bootstrap assumption (102) on the total norm $\Gamma$. For the third term, there holds
$\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}(\eta,\underline{\eta})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert}\lesssim 1.$
The result follows. ∎
Proposition 5.15.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there holds
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\chi}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\alpha]+1.$
Proof.
We begin by recalling the schematic equation
$\nabla_{4}\mbox{tr}\chi=\lvert\hat{\chi}\rvert^{2}+\psi_{g}\psi_{g}.$
Commuting with $i$ angular derivatives using 4.1, we obtain
(128) $\begin{split}\nabla_{4}\nabla^{i}\mbox{tr}\chi=&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\psi_{g}.\end{split}$
Passing to scale-invariant norms, we have
(129) $\begin{split}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\chi}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
From the first term, the most dangerous case is when $i_{4}$ falls on $\hat{\chi}$, so we only give details for that. We distinguish two cases:
  • If in the term $\nabla^{i_{1}}\psi_{g}^{i_{2}}$ there exists some $\psi_{g}$ whose derivative is of order >N+3>N+3, we bound that term in $L^{2}_{(sc)}$ and the rest of the terms in $L^{\infty}_{(sc)}$. Notice, crucially, that from Remark LABEL:remarkhat, we can bound $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{\frac{1}{2}}$ for small $i$. As a consequence, we have the bound
    $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}^{2}}{\lvert u\rvert}\lesssim 1.$
  • Otherwise, in the expression $\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}\hat{\chi}$, at most one index $i_{3},i_{4}$ is greater than $N+1$ (in which case, we cannot bound that term in $L^{\infty}_{(sc)}$). Say without loss of generality, that $i_{3}>N+1$. We bound $(a^{\frac{1}{2}})^{i_{3}-1}\nabla^{i_{3}}\hat{\chi}$ in $L^{2}_{(sc)}$ above by $\mathcal{R}[\alpha]+1$ and the rest of the terms in $L^{\infty}_{(sc)}$ above by $1$ (using the improvement mentioned in Remark LABEL:remarkhat), whence
    $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\left(\mathcal{R}[\alpha]+1\right)\cdot 1.$
The second term is handled as in the previous propositions.
Proposition 5.16.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), the following estimates hold:
$\frac{a}{\lvert u\rvert}\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1,\hskip 5.69054pt\frac{a}{|u|^{2}}\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1.$
Proof.
Notice that $\mbox{tr}\underline{\chi}$ satisfies the following structure equation:
$\nabla_{3}\mbox{tr}\underline{\chi}+\frac{1}{2}(\mbox{tr}\underline{\chi})^{2}=-\lvert\underline{\hat{\chi}}\rvert^{2}-2\underline{\omega}\mbox{tr}\underline{\chi}.$
Commuting with $i$ angular derivatives using 2, we get
(130) $\begin{split}\nabla_{3}\nabla^{i}\mbox{tr}\underline{\chi}+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\mbox{tr}\underline{\chi}=&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\underline{\hat{\chi}}\nabla^{i_{4}}\underline{\hat{\chi}}\\ +&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\underline{\omega}\nabla^{i_{4}}\mbox{tr}\underline{\chi}\\ +&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\mbox{tr}\underline{\chi}\\ +&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\mbox{tr}\underline{\chi}\\ +&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\mbox{tr}\underline{\chi}:=G_{i}.\end{split}$
Passing to scale invariant norms and using the weighted transport inequality from Proposition 5.6, we obtain
(131) $\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\hskip 1.42262pt$ $\frac{a}{\lvert u_{\infty}\rvert^{2}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}})}$
(132) $+\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}G_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}.$
We focus on $0\leq i\leq N+4$. For the first term in $G_{i}$, there holds
(133) $\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\underline{\hat{\chi}}\nabla^{i_{4}}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\cdot\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim 1.$
The second and third terms are handled in the same way. For the fourth term, there holds
(134) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\cdot\frac{\lvert u^{\prime}\rvert^{2}}{a}\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{2}}}\cdot\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert}\lesssim 1.\end{split}$
For the fifth term, we bound
(135) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\cdot\frac{\lvert u^{\prime}\rvert^{4}}{a^{2}}\cdot\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{\Gamma^{3}}{\lvert u\rvert}\lesssim 1.\end{split}$
For the sixth term, we can bound
(136) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ =&\int_{u_{\infty}}^{u}\frac{a^{\frac{5}{2}}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i-1}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}+1=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a^{\frac{5}{2}}}{\lvert u^{\prime}\rvert^{4}}\cdot\frac{\lvert u^{\prime}\rvert^{4}}{a^{2}}\cdot\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\hskip 1.42262pt\Gamma^{3}}{\lvert u\rvert}\lesssim 1.\end{split}$
Crucially, this implies that
(137) $\frac{a}{\lvert u\rvert^{2}}\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{I}+1,$
by the Sobolev embedding. This will prove useful in the estimates for $\widetilde{\mbox{tr}\underline{\chi}}$. ∎
We conclude this section with the corresponding estimate on $\underline{\eta}$ and its angular derivatives.
Proposition 5.17.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there holds
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\beta]+\underline{\mathcal{R}}[\underline{\beta}]+1.$
Proof.
The schematic equation for $\underline{\eta}$ is as follows:
(138) $\nabla_{3}\underline{\eta}+\frac{1}{2}\mbox{tr}\underline{\chi}\hskip 1.42262pt\underline{\eta}=\underline{\beta}+\mbox{tr}\underline{\chi}\eta+\psi_{g}\underline{\hat{\chi}}$.
$\nabla_{3}\nabla^{i}\underline{\eta}+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\underline{\eta}=$ $\nabla^{i}\underline{\beta}+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\underline{\beta}$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\underline{\eta}$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\underline{\eta}$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\underline{\eta}:=H_{i}.$
Calculating in scale-invariant norms, we arrive at
(139) $\frac{1}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{1}{\lvert u_{\infty}\rvert}\lVert(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}})}+\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}H_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}.$
We restrict attention to $0\leq i\leq N+4$. For the first term, there holds
(140) $\frac{1}{\lvert u_{\infty}\rvert}\lVert(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}})}\lesssim\frac{\mathcal{I}}{\lvert u_{\infty}\rvert}\lesssim\frac{\mathcal{I}}{\lvert u\rvert}$.
There holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\beta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$\lesssim$ $\left(\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\beta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\right)^{\frac{1}{2}}\left(\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{4}}\hskip 1.42262pt\text{d}u^{\prime}\right)^{\frac{1}{2}}$
(141) $\lesssim$ $\underline{\mathcal{R}}[\underline{\beta}]\cdot\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{3}{2}}}\lesssim\frac{\underline{\mathcal{R}}[\underline{\beta}]}{\lvert u\rvert}$.
For the next term, there holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\underline{\beta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(142) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\cdot\frac{a^{\frac{1}{2}}\hskip 1.42262pt\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{1}{\lvert u\rvert}$.
For the third term, there holds
(143) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{\mathcal{R}[\beta]+1}{\lvert u\rvert}$.
This is done by further taking into account that the schematic product appearing is actually $\psi_{g}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})=\underline{\hat{\chi}}\cdot\psi_{g}+\mbox{tr}\underline{\chi}\eta$. As such, we use the improvement obtained in Proposition 5.16 for the $\mbox{tr}\underline{\chi}$-term as well as the improvement obtained in Proposition 5.14 for $\eta$. The term $\underline{\hat{\chi}}$ is less anomalous than $\mbox{tr}\underline{\chi}$ and hence the above bound is easier to obtain. Continuing the estimates, it holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(144) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\cdot\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{2}}}\cdot\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{1}{\lvert u\rvert}$.
For the sixth term, there holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(145) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\cdot\frac{\lvert u^{\prime}\rvert^{2}}{a}\cdot\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{1}{\lvert u\rvert}$.
For the seventh term, there holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(146) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a^{\frac{3}{2}}}{\lvert u^{\prime}\rvert^{3}}\cdot\frac{\lvert u^{\prime}\rvert^{2}}{a}\cdot\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\hskip 1.42262pt\Gamma^{3}}{\lvert u\rvert^{2}}\lesssim\frac{1}{\lvert u\rvert}$.
Putting everything together, we have
$\frac{1}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{\mathcal{R}[\beta]+\underline{\mathcal{R}}[\underline{\beta}]+1}{\lvert u\rvert}$,
whence the result follows.
This concludes the estimates on Ricci coefficients.
6. $L^{2}(S_{u,\underline{u}})$ Estimates for the Weyl curvature components
Proposition 6.1.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there holds
$\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1.$
Proof.
Recall the Bianchi equation for $\alpha$:
$\nabla_{3}\alpha+\frac{1}{2}\mbox{tr}\underline{\chi}\alpha=$ $\nabla\hat{\otimes}\beta+4\underline{\omega}\alpha-3\left(\hat{\chi}\rho+\prescript{*}{}{\hat{\chi}}\sigma\right)+(\zeta+4\eta)\hat{\otimes}\beta.$
Schematically, the above rewrites as
$\nabla_{3}\alpha+\frac{1}{2}\mbox{tr}\underline{\chi}\alpha=$ $\nabla\beta+\psi_{g}\alpha+\hat{\chi}(\rho,\sigma)+\psi_{g}\beta.$
Commuting with $i$ angular derivatives using 2, we arrive at
$\nabla_{3}\nabla^{i}\alpha+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\alpha$
$=$ $\nabla^{i+1}\beta+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\beta+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\alpha$
$+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\rho,\sigma,\beta)$
$+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\alpha$
$+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\alpha$
(147) $+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\alpha:=T_{i}^{1}+\dots+T_{i}^{7}.$
Passing to scale-invariant norms and using the weighted transport equality from Proposition 5.6, we can estimate as follows.
(148) $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}$ $\lesssim\lVert(a^{\frac{1}{2}}\nabla)^{i}\alpha\rVert_{L^{2}_{(sc)}(S_{u_{\infty},u})}+\sum_{1\leq j\leq 12}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}.$
There holds:
(149) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\lVert{a^{\frac{i}{2}}T_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}+\lVert{a^{\frac{i}{2}}T_{i}^{5}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\right)\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}\Gamma^{2}}{\lvert u^{\prime}\rvert^{3}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim 1.$
Moreover, there holds
(150) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{6}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim 1.$
For the eighth term, we can bound
(151) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{7}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a\Gamma^{2}}{\lvert u^{\prime}\rvert^{3}}\lesssim 1.$
The result follows.
Proposition 6.2.
For $\Psi_{u}$ defined as in Section 4.6, there holds
$\sum_{0\leq i\leq N+3}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\alpha]+1,$
[…]
Proof.
Each of the $\Psi_{u}$ satisfies the following schematic equation:
(152) $\nabla_{4}\Psi_{u}=\nabla\left(\Psi_{u},\alpha\right)+(\psi,\underline{\hat{\chi}})\left(\Psi_{u},\alpha\right).$
Commuting with $i$ angular derivatives using 4.1, we obtain
$\nabla_{4}\nabla^{i}\Psi_{u}=$ $\nabla^{i+1}(\Psi_{u},\alpha)+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}(\Psi_{u},\alpha)$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi,\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{u},\alpha)$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}.$
Passing to scale-invariant norms and estimating, we have
$\hskip 19.91692pt\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}$
$\lesssim$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
We restrict attention to $0\leq i\leq N+3$. For the first term, we have,
(153) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{1/2}=\mathcal{R}[\alpha],$
by using Hölder’s inequality. For the second term, since the $\Psi_{u}$ are regular with respect to scaling, we conclude that
(154) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{1}{a^{\frac{1}{2}}}\mathcal{R}\lesssim\frac{1}{a^{\frac{1}{2}}}R\lesssim 1,$
by the bootstrap assumptions (102). For the third term, we have $i_{3}+1\leq i$, hence everything can be closed using the $\Gamma$ total norm. We have
(155) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\lesssim\frac{\Gamma^{2}}{\lvert u\rvert}\lesssim 1.$
The fourth term can be estimated by
(156) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{\lvert u\rvert}{a}\frac{\Gamma^{2}}{\lvert u\rvert}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim 1.$
Moreover, for $\Gamma(\mbox{tr}\chi)$, we make use of Proposition 5.16 to bound $\Gamma(\mbox{tr}\underline{\chi})\lesssim 1$. For the seventh term, there holds
(157) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{a^{\frac{1}{2}}\Gamma^{2}}{\lvert u\rvert}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim 1.$
For the eighth term, there holds
(158) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{a^{\frac{3}{2}}\Gamma^{3}}{\lvert u\rvert^{2}}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim 1.$
Putting all of the above together, the result follows. ∎
Proposition 6.3.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there holds
$\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1.$
Proof.
Recall the Bianchi equation for $\underline{\alpha}$:
(159) $\begin{split}\nabla_{4}\underline{\alpha}+\frac{1}{2}\mbox{tr}\chi\hskip 1.42262pt\underline{\alpha}=&-\nabla\hat{\otimes}\underline{\beta}+4\omega\hskip 1.42262pt\underline{\alpha}-3\left(\underline{\hat{\chi}}\rho-\prescript{*}{}{\underline{\hat{\chi}}}\sigma\right)+\left(\zeta-4\underline{\eta}\right)\hat{\otimes}\underline{\beta}.\end{split}$
Schematically, the above rewrites as
(160) $\nabla_{4}\underline{\alpha}=\nabla\Psi_{u}+\psi_{g}(\underline{\alpha},\underline{\beta}^{R})+\underline{\hat{\chi}}(\rho,\sigma).$
Commuting with $i$ angular derivatives, we arrive at
$\hskip 19.91692pt\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}$
$\lesssim$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{\underline{u}})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
In the above expression, all terms can be bounded above by $1$, in the same way as in the preceding Proposition. The result follows.
6.1. Energy Estimates for the Weyl/Riemann Curvature components
For $(\Psi_{1},\Psi_{2})\in\begin{Bmatrix}(\alpha,\beta^{R}),(\beta,(\rho,\sigma)),((\rho,\sigma),\underline{\beta}),(\underline{\beta}^{R},\underline{\alpha})\end{Bmatrix}$ the energy estimates are carried out in Bianchi pairs, via the aid of the following proposition:
Proposition 6.4.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), for a Bianchi pair $(\Psi_{1},\Psi_{2})$ satisfying
$\nabla_{3}\nabla^{i}\Psi_{1}+\left(\frac{i+1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\nabla^{i}\Psi_{1}-\mathcal{D}\nabla^{i}\Psi_{2}=P_{i},$
$\nabla_{4}\nabla^{i}\Psi_{2}-\prescript{*}{}{\mathcal{D}}\nabla^{i}\Psi_{1}=Q_{i},$
the following holds true:
(161) $\begin{split}&\int_{H_{u}^{(0,\underline{u})}}\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{\underline{H}_{\underline{u}}^{(u_{\infty},u)}}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{H_{u_{\infty}}^{(0,\underline{u})}}\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{\underline{H}_{0}^{(u_{\infty},u)}}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},0})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\\ &+\iint_{\mathcal{D}_{u,\underline{u}}}\frac{a}{\lvert u^{\prime}\rvert}\lVert{\nabla^{i}\Psi_{1}\cdot P_{i}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\iint_{\mathcal{D}_{u,\underline{u}}}\frac{a}{\lvert u^{\prime}\rvert}\lVert{\nabla^{i}\Psi_{2}\cdot Q_{i}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
Before embarking on the energy estimates, we provide a final helpful proposition, which can be found for example in [22].
Proposition 6.5.
Let $f(x,y),\hskip 1.42262ptg(x,y)$ be positive functions defined on a rectangle $U:=\begin{Bmatrix}0\leq x\leq x_{0},\hskip 1.42262pt0\leq y\leq y_{0}\end{Bmatrix}.$ Suppose there exist constants $J,c_{1},c_{2}$ such that
$f(x,y)+g(x,y)\lesssim J+c_{1}\int_{0}^{x}f(x^{\prime},y)\hskip 1.42262pt\text{d}x^{\prime}+\int_{0}^{y}g(x,y^{\prime})\hskip 1.42262pt\text{d}y^{\prime},$
for all $(x,y)\in U$. Then there holds
$\forall(x,y)\in U:\hskip 14.22636ptf(x,y)+g(x,y)\lesssim J\mathrm{e}^{c_{1}x+c_{2}y}.$
With this Proposition as the main tool, we begin with $(\alpha,\beta)$.
Proposition 6.6.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there holds, for all $0\leq i\leq N+4$:
(162) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\\ &\lesssim\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+1.\end{split}$
Proof.
We recall the (schematic) Bianchi equations for $\nabla^{i}\alpha,\nabla^{i}\beta$:
$\nabla_{3}\nabla^{i}\alpha+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\alpha-\mathcal{D}\nabla^{i}\beta$
$=$ $\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\beta^{R}+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\alpha$
$+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\rho,\sigma,\beta)$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\alpha$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\alpha$
(163) $+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\alpha:=P_{i}^{1}+\dots+P_{i}^{6}.$
Similarly, we have
$\nabla_{4}\nabla^{i}\beta-\prescript{*}{}{\mathcal{D}}\nabla^{i}\alpha=$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\beta,\alpha)$
(164) $+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\beta$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\beta:=Q_{i}^{1}+\dots+Q_{i}^{3}.$
Applying Proposition 6.4, we have
(165) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\\ \lesssim&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}\\ &+\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
By Hölder’s inequality, one has
(166) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \leq&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\sum_{j=1}^{10}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\cdot\sup_{u^{\prime}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\alpha\rVert_{L^{2}_{(sc)}(H_{u^{\prime}}^{(0,\underline{u})})},\end{split}$
Let us focus on the sum in the above line. For the first three terms, there holds
$\sum_{j=1}^{3}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\frac{a^{\frac{1}{2}}\Gamma\cdot R}{\lvert u^{\prime}\rvert}$.
For the fourth and fifth terms, there holds
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}+\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{5}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\frac{a^{\frac{1}{2}}\Gamma\cdot M}{\lvert u^{\prime}\rvert}$.
For the sixth term, there holds
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{6}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\Gamma^{3}.$
For the seventh term, there holds
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{7}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\frac{a^{\frac{1}{2}}\Gamma\cdot M}{\lvert u^{\prime}\rvert}$.
For the eighth term, there holds
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{8}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\Gamma^{2}+\Gamma R.$
For the ninth and tenth terms, there holds
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{9}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}+\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{10}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\Gamma^{3}+\Gamma^{2}R.$
Putting everything together, there holds
(167) $\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\Gamma^{3}+\Gamma^{2}R+\Gamma R+1.$
Similarly, for the analogous term involving $\beta$, there holds
(168) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \leq&\sum_{j=1}^{7}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\cdot\sup_{u^{\prime}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\beta\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}.\end{split}$
We estimate term by term. For the first term, there holds
(169) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma(R+\Gamma)}{\lvert u\rvert}$
For the second term, there holds
(170) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma(M+\Gamma)}{\lvert u\rvert}$
For the third and fourth terms, there holds
(171) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{3}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}+\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a\Gamma(M+\Gamma)}{\lvert u\rvert}\end{split}$
For the fifth term there holds
(172) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{5}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma(R+\Gamma)}{\lvert u\rvert}$
For the sixth term, we have
(173) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{6}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma^{2}(R+\Gamma)}{\lvert u\rvert^{2}}$
For the seventh term, we have
(174) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{7}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{\Gamma^{2}(R+\Gamma)}{\lvert u\rvert^{3}}$
Putting everything together, we have
(175) $\sum_{j=1}^{7}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma(R+M+\Gamma)}{\lvert u\rvert}+1.$
Combining (167) and (175), we have
(176) $\begin{split}&\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\\ \lesssim&\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}\\ &+\frac{1}{a^{\frac{1}{2}}}\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\frac{1}{a^{\frac{1}{2}}}\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\frac{1}{a^{\frac{1}{2}}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+\frac{1}{a^{\frac{1}{3}}}.\end{split}$
The claim follows. ∎
We now move on to energy estimates for the remaining pairs $(\beta,(\rho,\sigma)),((\rho,\sigma),\underline{\beta})$ and $(\underline{\beta},\underline{\alpha})$.
Proposition 6.7.
Let $(\Psi_{1},\Psi_{2})\in\begin{Bmatrix}a^{\frac{1}{2}}(\beta,(\rho,\sigma)),a^{\frac{1}{2}}((\rho,\sigma),\underline{\beta}),(\underline{\beta},\underline{\alpha})\end{Bmatrix}$. For $0\leq i\leq N+4$, there holds:
(177) $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}$
(178) $\lesssim$ $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+1.$
Proof.
The schematic equations for $\Psi_{1},\Psi_{2}$ are:
$\nabla_{3}\Psi_{1}+\left(\frac{1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\Psi_{1}-\mathcal{D}\Psi_{2}=$ $(\psi,\hat{\chi})\Psi,$
(179) $\begin{split}\nabla_{4}\Psi_{2}-\prescript{*}{}{\mathcal{D}}\Psi_{1}=(\psi,\underline{\hat{\chi}})(\Psi_{u},\alpha).\end{split}$
Commuting with $i$ angular derivatives, for $\Psi_{1}$, we have:
(180) $\begin{split}&\nabla_{3}\nabla^{i}\Psi_{1}+\left(\frac{i+1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\nabla^{i}\Psi_{1}-\mathcal{D}\nabla^{i}\Psi_{2}\\ =&\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}\Psi\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\Psi_{1}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\Psi_{1}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\Psi_{1}:=P_{i}.\end{split}$
Analogously, for $\Psi_{2}$, we have
(181) $\begin{split}&\nabla_{4}\nabla^{i}\Psi_{2}-\prescript{*}{}{\mathcal{D}}\nabla^{i}\Psi_{1}\\ =&\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{1}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi,\alpha)\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{2}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{2}:=Q_{i}.\end{split}$
Making use of Proposition 6.4 once again, we arrive at
(182) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}^{2}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}^{2}\\ \lesssim&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}^{2}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}^{2}\\ &+\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
For the first spacetime integral in the above, we estimate
(183) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\cdot\left(\int_{0}^{\underline{u}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}.\end{split}$
For the first term:
(184) $\begin{split}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime},\end{split}$
if $i_{3}+1\geq N+3$, we estimate
(185) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\sup_{0\leq\underline{u}^{\prime}\leq\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{(a^{\frac{1}{2}}\nabla)^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\lVert{(a^{\frac{1}{2}})^{i_{1}+i_{2}}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}}\rVert_{L^{\infty}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\end{split}$
and we can estimate $\lVert{(a^{\frac{1}{2}})^{i_{1}+i_{2}}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}}\rVert_{L^{\infty}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}$ by $\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{2}}}$ using the bootstrap assumption (102), to obtain
(186) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\sup_{0\leq\underline{u}^{\prime}\leq\underline{u}}\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{R}{\lvert u\rvert^{\frac{1}{2}}}\lesssim 1.\end{split}$
If, however, $i_{3}+1\leq N+2$, we can control the corresponding $L^{2}_{(sc)}(S)$ norm just by the bootstrap assumption (102) to get the bound
(187) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma^{2}}{\lvert u\rvert^{2}}\lesssim 1.\end{split}$
For the rest of the terms, we estimate using the same philosophy as appropriate. There holds
(188) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}\Psi}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a^{\frac{1}{2}}\Gamma(R+\Gamma)}{\lvert u\rvert}\lesssim 1.\end{split}$
Here in particular, when $i_{4}\geq N+3$, we treat the cases $\Psi=\Psi_{u}$ and $\Psi=\Psi_{\underline{u}}$ separately. For the sixth and seventh terms, we can bound them by one as in previous calculations. For the eighth term, using the fact that $i-2\leq N+2$ and the improvements from Proposition 6.2 and Proposition 6.6, we arrive at
(189) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a}{\lvert u\rvert}\Gamma[\mbox{tr}\underline{\chi}]\Gamma[\alpha^{F}]^{2}\lesssim\frac{a}{\lvert u\rvert}\Gamma[\eta,\underline{\eta}]\Gamma[\mbox{tr}\underline{\chi}]\Gamma[\Psi_{1}]\lesssim\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta^{R},\underline{\beta}^{R})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\cdot 1\cdot(\mathcal{R}[\alpha]+1)\\ \lesssim&(\mathcal{R}[\alpha]+1)^{2}\lesssim 1,\end{split}$
where in the last line we made use of Proposition 6.6. This completes the estimates for the first spacetime integral in (182). For the second and last one, a double application of Hölder’s inequality yields
(190) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{0}^{\underline{u}}\left(\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}Q_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\right)^{\frac{1}{2}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}\\ \lesssim&\left(\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}Q_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\left(\int_{0}^{\underline{u}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}\rVert^{2}_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\\ \lesssim&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}Q_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\frac{1}{4}\int_{0}^{\underline{u}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}\rVert^{2}_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\end{split}$
Define $B:=\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}Q_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$. We can then estimate $B$ as follows:
(191) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\psi_{g}\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\nabla\psi_{g}\nabla^{i-1}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\psi_{g}\hskip 1.42262pt\psi_{g}\nabla^{i-1}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\sum_{\begin{subarray}{c}i_{1}+i_{2}+i_{3}=i-1\\ i_{3}<i-2\end{subarray}}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&1,\end{split}$
where in the first three integrals we estimate $\psi_{g},\nabla\psi_{g}$ in $L^{\infty}_{(sc)}(S_{u,\underline{u}})$ and $\nabla^{i}\Psi_{1}$ in the hypersurface norm $L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})$ and in the last integral, since $i-2\leq N+2$, we can estimate $\nabla^{i}\Psi_{1}$ in $L^{2}_{(sc)}(S_{u,\underline{u}})$ using the bootstrap assumption on the norm $\Gamma$. For the second term, we similarly have
$\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\Psi,\alpha)}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
(192) $\lesssim$ $\left(\mathcal{R}[\alpha]+1\right)^{2}\lesssim 1,$
where we have used the improvements on $\hat{\chi}$ from Proposition 5.12 and the energy estimate from Proposition 6.6. The rest of the terms can also be bounded above by $1$, using the same approach. We finally arrive at an estimate of the form
(193) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&1+\frac{1}{4}\int_{0}^{\underline{u}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}\rVert^{2}_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
From here, collecting all the terms together and using Proposition 6.5, we arrive at the desired result. ∎
This completes the central part of this work. In particular, we prove that there exists an open set of initial data that is almost compatible with the scaling of [3] (with some differences handled by new scaling) and is capable of producing a semi-global characteristic development of the vacuum Einstein’s equations. Simultaneously, we also obtain concentrated generalized Yau mean curvature $c$ (as defined in Theorem 1.2) along the null hypersurface u=au=-a. The idea then is to show that the radius of the boundary $S_{-a,\epsilon}$ is large enough (compared to that of $S_{-a,0}$) so that the condition stated in the theorem 1.2 is met. The last part involves constructing the data on the Cauchy slice $\mathcal{M}_{t=-a}$ and evolving it for a short enough time $\epsilon>0$ (see the diagram 1 for clarity). The motivation behind solving the characteristic initial value problem is to concentrate the mean curvature along the incoming direction by means of large conjugate shear $\underline{\hat{\chi}}$ concentrated along the initial null-hypersurface $\underline{H}_{0}$ starting from dispersed data at past null-infinity.
7. The Cauchy problem on the slab $a,-a+\epsilon]\times\widehat{\mathcal{M}}$ and completion of the argument
In this section, we complete the proof of Theorem 3.1 by establishing the formation of an apparent horizon strictly to the future of the Cauchy hypersurface
$\widehat{\mathcal{M}}_{t=-a}\subset\widehat{\mathcal{M}}$.
We work in spacetime harmonic gauge and solve the vacuum Einstein equations on the slab
$a,-a+\epsilon]\times\widehat{\mathcal{M}},\qquad 0<\epsilon\ll 1,$
with initial data assembled by a gluing procedure. We then propagate the quasi-local boundary geometry forward in time, ultimately showing that the Schoen–Yau curvature radius condition is violated in the future domain, forcing the existence of an apparent horizon. In this section, we provide a sketch of the construction of the data. As mentioned previously, the detailed construction is executed in the sequel [28] since it is, by itself, a very nontrivial problem to solve.
7.1. Geometric setup and decomposition of the initial slice
Let (^,g^)($\widehat{\mathcal{M}}$$\widehat${g}) be the Lorentzian manifold constructed in previous sections. Let $u,\underline{u}$ be double-null coordinates associated to the canonical foliation, and let
$T:=\tfrac{1}{2}(e_{4}+e_{3})$
denote the future-directed unit timelike vector field orthogonal (in the induced sense) to the leaves $S_{u,\underline{u}}\subset\widehat{\mathcal{M}}_{t=u+\underline{u}}$.
The interior region $\widehat{\mathcal{M}}_{\mathrm{int}}\subset\widehat{\mathcal{M}}$ admits a foliation by topological 22–spheres outside a compact subset. Our analysis requires only the boundary geometry $\partial\widehat{\mathcal{M}}_{\mathrm{int}}$ and its associated Yau radius.
Fix times
$t_{1}=-a,\qquad t_{2}=-a+\epsilon,$
where $\epsilon>0$ is sufficiently small. Although we do not require the double-null foliation to describe the interior evolution, it is convenient near $\partial\widehat{\mathcal{M}}_{\mathrm{int}}$.
On the initial slice $\widehat{\mathcal{M}}_{t=-a}$ we introduce the decomposition
(194) $\widehat{\mathcal{M}}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
where
$\mathcal{M}_{1}:=\widehat{\mathcal{M}}_{\mathrm{int}},\qquad\mathcal{M}_{2}:=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon},\qquad\mathcal{M}_{3}:=\widehat{\mathcal{M}}_{t=-a}\setminus(\mathcal{M}_{1}\cup\mathcal{M}_{2}).$
7.2. Specification of the Cauchy data
The Cauchy data $(g,k)$ on $\widehat{\mathcal{M}}_{t=-a}$, consisting of the induced Riemannian metric $g$ and second fundamental form $k$, are specified as follows:
  1. (a)
    Interior data. On $\mathcal{M}_{1}$ we prescribe smooth $C^{\infty}$ data $(g,k)$ satisfying
    (195) $\operatorname{Rad}(\mathcal{M}_{1})<\frac{3\pi}{2\,c_{\partial\mathcal{M_{1}}}},$
    where $c_{\partial\mathcal{M}_{1}}$ denotes the generalized mean curvature coefficient of $\partial\mathcal{M}_{1}$ in the sense of Schoen–Yau. Existence of such data follows from the quasi-local analysis of [38, 37].
  2. (b)
    Matching data in the characteristic region. On $\mathcal{M}_{2}=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon}$, the data are induced from the characteristic development on the double-null slab $\mathcal{D}_{-a,\epsilon}$. Well-posedness of the characteristic problem on this region was established in the previous section.
  3. (c)
    Asymptotically Kerr exterior. On $\mathcal{M}_{3}$, we prescribe smooth data $(g,k)$ asymptotic to a Kerr slice with prescribed mass $m>0$ and angular momentum $J$. This ensures asymptotic flatness.
A gluing construction in the spirit of [20, 23, 10] yields a smooth, asymptotically flat vacuum Cauchy pair $(g,k)$ on $\widehat{\mathcal{M}}_{t=-a}$. Complete geometric–analytic details of this construction will appear in [28].
7.3. Setup, decomposition, constraints, and gauges
Let (^,g^)($\widehat{\mathcal{M}}$$\widehat${g}) be the semi-globally constructed Lorentzian manifold equipped with a canonical double–null foliation $(u,\underline{u})$ and Cauchy time $t=u+\underline{u}$. Fix $a\gg 1$ and a small $0<\epsilon\ll 1$. On the initial slice $\widehat{\mathcal{M}}_{t=-a}$ set
$\widehat{\mathcal{M}}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
with $\mathcal{M}_{1}=\widehat{\mathcal{M}}_{\mathrm{int}}$, $\mathcal{M}_{2}=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon}$ (the double–null slab), and $\mathcal{M}_{3}=\widehat{\mathcal{M}}_{t=-a}\setminus(\mathcal{M}_{1}\cup\mathcal{M}_{2})$. We must solve on $\widehat{\mathcal{M}}_{t=-a}$ the vacuum constraint system
(196) $\text{Scal}(g)-|k|_{g}^{2}+(\mbox{tr}_{g}k)^{2}=0,\qquad\mbox{div }_{g}k-d(\mbox{tr}_{g}k)=0,$
subject to (i) prescribed boundary/interface conditions across $\partial\mathcal{M}_{i}$, (ii) MOTS–exclusion on t=at=-a, and (iii) asymptotically Kerr behavior on $\mathcal{M}_{3}$.
Throughout we can use weighted $W^{2,p}_{-\delta}\times W^{1,p}_{-1-\delta}$ scales with $p>3$ and $0<\delta<1$.
7.4. Interior data on $\mathcal{M}_{1}$: strict Yau–radius subcriticality with boundary control
Let $\Omega:=\mathcal{M}_{1}$ and denote its outer boundary by $\partial\Omega=\partial\widehat{\mathcal{M}}_{\mathrm{int}}$. Let $H$ be the mean curvature of $\partial\Omega$ in the initial slice and $p\equiv k$ the slice’s second fundamental form in spacetime. Fix a strictly positive boundary gap
(197) $c_{\ast}:=\inf_{\partial\Omega}\bigl(H-|\mbox{tr}_{\partial\Omega}k|\bigr)>0,$
and choose a target radius bound $0<\text{Rad}(\Omega)\leq R_{\ast}<R_{\mathrm{crit}}(c_{\ast})$ strictly below Yau’s critical threshold (quantified below).
Boundary barrier via Yau’s brane functional.
Let $f>0$ be the first eigenfunction for the mixed operator $\Delta+\tfrac{1}{2}R-h$ with Robin data $\partial_{\nu}f+kf=0$ on $\partial\Omega$ (see [38] for details). For $\Gamma\subset\partial\Omega$ realizing $\text{Rad}(\Omega)$ up to $o(1)$, and a spanning disk $\Sigma\subset\Omega$ with $\partial\Sigma=\Gamma$, set
$\mathcal{L}_{f}(\Sigma)=\int_{\Sigma}f-c\int_{\Omega_{\Sigma}}f,\qquad c\in(0,c_{\ast}].$
[…]
$\min_{\partial\Omega}\bigl(H(\tilde{g})-|\mbox{tr}_{\partial\Omega}\tilde{k}|\bigr)\geq c_{\ast}>0,\qquad\text{Rad}(\Omega)\leq R_{\ast}<R_{\mathrm{crit}}(c_{\ast}),$
so $\mathcal{M}_{1}$ contains no apparent horizon on t=at=-a. Moreover, these inequalities are stable under the smoothing that we will do across the interfaces.
7.5. Characteristic matching on $\mathcal{M}_{2}$: importing the double–null data
On $\mathcal{M}_{2}=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon}$, take the data induced from the double–null development on the slab $\mathcal{D}_{-a,\epsilon}$, in the normalized frame
$e_{4}=\Omega^{-1}\partial_{\underline{u}},\qquad e_{3}=\Omega^{-1}(\partial_{u}+b^{A}\partial_{\theta^{A}}),\qquad g|_{S_{u,\underline{u}}}=\gamma_{AB}d\theta^{A}d\theta^{B}.$
Denote the null second fundamental forms by $\chi=\widehat{\chi}+\tfrac{1}{2}(\mbox{tr}\chi)\gamma$ and $\underline{\chi}=\widehat{\underline{\chi}}+\tfrac{1}{2}(\mbox{tr}\underline{\chi})\gamma$. We import the full set of Ricci and curvature components on $\mathcal{M}_{2}$, including the incoming shear $\widehat{\underline{\chi}}$, with the scale–critical bounds dictated by the construction of the characteristic initial value problem (on $H_{0}$ and $H_{u_{\infty}}$):
$\|\widehat{\chi}\|_{L^{\infty}(S_{u,\underline{u}})}\lesssim|u|^{-1},\qquad\|\widehat{\underline{\chi}}\|_{L^{\infty}(S_{u,\underline{u}})}\lesssim a^{1/2}|u|^{-2},$
$|\mbox{tr}\chi|\lesssim|u|^{-1},\quad|\mbox{tr}\underline{\chi}|\lesssim|u|^{-1},\quad\|\eta\|_{L^{\infty}}+\|\underline{\eta}\|_{L^{\infty}}\lesssim a^{1/2}|u|^{-2},\quad|\omega|+|\underline{\omega}|\lesssim|u|^{-1}.$
These are propagated from the double–null hierarchy and are compatible with the transport system
$\nabla_{3}\mbox{tr}\chi+\tfrac{1}{2}(\mbox{tr}\chi)^{2}=-|\widehat{\chi}|^{2}-2\underline{\omega}\,\mbox{tr}\chi,\qquad\nabla_{4}\mbox{tr}\underline{\chi}+\tfrac{1}{2}(\mbox{tr}\underline{\chi})^{2}=-|\widehat{\underline{\chi}}|^{2}-2\omega\,\mbox{tr}\underline{\chi},$
together with the $\nabla_{4}$$\nabla_{3}$ equations for $\widehat{\chi}$ and $\widehat{\underline{\chi}}$. In particular, on the interface sphere $S_{-a,\underline{u}}$ the combination
$H-|\kappa|=\tfrac{1}{2}(\mbox{tr}\chi-\mbox{tr}\underline{\chi})-\tfrac{1}{2}\bigl|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\bigr|$
is explicitly computable from these inputs, ensuring (via the chosen profile of incoming shear $\underline{\hat{\chi}}$ and the initial incoming expansion $\mbox{tr}\underline{\chi}$) that $c:=\min(H-|\kappa|)$ on the boundary of the interior piece at t=at=-a is below the Yau threshold, hence no MOTS appear there. The first and second fundamental forms of the slice and of the $S_{u,\underline{u}}$ leaves are matched in $C^{\infty}$ across $\partial\mathcal{M}_{1}\cap\partial\mathcal{M}_{2}$ by a standard partition–of–unity interpolation inside a thin collar, after which we re-solve the constraints by compactly supported corrections in the next section.
7.6. Asymptotically Kerr exterior on $\mathcal{M}_{3}$ and compactly supported correction
Choose a large coordinate sphere that encloses $\mathcal{M}_{1}\cup\mathcal{M}_{2}$ and fix asymptotically flat coordinates outside. Pick a member of the admissible Kerr family (mass $m>0$, angular momentum $J$ and Poincaré parameters) and carry out the Corvino–Schoen gluing: smoothly splice $(g,k)$ to the chosen Kerr slice in an annulus $A_{R}=\{R<r<2R\}$, then perturb by a compactly supported deformation inside $A_{R}$ to annihilate the ten–dimensional cokernel of the linearized constraint operator in $W^{2,p}_{-\delta}\times W^{1,p}_{-1-\delta}$. The surjectivity of the linearization in these scales and the degree–argument for the parameter map $\Theta$ produce a solution $(g^{\sharp},\pi^{\sharp})$ that equals Kerr outside $|x|\geq 2R$ and equals the original data inside $|x|\leq R$, while solving (196) globally.
7.7. Smoothing across interfaces and final properties on t=at=-a
Using a collar–supported partition of unity on the 2-spheres separating $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$, we ensure $C^{\infty}$ matching of the slice’s first and second fundamental forms, and re-solve the constraints by compactly supported corrections that do not alter $\mathcal{M}_{1}\cup\mathcal{M}_{2}$ nor the exterior Kerr end. The resulting global Cauchy data $(g,k)$ on $\widehat{\mathcal{M}}_{t=-a}$ satisfy:
  1. (1)
    $(g,k)$ are smooth and vacuum everywhere and asymptotically Kerr outside a compact set;
  2. (2)
    $\mathcal{M}_{1}$ is Yau–radius subcritical with boundary gap $c_{\ast}>0$; hence no MOTS occur on t=at=-a;
  3. (3)
    the characteristic shear content on $\mathcal{M}_{2}$ ($\widehat{\underline{\chi}}$) is large in the prescribed sense; the future slice $t=-a+\epsilon$ then satisfies a strengthened boundary–curvature inequality $c_{S_{-a,\epsilon}}>\tfrac{3\pi}{2\text{Rad}(\cdot)}$ for the interior domain, forcing an apparent horizon in its interior in the future development.
7.8. Proof notes for the boundary barrier and the Yau radius threshold
Let $f$ solve the mixed eigenvalue problem $\partial_{\nu}f+kf=0$ on $\partial\Omega$ as stated in [38]. For $\Gamma\subset\partial\Omega$ realizing $\text{Rad}(\Omega)$, minimizing $\mathcal{L}_{f}(\Sigma)$ among disks $\Sigma$ with $\partial\Sigma=\Gamma$ yields a stationary surface obeying a mean–curvature equation whose first and second variations control the concentrating competitors inside the tubular neighborhood of $\partial\Omega$. Using the barrier built from the distance function $d(x,\partial\Omega)$ and profiles $\varphi(d)$ with $\varphi^{\prime}(\varepsilon)=\infty$, one enforces that minimizers stay away from $\partial\Omega$, and the second variation produces the quantitative inequality linking $c_{\ast}$ and $\text{Rad}(\Omega)$. Choosing $R_{\ast}<R_{\mathrm{crit}}(c_{\ast})$ excludes MOTS on t=at=-a. Now note the following vital points that are made explicit.
  • The interior construction produces $(g,k)$ with a fixed positive gap $H-|\mbox{tr}_{\partial\Omega}k|\geq c_{\ast}$ and with $\text{Rad}(\Omega)$ prescribed strictly below the critical Yau threshold.
[…]
  • The exterior gluing gives an exact Kerr outside a compact set without spoiling the interior Yau gap or the characteristic content.
Now we briefly discuss the constraints. ∎
7.9. Constraint map and local deformation
Let $\Phi(g,\pi)=\big(\text{Scal}(g)+(\mbox{tr}_{g}\pi)^{2}-|\pi|_{g}^{2},\ \text{Div}_{g}\pi-d(\mbox{tr}_{g}\pi)\big)$ denote the vacuum constraint operator on a 33–manifold. Its linearization $D\Phi_{(g,\pi)}$ at an asymptotically flat (AF) seed and the formal $L^{2}$–adjoint $D\Phi_{(g,\pi)}^{*}$ are overdetermined–elliptic. On weighted Sobolev scales $W^{2,p}_{-\delta}\times W^{1,p}_{-1-\delta}$ ($p>3$, $0<\delta<1$), the adjoint has surjective range modulo a finite–dimensional obstruction space [25, 26]
$\mathcal{K}\cong\mathbb{R}^{10},$
canonically identified with Killing initial data corresponding to the Poincaré charges $(E,P,J,C)$ at infinity. Compactly supported solutions of
$D\Phi^{*}_{(g,\pi)}(u,X)=F\quad\text{on a domain }U\Subset M$
exist and yield compactly supported corrections to $(g,\pi)$ after a quadratic iteration, provided the $\mathcal{K}$–projection of the residual is killed by parameter variation in an admissible model family (Kerr) on the outer end.
7.10. Regionwise seeds
Fix times $t_{1}=-a$, $t_{2}=-a+\epsilon$, $\epsilon>0$ small. Decompose the initial slice
(198) $\widehat{\mathcal{M}}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
with interfaces along smooth 22–spheres.
(I) Interior seed on $\mathcal{M}_{1}$ with prescribed Yau radius as discussed previously.
Let $\Omega:=\mathcal{M}_{1}$ with outer boundary $\partial\Omega$. Choose a smooth Riemannian pair $(g^{\rm in},k^{\rm in})$ on $\Omega$ such that:
  1. (i)
    Boundary gap & Yau radius. On $\partial\Omega$,
    $c_{*}:=\min_{\partial\Omega}\big(H(g^{\rm in})-|\mbox{tr}_{\partial\Omega}k^{\rm in}|\big)>0,\qquad\text{Rad}(\Omega)=R_{*}<\frac{3\pi}{2\,c_{*}}.$
  2. (ii)
    Interface geometry. On the interface sphere $S_{-a,\underline{u}_{0}}=\partial\Omega\cap\partial\mathcal{M}_{2}$, set
    $\big(H(g^{\rm in}),\ \mbox{tr}_{S}k^{\rm in},\ \widehat{k}^{\rm in}\big)\big|_{S_{-a,\underline{u}_{0}}}=\big(H^{\rm dn},\ \mbox{tr}_{S}k^{\rm dn},\ \widehat{k}^{\rm dn}\big),$
    where the right side is imported from the double–null development (including the incoming shear $\widehat{\underline{\chi}}$ through the standard identification with tangential pieces of $k$).
No constraint is imposed yet in the interior of $\Omega$.
(II) Characteristic seed on $\mathcal{M}_{2}$ (from the double–null slab).
On $\mathcal{M}_{2}=\widehat{\mathcal{M}}_{t=-a}\cap\mathcal{D}_{-a,\epsilon}$ take the induced Cauchy data $(g^{\rm dn},k^{\rm dn})$ coming from the slab solution with scale–critical control on $(\widehat{\chi},\mbox{tr}\chi,\widehat{\underline{\chi}},\mbox{tr}\underline{\chi},\eta,\underline{\eta},\omega,\underline{\omega})$ and the shear budget $\int|\widehat{}\underline{\chi}_{0}|^{2}$ as prescribed in the characteristic problem. In particular, the transport system for $\widehat{\underline{\chi}}$ and $\mbox{tr}\underline{\chi}$ along $e_{4}$ is part of this data, hence the incoming focusing appears explicitly on $S_{-a,\underline{u}_{0}}$.
(III) Kerr seed on $\mathcal{M}_{3}$.
On the AF exterior choose a fixed Kerr slice $(g^{K}_{m,a},\pi^{K}_{m,a})$ with mass-angular momentum parameters $(m,J)$ and arbitrary Poincaré charges $(E,P,J,C)$ to be tuned later. Denote this by $(g^{\rm ex},k^{\rm ex})$.
7.11. Partition of unity and raw splice
Choose nested collars and annuli:
$\mathcal{C}_{1}\Subset\mathcal{M}_{1},\quad\mathcal{C}_{2}\Subset\mathcal{M}_{2},\quad A_{R}:=\{R<r<2R\}\Subset\mathcal{M}_{3},$
so that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are disjoint, and $A_{R}$ lies far out. Let $\{\xi_{1},\xi_{2},\xi_{3}\}$ be a smooth partition subordinate to $\{\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}\}$, with $\nabla\xi_{i}$ supported inside the chosen collars and annuli.
Define the splice (piecewise smooth) pair
$(g_{0},\pi_{0}):=\big(\xi_{1}g^{\rm in}+\xi_{2}g^{\rm dn}+\xi_{3}g^{\rm ex},\ \xi_{1}k^{\rm in}+\xi_{2}k^{\rm dn}+\xi_{3}k^{\rm ex}\big).$
By construction, $(g_{0},\pi_{0})$ agrees exactly with each seed away from the collars, and is smooth globally. Its constraint residual
$\mathcal{R}_{0}:=\Phi(g_{0},\pi_{0})$
is supported in $\operatorname{supp}\nabla\xi_{1}\cup\operatorname{supp}\nabla\xi_{2}\cup\operatorname{supp}\nabla\xi_{3}\subset\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup A_{R}$.
7.12. Compactly supported constraint correction (interior and interface)
Step A: kill $\mathcal{R}_{0}$ in $\mathcal{C}_{1}$ (purely interior).
Solve the projected adjoint equation on $\mathcal{C}_{1}$:
(199) $D\Phi^{*}_{(g_{0},\pi_{0})}(u_{1},X_{1})=\mathcal{T}_{1}(\mathcal{R}_{0})\quad\text{on }\mathcal{C}_{1},\qquad(u_{1},X_{1})\text{ compactly supported\ in }\mathcal{C}_{1},$
where $\mathcal{T}_{1}$ is a right–inverse onto the orthogonal complement of the local obstruction space. Solve the associated quadratic corrected system to get a compactly supported perturbation
$(\delta g_{1},\delta\pi_{1})=\mathcal{Q}_{1}(u_{1},X_{1})$
with $\Phi(g_{0}+\delta g_{1},\pi_{0}+\delta\pi_{1})=\mathcal{R}_{0}-\mathcal{P}_{1}(\mathcal{R}_{0})$ on $\mathcal{C}_{1}$, where $\mathcal{P}_{1}$ is the projection onto the local obstruction. Because $\mathcal{C}_{1}$ is strictly interior, no global charges are involved.
Step B: kill $\mathcal{R}_{0}$ in $\mathcal{C}_{2}$ (interface to double–null).
Repeat (199) on $\mathcal{C}_{2}$ to obtain a c.s. correction $(\delta g_{2},\delta\pi_{2})$ supported in $\mathcal{C}_{2}$, leaving the trace data on the common sphere $S_{-a,\underline{u}_{0}}$ unchanged. This guarantees that the double–null induced tangential geometry—hence the scalars built from $(\mbox{tr}\chi,\mbox{tr}\underline{\chi})$ and the shear $\widehat{\underline{\chi}}$—is preserved on the interface.
After Steps A and B we have
$\Phi(g_{1},\pi_{1})=\Phi\big((g_{0},\pi_{0})+(\delta g_{1}+\delta g_{2},\delta\pi_{1}+\delta\pi_{2})\big)=\mathcal{R}_{R}$
with $\operatorname{supp}\mathcal{R}_{R}\subset A_{R}$.
7.13. Exterior compactly supported correction with Kerr parameter tuning
On the large annulus $A_{R}$, perform the compactly supported exterior correction: solve
$D\Phi^{*}_{(g_{1},\pi_{1})}(u_{3},X_{3})=\mathcal{T}_{R}(\mathcal{R}_{R})\quad\text{on }A_{R},$
producing a compactly supported perturbation $(\delta g_{3},\delta\pi_{3})$ supported in $A_{R}$. The remaining obstruction lives in the global $10$-dimensional space $\mathcal{K}$; adjust the Kerr parameters $(E,P,J,C)$ so that the parameter map hits the obstruction and annihilates it. Concretely, choose $(E,P,J,C)$ so that
$\Pi_{\mathcal{K}}\big(\Phi(g_{1}+\delta g_{3},\pi_{1}+\delta\pi_{3})\big)=0,$
which is possible by the surjectivity argument. The outcome is a vacuum pair
$(g^{\sharp},\pi^{\sharp})$
with
$\Phi(g^{\sharp},\pi^{\sharp})=0,\qquad(g^{\sharp},\pi^{\sharp})=\begin{cases}(g^{\rm in},k^{\rm in})&\text{on }\mathcal{M}_{1}\setminus\mathcal{C}_{1},\\ (g^{\rm dn},k^{\rm dn})&\text{on }\mathcal{M}_{2}\setminus\mathcal{C}_{2},\\ (g^{K}_{m,a},\pi^{K}_{m,a})&\text{on }\{r\geq 2R\},\end{cases}$
and smooth across all interfaces.
7.14. Yau–radius prescription and MOTS/future AH behavior
By design, the boundary geometry on $\partial\Omega$ (of $\mathcal{M}_{1}$) is that of the double–null interface sphere, with
$c_{*}=\min_{\partial\Omega}\big(H-|\mbox{tr}_{\partial\Omega}k|\big)>0,\qquad\text{Rad}(\Omega)=R_{*}<\frac{3\pi}{2\,c_{*}}.$
The compactly supported corrections never alter $\partial\Omega$, hence the Yau barrier conditions and the prescribed Yau radius are retained on the initial slice. Consequently, there is no MOTS on t=at=-a. On the other hand, the shear content imported on $\mathcal{M}_{2}$ (including $\widehat{\underline{\chi}}$) guarantees MOTS formation in the future development (as per the scale-critical characteristic criteria) while the initial slice remains regular.
7.15. Local evolution in harmonic gauge
We now solve the vacuum Einstein equations
$\mathrm{Ric}[\widehat{g}]=0$
in the harmonic gauge
(200) $\Gamma^{\mu}:=\widehat{g}^{\alpha\beta}\Gamma[\widehat{g}]^{\mu}_{\alpha\beta}=0.$
Under (200) the field equations reduce to the quasilinear wave system
[…]
where the nonlinearity $\mathcal{N}$ is quadratic in $\partial\widehat{g}$. By the standard local well-posedness theorem in the spacetime harmonic gauge, there exists $\epsilon>0$ such that the Cauchy problem for (201) with initial data $(g,k)$ on $\widehat{\mathcal{M}}_{t=-a}$ admits a unique smooth solution on
$a,-a+\epsilon]\times\widehat{\mathcal{M}}$.
Let
$(g(-a+\epsilon),k(-a+\epsilon))$
denote the induced Cauchy data on $\widehat{\mathcal{M}}_{t=-a+\epsilon}$. Cauchy stability implies that $(g(-a+\epsilon),k(-a+\epsilon))$ is quasi-isometric to $(g,k)$.
7.16. Propagation of the radius and Yau curvature condition
Consider the future domain
$J^{+}(\mathcal{M}_{1})\cap\widehat{\mathcal{M}}_{t=-a+\epsilon}\subset\widehat{\mathcal{M}}_{t=-a+\epsilon}.$
By assumption, the null boundary $\{u=-a\}$ has positive outgoing expansion $\mbox{tr}\chi>0$. Integrating the first variation equation along the outgoing null generators yields
(202) $\operatorname{Rad}\!\bigl(\partial(J^{+}(\mathcal{M}_{1})\cap\widehat{\mathcal{M}}_{t=-a+\epsilon})\bigr)\geq\operatorname{Rad}(\partial\mathcal{M}_{1})+\epsilon^{1/2},$
for $\epsilon>0$ sufficiently small.
Next we compare the Schoen–Yau generalized mean curvature $H-|\operatorname{tr}_{\Sigma}K|$ on boundary 22–surfaces $\Sigma\subset\widehat{\mathcal{M}}_{t}$. By the double-null structure equations and the square-integrability of the incoming shear $\underline{\hat{\chi}}$, one has
(203) $\bigl(H-|\mbox{tr}_{\Sigma}K|\bigr)(-a,0)$ $=-\frac{2|u_{\infty}|}{a}\,\mbox{tr}\underline{\chi}(u_{\infty},0)+\frac{2}{a}\!\int_{u_{\infty}}^{-a}\!|u^{\prime}|\,|\underline{\hat{\chi}}(u^{\prime},0)|^{2}\,du^{\prime}-\frac{C(\mathcal{I}^{0})}{a^{2}},$
(204) $\bigl(H-|\mbox{tr}_{\Sigma}K|\bigr)(-a,\epsilon)$ $=-\frac{2|u_{\infty}|}{a}\,\mbox{tr}\underline{\chi}(u_{\infty},0)+\frac{2}{a}\!\int_{u_{\infty}}^{-a}\!|u^{\prime}|\,|\underline{\hat{\chi}}(u^{\prime},\epsilon)|^{2}\,du^{\prime}-\frac{C(\mathcal{I}^{0})}{a^{2}}.$
Property (b) of Theorem 3.1 ensures strict positivity of the difference when passing from t=at=-a to $t=-a+\epsilon$; combined with (202) this yields
$c_{S_{-a,\epsilon}}=C_{\partial J^{+}\bigl(J^{+}(\mathcal{M}_{1})\cap\widehat{\mathcal{M}}_{t=-a+\epsilon}\bigr)}>\frac{3\pi}{2\,\operatorname{Rad}\!\bigl(J^{+}(\mathcal{M}_{1})\cap\widehat{\mathcal{M}}_{t=-a+\epsilon}\bigr)}.$
The Schoen–Yau quasi-local inequality [38] now implies that
$J^{+}(\mathcal{M}_{1})\cap\widehat{\mathcal{M}}_{t=-a+\epsilon}$
must contain an apparent horizon in its interior. This completes the proof. ∎