### v4

# Dynamical Formation of Black Holes due to Boundary Effect in Vacuum

Abstract.
We prove that a marginally outer trapped surface (MOTS) can form as a result of Einsteinian evolution in pure vacuum spacetime starting from regular initial data free of MOTSs due to pure boundary effects. We adapt a Cauchy-double-null framework and use the boundary generalized mean curvature condition for the existence of an interior MOTS imposed by the author S-T Yau in [44]. In particular, we prove that the condition of [44] can be met dynamically starting from a configuration that does not verify the same through a focusing mechanism. In fact, a very mild incoming radiation can cause a large enough generalized boundary mean curvature of an isotropically large domain so that a MOTS exists in the interior. This is fundamentally different from black hole formation by standard “gravitational collapse" and can be interpreted as the dynamical realization of a long-suspected idea in GR: MOTS can form because of “global geometry", not just quasi-local concentration of gravity/matter.
1. Introduction and the Main Theorem
In this article, we consider the $3+1$ dimensional pure vacuum Einstein’s equation and investigate the issue of large data semi-global existence and dynamical existence of an MOTS. Consider a $3+1$ dimensional globally hyperbolic $C^{\infty}$ connected oriented Lorentzian manifold $(\mathcal{M},g)$. The vacuum Einstein equations correspond to the vanishing of the Ricci curvature of $(\mathcal{M},g)$
(1) $\text{Ric}[g]=0$
and therefore the free gravity is described by the Weyl curvature components of the spacetime Riemann curvature. In the context of these equations, the formation of a black hole is one of the central issues and deserves deep attention from a rigorous analytic perspective. We briefly recall the following historical note
1.1. Background
Marginally outer trapped surfaces (MOTSs) have played a central role in General Relativity since the mid–twentieth century. Although Schwarzschild’s solution was discovered in 1915, its global causal structure was understood only decades later, when it became clear that it contains a region $\mathcal{B}$ from which no signal can reach future null infinity $\mathcal{I}^{+}$, and in which all timelike observers encounter geodesic incompleteness in finite proper time. [^2] See Sbierski [37] for a sharp formulation showing divergence of tidal forces along incomplete timelike geodesics.
For many years, these features were widely regarded as artifacts of the high degree of symmetry of the Schwarzschild metric, rather than as generic consequences of the Einstein equations. [^3] At that stage, a rigorous formulation of the Einstein initial value problem was not yet available. This view was decisively overturned by Penrose’s incompleteness theorem in the 1960s, which showed that spacetime singularities arise under minimal and physically natural assumptions. Central to this breakthrough was Penrose’s introduction of trapped surfaces [34], a quasi-local geometric notion encoding the irreversible focusing of null geodesics. Modern developments place marginally outer trapped surfaces at the heart of this theory, providing the natural boundary between dispersive and trapped gravitational dynamics.
Definition 1.
Given a $(3+1)$- dimensional Lorentzian manifold $(\mathcal{M},g)$, a closed spacelike 22-surface $S$ is caled trapped if the following two fundamental forms $\chi$ and $\underline{\chi}$ have everywhere pointwise negative expansions on $S$:
$\chi(X,Y):=g(D_{X}L,Y),\hskip 5.69054pt\underline{\chi}(X,Y):=g(D_{X}\underline{L},Y).$
Here $D$ denotes the Levi-Civita connection of $g$, $L$ and $\underline{L}$ denote a null basis of the 2-dimensional orthogonal complement of $T_{p}S$ in $T_{p}\mathcal{M}$, extended as smooth vector fields and $X,Y$ are arbitrary $S$-tangent vector fields.
In other words, a surface is called trapped if both $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ are pointwise negative everywhere on $S$. These traces signify the infinitesimal changes in area along the null generators normal to $S$, whence one can interpret trapped surfaces as closed, spacelike 22-surfaces that infinitesimally decrease in area "along any possible future direction".
Closely related to the trapped surface is the notion of MOTS. The definition of an MOTS differs from that of a trapped surface by the fact that the trace $\mbox{tr}\chi$ of the null outgoing second fundamental form $\chi$ vanishes point-wise, while $\mbox{tr}\underline{\chi}$, the trace of the null incoming second fundamental form, is point-wise negative. The formal definition is as follows
Definition 2.
Given a $(3+1)$- dimensional Lorentzian manifold $(\mathcal{M},g)$, a closed spacelike 22-surface $S$ is caled MOTS if the fundamental forms $\chi$ and $\underline{\chi}$ have everywhere zero and negative expansions on $S$, respectively i.e.,
(2) $\mbox{tr}\chi=0,~\mbox{tr}\underline{\chi}<0~on~S.$
Formally, a MOTS can be interpreted as the outermost boundary of a domain containing closed trapped surfaces in a Cauchy slice.
The incompleteness theorem is now presented.
Theorem 1.1 (Penrose Incompleteness).
Let $(\mathcal{M},g)$ be a spacetime containing a non-compact Cauchy hypersurface. If $(\mathcal{M},g)$ moreover satisfies the null energy codition and contains a closed trapped surface, it is geodesically incomplete.
The existence of an MOTS (and trapped surfaces contained in it) is a stable feature in the context of dynamics. Indeed, sufficiently small perturbations of Schwarzschild initial data must also contain such surfaces, by Cauchy stability. As such, incompleteness is not an accident, but rather a recurring theme in the dynamics of the Einstein equations.
In this article, we focus on the study by S-T Yau [44] and proving that indeed an MOTS can form in a dynamical manner, starting from a regular configuration. This is a completely new idea in the sense that we want to obtain the existence of a MOTS as a result of boundary effects in a Cauchy slice that is a product of evolution according to the vacuum Einstein equations. The key idea is that if a three-dimensional manifold M has a boundary with strongly positive mean curvature, the effect of this mean curvature can influence the internal geometry of $M$. Let us make this idea precise. Let $\mathcal{M}_{t}$ be a domain with boundary $S_{t}$ in the Cauchy slice $\mathcal{M}_{t}$. On the Cauchy slice $\mathcal{M}_{t}$, the Einstein constraint equations
(3) $\text{Scal}[h]-|k|^{2}+(\mbox{tr}_{h}k)^{2}=0,~\mbox{div }_{h}(k-\mbox{tr}_{h}kh)=0$
are verified, where $h$ is the induced metric on the slice $\mathcal{M}_{t}$ and $k$ is second fundamental form of this slice. Here $\text{Scal}[h]$ denotes the scalar curvature of the metric $h$. The boundary $S_{t}$ of the domain $\mathcal{M}_{t}\subset\mathcal{M}_{t}$ is co-dimension $2$ in the spacetime $(\mathcal{M},g)$ and hence possesses a time-like and a spacelike second fundamental form. Let $(e_{T},e_{S})$ be the orthonormal pair spanning the tangent bundle of $S_{t}$ where $e_{T}$ is time-like and $e_{S}$ is spacelike. We define the spacelike and timelike second fundamental form of $S_{t}$ as
(4) $s_{AB}:=\langle\nabla_{e_{A}}e_{S},e_{B}\rangle,k_{AB}:=\langle\nabla_{e_{A}}e_{T},e_{B}\rangle,~A=1,2$
where $\{e_{A}\}_{A=1,2}$ are the orthonormal frame tangential to $S_{t}$. If the induced metric on $S_{t}$ is denoted by $\Sigma_{AB}$, then $H=s_{AB}(\Sigma^{-1})^{AB}$ is the spacelike mean curvature of $S_{t}$ while $\kappa:=k_{AB}(\Sigma^{-1})^{AB}$ is the time-like mean curvature of $S_{t}$ also same as the trace of the restriction of second fundamental form $k_{ij}$ of $\mathcal{M}_{t}$ to $S_{t}$. In addition, we also recall the notion of radius of $\mathcal{M}_{t}$ defined by [44]. The two entities that we are interested are $H-|\kappa|$ of $S_{t}$ and the $H$-radius or Schoen-Yau radius of a domain $\mathcal{M}_{t}$ with boundary $S_{t}$. In particular, there is a sharp threshold on $H-|\kappa|$ of $S_{t}:=\partial\mathcal{M}_{t}$ in terms of $\text{Rad}(\mathcal{M}_{t})$ that allows for an existence (and non-existence) of a MOTS inside $\widetilde{\mathcal{M}}_{1}$-see theorem 2.1 for the specifics. The vital question that arises “Can one realize this condition of [44] for the existence of MOTS in the interior of a domain in a dynamical manner?" We answer affirmatively to this question in this article. Let us state the main theorem that we prove in this article
Theorem 1.2 (Main Theorem).
There exists an open class of smooth, asymptotically flat vacuum initial data for the Einstein equations, containing no trapped surfaces and no marginally outer trapped surfaces, with the following property.
The corresponding maximal Cauchy development admits a spacelike hypersurface $\Sigma_{t_{*}}$, reached in finite proper time $t_{*}>0$, that contains a compact embedded two–sphere $S\subset\Sigma_{t_{*}}$ satisfying
$\mbox{tr}\chi(S)=0,\mbox{tr}\underline{\chi}(S)<0$
that is, $S$ is a marginally outer trapped surface.
The initial data are not assumed to be small in any global norm. The formation of $S$ occurs dynamically and prior to any breakdown of the spacetime, and is driven by a boundary–induced concentration mechanism: along a suitable interior region, the generalized mean curvature
$c:=\sup_{S}\bigg(H-|\kappa|\bigg)$
crosses the Yau threshold 2.1 while the intrinsic radius of the region remains uniformly large.
In particular, the marginally outer trapped surface is not generated by forcing the outgoing null expansion to become negative, but instead arises through a dynamic realization of the Yau boundary criterion 2.1 in vacuum general relativity.
Figure 1. Schematic of the boundary mechanism yielding a MOTS through the evolution of pure vacuum Einstein equations. The main idea is to maintain the Yau’s generalized mean curvature along u=au=-a while increasing the thickness of the interior $\mathcal{M}_{1}$ uniformly in its causal future under vacuum Einstein evolution. Roughly, the mass in the isotropically large ringed region in the Cauchy slice (top view) needs to be reasonably large. The Hawking mass of each sphere foliating this region is $\gg 1$ in our study. The remarkable fact is that the interior $\mathcal{M}_{1}$ is isotropically thick. It is natural to expect a mass-length relation for the black hole formation since, due to the scale invariance of the background (Minkowski), there is no lower bound on the mass for black hole formation. The well-known expectation is that the concentration of sufficiently large mass in a suitably defined small enough domain would lead to the formation of a black hole. The main point of this study is that a large mass can be distributed within an isotropically large domain and still form MOTS.
We begin by contrasting our result with the seminal work of Christodoulou [19]. In a double null framework, Christodoulou showed that sufficiently strong incoming gravitational radiation at past null infinity—quantified through the largeness of the shear $\hat{\chi}$ in a suitable norm—forces the formation of a trapped surface in the interior purely through vacuum Einstein evolution.
The mechanism considered in the present work is fundamentally different from the gravitational collapse via the concentration of gravitational radiation (e.g., the short pulse data technique of Christodoulou [19]). We do not assume largeness of the incoming shear in the sense of [19], nor do we seek trapped surfaces generated by driving the outgoing null expansion negative. Instead, we introduce a distinct class of initial data and a different dynamical mechanism leading to the formation of a marginally outer trapped surface. The construction exploits a boundary-induced concentration phenomenon, in which the generalized mean curvature crosses the Yau threshold while the intrinsic radius of the relevant region remains uniformly controlled.
From a technical perspective, the argument rests on three complementary components: the construction of a semi-global solution to a characteristic initial value problem (where focusing of the generalized Yau mean curvature of a topological 22-sphere is achieved), the local-in-time evolution of a Cauchy problem compatible with this characteristic development, and a quantitative control of the Schoen-Yau radius under the ensuing evolution. This combination of ideas appears to be new and does not arise in previous approaches to trapped surface or MOTS formation in vacuum general relativity. A corollary of the main theorem is stated below
Corollary 1.1 (Black hole formation without trapped-surface initial data).
There exists an open class of smooth, asymptotically flat vacuum initial data for the Einstein equations that contain no trapped surfaces and no marginally outer trapped surfaces, whose maximal Cauchy development nevertheless contains an apparent horizon.
More precisely, for such initial data, the spacetime admits a marginally outer trapped surface that forms dynamically in finite time, and hence the spacetime is causal future incomplete.
It is natural to expect a mass-length relation for the black hole formation since, due to the scale invariance of the background (Minkowski), there is no lower bound on the mass for black hole formation. This issue is deeply tied to the fact that $ADM$ mass is quite rough in terms of regularity, and hence the positive mass theorem is not stable in a smooth sense (one can have black holes with arbitrarily small masses). The well-known expectation is that the concentration of sufficiently large mass in a suitably defined small enough domain would lead to the formation of a black hole. The main point of this study is that a large mass can be distributed within an isotropically large domain and still form MOTS.
The formation of the marginally outer trapped surface occurs strictly prior to any breakdown of the spacetime and does not arise from driving the outgoing null expansion negative via concentration of curvature. Instead, the mechanism is geometric: along a dynamically evolving spacelike domain $\Omega_{t}\subset\Sigma_{t}$, the generalized boundary mean curvature $H-|\kappa|$ crosses the Schoen–Yau threshold (2.1) while the intrinsic radius $\text{Rad}(\Omega_{t})$ remains uniformly controlled.
This provides the first realization, within vacuum general relativity, of a horizon formation mechanism driven by boundary geometry rather than gravitational collapse in the sense of Christodoulou. In particular, the marginally outer trapped surface arises without requiring large incoming radiation, large null shear, or short-pulse initial data. The following three are the main points that are addressed through our study, where we take this different approach of dynamical formation of the apparent horizon
(a) It decouples horizon formation from collapse.
(b) It allows for quasi-stationary or mild dynamics to trigger a horizon.
(c) It provides a rigorous PDE realization of a long-suspected idea in GR: horizons can form because of global geometry, not just quasi-local gravity or matter energy density.
From a physical point of view, this study possesses the potential to explain an important open problem in physics: the formation of supermassive black holes (mass $\approx 10^{10}M_{\odot}$). In particular, the possibility of formation through short pulse gravitational collapse as in [19] is ruled out, and our mechanism provides a potential route for such formation.
2. Overview and strategy of the proof
The first part of our argument is to establish a semi-global development of regular characteristic initial data free of trapped surfaces or MOTS. Let $(\mathcal{M},g)$ be a smooth, time-oriented, four-dimensional Lorentzian manifold. We assume that $(\mathcal{M},g)$ admits a smooth double null foliation in the following sense: there exist smooth optical functions
$(u,\underline{u}):\mathcal{M}\to\mathbb{R}\times\mathbb{R}$
such that for each fixed value of $u$ (respectively $\underline{u}$), the level set
$H_{u}:=\{p\in\mathcal{M}\,:\,u(p)=\mathrm{const}\},\qquad\underline{H}_{\underline{u}}:=\{p\in\mathcal{M}\,:\,\underline{u}(p)=\mathrm{const}\}$
is a smooth null hypersurface, which we refer to as an outgoing (respectively incoming) null hypersurface.
For each pair $(u,\underline{u})$ for which $H_{u}\cap\underline{H}_{\underline{u}}\neq\emptyset$, we define
$S_{u,\underline{u}}:=H_{u}\cap\underline{H}_{\underline{u}}.$
We assume that $S_{u,\underline{u}}$ is a smooth, embedded, spacelike 22–surface diffeomorphic to $\mathbb{S}^{2}$, and we denote by $\not{\gamma}_{u,\underline{u}}$ the Riemannian metric on $S_{u,\underline{u}}$ induced by $g$. When convenient, we abbreviate $\Sigma_{u,\underline{u}}:=(S_{u,\underline{u}},\not{\gamma}_{u,\underline{u}})$. To define angular coordinates on each $S_{u,\underline{u}}$ in a smooth way, we begin by defining angular coordinates on $S_{u_{\infty},0}$. Since this is a standard 2-sphere in Minkowki space, we can use the stereographic projection coordinates $(\theta^{1},\hskip 1.42262pt\theta^{2})$ on $S_{u_{\infty},0}$. We first extend this coordinate to the whole of $\underline{H}_{0}$ by insisting that $\not{\mathcal{L}}_{\underline{L}}\theta^{A}=0$ on $\underline{H}_{0}$ for $A=1,\hskip 1.42262pt2$ and then to the whole spacetime by insisting that, for all $u$, $\not{\mathcal{L}}_{L}\theta^{A}=0$, where $L$ initially starts normal to some $S_{u,0}$. As such we have established a coordinate system $(u,\underline{u},\theta^{1},\theta^{2})$ in a neighbourhood of the initial sphere. In these coordinates, the vectors $e_{3},\hskip 1.42262pte_{4}$ become
$e_{3}=\Omega^{-1}\left(\frac{\partial}{\partial u}+b^{A}\hskip 1.42262pt\frac{\partial}{\partial\theta^{A}}\right),\hskip 5.69054pte_{4}=\Omega^{-1}\frac{\partial}{\partial\underline{u}}$
and the metric now takes the following form:
[…]
The section that maps $p\in\mathcal{M}\mapsto\left({\left.\kern-1.2pt\theta^{1}\right|_{p}},{\left.\kern-1.2pt\theta^{2}\right|_{p}}{\left.\kern-1.2pte_{3}\right|_{p}},{\left.\kern-1.2pte_{4}\right|_{p}}\right)$ is the double null gauge that we want. We begin by decomposing curvature components and Ricci coefficients with respect to the frame $(e_{1},e_{2},e_{3},e_{4})$. Let $A,B$ take values in $\begin{Bmatrix}1,2\end{Bmatrix}$.
Let us recall the definitions of the connection coefficients in this double null framework
$\chi_{AB}:=g(D_{A}e_{4},\hskip 1.42262pte_{B}),\hskip 5.69054pt\underline{\chi}_{AB}:=g(D_{A}e_{3},e_{B}),$
$\eta_{A}:=-\frac{1}{2}g(D_{A}e_{3},\hskip 1.42262pte_{4}),\hskip 5.69054pt\underline{\eta}_{A}:=-\frac{1}{2}g(D_{A}e_{4},e_{3}),$
$\omega:=-\frac{1}{4}g(D_{4}e_{3},\hskip 1.42262pte_{4}),\hskip 5.69054pt\underline{\omega}:=-\frac{1}{4}g(D_{3}e_{4},\hskip 1.42262pte_{3}),$
$\zeta_{A}:=\frac{1}{2}g(D_{A}e_{4},\hskip 1.42262pte_{3}).$
Moreover, if $\gamma$ denotes the induced metric on $S_{u,\underline{u}}$, we make the following further decomposition:
$\chi=\hat{\chi}+\frac{1}{2}\mbox{tr}\chi\gamma,\hskip 5.69054pt\underline{\chi}=\underline{\hat{\chi}}+\frac{1}{2}\mbox{tr}\underline{\chi}\hskip 1.42262pt\gamma.$
The Weyl curvature components read
$\alpha_{AB}:=W(e_{A},e_{4},e_{B},e_{4}),\hskip 5.69054pt\underline{\alpha}_{AB}:=W(e_{A},e_{3},e_{B},e_{3}),$
$\beta_{A}:=\frac{1}{2}W(e_{A},e_{4},e_{3},e_{4}),\hskip 5.69054pt\underline{\beta}_{A}:=\frac{1}{2}W(e_{A},e_{3},e_{3},e_{4}),$
$\rho:=\frac{1}{4}W(e_{3},e_{4},e_{3},e_{4}),\hskip 5.69054pt\sigma=\frac{1}{4}\prescript{*}{}{W}(e_{3},e_{4},e_{3},e_{4}).$
We prescribe characteristic initial data on two intersecting null hypersurfaces, namely on the incoming null hypersurface
$\underline{H}_{0}:=\{\underline{u}=0\}$
and the outgoing null hypersurface
$H_{u_{\infty}}:=\{u=u_{\infty}\},$
where $u_{\infty}\in\mathbb{R}$ is fixed. In addition, the Cauchy data is provided in the interior Cauchy slice $\mathcal{M}_{1}$ as in the diagram 2. The hierarchy and size of the data that we impose on $H_{u_{\infty}}\cup\underline{H}_{0}$ differ in an essential way from those considered in [1, 5, 6, 12, 13]. In particular, we allow large (nonperturbative) radiation along one null direction, while retaining only a degenerate smallness along the transverse direction. This constitutes the first principal new feature of the present work.
From these data, we construct a semi-global causal development, which we denote by
$D_{a,1}:=[u_{\infty},-a]\times[0,1]\times S,~D^{{}^{\prime}}:=[u_{\infty},-a-\frac{1}{a}]\times[0,1]\times S$
where $S\simeq\mathbb{S}^{2}$ is a fixed reference two–sphere, $a\gg 1$ is a large parameter, and $\epsilon>0$ is a small parameter. The notation above is to be interpreted as follows: the variables $(u,\underline{u})$ range in the rectangle
$u_{\infty}\leq u\leq-a,\qquad 0\leq\underline{u}\leq 1,$
and the angular variables range in $S$. Hence $D_{a,1}$ is the spacetime region covered by the portion of the double null foliation determined by these bounds. A schematic representation of $D_{a,1}$ is provided in Figure 2.
We now introduce the associated canonical spacelike foliation. Define the function
$t:=u+\underline{u}.$
For each constant $t\in\mathbb{R}$, we denote by
$\mathcal{M}_{t}:=\{p\in\mathcal{M}\,:\,u(p)+\underline{u}(p)=t\}$
the corresponding level set. We assume that, for the range of $t$ under consideration, $\mathcal{M}_{t}$ is a smooth spacelike Cauchy hypersurface for the relevant portion of $(\mathcal{M},g)$; in particular, $\{\mathcal{M}_{t}\}$ defines a spacelike foliation compatible with the above double null foliation.
Fix $a\gg 1$ as above and consider the slice $\mathcal{M}_{t=-a}$. We define the intersection of this slice with the semi-global development by
$\mathcal{M}_{2}:=\mathcal{M}_{t=-a}\cap D_{a,1}.$
It is convenient to decompose the entire Cauchy slice $\mathcal{M}_{t=-a}$ into three regions: an “interior” region $\mathcal{M}_{1}$, lying in the interior; the interaction region $\mathcal{M}_{-a}$ defined above; and an “exterior” region $\mathcal{M}_{\mathrm{ext}}$, lying to the outside of $D_{a,1}$. In particular, we write
(6) $\mathcal{M}_{t=-a}\;=\;\mathcal{M}_{1}\;\cup\;\mathcal{M}_{2}\;\cup\;\mathcal{M}_{\mathrm{ext}}.$
The decomposition (6) isolates the portion $\mathcal{M}_{2}$ of the Cauchy hypersurface on which the large characteristic data propagate and interact, separating it from the complementary interior and exterior regions of $\mathcal{M}_{t=-a}$. One of the main parts of the current work is the proof of this semi-global development with a new hierarchy of data and concentration of the generalized Yau mean curvature along u=au=-a (or u=a1/au=-a-1/a) null cone. Now we recall the second ingredient in the construction of [44], the $H$-radius of Schoen-Yau radius $\text{Rad}(\Omega)$ of a domain $\Omega$.
Definition 3.
Let $\Gamma$ be a simply closed curve in $\Omega$ that bounds a disk in $\Omega$. We let $N_{r}(\Gamma)$ denote the set of points within a distance $r$ of $\Gamma$. We define the $H$-radius or Schoen-Yau radius of $\Omega$ with respect to $\Gamma$ as
$\text{Rad}(\Omega,\Gamma):=\sup\{r:\text{dist}(\Gamma,\partial\Omega)>r,~\Gamma\text{does not bound a disk in}N_{r}(\Gamma)\}.$
We define the Schoen-Yau radius or H-radius of $\Omega$ denoted by $\text{Rad}(\Omega)$ as follows
$\text{Rad}(\Omega):=\sup\{\text{Rad}(\Omega,\Gamma):\Gamma\text{as above}\}.$
The guiding principle behind the introduction of this radius is to quantify, in a geometrically intrinsic manner, the effective interior thickness of a domain $\Omega$, as opposed to its global diameter or volume. More precisely, the radius is designed to discriminate between domains that are uniformly thick in all directions and those that exhibit pronounced geometric anisotropy, such as long, thin, tube-like regions.
This distinction is already apparent in basic model geometries. If $\Omega\subset\mathbb{R}^{3}$ is a Euclidean ball of radius $R$, then the $H$-radius or Schoen-Yau radius satisfies
$\operatorname{Rad}(\Omega)=\frac{R}{2},$
reflecting the fact that the domain possesses comparable thickness in every direction. By contrast, if $\Omega$ is a cylindrical region of the form $\mathbb{S}^{2}_{R}\times(-L,L)$, where $\mathbb{S}^{2}_{R}$ denotes the round sphere of radius $R$, then a direct computation yields
$\operatorname{Rad}(\Omega)=\min\!\left(\frac{\pi R}{2},\,L\right).$
In particular, in the highly elongated regime $L\gg R$, the HH–radius is governed entirely by the transverse scale $R$, and is insensitive to the longitudinal extent of the domain. This behavior precisely captures the intended geometric content: the radius detects the maximal scale on which the domain remains uniformly thick, while deliberately ignoring directions along which the geometry degenerates into a thin tube.
Accordingly, the $H$-radius provides a robust quantitative measure of interior thickness that sharply distinguishes globally round geometries from elongated or neck–like configurations, a feature that will play a crucial role in the subsequent geometric and analytic arguments. With these geometric-topological notions, [44] proves a remarkable theorem on the existence of an MOTS on a Cauchy slice. More precisely, the theorem is as follows
Theorem 2.1.
[44] Let $M$ be a space-like hypersurface in a spacetime. Let $g_{ij}$ be its induced metric and $k_{ij}$ be its second fundamental form. Assume that the spacelike mean curvature $H$ of $\partial M$ is strictly greater than its time-like mean curvature $\mbox{tr}_{\partial M}k=\kappa$. Let $c:=\min_{\partial M}\bigg({H-|\kappa|}\bigg)$. If $c\geq\frac{3\pi}{2\text{Rad}(M)}$, then $M$ must admit a MOTS in its interior.
Note that this is purely a boundary effect in the sense that if the (generalized) boundary mean curvature of a large domain (in the sense of large Schoen-Yau radius) in the Cauchy slice is significantly higher, then an MOTS must exist inside. This leads to the following natural question: Can one start from a regular configuration free of any MOTSs and form an MOTS in an evolutionary manner in finite time. This statement needs to be made more precise, such as exactly what it means by a regular configuration, etc. We will do this momentarily.
Figure 2. A detailed diagram depicting the mechanism of producing a MOTS in an evolutionary manner: The schematics of the current framework: concentration of the generalized mean curvature $H-|\kappa|>0$ while increasing the radius. Consider the domain $\mathcal{M}^{1/a}_{int}\subset\mathcal{M}_{-a}$. The choice of initial data on this interior and the characteristic evolution are designed to produce quasi-isometry between $(\Phi_{T}(M^{1/a}_{int}),(\Phi^{-1}_{T})^{*}g)$ and $(M^{1/a}_{int},g)$ for T=a1/a+3/4T=-a-1/a+3/4 up to $\frac{g}{a^{\frac{3}{2}}}$ while the extra collar $\bigg(J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{-a-1/a+\epsilon}\bigg)\setminus\Phi(\mathcal{M}^{1/a}_{int})_{-a-1/a+\epsilon}$ in the causal future contributes an additional $O(1)$ thickness gain while the generalized mean curvature $H-|\kappa|$ is preserved up to $O(a^{-5/2})$ along u=a1/au=-a-1/a up to $\underline{u}=3/4$ triggering the sign flip of the entity $(H-|\kappa|)-\frac{3\pi}{2\text{Rad}(\Omega)}$ from $\mathcal{M}^{1/a}_{int}$ to $J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{-a-1/a+3/4}$ and subsequent formation of a MOTS inside the later.
2.1. Yau [44] criterion for MOTS formation
By the semi-global existence result established in the previous section, the development $D_{a,1}$ induces canonical Cauchy data on the portion $\mathcal{M}_{-a}:=\mathcal{M}_{t=-a}\cap D_{a,1}$. On the complementary regions $\mathcal{M}_{1}$ and $\mathcal{M}_{\mathrm{ext}}$, one retains the freedom to prescribe data independently, subject only to the constraint and rigidity conditions imposed by the positive mass theorem [39, 40, 41]. In particular, the data on the exterior region $\mathcal{M}_{\mathrm{ext}}$ are far more constrained than those on the interior piece $\mathcal{M}_{1}$, and a natural choice is to attach a Kerr exterior with prescribed ADM mass and angular momentum via the Corvino–Schoen gluing construction. The objective is then to glue the data on the three components
$\mathcal{M}_{\mathrm{1}},\qquad\mathcal{M}_{2},\qquad\mathcal{M}_{\mathrm{ext}},$
so as to produce a smooth global Cauchy data set on $\mathcal{M}_{t=-a}$ satisfying the Einstein constraint equations. Note that in the context of dynamical trapped surface formation, [28] first constructed an initial data set free of trapped surfaces for the Cauchy problem compatible with Christodoulou’s short pulse data type. Later, [11] constructed an initial data set for the dynamical formation of the MOTS and addressing a spacetime Penrose inequality.
The guiding idea is to choose the initial data on the null cones $H_{u_{\infty}}$, $\underline{H}_{0}$, and on the interior Cauchy region $\mathcal{M}_{1}$ in such a way that the boundary of $\mathcal{M}_{1}$ fails to satisfy the MOTS condition of [44], namely,
(7) $c_{\partial\mathcal{M}_{1}}:=\min_{\partial\mathcal{M}_{1}}\big[H-|\kappa|\big]<\frac{3\pi}{2\,\mathrm{Rad}(\mathcal{M}_{1})}.$
The strict inequality (7) ensures that $\mathcal{M}_{1}$ contains no MOTS. This prescription is perfectly compatible with Causality since the data in the interior $\widetilde{\mathcal{M}}_{1}$ is not affected by that on $H_{u_{\infty}}$ and $\underline{H}_{0}$. After having proven that throughout the semi–global region $D_{a,1}$ one has
$\mbox{tr}\chi>0,\qquad\mbox{tr}\underline{\chi}<0,$
and the fact that the exterior Kerr portion $\mathcal{M}_{\mathrm{ext}}$ is also free of closed trapped surfaces, it follows that no MOTS occurs anywhere in the spacetime before or up to the slice t=at=-a.
The vital point to note here is that it is rather difficult to simultaneously control the generalized mean curvature $c$ and the radius $\text{Rad}(M)$ such that $c\cdot\text{Rad}(M)\geq\frac{3\pi}{2}$. This is precisely why we consider the two-step process. Let us revert our attention back to the diagram 2. The initial data is provided on the hypersurfaces $\underline{u}=0$ (or $\underline{H}_{0}$), $u=u_{\infty}$ (or $H_{u_{\infty}}$), and $\mathcal{M}_{1}$ (or $\widetilde{\mathcal{M}}_{1}$). The strategy that we adopt here is that first, we want to control the generalized mean curvature $c$ along u=au=-a. In the next step, we will control the radius of slice $\mathcal{M}_{1}$ (or in pactice we will consider a modified domain $\mathcal{M}^{1/a}_{int}$) as it evolves in the future. Notice that along u=au=-a, evolution from $s_{-a,0}$ to $S_{-a,\epsilon}$ indicates moving in the outgoing direction and hence naively expect a non-concentration of the generalized mean curvature $c$. The first idea is whether one can preserve the generalized mean curvature along u=au=-a up to error terms $O(a^{-5/2})$. Then, naturally, one can impose the initial condition on $\mathcal{M}_{1}$ in such a way that 7 holds, but since along u=au=-a, the radius of the bounding Cauchy slices increases, the condition as stated in the theorem 2.1 can be satisfied for $J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{t=-a-1/a+\epsilon}$. Therefore, the two main challenges in this study are the proof of the semi-global development with the choice of a new hierarchy of initial data and controlling the Cauchy evolution upto time $a+\epsilon$ for $\epsilon=O(1)$. We illustrate these steps next.
2.2. Focusing $H-|\mbox{tr}_{S}k|$
Let us now try to understand how this mechanism would unfold in a heuristic manner. First, recall that in the double null gauge that we are considering here
(8) $e_{4}=\Omega^{-1}\partial_{\underline{u}},~e_{3}=\Omega^{-1}\bigg(\partial_{u}+b^{A}\partial_{\theta^{A}}\bigg).$
The corresponding connection coefficients that appear in this study are $(\hat{\chi},\mbox{tr}\chi,\underline{\hat{\chi}},\mbox{tr}\underline{\chi},\omega,\underline{\omega},\eta,\underline{\eta})$. In this coordinate, one may explicitly compute the $H-|\kappa|$ for a topological sphere $S_{u,\underline{u}}$ constituting the double null foliation
(9) $H-|\kappa|=H-|\mbox{tr}_{S}K|=\frac{1}{2}\bigg(\mbox{tr}\chi-\mbox{tr}\underline{\chi}\bigg)-\frac{1}{2}\bigg|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\bigg|.$
This leads us to directly control $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ instead of working with separate evolution equations for $H$ and $\kappa$. The main equations that we focus on are the following
(10) $\nabla_{3}\mbox{tr}\underline{\chi}+\frac{1}{2}(\mbox{tr}\underline{\chi})^{2}=-|\underline{\hat{\chi}}|^{2}-2\underline{\omega}\mbox{tr}\underline{\chi},$
(11) $\nabla_{4}\mbox{tr}\chi+\frac{1}{2}(\mbox{tr}\chi)^{2}=-|\hat{\chi}|^{2}-2\omega\mbox{tr}\chi,$
(12) $\nabla_{4}\hat{\underline{\chi}}+\frac{1}{2}\mbox{tr}\chi\hat{\underline{\chi}}=\nabla\hat{\otimes}\underline{\eta}+2\omega\hat{\underline{\chi}}-\frac{1}{2}tr\underline{\chi}\hat{\chi}+\underline{\eta}\hat{\otimes}\underline{\eta}$
Now, the scaling hierarchy of norm that we will use, the lapse $\Omega$, the shift $b$, and the Ricci coefficients will verify the following estimates
[…]
(14) $||\eta,\underline{\eta}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|^{2}},~||\omega||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|^{2}},~||\hat{\chi}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{-\frac{1}{2}}}{|u|}.$
and the Weyl curvature scaling
(15) $|\alpha|=O(a^{-\frac{1}{2}}|u|^{-1}),~|\beta|=O(|u|^{-2}),~|\underline{\beta}|=O(a|u|^{-4}),~|(\rho,\sigma)|=O(a^{\frac{1}{2}}|u|^{-3}),~|\underline{\alpha}|=O(a^{\frac{3}{2}}|u|^{-5})$
Here, note importantly that we will have to prove that the involved constants in $\lesssim$ depend only on the initial scale-invariant data of the respective Ricci coefficients, and this hierarchy is propagated in the domain of spacetime constructed via semi-global development. With this scaling hierarchy, we can integrate $\mbox{tr}\underline{\chi}$ along the incoming direction, i.e., work with the following equation
(16) $\nabla_{3}\mbox{tr}\underline{\chi}+\frac{1}{2}(\mbox{tr}\underline{\chi})^{2}=-|\underline{\hat{\chi}}|^{2}-2\underline{\omega}\mbox{tr}\underline{\chi}.$
Now use the fact that $e_{3}=\Omega^{-1}\bigg(\partial_{u}+b^{A}\partial_{\theta^{A}}\bigg)$ and write
(17) $\nabla_{3}(|u|\mbox{tr}\underline{\chi})=-\bigg(\mbox{tr}\underline{\chi}+\frac{2}{|u|}\bigg)\frac{|u|\mbox{tr}\underline{\chi}}{2}-|u||\underline{\hat{\chi}}|^{2}-2|u|\underline{\omega}\mbox{tr}\underline{\chi}+(1-\Omega^{-1})\mbox{tr}\underline{\chi}$
which, upon integration and using the bounds (13) one obtains for every $\underline{u}\in[0,\epsilon]$ (provided one have the semi-global construction already)
(18) $\mbox{tr}\underline{\chi}(-a,\underline{u})=\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\underline{u})}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\underline{u})du^{{}^{\prime}}+\frac{1}{a}\int_{u_{\infty}}^{-a}\bigg(\mbox{tr}\underline{\chi}+\frac{2}{|u^{{}^{\prime}}|}\bigg)du^{{}^{\prime}}+\mathscr{E},$
where the error term $\mathscr{E}$ is $O(a^{-\frac{5}{2}})$. Now we need to control $\widetilde{\mbox{tr}\underline{\chi}}:=\mbox{tr}\underline{\chi}+\frac{2}{|u|}$. Recall the following equation verified by $\widetilde{\mbox{tr}\underline{\chi}}$
$\nabla_{3}\widetilde{\mbox{tr}\underline{\chi}}+\mbox{tr}\underline{\chi}\widetilde{\mbox{tr}\underline{\chi}}=\frac{2}{|u|^{2}}(\Omega^{-1}-1)+|\widetilde{\mbox{tr}\underline{\chi}}|^{2}+2\underline{\omega}\mbox{tr}\underline{\chi}-|\underline{\hat{\chi}}|^{2}$
and subsequently
$\nabla_{3}(|u|^{2}\widetilde{\mbox{tr}\underline{\chi}})=2(\Omega^{-1}-1)+2|u|^{2}\underline{\omega}\mbox{tr}\underline{\chi}-|u|^{2}|\underline{\hat{\chi}}|^{2}$
which yields
(19) $|u|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u,\underline{u})=|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},\underline{u})-\int_{u_{\infty}}^{u}|u^{{}^{\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime}}+O(\frac{a^{\frac{1}{2}}}{|u|}).$
Substitute (19) into (18) and obtain
(20) $\mbox{tr}\underline{\chi}(-a,\underline{u})=\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\underline{u})}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},\underline{u})}{a^{2}}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\underline{u})du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E},$
where the error term $\mathscr{E}=O(a^{-\frac{5}{2}})$ and negligible for sufficiently large $a\gg 1$. Now we control $\mbox{tr}\chi$ through integrating along $\nabla_{4}$ direction. Recall that the Raichoudhury equation
(21) $\nabla_{4}\mbox{tr}\chi+\frac{1}{2}(\mbox{tr}\chi)^{2}=-|\hat{\chi}|^{2}-2\omega\mbox{tr}\chi.$
But the presence of $|\hat{\chi}|^{2}$ term with a negative sign can reduce $\mbox{tr}\chi$. As it turns out, controlling the size of $\epsilon$ and the new hierarchy where $\hat{\chi}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}|u|^{-1}$, the reduction of $\mbox{tr}\chi$ can be controlled and this potentially dangerous term contributes to a negligible amount. Integration and using the bounds 13 yield for $0<\epsilon\leq 1$
(22) $\mbox{tr}\chi(u,\epsilon)=\mbox{tr}\chi(u,0)+\int_{0}^{\epsilon}\bigg(-\frac{1}{2}(\mbox{tr}\chi)^{2}-|\hat{\chi}|^{2}-2\omega\mbox{tr}\chi\bigg)du^{{}^{\prime}}\approx\mbox{tr}\chi(u,0)-\frac{4\epsilon}{|u|^{2}}.$
Here $\approx$ means up to terms that decay strictly faster than $|u|^{-2}$. Note that these terms can be made much smaller than the $O(|u|^{-2})$ terms and therefore can be ignored at a heuristic level. Now we are at a place to actually compute $H-|\kappa|$ at $(-a,\epsilon)$ and (a,0)(-a,0). Using (9) and absorbing $\frac{1}{2}$ to the left, we obtain
(23) $2\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=\bigg(\mbox{tr}\chi-\mbox{tr}\underline{\chi}\bigg)(-a,\epsilon)-\bigg|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\bigg|(-a,\epsilon)$
$=\bigg(\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}-\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},\epsilon)}{a^{2}}$
$+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}+\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg)$
$-\bigg|\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},\epsilon)}{a^{2}}$
$-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg|$
and similarly
(24) $2\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,0)$ $=\bigg(\mbox{tr}\chi(-a,0)-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}-\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg)$
$-\bigg|\mbox{tr}\chi(-a,0)+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg|.$
Integrating the $\nabla_{4}$ equation one can control $|\underline{\hat{\chi}}|^{2}(u,\epsilon)$ in terms of $|\underline{\hat{\chi}}|^{2}(u,0)$. Recall the following $\nabla_{4}$ equation verified by $\underline{\hat{\chi}}$
(25) $\nabla_{4}\hat{\underline{\chi}}+\frac{1}{2}\mbox{tr}\chi\hat{\underline{\chi}}=\nabla\hat{\otimes}\underline{\eta}+2\omega\hat{\underline{\chi}}-\frac{1}{2}\mbox{tr}\underline{\chi}\hat{\chi}+\underline{\eta}\hat{\otimes}\underline{\eta}$
implying
(26) $\nabla_{4}\bigg[|u||\underline{\hat{\chi}}|^{2}\bigg]=-|u|\mbox{tr}\chi|\underline{\hat{\chi}}|^{2}+2|u|\underline{\hat{\chi}}\nabla\otimes\underline{\eta}+4|u|\omega|\underline{\hat{\chi}}|^{2}-|u|\mbox{tr}\underline{\chi}\hat{\chi}\cdot\underline{\hat{\chi}}+2|u|\underline{\hat{\chi}}\underline{\eta}\otimes\underline{\eta}$
which, after integration and utilizing the estimates (13)
(27) $|u||\underline{\hat{\chi}}|^{2}(u,\epsilon)\approx|u||\underline{\hat{\chi}}|^{2}(u,0)+\frac{C\epsilon a^{\frac{3}{2}}}{|u|^{4}}+\frac{C\epsilon a^{\frac{3}{4}}}{|u|^{4}}+\frac{C\epsilon a^{\frac{3}{4}}}{|u|^{3}}+\frac{C\epsilon a^{\frac{7}{4}}}{|u|^{5}}.$
Therefore, $|u||\underline{\hat{\chi}}|^{2}(u,\epsilon)\approx|u||\underline{\hat{\chi}}|^{2}(u,0)$ up to a negligible error term at least $O(a^{-5/2})$ since each of the error terms decay much faster than $|u|^{-2}$. Therefore, we observe the following
(28) $2\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=$ $=\bigg(\mbox{tr}\chi(-a,0)-\frac{4\epsilon}{a^{2}}-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}-\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg)$
$-\bigg|\mbox{tr}\chi(-a,0)-\frac{4\epsilon}{a^{2}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg|$
Now, our goal is to obtain a strictly positive lower bound for $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)$. But the second term can be potentially problematic. Note on the other hand that we prescribe regular data on $H_{u_{\infty}}$. Given $\underline{\hat{\chi}}$ on $\underline{H}_{0}$, the remaining Ricci coefficients are determined on $\underline{H}_{0}$. Therefore, one ought to integrate the $\nabla_{3}$ transport equation for $\mbox{tr}\chi$ to estimate $\mbox{tr}\chi(-a,0)$ using the data $\mbox{tr}\chi(u_{\infty},0)$. Subsequently, we would want
(29) $\mbox{tr}\chi(-a,0)+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}-\frac{4\epsilon}{a^{2}}$
$-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}=O(a^{-\frac{5}{2}})$
which is compatible with the data choice. Now we ought to obtain $\mbox{tr}\chi(-a,0)$ in terms of its data $\mbox{tr}\chi(u_{\infty},0)$ since it is prescribed on $u=u_{\infty}$ hypersurface. Direct integration of the equation
$\nabla_{3}\mbox{tr}\chi+\frac{1}{2}tr\underline{\chi}\mbox{tr}\chi=2\underline{\omega}\mbox{tr}\chi+2\text{div}\eta+2|\eta|^{2}+2\rho-\hat{\chi}\cdot\hat{\underline{\chi}}$
yields
(30) $\mbox{tr}\chi(-a,0)=\frac{|u_{\infty}|\mbox{tr}\chi(u_{\infty},0)}{a}+2\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\bigg(\mbox{div }\eta+\rho-\frac{1}{2}\hat{\chi}\cdot\underline{\hat{\chi}}\bigg)(u^{{}^{\prime}},0)du^{{}^{\prime}}+O(a^{-\frac{5}{2}}).$
Importantly, $\mbox{div }\eta+\rho-\frac{1}{2}\hat{\chi}\cdot\underline{\hat{\chi}}$ is nothing but the negative of the mass aspect function $\mu$. Consequently
(31) $\mbox{tr}\chi(-a,0)=\frac{|u_{\infty}|\mbox{tr}\chi(u_{\infty},0)}{a}-\frac{2}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}+O(a^{-\frac{5}{2}}).$
Now this forces the choice of $\mbox{tr}\chi(u_{\infty},0)$ since we want to keep the data $\underline{\hat{\chi}}$ on $\underline{H}_{0}$ free. Substituting $\mbox{tr}\chi(-a,0)$ from (30) into (29) yields
$\frac{|u_{\infty}|}{a}\mbox{tr}\chi(u_{\infty},0)-\frac{4\epsilon}{a^{2}}-\frac{2}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},0)}{a^{2}}$
$\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}=\mathscr{E},$
where $\mathscr{E}=O(a^{-\frac{5}{2}})$. The idea here is that we prescribe the data on $H_{u_{\infty}}$ in terms of data on $H_{0}$ that is consistent with the null structure equations. This is because once we prescribe $\underline{\hat{\chi}}$ on $H_{0}$, the remaining Ricci coefficients are determined. This yields
(32) $\bigg(\mbox{tr}\chi(u_{\infty},0)+\mbox{tr}\underline{\chi}(u_{\infty},0)\bigg)-\frac{4\epsilon}{|u_{\infty}|a^{2}}-\frac{2}{|u_{\infty}|}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu du^{{}^{\prime}}+\frac{a}{|u_{\infty}|}\bigg[\frac{|u_{\infty}|^{2}}{a^{2}}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty})$
(33) $-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}\bigg]=\frac{a}{|u_{\infty}|}\mathscr{E}.$
One aims for the following for $\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty})$
(34) $\frac{|u_{\infty}|^{2}}{a^{2}}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty})=\frac{1}{10a}\bigg[\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}\bigg]$
since $\underline{\hat{\chi}}$ on $\underline{H}_{0}$ is free data and therefore
(35) $\bigg(\mbox{tr}\chi(u_{\infty},0)+\mbox{tr}\underline{\chi}(u_{\infty},0)\bigg)=\frac{4\epsilon}{|u_{\infty}|a^{2}}+\frac{2}{|u_{\infty}|}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{9}{10|u_{\infty}|}\bigg[\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}$
$+\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}\bigg]+\frac{a}{|u_{\infty}|}\mathscr{E}$
Now, in terms of the scaling (13)
(36) $\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}=O(a^{-\frac{1}{2}}),~\mathscr{E}=O(a^{-\frac{5}{2}}),~\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}+\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}=O(a^{-1}),$
which, in light of the estimates
(37) $|\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty})|=O(|u_{\infty}|^{-2})~i.e.,~\bigg|\mbox{tr}\underline{\chi}(u_{\infty})+\frac{2}{|u_{\infty}|}\bigg|=O(|u_{\infty}|^{-2})$
yields the following asymptotics of $\mbox{tr}\chi$
(38) $\bigg|\mbox{tr}\chi(u_{\infty})-\frac{2}{|u_{\infty}|}\bigg|=O(a^{-\frac{1}{2}}|u_{\infty}|^{-1})$
and a more precise value of $\mbox{tr}\chi(u_{\infty})$ would be
(39) $\mbox{tr}\chi(u_{\infty})=\frac{2}{|u_{\infty}|}+\frac{4\epsilon}{|u_{\infty}|a^{2}}+\frac{2}{|u_{\infty}|}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}|\mu(u^{{}^{\prime}},0)du^{{}^{\prime}}-\frac{9}{10|u_{\infty}|}\bigg[\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}$
$+\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}\bigg]$
up to an error term $O(a^{-\frac{3}{2}}|u_{\infty}|^{-1})$ and $a\gg 1$. Recall the following facts about the characteristic data on $\underline{H}_{0}$
  1. (1)
    The conformal class/metric on the corner sphere $S_{u_{\infty},0}$
  2. (2)
    The incoming shear $\underline{\hat{\chi}}$ prescribed on the null hypersurface $\underline{H}_{0}$.
  3. (3)
    A gauge normalization, for example: $\Omega$ fixed on $S_{u_{\infty},0}$, together with a choice of the shift vector $b$ along the initial null hypersurface $\underline{H}_{0}$.
There data are precisely provided on $\underline{H}_{0}$ and on the corner sphere $S_{u_{\infty},0}$. See chapter $2$ of [19]for the related concepts in their framework. This would lead to the following two expressions for the generalized mean curvature
(40) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,0)$ $=-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}$
(41) $+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}$
and
(42) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)$ $=-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}$
$+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}$
and the error terms $\mathscr{E}$ decay at least $O(a^{-\frac{5}{2}})$ and therefore negligible compared to $O(a^{-1})$ term $\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}$ and $O(a^{-2})$ terms involving $\underline{\hat{\chi}}$. The two main leading order terms that we are concerned with are the $O(a^{-1})$ term and $O(a^{-2})$ term. In addition note that these estimates are stable up to at least $O(a^{-\frac{5}{2}})$ upon perturbing $u$ up to a factor $1/a$.
2.3. Interior Cauchy development and compatibility with the characteristic region
A central component of the construction is the controlled Cauchy evolution of the interior region $J^{+}(\mathcal{M}_{1})$ over a fixed time interval of length $O(1)$, starting from initial data that are pointwise small but supported on a spatial domain of diameter $\sim a$. The data are arranged so as to be compatible, along a common interface sphere, with the characteristic development constructed in Section 3, while the far exterior is completed by a Kerr end with prescribed ADM parameters. We now formulate this precisely.
Let $\mathcal{M}_{-a}$ be a smooth spacelike Cauchy hypersurface, decomposed as
$\mathcal{M}_{-a}=\mathcal{M}_{1}\;\cup\;\mathcal{M}_{2}\;\cup\;\mathcal{M}_{\mathrm{ext}},\qquad\partial\mathcal{M}_{1}=S_{-a,0},$
where:
  • $\mathcal{M}_{1}$ is a connected interior region with $H$-radius or Schoen-Yau radius comparable to $a$;
  • $\mathcal{M}_{2}\subset D_{a,1}$ is the portion contained in the characteristic development region constructed earlier, foliated by a double–null optical pair $(u,\underline{u})$ satisfying $t=u+\underline{u}=-a$ on $\mathcal{M}_{-a}$;
  • $\mathcal{M}_{\mathrm{ext}}$ denotes the asymptotic exterior region.
Fix a truncated characteristic subdomain $D^{\prime}\subset D_{a,1}$ bounded by
$u=-a-\tfrac{1}{a},\qquad\underline{u}=0,\qquad\underline{u}=1,\qquad u=u_{\infty},$
and define the enlarged interior domain
$\mathcal{M}^{1/a}_{\mathrm{int}}:=\mathcal{M}_{1}\;\cup\;\bigl(\mathcal{M}_{2}\setminus(D^{\prime}\cap\mathcal{M}_{2})\bigr).$
We prescribe smooth vacuum Cauchy data $(g,k)$ on $\mathcal{M}^{1/a}_{\mathrm{int}}$ with the following quantitative bounds: for some fixed integer $s\leq 3$,
(43) $\|k\|_{L^{\infty}(\mathcal{M}^{1/a}_{\mathrm{int}})}+\|\partial(g-\delta)\|_{L^{\infty}(\mathcal{M}^{1/a}_{\mathrm{int}})}$ $\leq C\,a^{-3/2},$
(44) $\|k\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{\mathrm{int}})}+\|\partial(g-\delta)\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{\mathrm{int}})}$ $\leq C\,a^{-3/2},$
where $H^{m}_{\mathrm{ul}}$ denotes the uniformly local Sobolev norm defined by
$\|F\|_{H^{m}_{\mathrm{ul}}(\Omega)}:=\sup_{p\in\Omega}\|F\|_{H^{m}(B_{1}(p))},$
with $B_{1}(p)$ the unit geodesic ball in the induced metric. In addition, the lapse and shift associated with the ADM decomposition satisfy
(45) $\|N-1\|_{L^{\infty}}+\|\nabla N\|_{L^{\infty}}+\|X\|_{L^{\infty}}\leq C\,a^{-3/2}\qquad\text{on }\mathcal{M}^{1/a}_{\mathrm{int}}.$
The spacetime metric in a neighborhood of $\mathcal{M}_{-a}$ is expressed in spacetime harmonic gauge, $\Box_{g}x^{\mu}=0$, and in ADM form
$\mathsfbf g=-N^{2}dt^{2}+g_{ij}\,(dx^{i}+X^{i}dt)(dx^{j}+X^{j}dt),$
so that the vacuum Einstein equations reduce to a quasilinear hyperbolic system for $g_{\mu\nu}$. In this gauge, local existence and continuation criteria depend only on uniformly local Sobolev norms of the initial data and are stable under perturbations of size $O(a^{-3/2})$ (cf. [27]).
We impose a compatibility condition across the interface sphere $S_{-a,0}=\partial\mathcal{M}_{1}$ with the data induced from the characteristic development $D_{a,1}$. Precisely, letting $(g_{\mathrm{char}},k_{\mathrm{char}})$ denote the Cauchy data induced on $\mathcal{M}_{-a}\cap D_{a,1}$, we require:
  1. (1)
    The first and second fundamental forms induced on $S_{-a,0}$ by $(\mathcal{M}_{1},g,k)$ and by $(\mathcal{M}_{2},g_{\mathrm{char}},k_{\mathrm{char}})$ agree exactly.
  2. (2)
    All tangential covariant derivatives along $S_{-a,0}$ up to order $N$ (for some fixed $N\gg 1$) coincide.
To be completely rigorous, this is accomplished by further decomposing $\mathcal{M}_{1}$ into a compact subset $\widetilde{\mathcal{M}}_{1}\subset\mathcal{M}_{1}$ and the collar region $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$. In the compact set $\widetilde{\mathcal{M}}_{1}$, we provide the data $(g,k)$ which has $H$-radius or Schoen-Yau radius $\text{Rad}_{g}$ while the data in the collar $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ is the transition region that smoothly matches the data with the characteristic development ($D_{a,1}$) induced data on $\mathcal{M}_{2}$. In particular, the collar $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ has thickness $O(\widetilde{\epsilon}a),~\widetilde{\epsilon}\in[\frac{1}{1000},\frac{1}{100}]$ and $(g,k)$ on the compact subset $\widetilde{\mathcal{M}}_{1}$ has the estimate
(46) $\|\partial^{k}k\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\leq C\,a^{-k-3/2},~\|\partial^{k+1}g\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-1}$
$\|\partial^{k}k\|_{H^{s-1}_{\mathrm{ul}}(\widetilde{\mathcal{M}}_{1})}\leq C\,a^{-k-3/2},\|\partial^{k+1}g\|_{H^{s-1}_{\mathrm{ul}}(\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-1},~k\geq 0$
where on the collar $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ the estimate (same estimate for $K$) reads
(47) $\|\partial^{k}k\|_{L^{\infty}(\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-3/2},$
(48) $\|\partial^{k}k\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-3/2},~k\geq 0$
with $\partial g$ smoothly interpolated in the collar between $\widetilde{\mathcal{M}}_{1}$ (verifying (46)) and the estimate on $\mathcal{M}_{2}$ induced by the characteric development (where in harmonic coordinates, $\partial g||\lesssim a^{-\frac{3}{2}}$) along with the same estimates for the lapse and the shift. If these are understood, if there is no further confusion, we will not separately work with this further decomposition.
Remark 1.
Note that the largeness of the data is contained in the constants $C$ (which are independent of $a$ and are $O(1)$) and the smallness is encoded in the inverse power of $a$. For $a\gg 1$, the slice t=at=-a can be considered to be at a reverse late time, where one utilizes the decay ‘in time’ property of wave equations to build a large data theory where the non-linear terms are typically subdominant compared to the linear counterparts due to larger decay. This is the essential large-small control in the long time dynamics of Einstein equations.
Remark 2.
 [^4] Xuantao Chen brought this to our attention
Notice that this data verify the scale-critical estimate $\partial^{3/2}(\mathsfbf{g}-m)||_{L^{2}(\mathcal{M}_{1})}=O(1)$, $m$ being the Minkowski metric and $\mathsfbf{g}$ is the physical spacetime metric in harmonic gauge, since one derivative here costs $a^{-1}$ at best over the large domain of size $O(a)$ under study. Therefore the scale-critical norm $\dot{H}^{3/2}$ is $O(1)$ in our study. An interesting question would be to understand how small the $\dot{H}^{3/2}$ norm of $\mathsfbf{g}-m$ can be in order for the MOTS formation result to persist. This would help one understand if Minkowski space is stable under scale-critical small perturbations.
[…]
$\mathcal{M}_{1}\cup\mathcal{M}_{2}.$
Finally, we complete the data in the far exterior region $\mathcal{M}_{\mathrm{ext}}$ by a Corvino–Schoen type gluing to a Kerr initial data set outside a compact set. The glued end has ADM parameters satisfying
$m_{\mathrm{ADM}}\sim a^{1/2},\qquad|J|=O(a),$
[…]
The two analytic tasks are therefore separated as follows: (i) a controlled quasilinear hyperbolic evolution of the interior data over a time interval of length $O(1)$ (independent of $a$) in spacetime harmonic gauge, with deformation tensor and ADM variables remaining uniformly small, and (ii) a quantitative lower bound for the HH–radius along the evolved interior slices, treated in the next section.
2.4. Radius Comparison
In the previous section we showed that along the outgoing null hypersurface $H_{-a}$ the generalized mean curvature quantity $H-|\mbox{tr}_{\Sigma}K|$ is preserved up to errors of size $O(|u_{\infty}|^{-2}+a^{-5/2})$, which are negligible in the large–$a$ regime under consideration. Although monotonic decay of this quantity along outgoing null directions might be expected a priori, the admissible open class of characteristic initial data constructed here ensures that it remains effectively constant along $H_{-a}$.
The subsequent objective is to obtain quantitative control of the Schoen–Yau (or HH–) radius of the interior region of the Cauchy slice. This constitutes the central geometric step in the argument. More precisely, one must prove that the Schoen–Yau radius of the evolved interior domain
$J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}$
strictly exceeds the initial Schoen–Yau radius of $\mathcal{M}^{1/a}_{\mathrm{int}}$, and moreover to derive a sharp lower bound for this increase in terms of the large parameter $a$. We outline the geometric mechanism at a heuristic level here; the complete quantitative argument is given later, see in particular Proposition 7.2.
We refer to Figure 2. Consider first the compact manifold with boundary $\mathcal{M}_{1}$, whose boundary is the interface sphere $S_{-a,0}$. One must prescribe Cauchy data $(g,k)$ on $\mathcal{M}_{1}$ that are compatible across $S_{-a,0}$ with the data induced on the adjacent region $\mathcal{M}_{2}$ by the characteristic development in the slab $D_{a,1}$. The resulting data are then matched in the exterior region $\mathcal{M}_{3}$ with a Kerr end. This matching is achieved by a localized Corvino–Schoen type gluing construction, carried out in Section 6, which produces a global Cauchy data set on the slice $\mathcal{M}_{-a}$ satisfying the constraint equations and agreeing with the characteristic data and the Kerr data in their respective domains.
The glued initial data are then evolved from $\mathcal{M}_{-a}$ to the nearby slice $\mathcal{M}_{-a-1/a+\epsilon}$. For $\epsilon>0$ sufficiently small, depending only on the constructed initial norms, local well–posedness yields a unique spacetime development in this time interval. The data are pointwise small in the precise sense required for compatibility with the characteristic solution: the second fundamental form, shift vector field, and lapse gradient satisfy
$|k|+|X|+|\nabla N|=O(a^{-3/2})$
on $\mathcal{M}_{-a}$, while the associated ADM mass is of order $a^{1/2}$ and therefore large in the regime $a\gg 1$. In view of the quasi-linear wave equations, the allowable time of evolution may in fact be chosen uniformly of order one for example, in spacetime harmonic gauge.
Let $\Phi_{t}$ denote the flow generated by the time vector field $\partial_{t}$, which is globally defined by global hyperbolicity. Our objective is to obtain quantitative control of the transported Riemannian metric $(\Phi_{t}^{-1})^{*}g$ on the interior future domain
$J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon},$
which will be used to estimate the geometric radius and related interior quantities on the evolved slice.
During the Cauchy evolution, it is not a priori excluded that the transported metric $(\Phi_{t}^{-1})^{*}g$ on the future interior domain
$J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}$
may decrease relative to its initial size, even though $(\Phi_{t}^{-1})^{*}g$ remains quasi–isometric to the initial metric on $\mathcal{M}^{1/a}_{\mathrm{int}}$ under the smallness assumptions on the deformation tensor. At the same time, the evolved image $\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$ is a proper subset of $J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}$. Consequently, the collar region
$\Big(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}\Big)\setminus\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$
has strictly positive thickness and therefore contributes a positive amount to the geometric size measured by the Schoen–Yau (HH–) radius. The objective of this section is to show that this collar contribution dominates any possible metric contraction in the interior region; see Figure 4 for a schematic representation.
The argument relies on precise pointwise control of the second fundamental form $k$ on $\mathcal{M}_{1}$, dictated by the matching requirements with the data on $\mathcal{M}_{2}$ arising from the characteristic development in $D_{a,1}$. In the adapted frame $(T,S,e_{A})$ one has
$k_{ST}=0,\qquad|k_{SA}|+|k_{AT}|+|k_{AB}|+|k_{SS}|\;\lesssim\;a^{-3/2},$
and hence
$|k|\lesssim a^{-3/2}.$
The critical component is the tangential block $k_{AB}$, which admits the null decomposition
[…]
Individually, the traces $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ are of size $O(a^{-1})$ near the interface sphere $S_{-a,0}$ and therefore represent potentially dangerous terms. A naive estimate based solely on these trace bounds would yield $|k|=O(a^{-1})$ on $\mathcal{M}_{-a}$. Standard short–time existence theory for quasilinear wave systems would then permit a metric variation of the same order, comparable to the geometric gain coming from the collar region. Therefore, we need to turn to refined estimates.
A priori, quasilinear wave energy estimates together with direct integration of the metric transport equation along timelike and null directions suggests the comparison bound
(49) $g\Big(1-\frac{C\epsilon}{a}\Big)\;\leq\;(\Phi_{t}^{-1})^{*}g\;\leq\;g\Big(1+\frac{C\epsilon}{a}\Big)$
on the transported interior region $\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$, for a universal numerical constant $C>0$. Correspondingly, the geometric thickness contributed by the collar region
$\Big(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}\Big)\setminus\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$
is expected to satisfy a lower bound of the form
(50) $\text{Rad}_{\mathrm{collar}}\;\geq\;\text{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})\Big(1+\frac{C\epsilon}{a}\Big),$
possibly with a different constant $C$. At this level of precision, the potential metric contraction in (49) and the collar gain in (50) occur at the same order, and no effective lower bound for the radius of $J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}$ can be deduced.
The decisive structural improvement arises from the refined behavior of the combined null expansion $\mbox{tr}\chi+\mbox{tr}\underline{\chi}$. Owing to the matching construction and the null structure equations (see (29)), a cancellation occurs near the interface sphere $S_{-a,0}$ which yields the sharper estimate
$|\mbox{tr}\chi+\mbox{tr}\underline{\chi}|=O\!\big(a^{-5/2}\big),$
rather than the individually expected $O(a^{-1})$ bounds. This strengthened control is precisely what enforces the required focusing of the generalized mean curvature $H-|\mbox{tr}_{\Sigma}K|$ and breaks the apparent balance between interior metric loss and collar gain.
The cancellation mechanism described above yields the sharpened bound $|k|\lesssim a^{-3/2}$, consistent with the estimates for the remaining Ricci coefficients and stable under the matching construction. As a consequence, the metric transport estimates improve to
(51) $g\Big(1-\frac{C\epsilon}{a^{3/2}}\Big)\;\leq\;(\Phi_{t}^{-1})^{*}g\;\leq\;g\Big(1+\frac{C\epsilon}{a^{3/2}}\Big)$
on the evolved interior image $\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$, for a universal numerical constant $C>0$. In contrast, the geometric gain contributed by the collar region
$\Big(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}\Big)\setminus\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$
remains of order $O(1)$ and therefore satisfies the lower bound
(52) $\text{Rad}_{\mathrm{collar}}\;\geq\;\text{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})\Big(1+\frac{C\epsilon}{a}\Big).$
Since $a^{-3/2}\ll a^{-1}$ for $a\gg 1$, the collar contribution dominates the possible metric contraction in the interior. It follows that for sufficiently large $a$ one obtains a strict radius increase,
(53) $\text{Rad}\!\Big(J^{+}(\mathcal{M}_{1})\cap\mathcal{M}_{-a-1/a+\epsilon}\Big)\;\geq\;\text{Rad}(\mathcal{M}_{1})\,\sqrt{1+\frac{1}{10a}},$
after adjusting constants.
Invoking the Schoen–Yau barrier criterion [44], the existence of a MOTS in $J^{+}(\mathcal{M}_{1})\cap\mathcal{M}_{-a-1/a+\epsilon}$ together with the absence of trapped surfaces in $\mathcal{M}_{1}$ is reduced to verifying the sufficient condition
(54) $\frac{3\pi}{2\,\text{Rad}(\mathcal{M}_{1})\sqrt{1+\frac{1}{a}}}\;$ $<\;-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{\prime}|\,|\hat{\underline{\chi}}|^{2}(u^{\prime},\epsilon)\,du^{\prime}$
$\quad+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{\prime}|^{2}}\int_{u_{\infty}}^{u^{\prime}}|u^{\prime\prime}|^{2}\,|\hat{\underline{\chi}}|^{2}(u^{\prime\prime})\,du^{\prime\prime}\,du^{\prime}\;<\;\frac{3\pi}{2\,\text{Rad}(\mathcal{M}_{1})}.$
In particular, if the initial outgoing shear and the magnitude of the incoming expansion at past null infinity $u=u_{\infty}$ are sufficiently large, then the boundary sphere
$\partial\!\Big(J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{t=-a-1/a+\epsilon}\Big)=S_{-a-1/a,3/4}$
acquires sufficiently large generalized mean curvature $H-|\kappa|$. The Schoen–Yau criterion therefore guarantees the existence of a MOTS in the interior of $J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{t=-a-1/a+\epsilon}$.
Let us define the following entity
(55) $\mathsfbf{H}:=-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}$
which is nothing but the generalized mean curvature $(H-|\mbox{tr}_{\Sigma}K|)(-a-1/a,\epsilon)$ for $\epsilon\in[0,1]$ upto negligible $O(a^{-5/2})$ error terms. Before stating the technical version of the main theorem, let us define the locally uniform Sobolev norms.Fix a smooth cutoff $\chi\in C_{c}^{\infty}(B_{2}(0))$ with $\chi\equiv 1$ on $B_{1}(0)$, and for each $y\in\mathcal{M}^{1/a}_{int}$ define $\chi_{y}(x):=\chi(x-y)$. For an integer $s\geq 4$ set
$\|u(t)\|_{H^{s}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int}))}:=\sup_{y\in\mathcal{M}^{1/a}_{int}}\ \|\chi_{y}\,u(t)\|_{H^{s}(\mathcal{M}^{1/a}_{int})}$
(and similarly for $L^{2}_{\mathrm{ul}},\,L^{\infty}_{\mathrm{ul}}$).
Theorem 2.2 (Main Theorem: Technical Version).
Fix a smooth double null foliation of a four–dimensional Lorentzian manifold $(\mathcal{M},g)$ with optical functions $(u,\underline{u})$ such that the level sets $H_{u}$ and $H_{\underline{u}}$ are outgoing and incoming null hypersurfaces, respectively, and
$S_{u,\underline{u}}:=H_{u}\cap H_{\underline{u}}$
denotes the corresponding two–sphere sections equipped with the induced metric $\gamma$. Let $\nabla,$ $\nabla_{3}$, and $\nabla_{4}$ denote the angular and null derivatives associated with the null frame
$\{e_{1},e_{2},e_{3},e_{4}\}.$
(a) Semi–Global Existence of a Characteristic initial value problem: For every sufficiently large fixed constant $\mathcal{I}>0$, there exists $a_{0}=a_{0}(\mathcal{I})\gg 1$ such that the following holds. Let $a>a_{0}$ and let $N\in\mathbb{N}$ be sufficiently large. Suppose the initial data set is prescribed along the characteristic hypersurfaces $\underline{H}_{0}$, where $u_{\infty}<0$ is large in magnitude with
[…]
for sufficiently large $\mathcal{I}^{0}$ and moreover dispersive data is prescribed on the initial null hypersurface $H_{u_{\infty}}$ compatible with the following smooth data
[…]
$\mbox{tr}\chi-\frac{2}{|u|_{\infty}}=O(a^{-\frac{1}{2}}|u_{\infty}|^{-1}),~\hat{\chi}=O(a^{-\frac{1}{2}}|u_{\infty}|^{-1}),~\mbox{tr}\underline{\chi}+\frac{2}{|u_{\infty}|}=O(|u_{\infty}|^{-2}),~\underline{\hat{\chi}}=O(a^{\frac{1}{2}}|u_{\infty}|^{-2}),$
$\omega=O(a^{\frac{1}{2}}|u_{\infty}|^{-2}),~\underline{\omega}=O(a^{\frac{1}{2}}|u_{\infty}|^{-3}),~(\eta,\underline{\eta})=O(a^{\frac{1}{2}}|u_{\infty}|^{-2})$
$\alpha=O(a^{-\frac{1}{2}}|u_{\infty}|^{-1}),~\beta=O(|u_{\infty}|^{-2}),~\underline{\beta}=O(a|u_{\infty}|^{-4}),~(\rho,\sigma)=O(a^{\frac{1}{2}}|u_{\infty}|^{-3}),~\underline{\alpha}=O(a^{\frac{3}{2}}|u_{\infty}|^{-5}),$
where $\gamma_{0}$ is the standard unit round metric in a topological 22-sphere. Then there exists a unique smooth solution $g$ of the vacuum Einstein equations in the region
$D_{a,1}:=\big\{(u,\underline{u})\mid u_{\infty}\leq u\leq-a,\;\;0\leq\underline{u}\leq 1\big\},$
satisfying the constraint and transport equations, with all geometric quantities $\psi\in\{\mbox{tr}\chi,\hat{\chi},\underline{\hat{\chi}},\mbox{tr}\underline{\chi},\eta,\underline{\eta},\alpha,\underline{\alpha},\underline{\beta},\rho,\sigma\}$ remaining uniformly controlled by constants depending only on the initial data.
(b) Local Cauchy Evolution: Fix integers s,Ns,~N large enough and let $a\gg 1$. Let $\mathcal{M}_{-a}$ be a smooth spacelike Cauchy hypersurface decomposed as
$\mathcal{M}_{-a}=\mathcal{M}_{1}\;\cup\;\mathcal{M}_{2}\cup\mathcal{M}_{ext},\qquad\partial\mathcal{M}_{1}=S_{-a,0},$
where $\mathcal{M}_{2}\subset D_{a,1}$ is generated by a canonical characteristic development from past null infinity and is foliated by a double–null optical function pair $(u,\underline{u})$ with $t=u+\underline{u}=-a$.
Let $(g,k)$ be smooth vacuum Cauchy data on a slightly bigger domain $\mathcal{M}^{1/a}_{int}:=\mathcal{M}_{1}\cup(\mathcal{M}_{2}\setminus(D^{{}^{\prime}}\cap\mathcal{M}_{2})$ with $D^{{}^{\prime}}\subset D_{a,1}$ and $D^{{}^{\prime}}$ is bounded by $u=-a-\frac{1}{a},\underline{u}=0,\underline{u}=1,$ and $u=u_{\infty}$, and assume that on $\widetilde{\mathcal{M}}_{1}\subset\mathcal{M}^{1/a}_{\mathrm{int}}$ the initial data satisfy the smallness bounds
(56) $\|\partial^{k+1}g(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-1},~\|\partial^{k}k(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-3/2},$
[…]
(57) $\|\partial^{k+1}g(-a)\|_{H^{s-1}_{ul}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-1},~\|\partial^{k}k(-a)\|_{H^{s-1}_{ul}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-3/2},~k\geq 0,$
where on the collar $\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1}$ of thickness $O(\widetilde{\epsilon}a),~\widetilde{\epsilon}\in[\frac{1}{1000},\frac{1}{100}]$ the estimate reads
(58) $\|\partial^{k}k\|_{L^{\infty}(\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-\frac{3}{2}-k},$
[…]
(60) $\|\partial^{k}k\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-3/2},~k\geq 0,$
and $(k,\partial g)$ smoothly interpolates between the data on $\widetilde{\mathcal{M}}_{1}$ and the data on $\Sigma_{int}=\mathcal{M}_{-1}\cap D_{a,1}$ in spacetime harmonic coordinates. Assume moreover that the harmonic gauge constraints and the Einstein constraint equations hold initially.
Assume that $(g,k)$ is smoothly matched across $S_{-a,0}$ with the Cauchy data induced on $\mathcal{M}_{2}$ by the characteristic development $D_{a,1}$, in the sense that:
(1) the induced metric and second fundamental form on $S_{-a,0}$ agree when computed from $(\mathcal{M}_{1},g)$ and from the characteristic spacetime $(\mathcal{M}_{\mathrm{ext}}\cap D_{a,1},g_{\mathrm{char}})$,
(2) all tangential derivatives up to order $N$ coincide on $S_{-a,0}$.
[…]
$m_{\mathrm{ADM}}\sim a^{\frac{1}{2}},$
and angular momentum $J=O(a)$, preserving the vacuum constraints and the above Sobolev bounds. Let
[…]
$\varepsilon=\varepsilon\!\left(N,\,\mathrm{Rad}(\mathcal{M}_{1}),\,\|\tilde{g}\|_{H^{N+1}_{ul}},\,\|\tilde{k}\|_{H^{N}_{ul}}\right)>0$
such that the vacuum Einstein equations admit a unique local Cauchy development
$(\mathcal{M}\times[-a,-a+\varepsilon],\mathsfbf g)$
[…]
  1. (1)
    $\mathsfbf g\in C^{0}_{t}H^{N+1}_{x}\cap C^{1}_{t}H^{N}_{x}$;
  2. (2)
    the harmonic gauge constraints propagate on $a,-a+\varepsilon]$;
  3. (3)
    the solution coincides with the characteristic development $D_{a,1}$ in the overlap region.
Moreover, for all sufficiently large $a$, the lifespan $\varepsilon$ can be chosen uniformly, and in particular one may take
$\varepsilon=\frac{3}{4}.$
(c) Dynamical Formation of a MOTS: Suppose, in addition to the assumption of part $(a)$ and $(b)$, the initial Riemannian metric $g$ on $\mathcal{M}_{1}$ has $H$-radius $\text{Rad}(\mathcal{M}_{1})\gg 1$ and set $\text{Rad}(\mathcal{M}_{1})=\frac{3\pi}{4}(a-1)+O(a^{-1}),~a\gg 1$ and the initial data on $\underline{H}_{0}$ also verify the following isotropic condition
(61) $\frac{17}{9a}<\int_{u_{\infty}}^{-a}|u^{\prime}||\underline{\hat{\chi}}|^{2}(u^{\prime},\epsilon)\,du^{\prime}+\int_{u_{\infty}}^{-a}\frac{1}{|u^{\prime}|^{2}}\int_{u_{\infty}}^{u^{\prime}}|u^{\prime\prime}|^{2}|\underline{\hat{\chi}}|^{2}(u^{\prime\prime},\epsilon)\,du^{\prime\prime}\,du^{\prime}<\frac{19}{9a},$
then the Cauchy slice $\mathcal{M}_{-a}$ will not contain a MOTS and the interior Cauchy slice $J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{t=-a-1/a+\epsilon}\subset\mathcal{M}_{t=-a-1/a+\epsilon}$ at canonical double null time $t=-a-1/a+\epsilon,~\epsilon=\frac{3}{4}$ will contain a MOTS in its interior.
Remark 3.
Notice that given $\underline{\hat{\chi}}$ on $\underline{H}_{0}$, its geometry is determined once boundary condition in $S_{u_{\infty},0}$ is given. More precisely along with the conjugate shear, (a) the conformal class/metric on the corner sphere $S_{u_{\infty},0}$ and (b)a gauge normalization, for example: $\Omega$ fixed on $S_{u_{\infty},0}$, together with a choice of the shift vector $b$ along the initial null hypersurface $\underline{H}_{0}$. In addition, the decay rate for the Weyl curvature components are fixed by the peeling property [26], while the connection coefficients are less fundamental and dependent on the choice of frame.
Remark 4.
The scaling $a^{-3/2}$ is critical for compatibility with the Bel–Robinson energy hierarchy on time slabs of size $O(1)$.
The uniformity of $\varepsilon$ relies on the effective point-wise smallness of the initial k,|N1|k,|N-1|, and $\nabla N$; largeness is absorbed in the largeness of the Hawking mass of the spheres foliating the characteristic development.
Remark 5.
Notice that the condition on the strict upper bound on the Schoen-Yau radius of the interior $\mathcal{M}_{1}$ is very flexible in terms of the choice of the corresponding metrics in light of constructions of [44]. On the other hand, $a^{\frac{1}{2}}$ is roughly related to the ADM parameters of the initial slice $\mathcal{M}_{-a}$. This is precisely the idea of a large mass contained in an isotropically (in terms of the $H$-radius or Schoen-Yau radius) large domain. In particular, the interior domain $\mathcal{M}_{1}$ needs to be uniformly thick or isotropically large measured in terms of $H$-radius.
Remark 6.
Notice the vital importance of the interior estimate $|\partial g|=O(a^{-1})$ (in harmonic coordinates) in the interior $\widetilde{\mathcal{M}}_{1}$. This implies through fundamental theorem of calculus $|g-\delta|=O(1)$ in the interior which leads to the length estimate $|d_{g}(l)-d_{\delta}(l)|=O(a)$ on $\widetilde{\mathcal{M}}_{1}$ of a curve $l$ of $\delta$-length $O(a)$. This is necessary for $H$-radius to be $\text{Rad}(\mathcal{M}_{1})=\frac{3\pi}{4}(a-1)+O(a^{-1})$ in the interior-in particular the largeness of the interior $H$-radius originates form the largeness of the weighted entity $|a\partial g|$. Notice that if instead one had $|\partial g|=O(a^{-1-\delta})$ in the interior, then similarly one would have obtained $|g-\delta|=O(a^{-\epsilon})$ and $|d_{g}(l)-d_{\delta}(l)|=O(a^{1-\epsilon})$ which would be insufficient to satisfy the condition $\text{Rad}(\mathcal{M}_{1})=\frac{3\pi}{4}(a-1)+O(a^{-1})$ due to $a^{-\epsilon}$ loss.
2.5. Comparison with Previous Studies and Novelty
Following the singularity theorem of Penrose [34], the classical mechanism guaranteeing the presence of a black hole region requires the prior existence of a closed trapped surface. For a considerable period, the only available route to such a surface was to impose its existence directly at the level of the initial data set. This requirement is itself highly nontrivial, since the trapped surface condition is a nonlinear, fully geometric constraint involving both the intrinsic and extrinsic geometry of the initial slice.
A decisive advance at the level of initial data was achieved by Schoen–Yau [43], who established the first general existence theorem for marginally outer trapped surfaces (MOTS) under suitable geometric and energy conditions. Subsequently, Yau [44] discovered a substantially stronger principle: the existence of a MOTS can be forced purely by boundary geometry. More precisely, certain quantitative boundary convexity conditions imply the existence of a MOTS in the interior, independent of any positivity assumption on the matter density (and in particular allowing negative energy densities). This boundary–driven mechanism revealed a fundamentally new geometric effect and indicated that trapped surface formation is not exclusively tied to bulk matter concentration. This is the primary motivating point of this current study (See [7, 8] for studies related to dynamic and trapping horizon, later works on Jang’s equation and its relationship with MOTS).
These results naturally lead to a deeper dynamical question: are MOTSs genuinely evolutionary objects, in the sense that they arise from the Einstein evolution of regular initial data containing no trapped or marginally trapped surfaces? From both mathematical and physical perspectives, this issue is central. The geometric definition of a black hole region in general relativity derives its significance from its predictive and observational content; it must correspond to objects that can form dynamically from physically admissible configurations. In this sense, dynamical trapped surface formation constitutes a stringent consistency requirement linking the analytic theory of the Einstein equations with the physical interpretation of black holes. This is particularly relevant for astrophysical scenarios such as supermassive black holes, whose formation is not adequately modeled by short–pulse collapse mechanisms and therefore demands a large–scale, genuinely dynamical geometric theory of trapped surface formation.
The first results along this direction were obtained by Christodoulou for the Einstein equations coupled to a massless scalar field in spherical symmetry. Through a series of works [18], [16], [17], and [20], Christodoulou managed to not only prove trapped surface formation, but to understand the picture of gravitational collapse in its entirety for the given model and under the given symmetry. The breakthrough in the absence of symmetry came in [19] by the same author. In this work, Christodoulou introduced a hierarchy of small and large components in the initial data which (almost) persists under the evolution of the Einstein equations. He termed his method the short pulse method. After Christodoulou, another breakthrough work by Klainerman-Rodnianski [25] reduces the size of Christodoulou’s work from about 600 to approximately 120 pages by using a slightly different hierarchy. Moreover, it reduces the number of derivatives of curvature required to prove semi-global existence from two to one through refined trace estimates. A substantial extension of the result of Christodoulou (which required a uniform condition along all null geodesic generators instead) was executed by Klainerman-Luk-Rodnianski [24], proving a fully genuine anisotropic criterion for the formation of trapped surfaces in vacuum. More precisely, they provide local conditions on null data, concentrated in a neighborhood of a short null geodesic segment (possibly flat in all other directions) whose future development contains a trapped surface.
A few years later, An [3] introduces the signature for decay rates $s_{2}$ on his way to proving an extension of [25] from a finite region to a region close to past null infinity-this method proved to be very efficient in handling large data problems in general. In 2014, An and Luk [6] proved the first scale-critical trapped surface formation criterion for the vacuum equations in the absence of symmetry. While Christodoulou’s data in [19] were large in $\dot{H}^{1}(\mathbb{R}^{3})$, An and Luk give data which only have to be large in $\dot{H}^{\frac{3}{2}}(\mathbb{R}^{3})$, which is a scale-critical norm for the initial data. Taking advantage of the scale criticality in [6], An [2] constructs initial data that give rise not merely to trapped surfaces, but an apparent horizon, a smooth 3-dimensional hypersurface consisting of marginally outer trapped surfaces. In 2019, An [1] produces a 55-page proof of trapped surface formation for the vacuum equations, making use of the signature for decay rates and obtaining an existence result from a region close to past null infinity. In [4], An and Athanasiou extended [1] to the case of the Einstein-Maxwell system. Several other studies exist in the context of Einstein-Yang-Mills [12], Einstein-Vlasov system [23, 9, 10, 13], Einstein-Scalar field system, and Einstein-Spinor field system. Recently, [15], in a fundamental study, provided a short proof of the formation of a trapped surface in geodesic foliation.
The principal analytical difficulty of the present work lies in the prescription of the initial data, whether prescribed on a characteristic or a Cauchy hypersurface. One of the main novelties of this study and the difference with the previous studies is the delicate nature (cancellation) of the estimates associated with the transport equation. Arbitrary initial data, however, fail to exhibit such cancellation structures, and in fact, by the celebrated small-data result of Christodoulou–Klainerman [19], one already knows that for sufficiently small perturbations of Minkowski data (in global sense), no MOTS or trapped surface can form.
In the large-data regime, the situation is profoundly different: the system is fully nonlinear, no global small parameter exists, and long-time uniform control cannot, in general, be expected in the hyperbolic setting. As a result, one faces the possibility of geometric pathologies, including singularity formation and breakdown of the foliation. A central challenge, therefore, consists in constructing a hierarchy of large but controlled initial data $(g,k)$ that is compatible with a semi-global existence theorem for the Einstein vacuum equations, and whose evolution can be followed up to the onset of an MOTS.
From a structural viewpoint, the formation of a trapped surface is governed by monotonicity property of the Raychaudhuri equation for the null expansion $\mbox{tr}\chi$ along the outgoing null direction:
[…]
In the previous approach (cf. [18, 25, 1, 3]), one attempts to produce a negative expansion $\mbox{tr}\chi<0$ within a finite affine parameter time along the null generators. This mechanism is driven by the largeness of the incoming gravitational shear (radiation) $|\widehat{\chi}|^{2}$, which acts as a source in (62).
By contrast, the present work adopts a complementary perspective. Rather than directly forcing $\mbox{tr}\chi$ to become negative, we study the evolution of the generalized mean curvature quantity
(63) $c:=H-|\kappa|,$
as introduced in Section 2.1 and considered by Schoen-Yau [43] in the context of matter sourced gravity and by [44] in a more general framework that included the case of negative energy density (and pure vacuum in particular). Our objective is to obtain a strictly positive value of $c$ on an appropriate boundary hypersurface of an isotropically large (appropriately defined) domain (i.e., the $H$-radius or the Schoen-Yau radius is large), thus triggering the dynamical emergence of an MOTS from a previously horizon-free configuration. In the double-null gauge, one has the decomposition
(64) $H-|\kappa|=\tfrac{1}{2}\big(\mbox{tr}\chi-\mbox{tr}\underline{\chi}\big)-\tfrac{1}{2}\big|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\big|.$
The guiding principle is therefore to drive the difference $\mbox{tr}\chi-\mbox{tr}\underline{\chi}$ to become large and positive while simultaneously suppressing the absolute term $|\mbox{tr}\chi+\mbox{tr}\underline{\chi}|$ along the incoming null direction. This is essential not to simply control the generalized mean curvature $c=H-|\kappa|$, but also to control the radius of the domain under consideration (and gluing) as seen in the section 2.4 of introduction.
The construction is implemented by prescribing the incoming shear $\underline{\hat{\chi}}$ on the initial incoming null hypersurface $\underline{H}_{0}$ together with a normalized dispersive profile for the null expansions $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ on the distant outgoing hypersurface $H_{u_{\infty}}$. The data are arranged so as to satisfy a precise scale–invariant hierarchy which departs substantially from the Minkowskian regime while retaining dispersive decay in $|u|$. The parameter $a\gg 1$ measures the amplitude of the deviation from flat data, whereas inverse powers of $|u|$ encode null dispersion along the foliation.
More precisely, the Ricci coefficients are assumed to obey the asymptotic hierarchy
$\bigg|\mbox{tr}\chi-\frac{2}{|u|}\bigg|=O\!\left(a^{-\frac{1}{2}}|u|^{-1}\right),\qquad|\hat{\chi}|=O\!\left(a^{-\frac{1}{2}}|u|^{-1}\right),\qquad\bigg|\mbox{tr}\underline{\chi}+\frac{2}{|u|}\bigg|=O\!\left(|u|^{-2}\right),\qquad|\underline{\hat{\chi}}|=O\!\left(a^{\frac{1}{2}}|u|^{-2}\right),$
$|\omega|=O\!\left(a^{\frac{1}{2}}|u|^{-2}\right),\qquad|\underline{\omega}|=O\!\left(a^{\frac{1}{2}}|u|^{-3}\right),\qquad|\eta|+|\underline{\eta}|=O\!\left(a^{\frac{1}{2}}|u|^{-2}\right).$
This scaling is consistent with the null structure equations and is chosen so that the dominant large component is the incoming shear, while all remaining Ricci coefficients remain perturbative relative to their Minkowskian leading orders.
A principal analytic difficulty is that the null transport equations for the Ricci coefficients contain Weyl curvature components as source terms. These curvature components must satisfy decay estimates in $|u|$ compatible with the peeling behavior (cf. [26]) and, simultaneously, an amplitude hierarchy compatible with the above Ricci coefficient scaling. In particular, the curvature components are arranged to satisfy
$|\alpha|=O\!\left(a^{-\frac{1}{2}}|u|^{-1}\right),|\beta|=O\!\left(|u|^{-2}\right),|\underline{\beta}|=O\!\left(a|u|^{-4}\right),|\rho|+|\sigma|=O\!\left(a^{\frac{1}{2}}|u|^{-3}\right),|\underline{\alpha}|=O\!\left(a^{\frac{3}{2}}|u|^{-5}\right).$
The hierarchy is closed in the sense that, under the null structure and Bianchi equations, these weights are stable under propagation in the semi–global region of existence up to controlled losses. Ensuring this compatibility and propagation of scale is a central structural requirement in the argument.
Remark 7.
Note how the largeness of $a$ manifests. In simpler geometries such as an asymptotically Schwarchild case, $\rho=O(m|u|^{-3})$ and so $a^{\frac{1}{2}}$ in our context roughly behaves like mass-in fact the sphere $S_{-a,0}$ has Hawking mass $\approx a^{1/2}$. Thus, for $a\gg 1$, the data manifestly falls under the category of the large data or in the moduli space, $a^{\frac{1}{2}}$ essentially measures a notion of ‘distance’ from the Minkowski space. Contrast this scaling with the previous works [5, 4, 12], where $a$ appears instead of $a^{\frac{1}{2}}$ and the scaling of the radiation fields are different.
It is instructive to compare the incoming gravitational energy in the short–pulse framework of Christodoulou and its later refinements (see [19, 5, 4]) with the present construction. In the double–null formalism, the incoming gravitational energy flux through an outgoing null hypersurface $H_{u_{\infty}}$ over the slab $\underline{u}\in[0,1]$ is measured by
$E_{\infty}:=\int_{0}^{1}\!\int_{S_{u_{\infty},\underline{u}}}|\hat{\chi}|^{2}\,d\mu_{\gamma}\,d\underline{u},\qquad d\mu_{\gamma}=\sqrt{\det(\gamma_{AB})}\,.$
In the short–pulse regime of [5, 4], one prescribes
$|\hat{\chi}(u_{\infty},\underline{u})|\sim a^{1/2}|u_{\infty}|^{-1},\qquad\gamma(u_{\infty},\underline{u})\sim|u_{\infty}|^{2}\gamma_{0},$
so that $|S_{u_{\infty},\underline{u}}|\sim|u_{\infty}|^{2}$ and therefore
$E_{\infty}^{\mathrm{short\text{-}pulse}}\;\approx\;\int_{0}^{1}\!\int_{S_{u_{\infty},\underline{u}}}a\,|u_{\infty}|^{-2}\,d\mu_{\gamma}\;\sim\;a.$
Thus the incoming gravitational energy from past null infinity is large.
In contrast, in the present hierarchy one imposes the dispersive scaling
$|\hat{\chi}(u_{\infty},\underline{u})|\sim a^{-1/2}|u_{\infty}|^{-1},$
with the same area scale for $S_{u_{\infty},\underline{u}}$. Consequently,
$E_{\infty}^{\mathrm{present}}\;\approx\;\int_{0}^{1}\!\int_{S_{u_{\infty},\underline{u}}}a^{-1}|u_{\infty}|^{-2}\,d\mu_{\gamma}\;\sim\;a^{-1},$
which is small. The present construction therefore operates in a regime of weak incoming radiation from null infinity, in sharp contrast with the short–pulse mechanism.
We next examine the Hawking mass of the boundary sphere $S_{-a,0}\subset\underline{H}_{0}$. Recall that
$m_{H}(S)=\frac{r}{2}\left(1+\frac{1}{16\pi}\int_{S}\mbox{tr}\chi\,\mbox{tr}\underline{\chi}\,d\mu_{\gamma}\right),\qquad r=\Big(\frac{|S|}{4\pi}\Big)^{1/2}.$
On $S_{-a,0}$ we have $r\sim a$ and, by the Gauss equation,
$K=-\rho-\frac{1}{4}\mbox{tr}\chi\,\mbox{tr}\underline{\chi}+\frac{1}{2}\hat{\chi}\cdot\underline{\hat{\chi}}$.
Using $K\sim a^{-2}$ and $|\hat{\chi}\cdot\underline{\hat{\chi}}|\ll a^{-2}$ under the assumed hierarchy, one obtains
$\int_{S_{-a,0}}\mbox{tr}\chi\,\mbox{tr}\underline{\chi}\,d\mu_{\gamma}=-4\int_{S_{-a,0}}\rho\,d\mu_{\gamma}+O(1).$
With $|\rho|\sim a^{1/2}|u|^{-3}\sim a^{-5/2}$ at u=au=-a and $|S_{-a,0}|\sim a^{2}$, it follows that
$\int_{S_{-a,0}}\rho\,d\mu_{\gamma}\sim a^{-1/2},\qquad m_{H}(S_{-a,0})\sim a^{1/2}\gg 1.$
Thus the Hawking mass of the initial spheres foliating $\underline{H}_{0}$ is large despite the small incoming radiation from $H_{u_{\infty}}$.
This should be contrasted with the characteristic setup in Christodoulou’s short–pulse framework [19], where the incoming hypersurface $\underline{H}_{0}$ is taken to be exactly Minkowskian and all large effects arise from concentrated incoming radiation. The mechanism realized here is different: it is not a collapse driven by a high–energy short pulse, but a boundary–driven large–scale configuration in which substantial mass is distributed over an isotropically large domain while outgoing radiation remains mild.
This viewpoint is consistent with the boundary–driven geometric mechanisms established by Schoen and Yau [38] in the Riemannian setting, where quantitative lower bounds on boundary mean curvature impose rigid constraints on the admissible interior geometry and topology. In that context, sufficiently strong boundary convexity forces global interior consequences independent of any concentration of bulk energy. An analogous phenomenon was later identified by Yau in the Lorentzian setting, where appropriate boundary geometric conditions guarantee the existence of marginally outer trapped surfaces at the level of initial data, without requiring positive matter density assumptions.
The present analysis gives a genuinely dynamical realization of this boundary–effect principle within the Einstein vacuum equations. The mechanism developed here does not rely on short–pulse–type concentration of incoming gravitational radiation. Instead, the decisive inputs are the persistence of strong boundary mean curvature along the outgoing null hypersurface and the presence of large total mass distributed over an isotropically large spatial domain. Under these conditions, the null evolution amplifies the boundary geometry in a controlled manner and leads to the formation of a marginally outer trapped surface in the interior of the spacetime development.
In particular, this establishes that black hole formation is not restricted to collapse scenarios driven by highly concentrated energy flux. It also occurs in a complementary large–scale regime characterized by dispersed mass and mild radiation, thereby identifying a distinct and robust geometric pathway to trapped surface formation.
A key structural feature in our argument is the isotropic largeness of the strong–field region, quantified through a lower bound for the Schoen–Yau (HH–) radius. What is essential is not largeness of total mass alone, but largeness of the domain on which curvature and mean–convexity are simultaneously controlled in an essentially direction–independent manner.
This distinction is clarified by the localized gluing construction of Carlotto–Schoen [14]. Their theorem produces asymptotically flat, scalar–flat (time–symmetric) vacuum initial data sets with arbitrarily large ADM mass whose geometry agrees with a large–mass Schwarzschild end inside a prescribed cone, while remaining exactly Euclidean outside a slightly larger cone. In particular, the gravitational field can be made strong yet highly anisotropic, being confined to a narrow angular sector and completely shielded elsewhere.
Such examples show that large ADM mass by itself does not enforce any uniform, isotropic geometric control on large coordinate balls or quasi-round domains. In particular, one cannot deduce from mass alone the presence of a large mean–convex barrier region or a domain with large Schoen–Yau radius to which boundary–driven minimal or trapped surface arguments apply. The obstruction is geometric: curvature concentration that is strongly directional can be separated from large portions of the manifold by exactly flat regions.
This behavior stands in sharp contrast with the regime considered here. Our hypotheses impose quantitative mean–curvature and radius control on an isotropically large domain, ensuring that the dominant curvature and energy flux are not confined to a thin sector but are distributed over a region with uniform geometric thickness. It is precisely this combination — large scale together with strong positive boundary generalized mean curvature-that allows one to convert curvature concentration into a marginally outer trapped surface through evolution. The result, therefore, isolates a mechanism for MOTS formation that depends on global geometric size and boundary convexity, rather than on total mass alone or quasi-local energy concentration.
Figure 3. An equivalent alternative configuration yielding MOTS in an evolutionary manner. One may technically not need the characteristic evolution that is performed in this work. Recall that the data on middle part $\mathcal{M}_{2}$ is induced by the characteristic development. However, since we already know the type of data induced on $\mathcal{M}_{2}$, one can in principle start with this data on $\mathcal{M}_{2}$, the interior data on $\mathcal{M}_{1}$ and Kerr exterior on the outside $\mathcal{M}_{3}$ and initiate the time evolution. By uniqueness of the solution of the Cauchy problem of the Einstein’s equations one would have MOTS in the interior of isotropically large domain $\mathcal{M}^{{}^{\prime}}_{1}$.
2.6. Acknowledgements
This work is supported by the Beijing Institute of Mathematical Sciences and Applications of the Yau Mathematical Science Center at Tsinghua University and Beijing Municipal Govt.
3. Part 1: Solution of the Characteristic Part
This is an important aspect of this study. But the study of characteristic initial value problems and proving a semi-global development with appropriate smooth initial data is well-established. We will follow the signature of decay rate technology developed by An [5] that is designed to handle large data regime in a systematic way.
3.1. Construction of the double null gauge
Denote by $L\mathcal{M}$ the frame bundle of $\mathcal{M}$. We construct a double null gauge, meaning a smooth section of this bundle such that, through it, each point $p\in\mathcal{M}$ maps to a renormalized frame $(e_{1},e_{2},e_{3},e_{4})\in L\mathcal{M}$ with $g(e_{3},e_{4})=-2$, $g(e_{A},e_{B})=\delta_{AB}$ and $g(e_{3},e_{A})=g(e_{4},e_{A})=0$.
We begin with two null hypersurfaces $H_{u_{\infty}},\underline{H}_{0}$ and their intersection $S_{u_{\infty},0}$, a topological 2-sphere. For any point $q$ on this 22-sphere, the tangent space $T_{q}S_{u_{\infty},u}$ is 2-dimensional and admits a 22-dimensional orthogonal complement $T_{q}^{\text{Per}p}{S_{u_{\infty},0}}$, on which we can find two future-directed null vectors $L^{\prime}_{q}$ and $\underline{L}^{\prime}_{q}$, normalized so that
$g(L^{\prime}_{q},\underline{L}^{\prime}_{q})=-2.$
The pair $\begin{Bmatrix}L^{\prime}_{q},\underline{L}^{\prime}_{q}\end{Bmatrix}$ is uniquely determined up to a scaling factor $s>0$
$\begin{Bmatrix}L^{\prime}_{q},\underline{L}^{\prime}_{q}\end{Bmatrix}\mapsto\begin{Bmatrix}sL^{\prime}_{q},s^{-1}\underline{L}^{\prime}_{q}\end{Bmatrix}.$
Starting from $q$ and initially tangent to $L^{\prime}_{q}$, a unique geodesic is sent out. Call this geodesic, $l_{q}$. We extend the vectorfield $L^{\prime}$ along $l_{q}$ by parallel transport: $D_{L^{\prime}}L^{\prime}=0$. It then follows by simple calculation that $l_{q}$ is null, so that $g(L^{\prime},L^{\prime})=0$ along $l_{q}$. Gathering the $\begin{Bmatrix}l_{q}\end{Bmatrix}$ together we get a null hypersurface $H_{u_{\infty}}$. The null hypersurface $\underline{H}_{0}$ is obtained similarly. Note that, by construction, given a point $p$ on $H_{u_{\infty}}$ or $\underline{H}_{0}$, in the corresponding tangent spaces, there is a preferred null vector $L^{\prime}_{p}$ or $\underline{L}^{\prime}_{p}$.
We next choose a lapse function $\Omega$, which we define to be equal to 1 on $S_{u_{\infty},0}$ and then extend as a continuous function along both initial null hypersurfaces [^5] Indeed, there is a gauge freedom in choosing $\Omega$ on the initial hypersurfaces.. Define the vector fields
$L:=\Omega^{2}L^{\prime}\hskip 5.69054pt\text{along}\hskip 5.69054ptH_{u_{\infty}}\hskip 5.69054pt\text{and}\hskip 5.69054pt\underline{L}:=\Omega^{2}\underline{L}^{\prime}\hskip 5.69054pt\text{along}\hskip 5.69054pt\underline{H}_{0}.$
We use these vector fields to define two functions
$\underline{u}\hskip 5.69054pt\text{on}\hskip 5.69054ptH_{u_{\infty}}\hskip 5.69054pt\text{satisfying}\hskip 5.69054ptL\underline{u}=1\hskip 5.69054pt\text{on}\hskip 5.69054ptH_{u_{\infty}}\hskip 5.69054pt\text{and}\hskip 5.69054pt\underline{u}=0\hskip 5.69054pt\text{on}\hskip 5.69054ptS_{u_{\infty},0},$
$u\hskip 5.69054pt\text{on}\hskip 5.69054pt\underline{H}_{0}\hskip 5.69054pt\text{satisfying}\hskip 5.69054pt\underline{L}u=1\hskip 5.69054pt\text{on}\hskip 5.69054pt\underline{H}_{0}\hskip 5.69054pt\text{and}\hskip 5.69054ptu=0\hskip 5.69054pt\text{on}\hskip 5.69054ptS_{u_{\infty},0}.$
We now use these (so-called optical) functions to proceed further with the construction. Let $S_{u_{\infty},\underline{u}^{\prime}}$ be the embedded 22-surface on $H_{u_{\infty}}$ on which $\underline{u}=\underline{u}^{\prime}$ and define $S_{u,0}$ similarly. At each point $p\in S_{u_{\infty},\underline{u}^{\prime}}$ we have constructed a preferred null vector $L^{\prime}_{p}$. It follows that we can uniquely determine an incoming $g$-null vector $\underline{L}^{\prime}_{p}$ satisfying $g(L^{\prime}_{p},\underline{L}^{\prime}_{p})=-2$. Let $\underline{l}_{p}$ be the unique geodesic emanating from $p$ with tangent vector $\underline{L}^{\prime}_{p}$. We extend the definition of $\underline{L}^{\prime}$ along $\underline{l}_{q}$ by parallel transport, so that $D_{\underline{L}^{\prime}}\underline{L}^{\prime}=0$. Gathering all the $\begin{Bmatrix}\underline{l}_{p}\end{Bmatrix}$ on $S_{u_{\infty},\underline{u}^{\prime}}$, we thus obtain the null hypersurface $\underline{H}_{\underline{u}^{\prime}}$. We obtain the null hypersurface $H_{u^{\prime}}$ in an analogous way and define $S_{u,\underline{u}}:=H_{u}\cap\underline{H}_{\underline{u}}$. Having constructed the vector fields $L^{\prime}$ and $\underline{L}^{\prime}$ in all of the spacetime region, we extend the definition of the lapse function $\Omega$ by requiring, at each point $p\in S_{u,\underline{u}}$ that
$g(L^{\prime}_{p},\underline{L}^{\prime}_{p})=-2{\left.\kern-1.2pt\Omega^{-2}\right|_{p}}.$
The incoming null hypersurfaces $\begin{Bmatrix}\underline{H}_{\underline{u}}\end{Bmatrix}_{0\leq\underline{u}\leq 1}$ and outgoing null hypersurfaces $\begin{Bmatrix}H_{u}\end{Bmatrix}_{-a\leq u\leq u_{\infty}}$ along with their pairwise intersections $S_{u,\underline{u}}$ together define a double null foliation on the spacetime. On a given $S_{u,\underline{u}}$, we have $g(\Omega L^{\prime},\Omega\underline{L}^{\prime})=-2$ and hence the vectors
$e_{3}:=\Omega\underline{L}^{\prime},e_{4}:=\Omega L^{\prime}$
define a normalized null pair at each point on the sphere. We make the gauge choice $\Omega\equiv 1$ along both initial hypersurfaces.
3.2. Choice of coordinates and expression of the metric
To define angular coordinates on each $S_{u,\underline{u}}$ in a smooth way, we begin by defining angular coordinates on $S_{u_{\infty},0}$. Since this is a standard 2-sphere in Minkowki space, we can use the stereographic projection coordinates $(\theta^{1},\hskip 1.42262pt\theta^{2})$ on $S_{u_{\infty},0}$. We first extend this coordinate to the whole of $\underline{H}_{0}$ by insisting that $\not{\mathcal{L}}_{\underline{L}}\theta^{A}=0$ on $\underline{H}_{0}$ for $A=1,\hskip 1.42262pt2$ and then to the whole spacetime by insisting that, for all $u$, $\not{\mathcal{L}}_{L}\theta^{A}=0$, where $L$ initially starts normal to some $S_{u,0}$. As such we have established a coordinate system $(u,\underline{u},\theta^{1},\theta^{2})$ in a neighbourhood of the initial sphere. In these coordinates, the vectors $e_{3},\hskip 1.42262pte_{4}$ become
$e_{3}=\Omega^{-1}\left(\frac{\partial}{\partial u}+b^{A}\hskip 1.42262pt\frac{\partial}{\partial\theta^{A}}\right),\hskip 5.69054pte_{4}=\Omega^{-1}\frac{\partial}{\partial\underline{u}}$
and the metric now takes the following form:
[…]
The section that maps $p\in\mathcal{M}\mapsto\left({\left.\kern-1.2pt\theta^{1}\right|_{p}},{\left.\kern-1.2pt\theta^{2}\right|_{p}}{\left.\kern-1.2pte_{3}\right|_{p}},{\left.\kern-1.2pte_{4}\right|_{p}}\right)$ is the double null gauge we wanted to construct.
3.3. The vacuum Einstein equations expressed in the double null gauge
In this section we are going to express the Einstein equations in the double null gauge given above.
Before we are ready to present the equations, we introduce a few basic definitions. First of all, denote by $\nabla$ the covariant derivative operators induced by $D$ on $S_{u,\underline{u}}$. Let $\nabla_{3},\nabla_{4}$ denote the projections of the covariant derivatives $D_{3}$ and $D_{4}$ to $S_{u,\underline{u}}$. For two 11-forms $\phi_{A}^{1},\hskip 1.42262pt\phi_{A}^{2}$, we define
$(\phi_{1}\hat{\otimes}\phi_{2})_{AB}:=\phi_{A}^{1}\hskip 1.42262pt\phi_{B}^{2}+\phi_{B}^{1}\hskip 1.42262pt\phi_{A}^{2}-\gamma_{AB}\hskip 1.42262pt(\phi^{1}\cdot\phi^{2}),$
while for symmetric 22-tensors $\phi_{AB}^{1},\hskip 1.42262pt\phi_{AB}^{2}$, we define
$(\phi^{1}\wedge\phi^{2})_{AB}:=\not{\epsilon}^{AB}\hskip 1.42262pt(\gamma^{-1})^{CD}\hskip 1.42262pt\phi_{AB}^{1}\hskip 1.42262pt\phi_{CD}^{2}.$
Here $\not{\epsilon}$ is the volume form associated with the metric $\gamma$. Moreover, by $\phi^{1}\cdot\phi^{2}$ we mean an arbitrary contraction of the tensor product of $\phi^{1}$ and $\phi^{2}$ with respect to the metric $\gamma$. We also define suitable trace, divergence, and curl operators. For totally symmetric tensors $\phi$, we define these operators as follows:
$(\mbox{div }\phi)_{A_{1}\dots A_{r}}:=\nabla^{B}\phi_{BA_{1}\dots A_{r}},(\mbox{curl }\phi)_{A_{1}\dots A_{r}}:=\not{\epsilon}^{BC}\nabla_{B}\phi_{CA_{1}\dots A_{r}},(\mbox{tr}\phi)_{A_{1}\dots A_{r-1}}:=(\gamma^{-1})^{BC}\phi_{BCA_{1}\dots A_{r-1}}.$
[…]
$\prescript{*}{}{\phi}_{A}:=\gamma_{AC}\not{\epsilon}^{CB}\hskip 1.42262pt\phi_{B},$
$\prescript{*}{}{\phi}_{AB}:=\gamma_{BD}\hskip 1.42262pt\not{\epsilon}^{DC}\hskip 1.42262pt\phi_{AC}.$
Finally, on a 11-form $\phi$, the operator $\nabla\hat{\otimes}$ is defined as follows:
[…]
The vacuum Einstein equations take the following form in this double null gauge
(66) $\nabla_{4}\mbox{tr}\chi+\frac{1}{2}(\mbox{tr}\chi)^{2}=-|\hat{\chi}|^{2}_{\gamma}-2\omega\mbox{tr}\chi$
(67) $\nabla_{4}\hat{\chi}+\mbox{tr}\chi\hat{\chi}=-2\omega\hat{\chi}-\alpha$
(68) $\nabla_{3}tr\underline{\chi}+\frac{1}{2}(tr\underline{\chi})^{2}=-|\hat{\underline{\chi}}|^{2}_{\gamma}-2\underline{\omega}tr\underline{\chi}$
(69) $\nabla_{3}\hat{\underline{\chi}}+tr\underline{\chi}\hat{\underline{\chi}}=-2\underline{\omega}\hat{\underline{\chi}}-\underline{\alpha}$
(70) $\nabla_{4}\eta_{a}=-\chi\cdot(\eta-\underline{\eta})-\beta$
(71) $\nabla_{3}\underline{\eta}_{a}=-\underline{\chi}\cdot(\underline{\eta}-\eta)+\underline{\beta}$
(72) $\nabla_{4}\underline{\omega}=2\omega\underline{\omega}+\frac{3}{4}|\eta-\underline{\eta}|^{2}-\frac{1}{4}(\eta-\underline{\eta})\cdot(\eta+\underline{\eta})-\frac{1}{8}|\eta+\underline{\eta}|^{2}+\frac{1}{2}\rho$
(73) $\nabla_{3}\omega=2\omega\underline{\omega}+\frac{3}{4}|\eta-\underline{\eta}|^{2}+\frac{1}{4}(\eta-\underline{\eta})\cdot(\eta+\underline{\eta})-\frac{1}{8}|\eta+\underline{\eta}|^{2}+\frac{1}{2}\rho$
(74) $\nabla_{4}tr\underline{\chi}+\frac{1}{2}\mbox{tr}\chi tr\underline{\chi}=2\omega tr\underline{\chi}+2\text{div}\underline{\eta}+2|\underline{\eta}|^{2}_{\gamma}+2\rho-\hat{\chi}\cdot\hat{\underline{\chi}}$
(75) $\nabla_{3}\mbox{tr}\chi+\frac{1}{2}tr\underline{\chi}\mbox{tr}\chi=2\underline{\omega}\mbox{tr}\chi+2\text{div}\eta+2|\eta|^{2}+2\rho-\hat{\chi}\cdot\hat{\underline{\chi}}$
(76) $\nabla_{4}\hat{\underline{\chi}}+\frac{1}{2}\mbox{tr}\chi\hat{\underline{\chi}}=\nabla\hat{\otimes}\underline{\eta}+2\omega\hat{\underline{\chi}}-\frac{1}{2}tr\underline{\chi}\hat{\chi}+\underline{\eta}\hat{\otimes}\underline{\eta}$
(77) $\nabla_{3}\hat{\chi}+\frac{1}{2}tr\underline{\chi}\hat{\chi}=\nabla\hat{\otimes}\eta+2\underline{\omega}\hat{\chi}-\frac{1}{2}\mbox{tr}\chi\hat{\underline{\chi}}+\eta\hat{\otimes}\eta$
(78) $\text{div}\hat{\chi}=\frac{1}{2}\nabla\mbox{tr}\chi-\frac{1}{2}(\eta-\underline{\eta})\cdot(\hat{\chi}-\frac{1}{2}\mbox{tr}\chi\gamma)-\beta$
(79) $\text{div}\hat{\underline{\chi}}=\frac{1}{2}\nabla tr\underline{\chi}-\frac{1}{2}(\underline{\eta}-\eta)\cdot(\hat{\underline{\chi}}-\frac{1}{2}tr\underline{\chi}\gamma)-\underline{\beta}$
(80) $\text{curl}\eta=\hat{\underline{\chi}}\wedge\hat{\chi}+\sigma\epsilon=-\text{curl}\underline{\eta}$
(81) $K-\frac{1}{2}\hat{\chi}\cdot\hat{\underline{\chi}}+\frac{1}{4}\mbox{tr}\chi tr\underline{\chi}=-\rho.$
The Bianchi equations read in this gauge as follows
(82) $\begin{split}\nabla_{3}\alpha+\frac{1}{2}\mbox{tr}\underline{\chi}\alpha=&\nabla\hat{\otimes}\beta+4\underline{\omega}\alpha-3\left(\hat{\chi}\rho+\prescript{*}{}{\hat{\chi}}\sigma\right)+(\zeta+4\eta)\hat{\otimes}\beta\end{split}$
(83) $\nabla_{4}\beta+2\mbox{tr}\chi\beta=\text{div}\alpha-2\omega\beta+\left(\eta-2\zeta\right)\cdot\alpha,$
(84) $\nabla_{3}\beta+\mbox{tr}\underline{\chi}\beta=\nabla\rho+\prescript{*}{}{\nabla}\sigma+2\underline{\omega}\beta+2\hat{\chi}\cdot\underline{\beta}+3\left(\eta\rho+\prescript{*}{}{\eta}\sigma\right),$
(85) $\nabla_{4}\sigma+\frac{3}{2}\mbox{tr}\chi\sigma=-\text{div}\prescript{*}{}{\beta}+\frac{1}{2}\hskip 1.42262pt\underline{\hat{\chi}}\cdot\prescript{*}{}{\alpha}-(\zeta+2\underline{\eta})\cdot\prescript{*}{}{\beta},$
(86) $\nabla_{3}\sigma+\frac{3}{2}\mbox{tr}\underline{\chi}\sigma=-\text{div}\prescript{*}{}{\underline{\beta}}+\frac{1}{2}\hskip 1.42262pt\hat{\chi}\cdot\prescript{*}{}{\underline{\alpha}}-(\zeta+2\eta)\cdot\prescript{*}{}{\underline{\beta}},$
(87) $\nabla_{4}\rho+\frac{3}{2}\mbox{tr}\chi\rho=\text{\mbox{div }}\beta-\frac{1}{2}\underline{\hat{\chi}}\cdot\alpha+(\zeta+2\underline{\eta})\cdot\beta,$
(88) $\nabla_{3}\rho+\frac{3}{2}\mbox{tr}\underline{\chi}\rho=-\text{div}\underline{\beta}-\frac{1}{2}\hat{\chi}\cdot\underline{\alpha}+(\zeta-2\eta)\cdot\underline{\beta},$
(89) $\nabla_{4}\underline{\beta}+\mbox{tr}\chi\underline{\beta}=-\nabla\rho+\prescript{*}{}{\nabla}\sigma+2\omega\underline{\beta}+2\underline{\hat{\chi}}\cdot\beta-3\left(\underline{\eta}\rho-\prescript{*}{}{\underline{\eta}}\sigma\right),$
(90) $\nabla_{3}\underline{\beta}+2\mbox{tr}\underline{\chi}\underline{\beta}=-\text{div}\underline{\alpha}-2\underline{\omega}\underline{\beta}+\underline{\eta}\cdot\underline{\alpha},$
(91) $\begin{split}\nabla_{4}\underline{\alpha}+\frac{1}{2}\mbox{tr}\chi\hskip 1.42262pt\underline{\alpha}=&-\nabla\hat{\otimes}\underline{\beta}+4\omega\hskip 1.42262pt\underline{\alpha}-3\left(\underline{\hat{\chi}}\rho-\prescript{*}{}{\underline{\hat{\chi}}}\sigma\right)+\left(\zeta-4\underline{\eta}\right)\hat{\otimes}\underline{\beta}.\end{split}$
3.4. Integration
Let $U$ be a coordinate patch on a 22-sphere $S_{u,\underline{u}}$ and let $p_{U}$ be a partition of unity subordinate to $U$. For a function $\phi$, we define its integral on a 22-sphere as well as on the null hypersurfaces $H_{u}$ and $\underline{H}_{\underline{u}}$.
(92) $\int_{S_{u,\underline{u}}}\phi:=\sum_{U}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot p_{U}\cdot\sqrt{\text{det}\gamma}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2},$
(93) $\int_{H_{u}^{(0,\underline{u})}}:=\sum_{U}\int_{0}^{\underline{u}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot 2\hskip 1.42262ptp_{U}\cdot\Omega\cdot\sqrt{\text{det}\gamma}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime},$
(94) $\int_{\underline{H}_{\underline{u}}^{(u_{\infty},u)}}:=\sum_{U}\int_{u_{\infty}}^{u}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot 2\hskip 1.42262ptp_{U}\cdot\Omega\cdot\sqrt{\text{det}g}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2}\hskip 1.42262pt\text{d}u^{\prime},$
For a spacetime region $D_{u,\underline{u}}:=\begin{Bmatrix}\left(u^{\prime},\underline{u}^{\prime},\theta^{1},\theta^{2}\right)\hskip 1.42262pt\mid\hskip 1.42262ptu_{\infty}\leq u^{\prime}\leq u,0\leq\underline{u}^{\prime}\leq\underline{u}\end{Bmatrix}$, we define the spacetime integral
(95) $\int_{D_{u,\underline{u}}}\phi:=\sum_{U}\int_{u_{\infty}}^{u}\int_{0}^{\underline{u}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot p_{U}\cdot\Omega^{2}\cdot\sqrt{-\text{det}g}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\hskip 1.42262pt\text{d}u^{\prime}.$
We proceed with the definition of $L^{p}$ norms $(1\leq p<\infty)$ for an arbitrary tensorfield $\phi$:
(96) $\lVert\phi\rVert_{L^{p}(S_{u,\underline{u}})}^{p}:=\int_{S_{u,\underline{u}}}\langle\phi,\phi\rangle_{\gamma}^{\frac{p}{2}}$
(97) $\lVert\phi\rVert_{L^{p}(H_{u}^{(0,\underline{u})})}^{p}:=\int_{H_{u}^{(0,\underline{u})}}\langle\phi,\phi\rangle_{\gamma}^{\frac{p}{2}}$
(98) $\lVert\phi\rVert_{L^{p}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}^{p}:=\int_{\underline{H}_{\underline{u}}^{(u_{\infty},u)}}\langle\phi,\phi\rangle_{\gamma}^{\frac{p}{2}}.$
For the case $p=\infty$, we separately define
(99) $\lVert\phi\rVert_{L^{\infty}(S_{u,\underline{u}})}:=\underset{(\theta^{1},\theta^{2})\in S_{u,\underline{u}}}{\text{sup}}\langle\phi,\phi\rangle_{\gamma}^{\frac{1}{2}}(\theta^{1},\theta^{2}).$
3.5. Signature for decay rates and scale-invariant norms
Perhaps the most challenging aspect of trapped surface formation results, historically, has been the attempt to find initial data that are, in an appropriate sense, large (this is by necessity, as is implied by the monumental work of [21]) but also small enough to allow for an existence result of a spacetime region that gives trapped surfaces the time they would require to form. The first such initial data set, in the absence of symmetries, was given by [19]. Later contributions include [25], [6] and [2]. Moreover, one would have to construct norms that preserve, at least approximately, the hierarchy present in the initial data upon evolution of the Einstein equations. The signature for decay rates, which was first introduced in [3], is the tool we will use in the present paper to build scale-invariant norms. These will be norms that, upon evolution of the initial data, remain bounded above by a uniform constant (with the exception of a few anomalous terms). In particular, the dispersive estimates for the Ricci coefficients and the Weyl curvature components are restricted by their transport equations, and the signature of decay rates allows one to systematically obtain such estimates. For another application of this framework, see [4].
To each $\phi\in\begin{Bmatrix}\alpha,\underline{\alpha},\beta^{R},\underline{\beta}^{R},\rho,\sigma,\eta,\underline{\eta},\chi,\underline{\chi},\omega,\underline{\omega},\zeta,\gamma\end{Bmatrix}$ we associate its signature for decay rates $s_{2}(\phi)$:
$s_{2}(\phi)=0\cdot N_{4}(\phi)+\frac{1}{2}N_{A}(\phi)+1\cdot N_{3}(\phi)-1.$
Here $N_{\alpha}(\phi)$ $(\alpha=1,2,3,4)$ denotes the number of times $e_{\alpha}$ appears in the definition of $\phi$. We get the following tables of signatures:
$\phi$ $\alpha$ $\underline{\alpha}$ $\beta$ $\underline{\beta}$ $\rho$ $\sigma$ $\eta$ $\underline{\eta}$ $\chi$ $\underline{\chi}$ $\omega$ $\underline{\omega}$ $\zeta$ $\gamma$
$s_{2}(\phi)$ 0 2 0.5 1.5 1 1 0.5 0.5 0 1 0 1 0.5 0
Several properties of $s_{2}$ follow:
$s_{2}(\nabla_{4}\phi)=s_{2}(\phi),\hskip 5.69054pt$
$s_{2}(\nabla\phi)=s_{2}(\phi)+\frac{1}{2},\hskip 5.69054pt$
$s_{2}(\nabla_{3}\phi)=s_{2}(\phi)+1,\hskip 5.69054pt$
Finally, perhaps the most important property of $s_{2}$ is signature conservation:
(100) $s_{2}(\phi_{1}\cdot\phi_{2})=s_{2}(\phi_{1})+s_{2}(\phi_{2}),\hskip 5.69054pt.$
This allows for the (almost)-preservation of the scale-invariant norms upon evolution, as we shall see.
For any horizontal tensor-field $\phi$, we define the following norms:
(101) $\lVert{\phi}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}:=a^{-s_{2}(\phi)}\lvert u\rvert^{2s_{2}(\phi)+1}\lVert{\phi}\rVert_{L^{\infty}(S_{u,\underline{u}})},$
(102) $\lVert{\phi}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}:=a^{-s_{2}(\phi)}\lvert u\rvert^{2s_{2}(\phi)}\lVert{\phi}\rVert_{L^{2}(S_{u,\underline{u}})},$
(103) $\lVert{\phi}\rVert_{L^{1}_{(sc)}(S_{u,\underline{u}})}:=a^{-s_{2}(\phi)}\lvert u\rvert^{2s_{2}(\phi)-1}\lVert{\phi}\rVert_{L^{1}(S_{u,\underline{u}})},$
Notice the difference in the $u$-weights amongst the definitions.
[…]
(104) $\lVert{\mathrm{\Upsilon}_{1}\cdot\mathrm{\Upsilon}_{2}}\rVert_{L^{1}_{(sc)}(S_{u,\underline{u}})}\leq\frac{1}{\lvert u\rvert}\lVert{\mathrm{\Upsilon}_{1}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lVert{\mathrm{\Upsilon}_{2}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},$
(105) $\lVert{\mathrm{\Upsilon}_{1}\cdot\mathrm{\Upsilon}_{2}}\rVert_{L^{1}_{(sc)}(S_{u,\underline{u}})}\leq\frac{1}{\lvert u\rvert}\lVert{\mathrm{\Upsilon}_{1}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lVert{\mathrm{\Upsilon}_{2}}\rVert_{L^{1}_{(sc)}(S_{u,\underline{u}})},$
(106) $\lVert{\mathrm{\Upsilon}_{1}\cdot\mathrm{\Upsilon}_{2}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\leq\frac{1}{\lvert u\rvert}\lVert{\mathrm{\Upsilon}_{1}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lVert{\mathrm{\Upsilon}_{2}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}.$
Notice that this is possible partly thanks to the signature conservation property (100). In the region of study, the factor $\frac{1}{\lvert u\rvert}$ plays the role of measuring the smallness of the nonlinear terms. The above inequalities are the primary tools that will be used to close the bootstrap argument required for the existence part.
3.6. Norms
[…]
(107) $\Gamma_{i,\infty}(u,\underline{u}):=$ $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\psi_{g}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}$
(108) $+\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}$
(109) $+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\hat{\chi},\omega,\underline{\omega})}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}$
$\mathcal{R}_{i,\infty}(u,\underline{u}):=a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})},$
Furthermore, for $0\leq i\leq N+4$ and $0\leq j\leq N+4$, we define
(110) $\begin{split}\Gamma_{j,2}(u,\underline{u}):=&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ &+\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\hat{\chi},\omega,\underline{\omega})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},\end{split}$
(111) $\mathcal{R}_{i,2}(u,\underline{u}):=a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta,\underline{\beta},\rho,\sigma)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},$
Finally, for $0\leq i\leq N+4$, we define the norms along the null hypersurfaces:
(112) $\mathcal{R}_{i}(u,\underline{u}):=a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})},$
(113) $\underline{\mathcal{R}}_{i}(u,\underline{u}):=a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{\underline{u}}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}$
3.7. Commutation Formulae
Use the definition of the covariant derivatives and project it onto the topological 2-sphere $S_{u,\underline{u}}$ to yield
$[\nabla_{4},\nabla_{B}]\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{n}}=[D_{4},D_{B}]\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{n}}+(\nabla_{B}\log\Omega)\nabla_{4}\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{n}}$
$-\gamma^{CD}\chi_{BD}\nabla_{C}\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{n}}-\sum_{i=1}^{n}\gamma^{CD}\chi_{BD}\underline{\eta}_{A_{i}}\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\hat{A}_{i}C\cdot\cdot A_{n}}$
$+\sum_{i=1}^{n}\gamma^{CD}\chi_{A_{i}B}\underline{\eta}_{D}\mathcal{G}^{P}~_{QA_{1}A_{2}A_{3}\cdot\cdot\hat{A}_{i}C\cdot\cdot A_{n}}$
$[D_{4},D_{A}]\mathcal{G}^{P}~_{QA_{1}A_{2}\cdot\cdot\cdot\cdot A_{n}}=-\sum_{i}R(e_{C},e_{A_{i}},e_{4},e_{A})\mathcal{G}^{P}~_{QA_{1}\cdot\cdot\hat{A}_{i},\cdot\cdot A_{n}}+(\nabla_{A}\log\Omega)\nabla_{4}\mathcal{G}^{P}~_{QA_{1}A_{2}\cdot\cdot\cdot\cdot A_{n}}.$
Notice that the last term is redundant since it already appears in the previous expression. We need to take care of the curvature terms.
(114) $[\nabla_{4},\nabla_{A}]\mathcal{G}\sim\beta\mathcal{G}+(\eta+\underline{\eta})\nabla_{4}\mathcal{G}-\chi\nabla\mathcal{G}+\chi\underline{\eta}\mathcal{G}.$
For higher order commutation, we have the following lemma:
Lemma 3.1.
Suppose $\mathcal{G}$ is a section of the product vector bundle $^{k}\otimes T^{*}\mathbb{S}^{2}$, $k\geq 1$, that satisfies $\nabla_{4}\mathcal{G}=\mathcal{F}_{1}$ and $\nabla_{4}\nabla^{I}\mathcal{G}=\mathcal{F}^{I}_{1}$, then $\mathcal{F}^{I}_{1}$ verifies the following schematic expression:
(115) $\begin{split}\mathcal{F}^{I}_{1}\sim&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I-1}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\beta\nabla^{J_{4}}\mathcal{G}\\ &+\sum_{J_{1}+J_{2}+J_{3}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\mathcal{F}_{1}\\ +&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\hat{\nabla}^{J_{3}}\chi\nabla^{J_{4}}\mathcal{G}.\end{split}$
Similarly, for $\nabla_{3}\mathcal{G}=\mathcal{F}_{2}$, and $\nabla_{3}\nabla^{I}\mathcal{G}=\mathcal{F}^{I}_{2}$,
(116) $\begin{split}\mathcal{F}^{I}_{2}\sim&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I-1}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\underline{\beta}\nabla^{J_{4}}\mathcal{G}\\ +&\sum_{J_{1}+J_{2}+J_{3}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\mathcal{F}_{2}\\ +&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\hat{\nabla}^{J_{3}}\underline{\chi}\nabla^{J_{4}}\mathcal{G}.\end{split}$
Proof.
For $I=1$, this identity is clearly satisfied due to the calculations above. Assume it holds for J=I1J=I-1 and show that it holds for $J=I$. We omit the proof and refer to [4]. ∎
Remark 8.
By moving the top derivatives of $\mathcal{G}$ multiplied by $\mbox{tr}\underline{\chi}$ from the right-hand side to the left-hand side, one may also obtain
$\begin{split}\mathcal{F}^{I}_{2}+\frac{I}{2}\mbox{tr}\underline{\chi}\hat{\nabla}^{I}\mathcal{G}\sim&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I-1}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\beta\nabla^{J_{4}}\mathcal{G}\\ +&\sum_{J_{1}+J_{2}+J_{3}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\nabla^{J_{3}}\mathcal{F}_{2}\\ +&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}}\hat{\nabla}^{J_{3}}\widehat{\underline{\chi}}\nabla^{J_{4}}\mathcal{G}\\ +&\sum_{J_{1}+J_{2}+J_{3}+J_{4}=I-1}\nabla^{J_{1}}(\eta+\underline{\eta})^{J_{2}+1}\hat{\nabla}^{J_{3}}\mbox{tr}\underline{\chi}\nabla^{J_{4}}\mathcal{G}.\end{split}$
4. Preliminary estimates
4.1. Preliminary bootstrap assumptions
We derive the a priori estimates for the geometric and curvature norms by a bootstrap argument on the double–null development region
$D:=\Big\{(u,\underline{u},\theta^{1},\theta^{2})\ \Big|\ u_{\infty}\leq u\leq-a,\ \ 0\leq\underline{u}\leq 1\Big\}.$
Throughout, all implicit constants are universal and independent of $a$, depending only on fixed structural constants of the equations and on the number of derivatives $N$.
We employ the scale–invariant norms $\Gamma$ (Ricci coefficients) and $\mathcal{R}$ (curvature components) introduced in Section 3.6. These are defined as sums of weighted $L^{\infty}_{sc}$ and $L^{2}_{(sc)}$ norms of angular derivatives up to order $N+4$ and $N+3$, respectively. In particular, $\Gamma$ controls all Ricci coefficients
$\psi\in\{\mbox{tr}\chi,\hat{\chi},\mbox{tr}\underline{\chi},\underline{\hat{\chi}},\eta,\underline{\eta},\omega,\underline{\omega}\}$
in scale–invariant norms, while $\mathcal{R}$ controls the Weyl curvature components
$\Psi\in\{\alpha,\beta,\rho,\sigma,\underline{\beta},\underline{\alpha}\}$
through the corresponding null energy fluxes and supremum norms.
Initial bounds. Along the initial null hypersurfaces $H_{u_{\infty}}$ and $\underline{H}_{0}$, the data are prescribed so that the full hierarchy of Ricci coefficients and curvature components satisfies the dispersive scaling assumptions of Theorem 2.2. A direct analysis of the null constraint and transport equations along the initial hypersurfaces (cf. [6] and references therein) yields the quantitative bound
(117) $\Gamma_{0}+\mathcal{R}_{0}\;\lesssim\;\mathcal{I},$
where $\Gamma_{0},\mathcal{R}_{0}$ denote the initial values of the norms on $H_{u_{\infty}}\cup\underline{H}_{0}$ and $\mathcal{I}$ is the size of the prescribed data.
Bootstrap assumptions. We assume, on the spacetime region $D$, the bootstrap bounds
(118) $\Gamma\leq\Gamma_{\ast},\qquad\mathcal{R}\leq R_{\ast},$
for fixed constants $\Gamma_{\ast},R_{\ast}\geq 1$ to be chosen. These constants are taken sufficiently large so that
(119) $\mathcal{I}^{4}+\mathcal{I}^{2}+\mathcal{I}+1\;\ll\;\min\{\Gamma_{\ast},R_{\ast},M_{\ast}\},$
where $M_{\ast}$ denotes the corresponding metric norm bound, and at the same time satisfy the compatibility condition
(120) $(\Gamma_{\ast}+R_{\ast})^{20}\leq a^{1/16}.$
The smallness encoded in (120) ensures that all error terms produced by nonlinear interactions of Ricci and curvature components remain perturbative after integration in $u$ and $\underline{u}$.
The aim is to prove that the estimates implied by (118) can in fact be improved to
$\Gamma+\mathcal{R}\;\lesssim\;c(\mathcal{I}),\qquad c(\mathcal{I}):=\mathcal{I}^{4}+\mathcal{I}^{2}+\mathcal{I}+1,$
throughout $D$. By a standard continuity argument, this yields a closed a priori bound and therefore semi–global control of the solution.
Estimates on the metric components
We first control the metric quantities $\Omega$, the induced sphere metric $\gamma$, and the area radius of $S_{u,\underline{u}}$. We begin with the null lapse $\Omega$.
Proposition 4.1.
Under the assumptions of Theorem 2.2 and the bootstrap bounds (118), the null lapse satisfies, for every sphere $S_{u,\underline{u}}\subset D$,
$\|\Omega-1\|_{L^{\infty}(S_{u,\underline{u}})}\;\lesssim\;\frac{\Gamma_{\ast}\,a^{1/2}}{|u|^{2}}.$
Proof.
In double–null gauge the lapse obeys the transport equation
(121) $\nabla_{3}\log\Omega=-\,\underline{\omega}.$
On the initial outgoing hypersurface $H_{u_{\infty}}$ the data normalization gives
$\Omega(u_{\infty},\underline{u},\theta)=1\quad\text{for all }\underline{u}\in[0,1],$
hence $\log\Omega=0$ there. Integrating (121) along the incoming null generators from $u_{\infty}$ to $u$ at fixed $(\underline{u},\theta)$ yields
$\log\Omega(u,\underline{u},\theta)=-\int_{u_{\infty}}^{u}\underline{\omega}(u^{\prime},\underline{u},\theta)\,du^{\prime}.$
By the bootstrap bound on Ricci coefficients and the scale–invariant weights,
$|\underline{\omega}(u^{\prime},\underline{u},\theta)|\;\lesssim\;\frac{\Gamma_{\ast}\,a^{1/2}}{|u^{\prime}|^{3}}.$
Therefore,
$|\log\Omega(u,\underline{u},\theta)|\;\lesssim\;\Gamma_{\ast}a^{1/2}\int_{u_{\infty}}^{u}\frac{du^{\prime}}{|u^{\prime}|^{3}}\;\lesssim\;\frac{\Gamma_{\ast}a^{1/2}}{|u|^{2}}.$
Since the right–hand side is $\ll 1$ by (120), we conclude
$|\Omega(u,\underline{u},\theta)-1|\;\lesssim\;\frac{\Gamma_{\ast}a^{1/2}}{|u|^{2}}.$
Taking the supremum over $S_{u,\underline{u}}$ gives the stated estimate. ∎
Proposition 4.2.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), there exist two constants $c$ and $C$ depending only on the initial data such that the bounds
$c\leq\text{det}\hskip 1.42262pt\gamma\leq C.$
and
$\lvert\gamma_{AB}\rvert+\lvert\gamma^{-1}_{AB}\rvert\leq C$
hold throughout the slab of existence $D$.
Proposition 4.3.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), fix a point $(u,\theta)$ on the initial hypersurface $H_{\infty}$. Let $\Lambda(u)$ and $\lambda(u)$ be the largest and smallest eigenvalues of $\gamma^{-1}(u_{\infty},\underline{u},\theta)\hskip 1.42262pt\gamma(u,\underline{u},\theta)$ respectively, along the incoming null geodesics emanating from $(\underline{u},\theta)$. There holds
$\lvert\Lambda(u)-1\rvert+\lvert\lambda(u)-1\rvert\lesssim\frac{1}{a^{\frac{1}{2}}}.$
4.2. Estimates for transport equations
We shall be using two fundamental bounds on transport equations throughout this work.
Proposition 4.4.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), the following hold for an arbitrary $\mathcal{G}\in\Gamma(^{N}\otimes T^{*}S):$
(122) $\lVert{\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\lVert\mathcal{G}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime\prime}})}+\int_{\underline{u}^{\prime\prime}}^{\underline{u}}\lVert{\nabla_{4}\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
(123) $\lVert{\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\lVert\mathcal{G}\rVert_{L^{2}_{(sc)}(S_{u^{\prime\prime},\underline{u}})}+\int_{u^{\prime\prime}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla_{3}\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
There are, however, cases that are borderline and require more delicate control than what the above Proposition provides. These have to do with components $X$ satisfying an equation of the form $\nabla_{3}X=-\lambda\mbox{tr}\underline{\chi}X+\dots$, where $\lambda>0$. Keeping in mind that $\mbox{tr}\underline{\chi}$ is the worst Ricci coefficient in terms of peeling, one would hope to be able to get rid of its appearance and thus obtain stronger bounds regarding the peeling properties of $X$. The following weighted transport inequality achieves this.
Proposition 4.5.
Let $\mathcal{G},\mathcal{H}\in\Gamma(^{N}\otimes T^{*}S)$ and assume that the following equation holds:
$\nabla_{3}\mathcal{G}+\lambda_{0}\hskip 1.42262pt\mbox{tr}\underline{\chi}\hskip 1.42262pt\mathcal{G}=\mathcal{H}.$
Then, under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), the following is true:
$\lvert u\rvert^{\lambda_{1}}\lVert{\mathcal{G}}\rVert_{L^{2}(S_{u,\underline{u}})}\lesssim\lvert u_{\infty}\rvert^{\lambda_{1}}\lVert\mathcal{G}\rVert_{L^{2}(S_{u_{\infty},\underline{u}})}+\int_{u_{\infty}}^{u}\lvert u^{\prime}\rvert^{\lambda_{1}}\lVert\mathcal{H}\rVert_{L^{2}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
for $\lambda_{1}=2\lambda_{0}-1$.
Proof.
The variation of area formula for a scalar function $f$ reads:
(124) $\underline{L}\int_{S_{u,\underline{u}}}f=\int_{S_{u,\underline{u}}}\underline{L}f+\Omega\hskip 1.42262pt\mbox{tr}\underline{\chi}\hskip 1.42262ptf=\int_{S_{u,\underline{u}}}\Omega\hskip 1.42262pt\left(e_{3}(f)+\mbox{tr}\underline{\chi}\hskip 1.42262ptf\right).$
Plugging in $f=\lvert u\rvert^{2\lambda_{1}}\lvert\mathcal{G}\rvert_{\gamma}^{2}$, we calculate:
(125) $\begin{split}&\underline{L}\int_{S_{u,\underline{u}}}\lvert u\rvert^{2\lambda_{1}}\lvert\mathcal{G}\rvert_{\gamma}^{2}\\ =&\int_{S_{u,\underline{u}}}\Omega\left(-2\hskip 1.42262pt\lambda_{1}\lvert u\rvert^{2\hskip 1.42262pt\lambda_{1}-1}e_{3}(u)\lvert\mathcal{G}\rvert_{\gamma}^{2}+2\lvert u\rvert^{2\hskip 1.42262pt\lambda_{1}}\langle\mathcal{G},\nabla_{3}\mathcal{G}\rangle_{\gamma}+\mbox{tr}\underline{\chi}\hskip 1.42262pt\lvert u\rvert^{2\lambda_{1}}\hskip 1.42262pt\lvert\mathcal{G}\rvert_{\gamma}^{2}\right)\\ =&\int_{S_{u,\underline{u}}}\Omega\left(2\lvert u\rvert^{2\lambda_{1}}\langle\mathcal{G},\nabla_{3}\mathcal{G}+\lambda_{0}\mbox{tr}\underline{\chi}\mathcal{G}\rangle_{\gamma,\epsilon}\right)+\int_{S_{u,\underline{u}}}\Omega\lvert u\rvert^{2\lambda_{1}}\left(\frac{-2\lambda_{1}e_{3}(u)}{\lvert u\rvert}+(1-2\lambda_{0})\hskip 1.42262pt\mbox{tr}\underline{\chi}\right)\lvert\mathcal{G}\rvert_{\gamma}^{2}.\end{split}$
Notice that
(126) $\begin{split}&\frac{-2\lambda_{1}e_{3}(u)}{\lvert u\rvert}+(1-2\lambda_{0})\hskip 1.42262pt\mbox{tr}\underline{\chi}\\ =&\frac{-2\lambda_{1}(\Omega^{-1}-1)}{\lvert u\rvert}+(1-2\lambda_{0})(\mbox{tr}\underline{\chi}+\frac{2}{\lvert u\rvert})-\frac{2\lambda_{1}+2-4\lambda_{0}}{\lvert u\rvert}\\ \leq&\frac{\Gamma}{\lvert u\rvert^{2}},\end{split}$
where we have used the bootstrap assumption $\lVert{\mbox{tr}\underline{\chi}+\frac{2}{\lvert u\rvert}}\rVert_{L^{\infty}(S_{u,\underline{u}})}\leq\frac{\Gamma}{\lvert u\rvert^{2}}$ and the definition of $\lambda_{1}$. For the first term in the last line of (125) we then use Cauchy-Schwartz and for the second we apply Grönwall’s inequality to get:
(127) $\begin{split}&\lvert u\rvert^{2\lambda_{1}}\lVert{\mathcal{G}}\rVert_{L^{2}(S_{u,\underline{u}})}\\ \lesssim&\mathrm{e}^{\Gamma\lVert u^{-2}\rVert_{L_{u}^{1}}}\left(\lvert u_{\infty}\rvert^{\lambda_{1}}\lVert\mathcal{G}\rVert_{L^{2}(S_{u_{\infty},\underline{u}})}+\int_{u_{\infty}}^{u}\lvert u^{\prime}\rvert^{\lambda_{1}}\lVert\mathcal{H}\rVert_{L^{2}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\right)\\ \lesssim&\lvert u_{\infty}\rvert^{\lambda_{1}}\lVert\mathcal{G}\rVert_{L^{2}(S_{u_{\infty},\underline{u}})}+\int_{u_{\infty}}^{u}\lvert u^{\prime}\rvert^{\lambda_{1}}\lVert\mathcal{H}\rVert_{L^{2}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime},\end{split}$
where we have used the fact that $\mathrm{e}^{\Gamma\lVert u^{-2}\rVert_{L_{u}^{1}}}\lesssim\mathrm{e}^{\Gamma/a}\lesssim 1$.
4.3. Sobolev embedding
With the derived estimates for the metric $\gamma$, we can obtain a bound on the isoperimetric constant for a topological 22-sphere $S$:
(128) $I(S):=\sup_{U\subset S,~\partial U\in C^{1}}\frac{\min\left\{Area(U),Area(U^{c})\right\}}{[Perimeter(\partial U)]^{2}}.$
The following proposition yields an upper bound for $I(S)$.
Proposition 4.6 (Uniform control of the isoperimetric constant of $S_{u,\underline{u}}$).
Assume the initial data hypotheses and the bootstrap bounds (2.10) for the double–null development. Then for every sphere of the foliation
$S_{u,\underline{u}},\qquad u\in[u_{\infty},-a],\ \underline{u}\in[0,1],$
the isoperimetric constant satisfies the uniform estimate
(129) $I(S_{u,\underline{u}})\;\leq\;\frac{1}{\pi}.$
Here $I(S,\gamma)$ denotes the isoperimetric constant of the Riemannian 22–sphere $(S,\gamma)$, defined by
$I(S,\gamma):=\sup_{U\subset S}\frac{\min\{\text{Area}(U),\text{Area}(U^{c})\}}{\text{Per}(\partial U)^{2}},$
where the supremum ranges over all domains $U\subset S$ with smooth boundary.
Proof.
We first note that the isoperimetric constant is scale invariant: if $\gamma^{\prime}=\lambda^{2}\gamma$ on a surface $S$, then
$\frac{\min\{\text{Area}_{\gamma^{\prime}}(U),\text{Area}_{\gamma^{\prime}}(U^{c})\}}{\text{Per}_{\gamma^{\prime}}(\partial U)^{2}}=\frac{\lambda^{2}\min\{\text{Area}_{\gamma}(U),\text{Area}_{\gamma}(U^{c})\}}{\lambda^{2}\text{Per}_{\gamma}(\partial U)^{2}},$
hence $I(S,\gamma^{\prime})=I(S,\gamma)$. Therefore it suffices to prove the estimate for the renormalized metric
$\tilde{\gamma}_{u,\underline{u}}:=|u|^{-2}\gamma_{u,\underline{u}}\quad\text{on }S_{u,\underline{u}}.$
Suppose a metric $\gamma$ on $S^{2}$ satisfies, for some $\varepsilon\in(0,1)$,
(130) $(1-\varepsilon)\gamma_{0}\leq\gamma\leq(1+\varepsilon)\gamma_{0}$
as quadratic forms, where $\gamma_{0}$ is the unit round metric. Then for every smooth domain $U\subset S$,
$\text{Area}_{\gamma}(U)\leq(1+\varepsilon)\text{Area}_{\gamma_{0}}(U),\qquad\text{Area}_{\gamma}(U^{c})\leq(1+\varepsilon)\text{Area}_{\gamma_{0}}(U^{c}),$
and
$\text{Per}_{\gamma}(\partial U)\geq(1-\varepsilon)^{1/2}\text{Per}_{\gamma_{0}}(\partial U),\quad\Rightarrow\quad\text{Per}_{\gamma}(\partial U)^{2}\geq(1-\varepsilon)\text{Per}_{\gamma_{0}}(\partial U)^{2}.$
Hence
$\frac{\min\{\text{Area}_{\gamma}(U),\text{Area}_{\gamma}(U^{c})\}}{\text{Per}_{\gamma}(\partial U)^{2}}\leq\frac{1+\varepsilon}{1-\varepsilon}\frac{\min\{\text{Area}_{\gamma_{0}}(U),\text{Area}_{\gamma_{0}}(U^{c})\}}{\text{Per}_{\gamma_{0}}(\partial U)^{2}}.$
Taking the supremum gives
(131) $I(S,\gamma)\leq\frac{1+\varepsilon}{1-\varepsilon}I(S,\gamma_{0}).$
For the unit round sphere one has $I(S^{2},\gamma_{0})=\frac{1}{2\pi}$. Thus if $\varepsilon\leq\frac{1}{3}$, then
(132) $I(S,\gamma)\leq\frac{1+\varepsilon}{1-\varepsilon}\frac{1}{2\pi}\leq\frac{1}{\pi}.$
In the double–null foliation the induced metric satisfies the transport equations
$\nabla_{4}\gamma_{AB}=2\chi_{AB},\qquad\nabla_{3}\gamma_{AB}=2\underline{\chi}_{AB},$
that is,
$\nabla_{4}\gamma_{AB}=(\mbox{tr}\chi)\gamma_{AB}+2\hat{\chi}_{AB},\qquad\nabla_{3}\gamma_{AB}=(\mbox{tr}\underline{\chi})\gamma_{AB}+2\underline{\hat{\chi}}_{AB}.$
Define $\tilde{\gamma}_{AB}=|u|^{-2}\gamma_{AB}$. Using $\nabla_{4}u=0$ and the standard normalization of $\nabla_{3}u$, one obtains schematically
(133) $\nabla_{4}\tilde{\gamma}_{AB}$ $=\Big(\mbox{tr}\chi-\frac{2}{|u|}\Big)\tilde{\gamma}_{AB}+2|u|^{-2}\hat{\chi}_{AB},$
(134) $\nabla_{3}\tilde{\gamma}_{AB}$ $=\Big(\mbox{tr}\underline{\chi}+\frac{2}{|u|}\Big)\tilde{\gamma}_{AB}+2|u|^{-2}\underline{\hat{\chi}}_{AB}+\text{l.o.t.}$
By the initial data assumptions on $H_{u_{\infty}}$ we have
$\tilde{\gamma}(u_{\infty},0)=\gamma_{0}+O(a^{-1/2}|u_{\infty}|^{-3}).$
The bootstrap bounds (2.10) give, uniformly for $u\in[u_{\infty},-a]$, $\underline{u}\in[0,1]$,
$\Big|\mbox{tr}\chi-\frac{2}{|u|}\Big|\lesssim a^{-1/2}|u|^{-1},\qquad|\hat{\chi}|\lesssim a^{-1/2}|u|^{-1},$
$\Big|\mbox{tr}\underline{\chi}+\frac{2}{|u|}\Big|\lesssim|u|^{-2},\qquad|\underline{\hat{\chi}}|\lesssim a^{1/2}|u|^{-2}.$
Since $\underline{u}\in[0,1]$ and $|u|\geq a\gg 1$, Grönwall’s inequality applied to (133) along the $\nabla_{4}$ direction and to (134) along the $\nabla_{3}$ direction yields
(135) $\sup_{u\in[u_{\infty},-a],\ \underline{u}\in[0,1]}\|\tilde{\gamma}_{u,\underline{u}}-\gamma_{0}\|_{C^{0}(S_{u,\underline{u}})}\lesssim a^{-1/2}.$
For $a\geq a_{0}$ sufficiently large, this implies the bilipschitz comparison (130) with some $\varepsilon\leq\frac{1}{3}$, uniformly for all $S_{u,\underline{u}}$ in the slab.
By scale invariance,
$I(S_{u,\underline{u}},\gamma_{u,\underline{u}})=I(S_{u,\underline{u}},\tilde{\gamma}_{u,\underline{u}}).$
Combining (135) with (132) gives
$I(S_{u,\underline{u}})\leq\frac{1}{\pi}$
uniformly for all $u\in[u_{\infty},-a]$, $\underline{u}\in[0,1]$. This proves the proposition. ∎
Throughout this work, we will be using an $L^{2}-L^{\infty}$ Sobolev estimate. To obtain it, utilizing the basic estimates above, we may proceed to write down the following gauge-invariant Sobolev inequalities for the topological 22- sphere $S$.
Proposition 4.7 (Sobolev inequality for tensorfields in terms of the isoperimetric constant).
Let $(S,\gamma)$ be a smooth compact Riemannian 22–manifold. Denote by $\text{Area}(S)$ its $\gamma$–area and by $I(S)$ its isoperimetric constant
$I(S):=\sup_{U\subset S}\frac{\min\{\text{Area}(U),\text{Area}(U^{c})\}}{\text{Per}(\partial U)^{2}},$
where the supremum is taken over all domains $U\subset S$ with smooth boundary. Then for every $p\in(2,\infty)$ there exists a constant $C_{p}<\infty$ depending only on $p$ such that the following holds.
Let $\mathcal{G}$ be any smooth tensorfield of type $(0,N)$ on $S$ (in particular, $\mathcal{G}\in\Gamma(\otimes^{N}T^{*}S)$; more generally the same statement holds for any tensor bundle equipped with the metric induced by $\gamma$). Then
(136) $\text{Area}(S)^{-\frac{1}{p}}\,\|\mathcal{G}\|_{L^{p}(S)}\;\leq\;C_{p}\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\Big(\|\nabla\mathcal{G}\|_{L^{2}(S)}+\text{Area}(S)^{-\frac{1}{2}}\,\|\mathcal{G}\|_{L^{2}(S)}\Big).$
Proof.
We use the scalar Sobolev inequality on $(S,\gamma)$ in the formulation controlled by the isoperimetric constant: for each $p\in(2,\infty)$ there exists $C_{p}$ such that for every $f\in W^{1,2}(S)$,
(137) $\text{Area}(S)^{-\frac{1}{p}}\,\|f\|_{L^{p}(S)}\;\leq\;C_{p}\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\Big(\|\nabla f\|_{L^{2}(S)}+\text{Area}(S)^{-\frac{1}{2}}\,\|f\|_{L^{2}(S)}\Big).$
(See, e.g., the Federer–Fleming/Maz’ya isoperimetric–Sobolev inequality specialized to dimension $2$.)
To deduce (136) from (137), we apply (137) to the regularized norm of $\mathcal{G}$. For $\delta>0$ define the Lipschitz function
$f_{\delta}:=\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}\,,\qquad|\mathcal{G}|_{\gamma}^{2}:=\gamma^{A_{1}B_{1}}\cdots\gamma^{A_{N}B_{N}}\,\mathcal{G}_{A_{1}\cdots A_{N}}\,\mathcal{G}_{B_{1}\cdots B_{N}}.$
Since $f_{\delta}\in W^{1,2}(S)$ and $S$ is compact, (137) applies to $f_{\delta}$.
Estimate of $\nabla f_{\delta}$. By metric compatibility $\nabla\gamma=0$ and the Leibniz rule,
$\nabla(|\mathcal{G}|_{\gamma}^{2})=2\langle\mathcal{G},\nabla\mathcal{G}\rangle_{\gamma},$
where $\langle\cdot,\cdot\rangle_{\gamma}$ denotes the pointwise inner product induced by $\gamma$ on the relevant tensor bundle. Hence, in the weak sense (and pointwise a.e. since $f_{\delta}$ is Lipschitz),
$\nabla f_{\delta}=\frac{1}{2}(|\mathcal{G}|_{\gamma}^{2}+\delta)^{-1/2}\,\nabla(|\mathcal{G}|_{\gamma}^{2})=\frac{\langle\mathcal{G},\nabla\mathcal{G}\rangle_{\gamma}}{\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}}.$
By Cauchy–Schwarz,
$|\nabla f_{\delta}|\leq\frac{|\mathcal{G}|_{\gamma}\,|\nabla\mathcal{G}|_{\gamma}}{\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}}\leq|\nabla\mathcal{G}|_{\gamma},$
and therefore
(138) $\|\nabla f_{\delta}\|_{L^{2}(S)}\leq\|\nabla\mathcal{G}\|_{L^{2}(S)}.$
Moreover, since $f_{\delta}\geq|\mathcal{G}|_{\gamma}$ and $f_{\delta}\downarrow|\mathcal{G}|_{\gamma}$ pointwise as $\delta\to 0$,
(139) $\|f_{\delta}\|_{L^{2}(S)}\to\|\mathcal{G}\|_{L^{2}(S)},\qquad\|f_{\delta}\|_{L^{p}(S)}\to\|\mathcal{G}\|_{L^{p}(S)}\quad\text{for every }p\in[1,\infty),$
by monotone convergence (or dominated convergence on the compact manifold $S$).
Applying (137) to $f_{\delta}$ and using (138) gives
$\text{Area}(S)^{-\frac{1}{p}}\,\|f_{\delta}\|_{L^{p}(S)}\;\leq\;C_{p}\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\Big(\|\nabla\mathcal{G}\|_{L^{2}(S)}+\text{Area}(S)^{-\frac{1}{2}}\,\|f_{\delta}\|_{L^{2}(S)}\Big).$
Letting $\delta\to 0$ and using (139) yields (136). ∎
Proposition 4.8 ($L^{\infty}$ Sobolev embedding in terms of the isoperimetric constant).
Let $(S,\gamma)$ be a smooth compact Riemannian 22–manifold, with isoperimetric constant $I(S)$ and area $\text{Area}(S)$. Fix $p\in(2,\infty)$. Then there exists a constant $C_{p}<\infty$, depending only on $p$, such that the following holds.
For every smooth tensorfield $\mathcal{G}$ of type $(0,N)$ on $S$ (in particular $\mathcal{G}\in\Gamma(\otimes^{N}T^{*}S)$, or more generally any tensor bundle endowed with the norm induced by $\gamma$),
(140) $\|\mathcal{G}\|_{L^{\infty}(S)}\;\leq\;C_{p}\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\Big(\|\nabla\mathcal{G}\|_{L^{p}(S)}+\text{Area}(S)^{-\frac{1}{2}}\,\|\mathcal{G}\|_{L^{p}(S)}\Big).$
Proof.
We reduce to the scalar inequality and then pass to tensorfields by a gauge–invariant regularization argument, as in Proposition 4.7.
First control scalar $W^{1,p}\hookrightarrow L^{\infty}$ with isoperimetric control in 22-dimensions. Let $f\in W^{1,p}(S)$ with $p>2$. By the isoperimetric inequality, the $L^{1}$–Sobolev inequality holds with constant controlled by $\max\{1,I(S)\}^{1/2}$:
(141) $\|f-\bar{f}\|_{L^{2}(S)}\;\leq\;C\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\|\nabla f\|_{L^{1}(S)},\qquad\bar{f}:=\text{Area}(S)^{-1}\int_{S}f.$
Interpolating between $L^{1}$ and $L^{p}$ norms of $\nabla f$ and using Hölder’s inequality yields, with a constant depending only on $p$,
(142) $\|f-\bar{f}\|_{L^{2}(S)}\;\leq\;C_{p}\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\,\|\nabla f\|_{L^{p}(S)}.$
Next, using the standard Morrey embedding in dimension $2$ on compact manifolds (which can be proved by covering $S$ with finitely many coordinate charts and reducing to the Euclidean Morrey inequality), we have
(143) $\|f-\bar{f}\|_{L^{\infty}(S)}\;\leq\;C_{p}\Big(\|\nabla f\|_{L^{p}(S)}+\text{Area}(S)^{-1/2}\|f-\bar{f}\|_{L^{2}(S)}\Big).$
Combining (142) and (143) gives
(144) $\|f-\bar{f}\|_{L^{\infty}(S)}\;\leq\;C_{p}\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\,\|\nabla f\|_{L^{p}(S)}.$
Finally, control $\bar{f}$ by Hölder:
$|\bar{f}|\leq\text{Area}(S)^{-1}\|f\|_{L^{1}(S)}\leq\text{Area}(S)^{-1/p}\|f\|_{L^{p}(S)}.$
Thus, from (144),
(145) $\|f\|_{L^{\infty}(S)}\;\leq\;C_{p}\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\Big(\|\nabla f\|_{L^{p}(S)}+\text{Area}(S)^{-\frac{1}{2}}\|f\|_{L^{p}(S)}\Big).$
Now let $\mathcal{G}$ be a smooth tensorfield. For $\delta>0$ define
$f_{\delta}:=\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}\in W^{1,p}(S).$
As in the proof of Proposition 4.7, metric compatibility and the chain rule yield (a.e.)
$\nabla f_{\delta}=\frac{\langle\mathcal{G},\nabla\mathcal{G}\rangle_{\gamma}}{\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}},\qquad|\nabla f_{\delta}|\leq|\nabla\mathcal{G}|_{\gamma},$
hence
(146) $\|\nabla f_{\delta}\|_{L^{p}(S)}\leq\|\nabla\mathcal{G}\|_{L^{p}(S)}.$
Moreover, $f_{\delta}\downarrow|\mathcal{G}|_{\gamma}$ pointwise and monotonically as $\delta\to 0$, so
(147) $\|f_{\delta}\|_{L^{p}(S)}\to\|\mathcal{G}\|_{L^{p}(S)},\qquad\|f_{\delta}\|_{L^{\infty}(S)}\to\|\mathcal{G}\|_{L^{\infty}(S)}.$
Apply the scalar inequality (145) to $f_{\delta}$ and use (146):
$\|f_{\delta}\|_{L^{\infty}(S)}\;\leq\;C_{p}\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\Big(\|\nabla\mathcal{G}\|_{L^{p}(S)}+\text{Area}(S)^{-\frac{1}{2}}\|f_{\delta}\|_{L^{p}(S)}\Big).$
Letting $\delta\to 0$ and using (147) yields (140). ∎
The two inequalities above, together with Propositions 4.1-LABEL:34, allow us to control the $L^{2}$-norm of $\mathcal{G}$ in terms of its $H^{2}$-norm. Following the area estimates, we have $Area(S_{u,\underline{u}})\approx u^{2}$. Therefore, we obtain the following important inequality.
Proposition 4.9 (Sobolev embedding on $S_{u,\underline{u}}$).
Assume the hypotheses of Theorem 2.2 and the bootstrap assumptions (118). Let
$S_{u,\underline{u}},\qquad u\in[u_{\infty},-a],\ \underline{u}\in[0,1],$
be the 22–spheres of the canonical double–null foliation with induced metric $\gamma=\gamma(u,\underline{u})$, Levi–Civita connection $\nabla$, and area form $d\mu_{\gamma}$. Then for every smooth tensorfield $\mathcal{G}\in\Gamma(\otimes^{N}T^{*}S_{u,\underline{u}})$ one has the uniform Sobolev estimate
(148) $\|\mathcal{G}\|_{L^{\infty}(S_{u,\underline{u}})}\;\lesssim\;\sum_{I=0}^{2}\big\||u|^{I-1}\nabla^{I}\mathcal{G}\big\|_{L^{2}(S_{u,\underline{u}})},$
where the implicit constant is universal (independent of $u,\underline{u},a$) and depends only on the fixed bootstrap constants.
Equivalently, in the scale–invariant norms associated to the spherical scale $|u|$ (and the parameter $a$ fixed in Theorem 2.2)),
(149) $\|\mathcal{G}\|_{L^{\infty}_{sc}(S_{u,\underline{u}})}\;\lesssim\;\sum_{I=0}^{2}\big\|(a^{1/2}\nabla)^{I}\mathcal{G}\big\|_{L^{2}_{sc}(S_{u,\underline{u}})}.$
Proof sketch.
Fix $(u,\underline{u})$. Apply the $L^{\infty}$ Sobolev inequality of Proposition 4.8 to $(S_{u,\underline{u}},\gamma)$ with $p=4$:
$\|\mathcal{G}\|_{L^{\infty}(S_{u,\underline{u}})}\;\lesssim\;\big(\max\{1,I(S_{u,\underline{u}})\}\big)^{1/2}\,\text{Area}(S_{u,\underline{u}})^{1/4}\Big(\|\nabla\mathcal{G}\|_{L^{4}(S_{u,\underline{u}})}+\text{Area}(S_{u,\underline{u}})^{-1/2}\|\mathcal{G}\|_{L^{4}(S_{u,\underline{u}})}\Big).$
Next, estimate the $L^{4}$–terms by the $L^{2}$–Sobolev inequality of Proposition 4.7 with $p=4$, yielding control by $\|\nabla^{2}\mathcal{G}\|_{L^{2}}+\text{Area}(S_{u,\underline{u}})^{-1/2}\|\nabla\mathcal{G}\|_{L^{2}}+\text{Area}(S_{u,\underline{u}})^{-1}\|\mathcal{G}\|_{L^{2}}$. Finally, use the geometric bounds provided by Proposition 4.3 (area comparability $\text{Area}(S_{u,\underline{u}})\sim|u|^{2}$) and Proposition 4.6 (uniform isoperimetric control $I(S_{u,\underline{u}})\lesssim 1$) to rewrite the resulting estimate in the form (148). The scale-invariant formulation (149) is the same inequality expressed in the rescaled norms by factoring out the natural spherical scale $|u|$ (equivalently, the $a^{1/2}\nabla$ normalization used throughout the bootstrap). ∎
4.4. Estimates on the Ricci coefficients
Proposition 4.10 (Scale–invariant $L^{2}$ control of $\nabla\omega$).
Assume the hypotheses of Theorem 2.2 and the bootstrap assumptions (118). Then, for all
$u\in[u_{\infty},-a],\qquad\underline{u}\in[0,\epsilon],$
the lapse/vorticity coefficient $\omega$ satisfies the scale–invariant estimate
(150) $\sum_{i\leq N+4}\,\lVert{(a^{\frac{1}{2}}\nabla)^{i}\omega}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\;\lesssim\;\frac{a^{1/2}}{|u|}\;+\;\frac{a^{1/2}}{|u|}\,\underline{\mathcal{R}}[\rho].$
The implicit constant depends only on the fixed bootstrap constants and on $N$.
Proof.
We begin by recalling that $\omega$ satisfies the schematic equation
$\nabla_{3}\omega=\frac{1}{2}\rho+\psi_{g}\psi_{g}.$
Using the commutation formula 8 and the notation of Section 3.6, we have, for a general $i$:
(151) $\begin{split}\nabla_{3}\nabla^{i}\omega+\frac{i}{2}\mbox{tr}\underline{\chi}\nabla^{i}\omega=&\nabla^{i}\rho+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega\end{split}$
Note that since $\omega$ is not a section of the vector bundle $^{k}\otimes T^{*}S\otimes P_{Ad,\mathfrak{g}}$, the last term on the right-hand side of 8 does not appear. Passing to scale-invariant norms, we get
$\lVert{(a^{\frac{1}{2}}\nabla)^{i}\omega}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}$
$\lesssim$ $\lVert(a^{\frac{1}{2}}\nabla)^{i}\omega\rVert_{L^{2}_{(sc)}(S_{u_{\infty}},0)}+\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\rho}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}=i-1}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}.$
For $0\leq i\leq N+4$, the first term, by virtue of the fact that $\Omega\equiv 1$ initially, vanishes. The second term can be bounded, using Hölder’s inequality, by $\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}\underline{\mathcal{R}}[\rho].$ The third, fourth, and fifth terms can be bounded above by
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\hskip 1.42262pt\text{d}u^{\prime}\leq\frac{a^{\frac{3}{2}}\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert^{3}}.$
The sixth term is controlled as follows:
(152) $\begin{split}&\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{2}}}\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert^{3}}.\end{split}$
The seventh term is controlled as follows:
$\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(153) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\lvert u^{\prime}\rvert^{2}}{a}\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma^{3}}{\lvert u\rvert^{2}}.$
For the eighth and most borderline term, we estimate:
$\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$=$ $\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\frac{a^{\frac{3}{2}}}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(154) $\lesssim$ $\int_{u_{\infty}}^{u}\int_{u_{\infty}}^{u}\frac{a^{2}}{|u^{{}^{\prime}}|^{2}}\frac{1}{|u^{{}^{\prime}}|^{2}}\frac{|u^{{}^{\prime}}|}{a^{\frac{1}{2}}}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\Gamma^{3}du^{{}^{\prime}}\lesssim\frac{a^{2}\Gamma^{3}}{|u|^{3}}.$
Therefore collecting every terms and by boot-strap
$\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\omega}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{\lvert u\rvert}+\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\underline{\mathcal{R}}[\rho].$
This completes the proof. ∎
Proposition 4.11.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), there hold
$\sum_{i\leq N+4}\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1,\hskip 5.69054pt\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{1}{a^{\frac{1}{2}}}(\mathcal{R}[\alpha]+1).$
Proof.
The proof is exactly similar to that of [12]. The point to note here is that $\hat{\chi}$ and $\alpha$ have same scaling. ∎
The estimates for $\underline{\omega}$ are, in a sense, dual to those for $\omega$.
Proposition 4.12.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), there holds
$\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\leq\frac{1}{a^{\frac{1}{2}}}\mathcal{R}[\rho]+\frac{1}{a^{\frac{1}{2}}}.$
Proof.
We have, schematically,
$\nabla_{4}\underline{\omega}=\frac{1}{2}\rho+\psi_{g}\psi_{g}.$
As before, the schematic product of the Yang-Mills components is gauge-invariant. Using the commutation formula 3.1, we have, for a general $i$:
(155) $\begin{split}\nabla_{4}\nabla^{i}\underline{\omega}=&\nabla^{i}\rho+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}(\hat{\chi},\omega)\\ &++\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}\end{split}$
Passing to scale-invariant norms, we have,
(156) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ \lesssim&\lVert(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}\rVert_{L^{2}_{(sc)}(S_{u,0})}+\int_{0}^{\underline{u}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\rho}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
For $0\leq i\leq N+4$, the first four terms are controlled as in Proposition 4.10 and are bounded above by $\mathcal{R}[\rho]+1$. For the next terms, there holds
(157) $\begin{split}&\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{0}^{\underline{u}}a^{\frac{1}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\frac{\Gamma^{2}}{\lvert u\rvert^{2}}.\end{split}$
Working similarly, there hold
(158) $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{a\hskip 1.42262pt\Gamma^{3}}{\lvert u\rvert^{2}}.$
The claim follows. ∎
We move on to estimates for $\eta$.
Proposition 4.13.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), there holds
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\eta}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}\mathcal{R}[\beta]+1.$
Proof.
We begin with the schematic structure equation for $\eta$:
$\nabla_{4}\eta=\beta+(\hat{\chi},\mbox{tr}\chi)\cdot(\eta,\underline{\eta}).$
Using the commutation formula 3.1 for the $\nabla_{4}$-direction, we have
(159) $\begin{split}\nabla_{4}\nabla^{i}\eta=&\nabla^{i}\beta+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\beta^{R}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}(\eta,\underline{\eta})\end{split}$
Estimating in scale-invariant norms, we have
[…]
For $0\leq i\leq N+4$, the first term is bounded by $a^{-\frac{1}{2}}\mathcal{R}[\beta]$. The second term is bounded by
[…]
Notice that, since $i_{3}\leq i-1\leq N+3$, we can bound $i_{3}$ derivatives of $\beta$ using the bootstrap assumption (118) on the total norm $\Gamma$. For the third term, there holds
$\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}(\eta,\underline{\eta})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert}$.
The result follows. ∎
Proposition 4.14.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), there holds
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\chi}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|}\mathcal{R}[\alpha]+1.$
Proof.
We begin by recalling the schematic equation
$\nabla_{4}\mbox{tr}\chi=\lvert\hat{\chi}\rvert^{2}+\psi_{g}\psi_{g}.$
Commuting with $i$ angular derivatives using 3.1, we obtain
(162) $\begin{split}\nabla_{4}\nabla^{i}\mbox{tr}\chi=&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\psi_{g}.\end{split}$
Passing to scale-invariant norms, we have
[…]
From the first term, the most dangerous case is when $i_{4}$ falls on $\hat{\chi}$, so we only give details for that. We distinguish two cases:
  • If in the term $\nabla^{i_{1}}\psi_{g}^{i_{2}}$ there exists some $\psi_{g}$ whose derivative is of order >N+3>N+3, we bound that term in $L^{2}_{(sc)}$ and the rest of the terms in $L^{\infty}_{(sc)}$. Notice, crucially, we can bound $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{1}{a^{\frac{1}{2}}}$ for small $i$ in light of embedding theorems. As a consequence, we have the bound
    $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}^{2}}{\lvert u\rvert}\lesssim\frac{1}{a|u|}.$
  • Otherwise, in the expression $\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}\hat{\chi}$, at most one index $i_{3},i_{4}$ is greater than $N+1$ (in which case, we cannot bound that term in $L^{\infty}_{(sc)}$). Say without loss of generality, that $i_{3}>N+1$. We bound $(a^{\frac{1}{2}})^{i_{3}-1}\nabla^{i_{3}}\hat{\chi}$ in $L^{2}_{(sc)}$ above by $\mathcal{R}[\alpha]$ and the rest of the terms in $L^{\infty}_{(sc)}$ above by $1$, whence
    $\int_{0}^{\underline{u}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\hat{\chi}\nabla^{i_{4}}(\hat{\chi},\mbox{tr}\chi)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\left(\mathcal{R}[\alpha]+1\right)\cdot\frac{a^{\frac{1}{2}}}{|u|}.$
The second term is handled as in the previous propositions.
Proposition 4.15.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), the following estimates hold:
$\frac{a}{\lvert u\rvert}\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1,\hskip 5.69054pt\frac{a}{|u|^{2}}\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1.$
Proof.
Notice that $\mbox{tr}\underline{\chi}$ satisfies the following structure equation:
$\nabla_{3}\mbox{tr}\underline{\chi}+\frac{1}{2}(\mbox{tr}\underline{\chi})^{2}=-\lvert\underline{\hat{\chi}}\rvert^{2}-2\underline{\omega}\mbox{tr}\underline{\chi}.$
Commuting with $i$ angular derivatives using 8, we get
(164) $\begin{split}\nabla_{3}\nabla^{i}\mbox{tr}\underline{\chi}+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\mbox{tr}\underline{\chi}=&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\underline{\hat{\chi}}\nabla^{i_{4}}\underline{\hat{\chi}}\\ +&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\underline{\omega}\nabla^{i_{4}}\mbox{tr}\underline{\chi}\\ +&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\mbox{tr}\underline{\chi}\\ +&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\mbox{tr}\underline{\chi}\\ +&\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\mbox{tr}\underline{\chi}:=G_{i}.\end{split}$
Passing to scale invariant norms and using the weighted transport inequality from Proposition 4.5, we obtain
(165) $\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\hskip 1.42262pt$ $\frac{a}{\lvert u_{\infty}\rvert^{2}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}})}$
(166) $+\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}G_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}.$
We focus on $0\leq i\leq N+4$. For the first term in $G_{i}$, there holds
(167) $\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\underline{\hat{\chi}}\nabla^{i_{4}}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\cdot\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim 1.$
The second and third terms are handled in the same way. For the fourth term, there holds
(168) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\cdot\frac{\lvert u^{\prime}\rvert^{2}}{a}\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{2}}}\cdot\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert}\lesssim 1.\end{split}$
For the fifth term, we bound
(169) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\cdot\frac{\lvert u^{\prime}\rvert^{4}}{a^{2}}\cdot\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{\Gamma^{3}}{\lvert u\rvert}\lesssim 1.\end{split}$
For the sixth term, we can bound
(170) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a^{2}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ =&\int_{u_{\infty}}^{u}\frac{a^{\frac{5}{2}}}{\lvert u^{\prime}\rvert^{4}}\lVert{a^{\frac{i-1}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}+1=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a^{\frac{5}{2}}}{\lvert u^{\prime}\rvert^{4}}\cdot\frac{\lvert u^{\prime}\rvert^{4}}{a^{2}}\cdot\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\hskip 1.42262pt\Gamma^{3}}{\lvert u\rvert}\lesssim 1.\end{split}$
Crucially, this implies that
(171) $\frac{a}{\lvert u\rvert^{2}}\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{I}+1,$
by the Sobolev embedding. This will prove useful in the estimates for $\widetilde{\mbox{tr}\underline{\chi}}$. ∎
We conclude this section with the corresponding estimate on $\underline{\eta}$ and its angular derivatives.
Proposition 4.16.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), there holds
$\sum_{0\leq i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}\bigg(\mathcal{R}[\beta]+\underline{\mathcal{R}}[\underline{\beta}]\bigg)+1.$
Proof.
The schematic equation for $\underline{\eta}$ is as follows:
(172) $\nabla_{3}\underline{\eta}+\frac{1}{2}\mbox{tr}\underline{\chi}\hskip 1.42262pt\underline{\eta}=\underline{\beta}+\mbox{tr}\underline{\chi}\eta+\psi_{g}\underline{\hat{\chi}}$.
$\nabla_{3}\nabla^{i}\underline{\eta}+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\underline{\eta}=$ $\nabla^{i}\underline{\beta}+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\underline{\beta}$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\underline{\eta}$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\underline{\eta}$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\underline{\eta}:=H_{i}.$
Calculating in scale-invariant norms, we arrive at
(173) $\frac{1}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{1}{\lvert u_{\infty}\rvert}\lVert(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}})}+\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}H_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}.$
We restrict attention to $0\leq i\leq N+4$. For the first term, there holds
(174) $\frac{1}{\lvert u_{\infty}\rvert}\lVert(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}})}\lesssim\frac{\mathcal{I}}{\lvert u_{\infty}\rvert}\lesssim\frac{\mathcal{I}}{\lvert u\rvert}$.
There holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\beta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
$\lesssim$ $\left(\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\beta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\right)^{\frac{1}{2}}\left(\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{4}}\hskip 1.42262pt\text{d}u^{\prime}\right)^{\frac{1}{2}}$
(175) $\lesssim$ $\underline{\mathcal{R}}[\underline{\beta}]\cdot\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{3}{2}}}\lesssim\frac{\underline{\mathcal{R}}[\underline{\beta}]}{\lvert u\rvert}$.
For the next term, there holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\underline{\beta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(176) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\cdot\frac{a^{\frac{1}{2}}\hskip 1.42262pt\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{1}{\lvert u\rvert}$.
For the third term, there holds
(177) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{\mathcal{R}[\beta]+1}{\lvert u\rvert}$.
This is done by further taking into account that the schematic product appearing is actually $\psi_{g}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})=\underline{\hat{\chi}}\cdot\psi_{g}+\mbox{tr}\underline{\chi}\eta$. As such, we use the improvement obtained in Proposition 4.15 for the $\mbox{tr}\underline{\chi}$-term as well as the improvement obtained in Proposition 4.13 for $\eta$. The term $\underline{\hat{\chi}}$ is less anomalous than $\mbox{tr}\underline{\chi}$ and hence the above bound is easier to obtain. Continuing the estimates, it holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(178) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\cdot\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{2}}}\cdot\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{1}{\lvert u\rvert}$.
For the sixth term, there holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(179) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\cdot\frac{\lvert u^{\prime}\rvert^{2}}{a}\cdot\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{1}{\lvert u\rvert}$.
For the seventh term, there holds
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
(180) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a^{\frac{3}{2}}}{\lvert u^{\prime}\rvert^{3}}\cdot\frac{\lvert u^{\prime}\rvert^{2}}{a}\cdot\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\hskip 1.42262pt\Gamma^{3}}{\lvert u\rvert^{2}}\lesssim\frac{1}{\lvert u\rvert}$.
Putting everything together, we have
$\frac{1}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\eta}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{a^{-\frac{1}{2}}(\mathcal{R}[\beta]+\underline{\mathcal{R}}[\underline{\beta}])+1}{\lvert u\rvert}$,
whence the result follows.
This concludes the estimates on Ricci coefficients.
5. $L^{2}(S_{u,\underline{u}})$ Estimates for the Weyl curvature components
Proposition 5.1.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), there holds
$\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{1}{a^{\frac{1}{2}}}.$
Proof.
Recall the Bianchi equation for $\alpha$:
$\nabla_{3}\alpha+\frac{1}{2}\mbox{tr}\underline{\chi}\alpha=$ $\nabla\hat{\otimes}\beta+4\underline{\omega}\alpha-3\left(\hat{\chi}\rho+\prescript{*}{}{\hat{\chi}}\sigma\right)+(\zeta+4\eta)\hat{\otimes}\beta.$
Schematically, the above rewrites as
$\nabla_{3}\alpha+\frac{1}{2}\mbox{tr}\underline{\chi}\alpha=$ $\nabla\beta+\psi_{g}\alpha+\hat{\chi}(\rho,\sigma)+\psi_{g}\beta.$
Commuting with $i$ angular derivatives using 8, we arrive at
$\nabla_{3}\nabla^{i}\alpha+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\alpha$
$=$ $\nabla^{i+1}\beta+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\beta+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\alpha$
$+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\rho,\sigma,\beta)$
$+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\alpha$
$+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\alpha$
(181) $+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\alpha:=T_{i}^{1}+\dots+T_{i}^{7}.$
Passing to scale-invariant norms and using the weighted transport equality from Proposition 4.5, we can estimate as follows.
(182) $\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}$ $\lesssim\lVert(a^{\frac{1}{2}}\nabla)^{i}\alpha\rVert_{L^{2}_{(sc)}(S_{u_{\infty},u})}+\sum_{1\leq j\leq 12}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}.$
There holds:
(183) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\lVert{a^{\frac{i}{2}}T_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}+\lVert{a^{\frac{i}{2}}T_{i}^{5}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\right)\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}\Gamma^{2}}{\lvert u^{\prime}\rvert^{3}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma^{2}}{|u|^{2}}.$
Moreover, there holds
(184) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{6}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma^{3}}{|u|^{2}}\lesssim\frac{a^{\frac{1}{2}}}{|u|}.$
For the eighth term, we can bound
(185) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}T_{i}^{7}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\int_{u_{\infty}}^{u}\frac{a\Gamma^{2}}{\lvert u^{\prime}\rvert^{3}}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\lesssim\frac{a^{\frac{3}{2}}\Gamma^{2}}{|u|^{3}}.$
The result follows.
Proposition 5.2.
For $\Psi_{u}=(\beta,\underline{\beta},\rho,\sigma)$ defined as in Section 3.6, there holds
$\sum_{0\leq i\leq N+3}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta,\underline{\beta})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}(\mathcal{R}[\alpha]+1),$
[…]
Proof.
Each of the $\Psi_{u}$ satisfies the following schematic equation:
(186) $\nabla_{4}\Psi_{u}=\nabla\left(\Psi_{u},\alpha\right)+(\psi,\underline{\hat{\chi}})\left(\Psi_{u},\alpha\right).$
Commuting with $i$ angular derivatives using 3.1, we obtain
$\nabla_{4}\nabla^{i}\Psi_{u}=$ $\nabla^{i+1}(\Psi_{u},\alpha)+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}(\Psi_{u},\alpha)$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi,\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{u},\alpha)$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}.$
Passing to scale-invariant norms and estimating, we have
[…]
$+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
We restrict attention to $0\leq i\leq N+3$. For the first term, we have,
(187) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{1/2}=\frac{1}{a}\mathcal{R}[\alpha],$
by using Hölder’s inequality. For the second term, since the $\Psi_{u}$ are regular with respect to scaling, we conclude that
(188) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{1}{a^{\frac{1}{2}}}\mathcal{R}+\frac{1}{a}\mathcal{R}[\alpha],$
by the bootstrap assumptions (118). For the third term, we have $i_{3}+1\leq i$, hence everything can be closed using the $\Gamma$ total norm. We have
(189) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\lesssim\frac{\Gamma^{2}}{\lvert u\rvert}\lesssim a^{-\frac{1}{2}}.$
The fourth term can be estimated by
(190) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{\lvert u\rvert}{a}\frac{\Gamma^{2}}{\lvert u\rvert}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim a^{-\frac{1}{2}}.$
Moreover, for $\Gamma(\mbox{tr}\chi)$, we make use of Proposition 4.15 to bound $\Gamma(\mbox{tr}\underline{\chi})\lesssim 1$. For the seventh term, there holds
(191) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{\Gamma^{2}}{\lvert u\rvert}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim a^{-\frac{1}{2}}.$
For the eighth term, there holds
(192) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{a\Gamma^{3}}{\lvert u\rvert^{2}}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim a^{-\frac{1}{2}}.$
Putting all of the above together, the result follows. ∎
Proposition 5.3.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), there holds
$\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}.$
Proof.
Recall the Bianchi equation for $\underline{\alpha}$:
(193) $\begin{split}\nabla_{4}\underline{\alpha}+\frac{1}{2}\mbox{tr}\chi\hskip 1.42262pt\underline{\alpha}=&-\nabla\hat{\otimes}\underline{\beta}+4\omega\hskip 1.42262pt\underline{\alpha}-3\left(\underline{\hat{\chi}}\rho-\prescript{*}{}{\underline{\hat{\chi}}}\sigma\right)+\left(\zeta-4\underline{\eta}\right)\hat{\otimes}\underline{\beta}.\end{split}$
Schematically, the above rewrites as
(194) $\nabla_{4}\underline{\alpha}=\nabla\Psi_{u}+\psi_{g}(\underline{\alpha},\underline{\beta}^{R})+\underline{\hat{\chi}}(\rho,\sigma).$
Commuting with $i$ angular derivatives, we arrive at
[…]
$+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{\underline{u}})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
$+$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
In the above expression, all terms can be bounded above by $a^{-\frac{1}{2}}$, in the same way as in the preceding Proposition. The result follows.
5.1. Energy Estimates for the Weyl/Riemann Curvature components
For $(\Psi_{1},\Psi_{2})\in\begin{Bmatrix}(\alpha,\beta^{R}),(\beta,(\rho,\sigma)),((\rho,\sigma),\underline{\beta}),(\underline{\beta}^{R},\underline{\alpha})\end{Bmatrix}$ the energy estimates are carried out in Bianchi pairs, via the aid of the following proposition:
Proposition 5.4.
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), for a Bianchi pair $(\Psi_{1},\Psi_{2})$ satisfying
$\nabla_{3}\nabla^{i}\Psi_{1}+\left(\frac{i+1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\nabla^{i}\Psi_{1}-D\nabla^{i}\Psi_{2}=P_{i},$
$\nabla_{4}\nabla^{i}\Psi_{2}-\prescript{*}{}{D}\nabla^{i}\Psi_{1}=Q_{i},$
the following holds true:
(195) $\begin{split}&\int_{H_{u}^{(0,\underline{u})}}\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{\underline{H}_{\underline{u}}^{(u_{\infty},u)}}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{H_{u_{\infty}}^{(0,\underline{u})}}\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{\underline{H}_{0}^{(u_{\infty},u)}}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},0})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\\ &+\iint_{D_{u,\underline{u}}}\frac{a}{\lvert u^{\prime}\rvert}\lVert{\nabla^{i}\Psi_{1}\cdot P_{i}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\iint_{D_{u,\underline{u}}}\frac{a}{\lvert u^{\prime}\rvert}\lVert{\nabla^{i}\Psi_{2}\cdot Q_{i}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
Before embarking on the energy estimates, we provide a final helpful proposition, which can be found for example in [25].
Proposition 5.5 (Two–parameter Grönwall inequality on a rectangle).
Let $x_{0},y_{0}>0$ and set
$U:=\{(x,y)\in\mathbb{R}^{2}:\;0\leq x\leq x_{0},\;0\leq y\leq y_{0}\}.$
Let $f,g:U\to[0,\infty)$ be measurable functions such that for each fixed $y$ the map $x\mapsto f(x,y)$ is integrable on $0,x_{0}]$ and for each fixed $x$ the map $y\mapsto g(x,y)$ is integrable on $0,y_{0}]$. Assume there exist constants $J,c_{1},c_{2}\geq 0$ and $C_{0}\geq 1$ such that for all $(x,y)\in U$ one has
(196) $f(x,y)+g(x,y)\leq C_{0}\!\left(J+c_{1}\int_{0}^{x}f(x^{\prime},y)\,dx^{\prime}+c_{2}\int_{0}^{y}g(x,y^{\prime})\,dy^{\prime}\right).$
Then for every $(x,y)\in U$ there holds the exponential bound
(197) $f(x,y)+g(x,y)\leq C\,J\,\exp\!\big(C(c_{1}x+c_{2}y)\big),$
where $C\geq 1$ depends only on $C_{0}$ (in particular, it is independent of $x_{0},y_{0},J,c_{1},c_{2}$ and of the functions $f,g$).
With this Proposition as the main tool, we begin with $(\alpha,\beta)$.
Proposition 5.6 (Top–order curvature flux estimate for $(\alpha,\beta)$).
Let $(\mathcal{M},g)$ be a smooth vacuum spacetime endowed with a regular double–null foliation $(u,\underline{u})$ and associated null frame $(e_{3},e_{4},e_{A})$. Assume the hypotheses of Theorem 2.2 and the bootstrap bounds (118) hold on the spacetime region under consideration. Let $N\geq 0$ denote the maximal derivative level in the energy hierarchy.
Then for every integer $0\leq i\leq N+4$, the following renormalized curvature flux estimate holds for the extreme null Weyl component $\alpha$ and the adjacent component $\beta$:
(198) $a^{\frac{1}{2}}\,\lVert{\nabla^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}\;+\;a^{\frac{1}{2}}\,\lVert{\nabla^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\;\lesssim\;a^{\frac{1}{2}}\,\lVert{\nabla^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}\;+\;a^{\frac{1}{2}}\,\lVert{\nabla^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}\;+\;1.$
The implicit constant depends only on $N$, and is uniform in $a$, $u$, and $\underline{u}$.
In particular, the scale–renormalized outgoing flux of $\alpha$ along $H_{u}$ together with the incoming flux of $\beta$ along $\underline{H}_{\underline{u}}$ are controlled solely by the corresponding initial null fluxes on $H_{u_{\infty}}$ and $\underline{H}_{0}$, up to a universal error bound generated by lower–order curvature components and Ricci coefficients.
Proof.
We recall the (schematic) Bianchi equations for $\nabla^{i}\alpha,\nabla^{i}\beta$:
$\nabla_{3}\nabla^{i}\alpha+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\alpha-D\nabla^{i}\beta$
$=$ $\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\beta^{R}+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\alpha$
$+$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\rho,\sigma,\beta)$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\alpha$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\alpha$
(199) $+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\alpha:=P_{i}^{1}+\dots+P_{i}^{6}.$
Similarly, we have
$\nabla_{4}\nabla^{i}\beta-\prescript{*}{}{D}\nabla^{i}\alpha=$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\beta,\alpha)$
(200) $+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\beta$
$+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\beta:=Q_{i}^{1}+\dots+Q_{i}^{3}.$
Applying Proposition 5.4, we have
(201) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\\ \lesssim&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}\\ &+\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
By Hölder’s inequality, one has
(202) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \leq&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\sum_{j=1}^{10}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\cdot\sup_{u^{\prime}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\alpha\rVert_{L^{2}_{(sc)}(H_{u^{\prime}}^{(0,\underline{u})})},\end{split}$
Let us focus on the sum in the above line. For the first three terms, there holds
$\sum_{j=1}^{3}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\frac{a^{\frac{1}{2}}\Gamma\cdot R}{\lvert u^{\prime}\rvert}$.
For the fourth and fifth terms, there holds
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}+\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{5}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\frac{a^{\frac{1}{2}}\Gamma\cdot M}{\lvert u^{\prime}\rvert}$.
For the sixth term, there holds
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{7}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\frac{a^{\frac{1}{2}}\Gamma\cdot M}{\lvert u^{\prime}\rvert}$.
For the eighth term, there holds
Putting everything together, there holds
(203) $\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{a^{\frac{1}{2}}}{|u|}\bigg(\Gamma^{3}+\Gamma^{2}R+\Gamma R+1\bigg).$
Similarly, for the analogous term involving $\beta$, there holds
(204) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \leq&\sum_{j=1}^{7}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\cdot\sup_{u^{\prime}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\beta\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}.\end{split}$
We estimate term by term. For the first term, there holds
(205) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma(R+\Gamma)}{\lvert u\rvert}$
For the second term, there holds
(206) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma(M+\Gamma)}{\lvert u\rvert}$
For the third and fourth terms, there holds
(207) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{3}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}+\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a^{\frac{1}{2}}\Gamma(M+\Gamma)}{\lvert u\rvert}\end{split}$
For the fifth term there holds
(208) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{5}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma(R+\Gamma)}{\lvert u\rvert}$
For the sixth term, we have
(209) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{6}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma^{2}(R+\Gamma)}{\lvert u\rvert^{2}}$
For the seventh term, we have
(210) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{7}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{\Gamma^{2}(R+\Gamma)}{\lvert u\rvert^{3}}$
Putting everything together, we have
(211) $\sum_{j=1}^{7}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}}{|u|}\bigg((\Gamma(R+M+\Gamma)+1\bigg).$
Combining (203) and (211), we have
(212) $\begin{split}&a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\\ \lesssim&a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}\\ &+a^{\frac{1}{2}}\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+a^{\frac{1}{2}}\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+\frac{1}{a^{\frac{1}{2}}}.\end{split}$
The claim follows. ∎
We now move on to energy estimates for the remaining pairs $(\beta,(\rho,\sigma)),((\rho,\sigma),\underline{\beta})$ and $(\underline{\beta},\underline{\alpha})$.
Proposition 5.7.
Let $(\mathcal{M},g)$ be a vacuum spacetime equipped with a regular double–null foliation $(u,\underline{u})$ and associated null frame $(e_{3},e_{4},e_{A})$, and assume the bootstrap assumptions (118) and the curvature and Ricci coefficient bounds stated in Theorem 2.2. Let $N\geq 0$ be the top derivative order in the energy hierarchy.
Consider any adjacent pair of Weyl curvature components
$(\Psi_{1},\Psi_{2})\in\Big\{a^{\frac{1}{2}}(\beta,(\rho,\sigma)),\;a^{\frac{1}{2}}((\rho,\sigma),\underline{\beta}),\;a^{\frac{1}{2}}(\underline{\beta},\underline{\alpha})\Big\},$
where the factor $a^{1/2}$ denotes the renormalized scaling used in the curvature energy norms. Then for every integer $0\leq i\leq N+4$ there holds the flux bound
(213) $\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\;\lesssim\;\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+1.$
The implicit constant depends only on $N$ and is independent of $a$, $u$, and $\underline{u}$.
Proof.
The schematic equations for $\Psi_{1},\Psi_{2}$ are:
$\nabla_{3}\Psi_{1}+\left(\frac{1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\Psi_{1}-D\Psi_{2}=$ $(\psi,\hat{\chi})\Psi,$
(214) $\begin{split}\nabla_{4}\Psi_{2}-\prescript{*}{}{D}\Psi_{1}=(\psi,\underline{\hat{\chi}})(\Psi_{u},\alpha).\end{split}$
Commuting with $i$ angular derivatives, for $\Psi_{1}$, we have:
(215) $\begin{split}&\nabla_{3}\nabla^{i}\Psi_{1}+\left(\frac{i+1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\nabla^{i}\Psi_{1}-D\nabla^{i}\Psi_{2}\\ =&\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}\Psi\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\Psi_{1}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\Psi_{1}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\Psi_{1}:=P_{i}.\end{split}$
Analogously, for $\Psi_{2}$, we have
(216) $\begin{split}&\nabla_{4}\nabla^{i}\Psi_{2}-\prescript{*}{}{D}\nabla^{i}\Psi_{1}\\ =&\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{1}+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi,\alpha)\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{2}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{2}:=Q_{i}.\end{split}$
Making use of Proposition 5.4 once again, we arrive at
(217) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}^{2}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}^{2}\\ \lesssim&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}^{2}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}^{2}\\ &+\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ &+\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
For the first spacetime integral in the above, we estimate
(218) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\cdot\left(\int_{0}^{\underline{u}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}.\end{split}$
For the first term:
(219) $\begin{split}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime},\end{split}$
if $i_{3}+1\geq N+3$, we estimate
(220) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\sup_{0\leq\underline{u}^{\prime}\leq\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{3}}\lVert{(a^{\frac{1}{2}}\nabla)^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\lVert{(a^{\frac{1}{2}})^{i_{1}+i_{2}}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}}\rVert_{L^{\infty}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\end{split}$
and we can estimate $\lVert{(a^{\frac{1}{2}})^{i_{1}+i_{2}}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}}\rVert_{L^{\infty}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}$ by $\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{2}}}$ using the bootstrap assumption (118), to obtain
(221) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\sup_{0\leq\underline{u}^{\prime}\leq\underline{u}}\int_{u_{\infty}}^{u}\frac{a^{\frac{1}{2}}}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{R}{\lvert u\rvert^{\frac{1}{2}}}\lesssim 1.\end{split}$
If, however, $i_{3}+1\leq N+2$, we can control the corresponding $L^{2}_{(sc)}(S)$ norm just by the bootstrap assumption (118) to get the bound
(222) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma^{2}}{\lvert u\rvert^{2}}\lesssim 1.\end{split}$
For the rest of the terms, we estimate using the same philosophy as appropriate. There holds
(223) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}\Psi}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a^{\frac{1}{2}}\Gamma(R+\Gamma)}{\lvert u\rvert}\lesssim 1.\end{split}$
Here in particular, when $i_{4}\geq N+3$, we treat the cases $\Psi=\Psi_{u}$ and $\Psi=\Psi_{\underline{u}}$ separately. For the sixth and seventh terms, we can bound them by one as in previous calculations. For the eighth term, using the fact that $i-2\leq N+2$ and the improvements from Proposition 5.2 and Proposition 5.6, we arrive at
(224) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a}{\lvert u\rvert}\Gamma[\mbox{tr}\underline{\chi}]\Gamma[\alpha]^{2}\lesssim\frac{a}{\lvert u\rvert}\Gamma[\eta,\underline{\eta}]\Gamma[\mbox{tr}\underline{\chi}]\Gamma[\Psi_{1}]\lesssim\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta^{R},\underline{\beta}^{R})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\cdot 1\cdot(a^{-\frac{1}{2}}\mathcal{R}[\alpha]+1)\\ \lesssim&(a^{-\frac{1}{2}}\mathcal{R}[\alpha]+1)^{2}\lesssim 1,\end{split}$
where in the last line we made use of Proposition 5.6. This completes the estimates for the first spacetime integral in (217). For the second and last one, a double application of Hölder’s inequality yields
(225) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{0}^{\underline{u}}\left(\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}Q_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\right)^{\frac{1}{2}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}\\ \lesssim&\left(\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}Q_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\left(\int_{0}^{\underline{u}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}\rVert^{2}_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\\ \lesssim&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}Q_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\frac{1}{4}\int_{0}^{\underline{u}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}\rVert^{2}_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\end{split}$
Define $B:=\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}Q_{i}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$. We can then estimate $B$ as follows:
(226) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\psi_{g}\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\nabla\psi_{g}\nabla^{i-1}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\psi_{g}\hskip 1.42262pt\psi_{g}\nabla^{i-1}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ +&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\sum_{\begin{subarray}{c}i_{1}+i_{2}+i_{3}=i-1\\ i_{3}<i-2\end{subarray}}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&1,\end{split}$
where in the first three integrals we estimate $\psi_{g},\nabla\psi_{g}$ in $L^{\infty}_{(sc)}(S_{u,\underline{u}})$ and $\nabla^{i}\Psi_{1}$ in the hypersurface norm $L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})$ and in the last integral, since $i-2\leq N+2$, we can estimate $\nabla^{i}\Psi_{1}$ in $L^{2}_{(sc)}(S_{u,\underline{u}})$ using the bootstrap assumption on the norm $\Gamma$. For the second term, we similarly have
$\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\Psi,\alpha)}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
(227) $\lesssim$ $\left(\mathcal{R}[\alpha]+1\right)^{2}\lesssim 1,$
where we have used the improvements on $\hat{\chi}$ from Proposition 4.11 and the energy estimate from the Proposition 5.6. The rest of the terms can also be bounded above by $1$, using the same approach. We finally arrive at an estimate of the form
(228) $\begin{split}&\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}Q_{i}\cdot(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\\ \lesssim&1+\frac{1}{4}\int_{0}^{\underline{u}}\lVert(a^{\frac{1}{2}}\nabla)^{i}\Psi_{2}\rVert^{2}_{L^{2}_{(sc)}(\underline{H}_{\underline{u}^{\prime}}^{(u_{\infty},u)})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
From here, collecting all the terms we arrive at the desired result. This concludes the characteristic development. ∎
6. Cauchy Development and Completion of the argument
In this section, the remaining two major tasks are executed. First, we prove a local Cauchy development of the initial data on $\mathcal{M}_{-a}:=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{ext}$. Most importantly, the task is prove that the interior estimates on $\mathcal{M}_{1}$ can be propagated up to $O(1)$ time without substantial distortion. In particular, the interior thickness should not be reduced too much while measured in terms of the $H$-radius. This is due to the choice of initial data in the interior. The remarkable point to note here is that this interior data is chosen to be consistent with the induced Cauchy data on $\mathcal{M}_{2}$ by the characteristic development $D_{a,1}$. This choice causes the minimal distortion of the induced Riemannian geometry on $\mathcal{M}_{1}$ under evolution for $O(1)$ time. In effect, the choice of initial data for the Characteristic development is vital in the full analysis. Nevertheless, once one understands the Characteristic development completely, one can in a sense forget about the characteristic problem and start prescribing data on the Cauchy slice $\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{ext}$ altogether with appropriate asymptotically flat solution for the $\mathcal{M}_{ext}$ end.
6.1. Geometric setup and decomposition of the initial slice
Let (,g^)($\mathcal{M}$$\widehat${g}) be the Lorentzian manifold constructed in previous sections. Let $u,\underline{u}$ be double-null coordinates associated to the canonical foliation, and let
$T:=\tfrac{1}{2}(e_{4}+e_{3})$
denote the future-directed unit timelike vector field orthogonal (in the induced sense) to the leaves $S_{u,\underline{u}}\subset\mathcal{M}_{t=u+\underline{u}}$.
The interior region $\mathcal{M}_{1}\subset\mathcal{M}$ admits a foliation by topological 22–spheres outside a compact subset. Our analysis requires only the boundary geometry $\partial\mathcal{M}_{1}$ and its associated Yau radius.
Fix times
$t_{1}=-a,\qquad t_{2}=-a-1/a+\epsilon,$
where $\epsilon$ will be chosen to be $\frac{3}{4}$. Although we do not require the double-null foliation to describe the interior evolution (and not well defined for the interior), it is convenient near $\partial\mathcal{M}_{1}$.
On the initial slice $\mathcal{M}_{t=-a}$ we introduce the decomposition
(229) $\mathcal{M}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
where
$\mathcal{M}_{1}:=\mathcal{M}_{1},\qquad\mathcal{M}_{2}:=\mathcal{M}_{t=-a}\cap D_{a,1},\qquad\mathcal{M}_{3}:=\mathcal{M}_{t=-a}\setminus(\mathcal{M}_{1}\cup\mathcal{M}_{2})=\mathcal{M}_{ext}.$
6.2. Specification of the Cauchy data
The Cauchy data $(g,k)$ on $\mathcal{M}_{t=-a}$, consisting of the induced Riemannian metric $g$ and second fundamental form $k$, are specified as follows:
  1. (a)
    Interior data.
  2. (b)
    Matching data in the characteristic region. On $\mathcal{M}_{2}=\mathcal{M}_{t=-a}\cap D_{a,1}$, the data are induced from the characteristic development on the double-null slab $D_{a,1}$. Well-posedness of the characteristic problem on this region was established in the previous section.
  3. (c)
    Asymptotically Kerr exterior. On $\mathcal{M}_{3}$, we prescribe smooth data $(g,k)$ asymptotic to a Kerr slice with prescribed mass $m=O(a^{1/2})$ and angular momentum $J=a$. This ensures asymptotic flatness.
A gluing construction in the spirit of [22, 28, 11] yields a smooth, asymptotically flat vacuum Cauchy pair $(g,k)$ on $\mathcal{M}_{t=-a}$.
6.3. Setup, decomposition, constraints, and gauges
Let (,g^)($\mathcal{M}$$\widehat${g}) be the semi-globally constructed Lorentzian manifold equipped with a canonical double–null foliation $(u,\underline{u})$ and Cauchy time $t=u+\underline{u}$. Fix $a\gg 1$. On the initial slice $\mathcal{M}_{t=-a}$ set
$\mathcal{M}_{t=-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},$
with $\mathcal{M}_{1}=\mathcal{M}_{1}$, $\mathcal{M}_{2}=\mathcal{M}_{t=-a}\cap\mathcal{D}_{-a,1}$ (the double–null slab), and $\mathcal{M}_{3}=\mathcal{M}_{t=-a}\setminus(\mathcal{M}_{1}\cup\mathcal{M}_{2})$. We must solve on $\mathcal{M}_{t=-a}$ the vacuum constraint system
(230) $\text{Scal}(g)-|k|_{g}^{2}+(\mbox{tr}_{g}k)^{2}=0,\qquad\mbox{div }_{g}k-d(\mbox{tr}_{g}k)=0,$
subject to (i) prescribed boundary/interface conditions across $\partial\mathcal{M}_{i}$, (ii) MOTS–exclusion on t=at=-a, and (iii) asymptotically Kerr behavior on $\mathcal{M}_{3}$.
Throughout we can use weighted $W^{2,p}_{-\delta}\times W^{1,p}_{-1-\delta}$ scales with $p>3$ and $0<\delta<1$.
On $\mathcal{M}_{2}=\mathcal{M}_{t=-a}\cap D_{a,1}$, take the data induced from the double–null development on the slab $\mathcal{D}_{-a,\epsilon}$, in the normalized frame
$e_{4}=\Omega^{-1}\partial_{\underline{u}},\qquad e_{3}=\Omega^{-1}(\partial_{u}+b^{A}\partial_{\theta^{A}}),\qquad g|_{S_{u,\underline{u}}}=\gamma_{AB}d\theta^{A}d\theta^{B}.$
Denote the null second fundamental forms by $\chi=\widehat{\chi}+\tfrac{1}{2}(\mbox{tr}\chi)\gamma$ and $\underline{\chi}=\widehat{\underline{\chi}}+\tfrac{1}{2}(\mbox{tr}\underline{\chi})\gamma$. We import the full set of Ricci and curvature components on $\mathcal{M}_{2}$, including the incoming shear $\widehat{\underline{\chi}}$, with the scale–critical bounds dictated by the construction of the characteristic initial value problem (on $H_{0}$ and $H_{u_{\infty}}$):
$\|\widehat{\chi}\|_{L^{\infty}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}|u|^{-1},\qquad\|\widehat{\underline{\chi}}\|_{L^{\infty}(S_{u,\underline{u}})}\lesssim a^{\frac{1}{2}}|u|^{-2},$
$|\mbox{tr}\chi|\lesssim|u|^{-1},\quad|\mbox{tr}\underline{\chi}|\lesssim|u|^{-1},\quad\|\eta\|_{L^{\infty}}+\|\underline{\eta}\|_{L^{\infty}}\lesssim a^{\frac{1}{2}}|u|^{-2},\quad|\omega|+|\underline{\omega}|\lesssim a^{\frac{1}{2}}|u|^{-2}.$
In particular, on the interface sphere $S_{-a,\underline{u}}$ the combination
$H-|\kappa|=\tfrac{1}{2}(\mbox{tr}\chi-\mbox{tr}\underline{\chi})-\tfrac{1}{2}\bigl|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\bigr|$
is explicitly computable from these inputs, ensuring (via the chosen profile of incoming shear $\underline{\hat{\chi}}$ and the initial incoming expansion $\mbox{tr}\underline{\chi}$) that $c:=\min(H-|\kappa|)$ on the boundary of the interior piece at t=at=-a is below the Yau threshold, hence no MOTS appear there. The first and second fundamental forms of the slice and of the $S_{u,\underline{u}}$ leaves are matched in $C^{\infty}$ across $\partial\mathcal{M}_{1}\cap\partial\mathcal{M}_{2}$ by a standard partition–of–unity interpolation inside a thin collar, after which we re-solve the constraints by compactly supported corrections in the next section.
In this section, we complete the proof of Theorem 2.2 by establishing the formation of an MOTS strictly to the future of the Cauchy hypersurface
$\mathcal{M}_{t=-a}\subset\mathcal{M}.$
We work in the spacetime harmonic gauge and solve the vacuum Einstein equations on the slab
$a,-a-1/a+\epsilon]\times\mathcal{M},\qquad 0<\epsilon\ll 1,$
with initial data assembled by a gluing procedure. We then propagate the quasi-local boundary geometry forward in time, ultimately showing that the Schoen–Yau curvature radius condition is violated in the future domain, forcing the existence of an MOTS. In this section, we provide a sketch of the construction of the data.
6.4. Interior data on $\mathcal{M}_{1}$: strict Yau–radius subcriticality with boundary control
This is achieved by a refined geometric decomposition and constraint–compatible gluing of the interior Cauchy data. We decompose
$\mathcal{M}_{1}=\widetilde{\mathcal{M}}_{1}\cup\Big(\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}\Big),$
where $\widetilde{\mathcal{M}}_{1}\Subset\mathcal{M}_{1}$ is a compact subdomain with smooth boundary and the complement $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ is a collar (transition) region of uniformly bounded thickness $O(1)$ measured with respect to the background Euclidean metric. The decomposition is chosen so that $\partial\widetilde{\mathcal{M}}_{1}$ lies a fixed positive distance away from the outer interface $\partial\mathcal{M}_{1}$ and from the gluing interface with the characteristic region $D_{a,1}$.
On $\widetilde{\mathcal{M}}_{1}$ we prescribe vacuum constraint data $(g,k)$ satisfying the Einstein constraint equations and the quantitative smallness bounds
(231) $\|k\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\leq C\,a^{-3/2},~\|\partial g\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}$ $\leq Ca^{-1},$
(232) $\|k\|_{H^{s-1}_{\mathrm{ul}}(\widetilde{\mathcal{M}}_{1})}\leq C\,a^{-3/2},~\|\partial g\|_{H^{s-1}_{\mathrm{ul}}(\widetilde{\mathcal{M}}_{1})}$ $\leq Ca^{-1},$
for some fixed $s\gg 1$, where $H^{s-1}_{\mathrm{ul}}$ denotes the uniformly–local Sobolev space (defined via unit–scale coordinate balls). The same bounds hold for the lapse $N$ and shift $X$,
(233) $\|N-1\|_{L^{\infty}}+\|X\|_{L^{\infty}}+\|N-1\|_{H^{s}_{\mathrm{ul}}}+\|X\|_{H^{s}_{\mathrm{ul}}}\leq Ca^{-3/2},$
so that the data are a quantitatively small perturbation of Euclidean data (in homogeneous Sobolev norm) on each unit ball, with constants independent of the total $H$-radius $O(a)$ of $\mathcal{M}_{1}$. In particular, the local constraint norms remain $O(a^{-3/2})$ while global Sobolev norms may be $O(1)$ due to volume growth; all arguments below are formulated in uniformly–local norms to avoid this scaling loss.
The collar region $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ is used as a gluing zone in which $(g,k)$ is constructed by a constraint–preserving interpolation between the interior data on $\widetilde{\mathcal{M}}_{1}$ and the induced Cauchy data coming from the characteristic development $D_{a,1}$ on $\mathcal{M}_{2}$. Using a Corvino–Schoen type localized deformation together with a partition of unity and solvability of the linearized constraint operator with compact support, one obtains smooth vacuum data on $\mathcal{M}_{1}$ that agree exactly with the interior data on $\widetilde{\mathcal{M}}_{1}$ and with the characteristic–induced data near $\partial\mathcal{M}_{1}$, with all corrections supported strictly inside the collar (a relevant proposition 6.2 and a sketch of the proof. Since the collar has thickness $O(1)$, these localized corrections do not alter any geometric radius quantity of order $a$ except by an additive $O(1)$ error. Consequently, if $\text{Rad}_{g}(\widetilde{\mathcal{M}}_{1})$ denotes the Schoen–Yau (HH–)radius computed with respect to the interior metric $g$, then
(234) $\text{Rad}(\mathcal{M}_{1})=\text{Rad}_{g}(\widetilde{\mathcal{M}}_{1})+O(1),$
with the $O(1)$ constant depending only on the collar geometry and the fixed gluing profile.
Set $\Omega:=\mathcal{M}_{1}$ and denote its outer boundary by $\partial\Omega=\partial\mathcal{M}_{1}$. Let $H$ be the mean curvature of $\partial\Omega$ in the initial slice $(\Omega,g)$ with respect to the outward unit normal, and let $k$ be the second fundamental form of the slice in spacetime. Because the gluing corrections are supported away from $\partial\Omega$, the induced pair $(\gamma,k^{\top})$ on $\partial\Omega$ coincides with that coming from the double–null interface data. In particular, there is a strictly positive barrier gap
(235) $c_{*}:=\inf_{\partial\Omega}\Big(H-|\mbox{tr}_{\partial\Omega}k|\Big)>0.$
Fix a target radius bound
(236) $0<\text{Rad}(\Omega)\leq R_{*}<R_{\mathrm{crit}}(c_{*}),$
strictly below the Schoen–Yau critical radius associated to the gap $c_{*}$.
Following Schoen–Yau, consider the mixed boundary value problem for the operator
$\mathcal{L}:=-\Delta+\tfrac{1}{2}R-h$
on $\Omega$, with Robin boundary condition
$\partial_{\nu}f+(\mbox{tr}_{\partial\Omega}k)\,f=0\quad\text{on }\partial\Omega,$
where $R$ is the scalar curvature of $(\Omega,g)$ and $h$ is the quadratic form in $k$ appearing in the stability operator for marginally outer trapped surfaces (see [44] for the precise formula). Standard elliptic theory yields a strictly positive first eigenfunction $f>0$.
Let $\Gamma\subset\partial\Omega$ be a boundary curve realizing $\text{Rad}(\Omega)$ up to $o(1)$, and let $\Sigma\subset\Omega$ be a spanning disk with $\partial\Sigma=\Gamma$. Define the weighted functional
$\mathcal{L}_{f}(\Sigma)=\int_{\Sigma}f\,d\mu_{\Sigma}-c\int_{\Omega_{\Sigma}}f\,d\mu_{g},\qquad c\in(0,c_{*}],$
where $\Omega_{\Sigma}$ is the region enclosed by $\Sigma$ and $\Gamma$. Using first and second variation formulas for $\mathcal{L}_{f}$, together with the strict boundary barrier (235) and the distance–to–boundary foliation with explicit supersolution/subsolution profiles $\varphi(d)$, one obtains existence of a minimizing surface with $\partial\Omega$ acting as a strict barrier and derives the quantitative curvature–radius inequality of Schoen–Yau type. In particular, if $\text{Rad}(\Omega)<R_{\mathrm{crit}}(c_{*})$, then no marginally outer trapped surface can be contained in $\Omega$.
[…]
$\min_{\partial\Omega}\Big(H(\tilde{g})-|\mbox{tr}_{\partial\Omega}\tilde{k}|\Big)\geq c_{*}>0,\qquad\text{Rad}(\Omega)\leq R_{*}<R_{\mathrm{crit}}(c_{*}),$
and therefore $\mathcal{M}_{1}$ contains no MOTS on the initial slice t=at=-a. Since all subsequent smoothing and interface adjustments are performed by compactly supported deformations away from $\partial\Omega$ with size $O(a^{-3/2})$ in uniformly–local norms, both the boundary gap and the radius bound are stable under these operations.
It remains to ensure that the same nonexistence property holds on the full initial slice $\mathcal{M}_{-a}$. This is provided by the following proposition, which shows that the Schoen–Yau barrier–radius inequality persists across the glued regions and excludes trapped or marginally outer trapped surfaces on $\mathcal{M}_{-a}$.
Proposition 6.1 (Absence of MOTS on the initial slice).
Let $\mathcal{M}_{-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}$ be the initial Cauchy slice constructed as follows: (i) $\mathcal{M}_{1}$ is the interior region with boundary sphere $S_{-a,0}=\partial\mathcal{M}_{1}$ carrying the prescribed boundary geometry from the double–null interface; (ii) $\mathcal{M}_{2}=\mathcal{M}_{-a}\cap D_{a,1}$ is the characteristic gluing region with induced data from the characteristic development $D_{a,1}$; (iii) $\mathcal{M}_{3}=\mathcal{M}_{ext}$ is an exterior Kerr slice (in the domain of outer communication). Assume that the boundary data on $S_{-a,0}$ satisfy the strict Yau–barrier inequality in the form of condition (c) of Theorem 2.2, and that $\text{Rad}(\mathcal{M}_{1})$ is chosen by
(237) $\text{Rad}(\mathcal{M}_{1})=\frac{3\pi}{4}(a-1)+O(a^{-1}),~a\gg 1$
Then $\mathcal{M}_{-a}$ contains no closed marginally outer trapped surface and no closed trapped surface.
Proof.
With the choice (237), a direct expansion gives
(238) $\frac{3\pi}{2\text{Rad}(\mathcal{M}_{1})}-\frac{2}{a}=\frac{2}{a^{2}}+O(a^{-3}),$
and likewise
(239) $\frac{3\pi}{2\text{Rad}(\mathcal{M}_{1})\sqrt{1+\frac{1}{10a}}}-\frac{2}{a}=\frac{8}{5a^{2}}+O(a^{-3}).$
Under condition (c) of Theorem 2.2, i.e.
(240) $\frac{17}{10a^{2}}<\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{\prime}||\underline{\hat{\chi}}|^{2}(u^{\prime},\epsilon)\,du^{\prime}+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{\prime}|^{2}}\int_{u_{\infty}}^{u^{\prime}}|u^{\prime\prime}|^{2}|\underline{\hat{\chi}}|^{2}(u^{\prime\prime},\epsilon)\,du^{\prime\prime}\,du^{\prime}<\frac{19}{10a^{2}},$
the computation carried out in the proof of Theorem 2.2 yields the quantitative defect
(241) $\Big(H-|\kappa|\Big)\big|_{S_{-a,0}}-\frac{3\pi}{2\text{Rad}(\mathcal{M}_{1})}\leq-\frac{1}{10a^{2}}+O(a^{-5/2}).$
In particular, for $a\gg 1$ the right-hand side is strictly negative. Hence $S_{-a,0}$ violates the Yau criterion (cf. [44] and the formulation 2.1) with a strict margin. By the Yau barrier theorem, $\mathcal{M}_{1}$ contains no closed MOTS (and thus no closed trapped surface).
By construction, $\mathcal{M}_{3}$ is a Kerr slice lying in the domain of outer communication; in particular it contains no closed trapped surface and no closed MOTS.
Let $S\subset\mathcal{M}_{2}$ be any embedded $2$-sphere arising as a section of the outgoing null foliation in $D_{a,1}$. Along each incoming null generator, $\mbox{tr}\chi$ satisfies the Raychaudhuri transport equation
(242) $\nabla_{3}\mbox{tr}\chi+\frac{1}{2}\mbox{tr}\underline{\chi}\,\mbox{tr}\chi=2\underline{\omega}\,\mbox{tr}\chi+2\mbox{div }\eta+2|\eta|^{2}+2\rho-\hat{\chi}\cdot\underline{\hat{\chi}}$.
Using the bootstrap bounds available in $D_{a,1}$ (in particular the standard estimates on $\underline{\omega},\eta,\rho,\hat{\chi},\underline{\hat{\chi}}$ inherited from the characteristic construction) and Grönwall along the $\nabla_{3}$–flow, one obtains the quantitative perturbative estimate
(243) $\Big|\mbox{tr}\chi-\frac{2}{|u|}\Big|\lesssim a^{-1/2}|u|^{-1}\qquad\text{throughout }\mathcal{M}_{2}.$
On $\mathcal{M}_{2}$ one has $|u|\sim a$ (since $\mathcal{M}_{2}=\mathcal{M}_{-a}\cap D_{a,1}$ is localized near the interface), hence (243) improves to
(244) $\Big|\mbox{tr}\chi-\frac{2}{|u|}\Big|\lesssim a^{-3/2},\qquad\text{on }\mathcal{M}_{2}.$
Therefore, for $a\gg 1$, $0>u\in[u_{\infty},-a]$
$\mbox{tr}\chi\geq\frac{2}{|u|}-Ca^{-3/2}\geq ca^{-1}>0\qquad\text{on every section }S\subset\mathcal{M}_{2},$
with $c>0$ universal. In particular, no $S\subset\mathcal{M}_{2}$ can satisfy $\mbox{tr}\chi=0$, so $\mathcal{M}_{2}$ contains no closed MOTS; moreover, $\mbox{tr}\chi$ cannot be negative on such a closed section, so there are no closed trapped surfaces in $\mathcal{M}_{2}$. This completes the proof. ∎
We provide the following proposition sketching the gluing mechanism to construct data on the slice $\mathcal{M}_{-a}$. This is a direct application of the Corvino-Schoen gluing technique [22]. This is slightly less important in the sense that we are ultimately interested in the interior dynamics and dynamical formation of MOTS. This study deserves a separate rigorous attention which we leave it for future work. But we provide a sketch of the proof for completeness.
Proposition 6.2 (Construction of constraint data with interior–characteristic–Kerr matching).
Fix an integer $N\geq 4$, $s\leq 3$ for suffiicnetly large $\mathbb{N}$, and $a\gg 1$. There exist a smooth Riemannian metric $\widetilde{g}$ and a smooth symmetric 22–tensor $k$ on the Cauchy slice
$\mathcal{M}_{-a}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{ext},$
such that $(\widetilde{g},k)$ solve the vacuum Einstein constraint equations on $\mathcal{M}_{-a}$, i.e.
[…]
and enjoy the following matching properties.
(i) There exists an open subset $\widetilde{\mathcal{M}}_{1}\Subset\mathcal{M}_{1}$ such that
[…]
where $g$ is the prescribed interior metric with $H$-radius or Schoen-Yau radius $\text{Rad}_{g}(\widetilde{\mathcal{M}}_{1})\gg 1$. Moreover, in a fixed harmonic coordinate chart on $\widetilde{\mathcal{M}}_{1}$,
(247) $\|\partial^{k+1}g(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-1},~\|\partial^{k}k(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-3/2},$
[…]
(248) $\|\partial^{k+1}g(-a)\|_{H^{s-1}_{ul}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-1},~\|\partial^{k}k(-a)\|_{H^{s-1}_{ul}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-3/2},~k\geq 0,$
where on the collar $\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1}$ of thickness $O(\widetilde{\epsilon}a),~\widetilde{\epsilon}\in[\frac{1}{1000},\frac{1}{100}]$ the estimate reads
(249) $\|\partial^{k}k\|_{L^{\infty}(\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-\frac{3}{2}-k},$
[…]
(251) $\|\partial^{k}k\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-3/2},~k\geq 0,$
and $\partial g$ smoothly interpolates between the data on $\widetilde{\mathcal{M}}_{1}$ and the data on $\Sigma_{int}=\mathcal{M}_{-1}\cap D_{a,1}$. Here $H^{s-1}_{ul}$ denotes the uniformly local Sobolev space (defined in the technical version of the main theorem 2.2), $\delta$ is the background metric in the chosen harmonic chart, and $C$ is independent of $a$.
(ii) On $\mathcal{M}_{ext}$, the data $(\widetilde{g},k)$ are isometric to a $t=\operatorname{const}$ slice of the Kerr family with parameters $(m,J)$, and the parameters satisfy
(252) $|m-a^{\frac{1}{2}}|+|J-a|\lesssim a^{-3/2}.$
(iii) On $\mathcal{M}_{2}$, the data $(\widetilde{g},k)$ coincide with the induced Cauchy data on $\mathcal{M}_{-a}\cap D_{a,1}$ arising from the characteristic development $D_{a,1}$.
Sketch of the Proof: Write the vacuum constraints in the $(g,\pi)$-formulation used by Corvino–Schoen:
$\Phi(g,\pi)=\Big(R(g)+\tfrac{1}{2}(\mbox{tr}_{g}\pi)^{2}-|\pi|_{g}^{2},\ \mbox{div }_{g}\pi\Big),\qquad\pi:=k-(\mbox{tr}_{g}k)\,g,$
so that $\Phi(g,\pi)=0$ is equivalent to the standard constraint system $R(g)-|k|_{g}^{2}+(\mbox{tr}_{g}k)^{2}=0$, $\mbox{div }_{g}k-d(\mbox{tr}_{g}k)=0$. Let $s\geq\mathbb{N}$ be fixed.
We give a Corvino-Schoen [22] type localized deformation proof, in the form adapted to interior-characteristic–Kerr matching as in the Cauchy-data constructions of [28], for example (even though their framework is fundamentally different from ours). Throughout, $\Sigma:=\mathcal{M}_{t=-a}$, $s\geq 4$ is fixed, and $a\gg 1$.
Recall the decomposition
$\Sigma=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3},\qquad\mathcal{M}_{2}=\Sigma\cap\mathcal{D}_{-a,1},\qquad\mathcal{M}_{3}=\Sigma\setminus(\mathcal{M}_{1}\cup\mathcal{M}_{2}),$
and $\mathcal{M}_{1}=\widetilde{\mathcal{M}}_{1}\cup(\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1})$ with $\widetilde{\mathcal{M}}_{1}\Subset\mathcal{M}_{1}$ and $O(1)$ collar $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$. Let $(g,k)$ denote the prescribed interior constraint data on $\widetilde{\mathcal{M}}_{1}$ satisfying (276)–(278) . Let $(g_{\mathrm{char}},k_{\mathrm{char}})$ be the Cauchy data induced by (,g^)($\mathcal{M}$$\widehat${g}) on $\mathcal{M}_{2}$; by the stated double–null coefficient bounds on $D_{a,1}$, the restriction of $(g_{\mathrm{char}},k_{\mathrm{char}})$ to any fixed–width collar in $\mathcal{M}_{2}$ satisfies the same scale of smallness (in uniformly local Sobolev norms) as the interior collar data i.e., while expressed in the double null frame
[…]
$||\eta,\underline{\eta}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|^{2}},~||\omega||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|^{2}},~||\hat{\chi}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{-\frac{1}{2}}}{|u|}.$
Finally, for each parameter $m>0$ and with angular momentum fixed as $J:=a$, let
$(g_{\mathrm{Kerr}}(m),k_{\mathrm{Kerr}}(m)):=(g_{\mathrm{Kerr}}(m,J=a),k_{\mathrm{Kerr}}(m,J=a))$
denote the Kerr $t=\operatorname{const}$ Cauchy data on the exterior end. Choose open subsets
$\widetilde{\mathcal{M}}_{3}\Subset\mathcal{M}_{3},\qquad\widetilde{\mathcal{M}}_{1}\Subset\mathcal{M}_{1}$
and fix a gluing annulus $\mathcal{A}\Subset\Sigma$ such that
(253) $\mathcal{A}\cap\big(\widetilde{\mathcal{M}}_{1}\cup\mathcal{M}_{2}\cup\widetilde{\mathcal{M}}_{3}\big)=\varnothing,\qquad\Sigma\setminus\big(\widetilde{\mathcal{M}}_{1}\cup\mathcal{M}_{2}\cup\widetilde{\mathcal{M}}_{3}\big)\subset\mathcal{A},$
and $\mathcal{A}$ is contained in the union of the interior collar $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ and a compact collar region in $\mathcal{M}_{3}$.
Define a smooth approximate pair $(g_{0}(m),k_{0}(m))$ on $\Sigma$ by patching with a partition of unity:
$(g_{0}(m),k_{0}(m))=\begin{cases}(g,k)&\text{on }\widetilde{\mathcal{M}}_{1},\\ (g_{\mathrm{char}},k_{\mathrm{char}})&\text{on }\mathcal{M}_{2},\\ (g_{\mathrm{Kerr}}(m),k_{\mathrm{Kerr}}(m))&\text{on }\widetilde{\mathcal{M}}_{3},\end{cases}$
and interpolate smoothly inside $\mathcal{A}$. Each piece satisfies the constraints on its region, hence the constraint defect
$\mathcal{E}(m):=\Phi(g_{0}(m),k_{0}(m))$
is supported in $\mathcal{A}$. Moreover, since the collars have thickness $O(1)$ and the smallness in the sense the double–null bounds stated previously, We have the quantitative estimate
(254) $\|\mathcal{E}(m)\|_{H^{s-2}(\Sigma)}\lesssim a^{-1},$
uniformly for $m$ in a fixed neighborhood of $a^{1/2}$.
Now let $\Phi$ be the vacuum constraint map
$\Phi(g,k)=\big(R[g]-|k|_{g}^{2}+(\mbox{tr}_{g}k)^{2},\ \mbox{div }_{g}k-d(\mbox{tr}_{g}k)\big).$
Fix $m$ and abbreviate $(g_{0},k_{0}):=(g_{0}(m),k_{0}(m))$. Following Corvino–Schoen, we seek a compactly supported correction $(h,p)$ with
$\operatorname{supp}(h,p)\subset\mathcal{A},\qquad\Phi(g_{0}+h,k_{0}+p)=0.$
The key device is to impose the adjoint–generated form
(255) $(h,p)=D\Phi^{*}_{(g_{0},k_{0})}(\varphi,Y),$
where $D\Phi^{*}_{(g_{0},k_{0})}$ denotes the formal $L^{2}$–adjoint of the linearized constraint operator, and $(\varphi,Y)$ is a scalar-vector field pair supported in $\mathcal{A}$ (typically with homogeneous boundary conditions on $\partial\mathcal{A}$). Substituting (255) into the linearization yields an elliptic system for $(\varphi,Y)$:
$D\Phi_{(g_{0},k_{0})}\,D\Phi^{*}_{(g_{0},k_{0})}(\varphi,Y)=-\mathcal{E}(m)+\text{(quadratic terms)}.$
Denote the elliptic operator
$\mathcal{L}_{(g_{0},k_{0})}:=D\Phi_{(g_{0},k_{0})}\,D\Phi^{*}_{(g_{0},k_{0})}.$
This is the standard Corvino–Schoen reduction: $\mathcal{L}_{(g_{0},k_{0})}$ is elliptic on $\mathcal{A}$, and solving for $(\varphi,Y)$ yields a correction $(h,p)$ supported in $\mathcal{A}$.
The obstruction to solving $\mathcal{L}_{(g_{0},k_{0})}(\varphi,Y)=f$ on $\mathcal{A}$ is precisely the finite–dimensional space of ‘Killing initial data’ (KID) (kernel of $D\Phi^{*}$), which is the cokernel of $D\Phi$ and hence of $\mathcal{L}$. More concretely, let
$\mathcal{K}:=\ker\big(D\Phi^{*}_{(g_{0},k_{0})}\big)\subset H^{s}(\mathcal{A})\times H^{s}(\mathcal{A};T\Sigma)$
(with the boundary conditions used to define $\mathcal{L}$). Then $\mathcal{L}$ is invertible on $\mathcal{K}^{\text{Per}}p$, and solvability requires the standard orthogonality conditions
(256) $\langle\mathcal{E}(m),\ \kappa\rangle_{L^{2}(\mathcal{A})}=0\qquad\text{for all }\kappa\in\mathcal{K}.$
In our setting (Kerr end, fixed center, and with $J$ fixed), the only remaining balancing mode is the time–translation KID, i.e. the mass mode; the angular momentum mode is eliminated by the choice $J=a$.
Li-Yu construction [28] shows that the balancing functional in the mass direction depends nondegenerately on the Kerr mass parameter $m$: there exists a smooth scalar function $\mathfrak{B}(m)$ (the mass balance) such that (256) holds if and only if $\mathfrak{B}(m)=0$, and
(257) $\frac{d}{dm}\mathfrak{B}(m)\neq 0\qquad\text{for }m\text{ near }a^{1/2}.$
Moreover, by (254) the imbalance is small:
$|\mathfrak{B}(a^{1/2})|\lesssim a^{-1}.$
Hence, by the one–dimensional implicit function theorem using (257), there exists
[…]
We now fix this choice of $m=m(a)$ and keep $J=a$. Set $(g_{0},k_{0}):=(g_{0}(m(a)),k_{0}(m(a)))$ and $\mathcal{E}:=\mathcal{E}(m(a))$. By (256) (equivalently $\mathfrak{B}(m(a))=0$), the source $\mathcal{E}$ lies in the range of $D\Phi$ on $\mathcal{A}$ and hence in the range of $\mathcal{L}$ on $\mathcal{K}^{\text{Per}}p$.
We solve the nonlinear problem by a contraction/implicit function argument in Sobolev spaces. Let $\mathcal{Q}(h,p)$ denote the quadratic remainder in the Taylor expansion:
$\Phi(g_{0}+h,k_{0}+p)=\Phi(g_{0},k_{0})+D\Phi_{(g_{0},k_{0})}(h,p)+\mathcal{Q}(h,p).$
Using (255), we seek $(\varphi,Y)$ such that
(259) $\mathcal{L}_{(g_{0},k_{0})}(\varphi,Y)=-\mathcal{E}-\mathcal{Q}\big(D\Phi^{*}_{(g_{0},k_{0})}(\varphi,Y)\big).$
Since $\mathcal{L}$ is elliptic and invertible on $\mathcal{K}^{\text{Per}}p$, let $\mathcal{G}$ be its bounded inverse there (with the chosen boundary conditions). Define the map
$\mathcal{T}(\varphi,Y):=\mathcal{G}\Big(-\mathcal{E}-\mathcal{Q}\big(D\Phi^{*}(\varphi,Y)\big)\Big).$
Standard product estimates for the constraint nonlinearity (valid for $s\geq\mathbb{N}$ for sufficiently large $\mathbb{N}$ on the compact annulus $\mathcal{A}$) give
$\|\mathcal{Q}(D\Phi^{*}(\varphi,Y))\|_{H^{s-2}(\mathcal{A})}\lesssim\|(\varphi,Y)\|_{H^{s}(\mathcal{A})}^{2},$
and hence, using (254),
$\|\mathcal{T}(\varphi,Y)\|_{H^{s}(\mathcal{A})}\lesssim\|\mathcal{E}\|_{H^{s-2}(\mathcal{A})}+\|(\varphi,Y)\|_{H^{s}(\mathcal{A})}^{2}\lesssim a^{-1}+\|(\varphi,Y)\|_{H^{s}(\mathcal{A})}^{2}.$
For $a\gg 1$, $\mathcal{T}$ is a contraction on a ball of radius $Ca^{-3/2}$ in $H^{s}(\mathcal{A})$. Thus there exists a unique fixed point $(\varphi,Y)$ solving (259). Define $(h,p)$ by (255) and set
$(\widetilde{g},\widetilde{k}):=(g_{0}+h,k_{0}+p).$
Then $\Phi(\widetilde{g},\widetilde{k})=0$ on $\Sigma$. Elliptic regularity bootstraps $(\varphi,Y)$ (hence $(h,p)$) to smoothness since the coefficients are smooth and $\mathcal{A}$ is compact. By construction, $(h,p)=D\Phi^{*}(\varphi,Y)$ is supported in $\mathcal{A}$ and $\mathcal{A}$ is disjoint from the protected regions (253). Therefore
[…]
and on the far exterior region $\widetilde{\mathcal{M}}_{3}$,
$(\widetilde{g},\widetilde{k})\equiv(g_{\mathrm{Kerr}}(m(a),J=a),k_{\mathrm{Kerr}}(m(a),J=a)).$
This proves (i) and (iii), and yields (ii) with $J=a$ and $|m(a)-a^{1/2}|\lesssim a^{-3/2}$ from (258). The interior $L^{\infty}$ and $H^{s-1}_{\mathrm{ul}}$ bounds remain valid on $\widetilde{\mathcal{M}}_{1}$ since the deformation vanishes there. This completes the sketch of the proof. Notice that we provided a sketch of the proof and not the full technical details since that is tangential to the purpose of this article.
7. Control of the Second Fundamental Form and the H-radius
This completes the central part of this work. In particular, we prove that there exists an open set of initial data that is almost compatible with the scaling of [3] (with some differences handled by new scaling) and is capable of producing a semi-global characteristic development of the vacuum Einstein’s equations. Simultaneously, we also obtain concentrated generalized Yau mean curvature $c$ (as defined in Theorem 2.1) along the null hypersurface u=au=-a. The idea then is to show that the radius of the boundary $S_{-a,\epsilon}$ is large enough (compared to that of $S_{-a,0}$) so that the condition stated in the theorem 2.1 is met. The last part involves constructing the data on the Cauchy slice $\mathcal{M}_{t=-a}$ and evolving it for a short enough time $\epsilon>0$ (see the diagram 2 for clarity). The motivation behind solving the characteristic initial value problem is to concentrate the mean curvature along the incoming direction by means of large conjugate shear $\underline{\hat{\chi}}$ concentrated along the initial null-hypersurface $\underline{H}_{0}$ starting from dispersed data at past null-infinity.
7.1. Norms
In this section, we control the interior geometry. In particular, we prove uniform estimates for the spacetime Weyl curvature through the use of Bel-Robinson energy and its higher-order analogue. We provide the initial data on the interior which is smoothly matched with the data on the exterior domain. Then we prove the interior development $J^{+}(\mathcal{M}_{1})$ up to $O(1)$ time where the spacetime geometry is uniformly controlled. In practice, we consider a slightly bigger domain $\mathcal{M}^{1/a}_{int}:=\mathcal{M}_{1}\cup$
Let $(\mathcal{M}^{3+1},\mathsfbf g)$ be a smooth, time-oriented Lorentzian manifold with Levi-Civita connection $\mathsfbf D$ and volume form $dV$. Assume vacuum:
(260) $\text{Ric}(\mathsfbf g)=0.$
Let $W$ denote the Weyl tensor of $\mathsfbf g$; under (260), $W=\text{Riem}(\mathsfbf g)$ and the vacuum Bianchi identities are
(261) $\mathsfbf D^{\alpha}W_{\alpha\beta\gamma\delta}=0,\qquad\mathsfbf D^{\alpha}{}^{\star}W_{\alpha\beta\gamma\delta}=0.$
Fix $p\in\mathcal{M}$ and let $\mathcal{N}^{-}(p)$ be the past null cone from $p$. For $0<\tau\leq 1$ define the cone domain
(262) $\mathcal{D}_{\tau}(p):=J^{-}(p)\cap J^{+}(\Sigma_{t(p)-\tau}),\qquad\partial\mathcal{D}_{\tau}(p)=\Sigma_{t(p)-\tau}\cup\mathcal{N}^{-}_{\tau}(p),$
where $\Sigma_{t(p)-\tau}$ is a spacelike cut of the cone and $\mathcal{N}^{-}_{\tau}(p)$ is the portion of $\mathcal{N}^{-}(p)$ between $\Sigma_{t(p)-\tau}$ and $p$.
On $\mathcal{N}^{-}_{\tau}(p)$ fix a null frame $(e_{3},e_{4},e_{A})$, $A=1,2$, with
(263) $e_{4}=L\ \text{tangent to generators of }\mathcal{N}^{-}(p),\qquad\mathsfbf g(e_{3},e_{4})=-2,\qquad\mathsfbf g(e_{3},e_{3})=\mathsfbf g(e_{4},e_{4})=0,\qquad e_{A}\in TS,$
and define the canonical timelike vector
(264) $T:=\frac{1}{2}(e_{3}+e_{4}).$
Let $n$ denote the future unit normal to $\Sigma_{t(p)-\tau}$ and $d\mu_{\Sigma}$ its induced volume form. Let $d\mu_{\mathcal{N}}$ denote the induced null measure on $\mathcal{N}^{-}_{\tau}(p)$ (any fixed normalization consistent with (263); the identities below are independent of the parametrization).
7.2. Estimates on the metric
Under the bootstrap assumption, we will prove the uniform control of the metric on the time interval $a,-a+\frac{3}{4}]$ in this section. This is necessary since we will define the norm with respect to the dynamical metric $g$.
Proposition 7.1.
Let $a,-a+1]\times\mathcal{M}$ be a globally hyperbolic spacetime foliated by constant time slices $\mathcal{M}_{T=\text{constant}}$ and let $g(x,T)$ be the Riemannian metric on $\mathcal{M}_{T}$. Denote $\lambda_{1}(T)$ and $\lambda_{2}(T)$ by the maximal and minimal eigenvalues, respectively, of the symmetric bilinear form $g(x,T)$ with respect to the initial metric $g_{-a}$. Then, under the bootstrap assumption (280-282), we have the estimate
(265) $|\lambda_{1}(T)-1|+|1-\lambda_{2}(T)|\lesssim a^{-\frac{3}{2}+\delta}~on~[-a,-a+1]$
Proof.
Fix a point $x\in\mathcal{M}$, and consider the symmetric bilinear form $g(T,x)$ on $T_{x}M$. Define the maximal and minimal eigenvalues of $g(T,x)$ with respect to g(a,x)g(-a,x) at the point $x$ as
(266) $\lambda_{1}(x,T):=\sup_{0\neq Y\in T_{x}M}\frac{g(T,x)(Y,Y)}{g(-a,x)(Y,Y)},\quad\lambda_{2}(x,T):=\inf_{0\neq Y\in T_{x}M}\frac{g(T,x)(Y,Y)}{g(-a,x)(Y,Y)}.$
We aim to estimate $|\lambda_{1}(x,T)-1|$ and $|1-\lambda_{2}(x,T)|$ uniformly in $x$ and $T\in[-a,-a+1]$. Recall the evolution equation for the metric
$\partial_{t}g_{ij}=-2Nk_{ij}+(\mathscr{L}_{X}g)_{ij},$
where $N$ is the lapse function, $k_{ij}$ is the second fundamental form of the constant time slice, and $X$ is the shift vector field. For fixed tangent vector $Y\in T_{x}M$, one observes
$\partial_{t}g(Y,Y)=-2Nk(Y,Y)+(\mathscr{L}_{X}g)(Y,Y).$
To estimate the right-hand side, we begin with the bound
$|k(Y,Y)|\leq\sqrt{g(-a,x)^{ik}g(-a,x)^{jl}k_{ij}k_{kl}}\cdot g(-a,x)(Y,Y)$
and therefore, under the bootstrap assumption (280-282)
$|k(Y,Y)|\lesssim a^{-\frac{3}{2}+\delta}g(-a,x)(Y,Y)$
and
$(\mathscr{L}_{X}g)(Y,Y)\lesssim a^{-2+\delta}g(-a,x)(Y,Y)$
and so
$|g(T,x)(Y,Y)-g(-a,x)(Y,Y)|\lesssim a^{-\frac{3}{2}+\delta}g(-a,x)(Y,Y).$
To improve this to an estimate of relative deviation, we write
$\left|\frac{g(T,x)(Y,Y)}{g(-a,x)(Y,Y)}-1\right|\lesssim a^{-\frac{3}{2}+\delta}~\forall T\in[-a,-a+1].$
This completes the proof. ∎
Now let us focus on obtaining interior estimates. For any $(0,4)$-tensor $U$ with the algebraic symmetries of a Weyl tensor, define the Bel-Robinson tensor
(267) $Q_{\alpha\beta\gamma\delta}[U]:=U_{\alpha\mu\gamma\nu}U_{\beta}{}^{\mu}{}_{\delta}{}^{\nu}+{}^{\star}U_{\alpha\mu\gamma\nu}\,{}^{\star}U_{\beta}{}^{\mu}{}_{\delta}{}^{\nu}$.
Null components.
For $U=W$, recall the definition of standard null curvature components (relative to (263)):
(268) $\alpha_{AB}$ $:=W(e_{4},e_{A},e_{4},e_{B}),$
(269) $\beta_{A}$ $:=\tfrac{1}{2}W(e_{4},e_{A},e_{4},e_{3}),$
(270) $\rho$ $:=\tfrac{1}{4}W(e_{4},e_{3},e_{4},e_{3}),\qquad\sigma:=\tfrac{1}{4}{}^{\star}W(e_{4},e_{3},e_{4},e_{3}),$
(271) $\underline{\beta}_{A}$ $:=\tfrac{1}{2}W(e_{3},e_{A},e_{3},e_{4}),\qquad\underline{\alpha}_{AB}:=W(e_{3},e_{A},e_{3},e_{B}).$
Also recall the definition of the electric and magnetic parts (adapted to the Cauchy slice $T=u+\underline{u}$) of the Weyl curvature
$E:=W(T,\cdot,T,\cdot),~B:=^{*}W(T,\cdot,T,\cdot)$
Then (by direct contraction of (267) with (264) and (263)) one has
(272) $Q[W](e_{4},T,T,T)=\frac{1}{4}|\alpha|^{2}+\frac{1}{2}|\beta|^{2}+\frac{1}{2}(\rho^{2}+\sigma^{2})+\frac{1}{2}|\underline{\beta}|^{2},$
and in terms of the electric and magnetic fields
(273) $|E|^{2}+|B|^{2}\lesssim Q[W](T,T,T,T)\lesssim|E|^{2}+|B|^{2}$
Let us now define the locally uniform Sobolev norms for the interior $\mathcal{M}^{1/a}_{int}$. Fix a smooth cutoff $\chi\in C_{c}^{\infty}(B_{2}(0))$ with $\chi\equiv 1$ on $B_{1}(0)$, and for each $y\in\mathcal{M}^{1/a}_{int}$ define $\chi_{y}(x):=\chi(x-y)$. For an integer $s$ sufficiently large, set
$\|u(t)\|_{H^{s}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int}))}:=\sup_{y\in\mathcal{M}^{1/a}_{int}}\ \|\chi_{y}\,u(t)\|_{H^{s}(\mathcal{M}^{1/a}_{int})}$
(and similarly for $L^{2}_{\mathrm{ul}},\,L^{\infty}_{\mathrm{ul}}$).
Proposition 7.2 (Propagation of interior estimates up to unit time).
Fix integers $s,N$ sufficiently large. Let $(\mathcal{M},g)$ be a smooth vacuum spacetime written in spacetime harmonic gauge
$\Box_{g}x^{\mu}=0,$
so that the Einstein vacuum equations reduce to a quasilinear wave system
$\Box_{g}g_{\mu\nu}=Q_{\mu\nu}(\partial g,\partial g).$
Let $\mathcal{M}_{-a}$ be a Cauchy hypersurface with induced data $(g,k)$ satisfying the constraint equations, and suppose the metric is expressed in ADM form
$g=-N^{2}dt^{2}+g_{ij}(dx^{i}+X^{i}dt)(dx^{j}+X^{j}dt).$
Assume that $\mathcal{M}_{-a}$ admits a decomposition
$\mathcal{M}_{-a}=\mathcal{M}^{1/a}_{\mathrm{int}}\cup\mathcal{M}^{{}^{\prime}}_{\mathrm{ext}},$
where $\mathcal{M}^{1/a}_{\mathrm{int}}$ is a connected interior region of $H$-radius $O(a)$ ($\approx\frac{3\pi a}{4}$), whose future domain of dependence intersects the characteristic development domain $D_{a,1}$ constructed in Section 3. Assume that the induced data on the interface
$\Sigma_{\mathrm{int}}:=\mathcal{M}_{-a}\cap D_{a,1}$
agree with the data obtained from the characteristic development, to order $H^{s}$.
Define the uniformly local Sobolev norm by
$\|F\|_{H^{m}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{\mathrm{int}})}:=\sup_{p\in\mathcal{M}^{1/a}_{\mathrm{int}}}\|F\|_{H^{m}(B_{1}(p))},$
where $B_{1}(p)$ denotes the geodesic unit ball (with respect to $g_{ij}(-a)$).
Assume that on $\widetilde{\mathcal{M}}_{1}\subset\mathcal{M}^{1/a}_{\mathrm{int}}$ the initial data satisfy the smallness bounds
(274) $\|\partial^{k+1}g(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-1},~\|\partial^{k}k(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-3/2},$
[…]
(275) $\|\partial^{k+1}g(-a)\|_{H^{s-1}_{ul}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-1},~\|\partial^{k}k(-a)\|_{H^{s-1}_{ul}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-3/2},~k\geq 0,$
where on the collar $\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1}$ of thickness $O(\widetilde{\epsilon}a)$ for $\widetilde{\epsilon}\in[\frac{1}{1000},\frac{1}{100}]$, the estimate reads
(276) $\|\partial^{k}k\|_{L^{\infty}(\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-\frac{3}{2}-k},$
[…]
(278) $\|\partial^{k}k\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-3/2},~k\geq 0,$
and $\partial g$ smoothly interpolates between the data on $\widetilde{\mathcal{M}}_{1}$ and the data on $\Sigma_{int}=\mathcal{M}_{-1}\cap D_{a,1}$. Assume moreover that the harmonic gauge constraints and the Einstein constraint equations hold initially.
Then for every fixed $\delta\in(0,\tfrac{1}{2})$ there exists $a_{0}=a_{0}(\delta,s,C)$ such that for all $a\geq a_{0}$ the corresponding spacetime solution exists on the slab
$\mathcal{R}_{a}:=J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\{-a\leq t\leq-a+1\},$
and throughout $\mathcal{R}_{a}$ the following bounds hold:
(279) $\|k\|_{L^{\infty}(\mathcal{R}_{a})}$ $\leq C\,a^{-3/2+\delta},$
$\|X\|_{L^{\infty}(\mathcal{R}_{a})}$ $\leq C\,a^{-3/2+\delta},$
$\|N-1\|_{L^{\infty}(\mathcal{R}_{a})}$ $\leq C\,a^{-3/2+\delta}.$
The constant $C$ is a numerical and independent of $a$.
Proof.
We begin by recording the gauge–invariant curvature bounds which will be used, later, as an input in the spacetime harmonic gauge estimates for $(g_{ij},k_{ij},N,X)$. Although the local Cauchy development on the slab $\{-a\leq t\leq-a+1\}$ can be obtained directly from the reduced Einstein system in harmonic gauge, we include the Bel–Robinson argument both for completeness and because it yields a canonical, coordinate–free control of the Weyl tensor on truncated null cones.
Throughout the region
$\mathcal{R}_{a}:=J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\{-a\leq t\leq-a+1\},$
we assume the bootstrap bounds
(280) $\|k\|_{L^{\infty}(\mathcal{R}_{a})}+\|\partial g\|_{L^{\infty}(\mathcal{R}_{a})}$ $\leq C\,a^{-1+\delta},$
(281) $\|X\|_{L^{\infty}(\mathcal{R}_{a})}$ $\leq C\,a^{-3/2+\delta},$
(282) $\|N-1\|_{L^{\infty}(\mathcal{R}_{a})}$ $\leq C\,a^{-3/2+\delta}.$
The bootstrap will be closed by proving the corresponding improvements with the constant $\tfrac{1}{2}C$ on the right-hand side (after choosing $a\geq a_{0}(\delta)$ sufficiently large).
Fix a point $p\in\mathcal{R}_{a}$ and a parameter $\tau\in(0,\tau_{*}]$, where $\tau_{*}\leq 1$ is chosen so that the causal past $J^{-}(p)\cap\{t(p)-\tau\leq t\leq t(p)\}$ remains in $\mathcal{R}_{a}$. Let $D_{\tau}(p)$ denote the truncated past domain
$D_{\tau}(p):=J^{-}(p)\cap\{t(p)-\tau\leq t\leq t(p)\},$
with boundary decomposition
$\partial D_{\tau}(p)=\Sigma_{t(p)}(p)\ \cup\ \Sigma_{t(p)-\tau}(p)\ \cup\ \mathcal{N}^{-}_{\tau}(p),$
where $\Sigma_{t}(p):=\Sigma_{t}\cap J^{-}(p)$ and $\mathcal{N}^{-}_{\tau}(p)$ is the portion of the past null cone emanating from $p$ between $t=t(p)-\tau$ and $t=t(p)$.
Let $T=\partial_{t}$ be the time–translation vector field associated to the ADM foliation, and let $n$ be the future unit normal to $\Sigma_{t}$. The deformation tensor of $T$ is
$\pi_{T\,\alpha\beta}:=\mathsfbf{D}_{\alpha}T_{\beta}+\mathsfbf{D}_{\beta}T_{\alpha}.$
In ADM variables one has the schematic identity
(283) $|\pi_{T}|\ \lesssim\ |k|+|\nabla N|+|\nabla X|,$
and hence, under (280)–(282) together with the corresponding a priori control on $\nabla N$ and $\nabla X$ (proved later from the harmonic gauge system),
(284) $\|\pi_{T}\|_{L^{\infty}(\mathcal{R}_{a})}\ \leq\ C\,a^{-3/2+\delta}.$
Let $W$ denote the spacetime Weyl tensor. The Bel–Robinson tensor $Q[W]$ is defined by
$Q_{\alpha\beta\gamma\delta}[W]:=W_{\alpha\mu\gamma\nu}W_{\beta}{}^{\mu}{}_{\delta}{}^{\nu}+{}^{\ast}W_{\alpha\mu\gamma\nu}\,{}^{\ast}W_{\beta}{}^{\mu}{}_{\delta}{}^{\nu}$,
where $^{\ast}W$ is the Hodge dual. In vacuum, $W$ satisfies the Bianchi system $\mathsfbf{D}^{\alpha}W_{\alpha\beta\gamma\delta}=0$, and consequently
(285) $\mathsfbf{D}^{\alpha}Q_{\alpha\beta\gamma\delta}[W]=0.$
Define the Bel–Robinson current associated to $T$ by
$P^{\alpha}:=Q^{\alpha}{}_{\beta\gamma\delta}[W]\;T^{\beta}T^{\gamma}T^{\delta}.$
A direct computation using (285) and the symmetry of $Q$ in the first index pair yields the exact divergence identity
(286) $\mathsfbf{D}_{\alpha}P^{\alpha}=\frac{3}{2}\,Q_{\alpha\beta\gamma\delta}[W]\;\pi_{T}^{\alpha\beta}T^{\gamma}T^{\delta}.$
Applying the divergence theorem to $D_{\tau}(p)$ gives
(287) $\int_{D_{\tau}(p)}\mathsfbf{D}_{\alpha}P^{\alpha}\,dV=\int_{\partial D_{\tau}(p)}P\cdot\nu\,d\mu_{\partial},$
where $\nu$ is the future/outward normal density on each boundary component. On the spacelike pieces $\Sigma_{t}(p)$ one has $\nu=n$ and $d\mu_{\partial}=d\mu_{\Sigma_{t}}$, whereas on the null piece $\mathcal{N}^{-}_{\tau}(p)$ we take the standard normalization $\nu=L$ where $L$ is the null generator of $\mathcal{N}^{-}_{\tau}(p)$, and $d\mu_{\partial}=d\mu_{\mathcal{N}}$ is the induced null measure.
Combining (286) and (287) yields the exact energy–flux identity
$\int_{\Sigma_{t(p)}(p)}Q[W](n,T,T,T)\,d\mu_{\Sigma_{t(p)}}+\int_{\mathcal{N}^{-}_{\tau}(p)}Q[W](L,T,T,T)\,d\mu_{\mathcal{N}}$
(288) $\qquad=\int_{\Sigma_{t(p)-\tau}(p)}Q[W](n,T,T,T)\,d\mu_{\Sigma_{t(p)-\tau}}+\frac{3}{2}\int_{D_{\tau}(p)}Q_{\alpha\beta\gamma\delta}[W]\;\pi_{T}^{\alpha\beta}T^{\gamma}T^{\delta}\,dV.$
Define the curvature energy and null flux by
$E_{0}(\tau;p):=\int_{\Sigma_{t(p)-\tau}(p)}Q[W](n,T,T,T)\,d\mu_{\Sigma_{t(p)-\tau}},\qquad F_{0}(\tau;p):=\int_{\mathcal{N}^{-}_{\tau}(p)}Q[W](L,T,T,T)\,d\mu_{\mathcal{N}}.$
Then (7.2) becomes
(289) $E_{0}(0;p)+F_{0}(\tau;p)=E_{0}(\tau;p)+\frac{3}{2}\int_{D_{\tau}(p)}Q_{\alpha\beta\gamma\delta}[W]\;\pi_{T}^{\alpha\beta}T^{\gamma}T^{\delta}\,dV.$
We now estimate the bulk term. By the dominant property of $Q[W]$ and the Cauchy–Schwarz inequality for quadratic forms, there exists a universal numerical constant $C_{0}$ such that
(290) $\big|Q_{\alpha\beta\gamma\delta}[W]\;\pi_{T}^{\alpha\beta}T^{\gamma}T^{\delta}\big|\leq C_{0}\,|\pi_{T}|\,Q[W](T,T,T,T)\leq C_{0}\,|\pi_{T}|\,Q[W](n,T,T,T).$
Using $dV=N\,dt\,d\mu_{\Sigma_{t}}$ and (290), we infer from (289) that
(291) $E_{0}(0;p)+F_{0}(\tau;p)\leq E_{0}(\tau;p)+C\int_{0}^{\tau}\|\pi_{T}\|_{L^{\infty}(\Sigma_{t(p)-s}(p))}\,E_{0}(s;p)\,ds,$
for a numerical constant $C$. Since $F_{0}(\tau;p)\lesssim a^{-1}$, we obtain the standard Grönwall inequality
(292) $E_{0}(\tau;p)\leq E_{0}(0;p)\,\exp\!\left(C\int_{0}^{\tau}\|\pi_{T}\|_{L^{\infty}(\Sigma_{t(p)-s}(p))}\,ds\right)+Ca^{-1}$
In particular, using (284) and $\tau\leq 1$,
(293) $E_{0}(\tau;p)\ \leq\ E_{0}(0;p)\,\exp\!\big(Ca^{-3/2+\delta}\big)+Ca^{-1}\ \leq\ (1+o_{a\to\infty}(1))\,E_{0}(0;p)+Ca^{-1}$
uniformly for $\tau\leq 1$.
Finally, we recall the null–decomposition of the Weyl tensor relative to a null frame $\{e_{4}=L,e_{3}=\underline{L},e_{1},e_{2}\}$ adapted to $\mathcal{N}^{-}_{\tau}(p)$. With the standard normalization $g(L,\underline{L})=-2$, the integrand $Q[W](L,T,T,T)$ controls, up to universal weights depending only on the frame normalization, the square sum of null curvature components $(\alpha,\beta,\rho,\sigma,\underline{\beta},\underline{\alpha})$ along $\mathcal{N}^{-}_{\tau}(p)$. In particular, One has the coercivity estimate
(294) $\int_{\mathcal{N}^{-}_{\tau}(p)}Q[W](L,T,T,T)\,d\mu_{\mathcal{N}}\ \gtrsim\ \int_{\mathcal{N}^{-}_{\tau}(p)}\Big(|\alpha|^{2}+|\beta|^{2}+|\rho|^{2}+|\sigma|^{2}+|\underline{\beta}|^{2}\Big)\,d\mu_{\mathcal{N}},$
where the implied constant is numerical and depends only on the choice of normalization.
At this stage, the control of $E_{0}(0;p)$ and of the initial flux entering (294) is supplied by the characteristic development in $D_{a,1}$ (Section 3), which provides pointwise bounds for the null curvature components near u=au=-a and hence an a priori bound on the initial Bel–Robinson energy. Combined with (293)–(LABEL:eq:F0-control), this yields a uniform $L^{2}$–control of the Weyl curvature on truncated null cones within $\mathcal{R}_{a}$.
Figure 4. Shaded region is the domain of bulk integration
We next derive energy–flux estimates for covariant derivatives of the Weyl curvature. The key point is that, upon commuting the Bianchi system, the commuted Weyl fields remain algebraically Weyl but acquire nontrivial divergence currents which are quadratic in lower derivatives of $W$. These error currents enter the divergence of the Bel–Robinson tensor and must be treated as source terms in the cone energy identity.
We do this now. Let $U$ be a Weyl field, i.e. a (0,4)(0,4)–tensor satisfying the algebraic symmetries of the Weyl curvature (traceless, alternating in each index pair, symmetric under exchange of pairs, and satisfying the algebraic Bianchi identity). Suppose $U$ satisfies an inhomogeneous Bianchi system
(295) $\mathsfbf D^{\alpha}U_{\alpha\beta\gamma\delta}=J_{\beta\gamma\delta},\qquad\mathsfbf D^{\alpha}{}^{\star}U_{\alpha\beta\gamma\delta}={}^{\star}J_{\beta\gamma\delta},$
for some (0,3)(0,3)–tensors $J,{}^{\star}J$. Then, expanding $\mathsfbf D^{\alpha}Q_{\alpha\beta\gamma\delta}[U]$ directly from the definition of $Q[U]$ by the Leibniz rule and using (295), one obtains the exact identity
$\mathsfbf D^{\alpha}Q_{\alpha\beta\gamma\delta}[U]$ $=U_{\beta}{}^{\mu}{}_{\delta}{}^{\nu}\,J_{\gamma\mu\nu}+U_{\beta}{}^{\mu}{}_{\gamma}{}^{\nu}\,J_{\delta\mu\nu}+U_{\gamma}{}^{\mu}{}_{\delta}{}^{\nu}\,J_{\beta\mu\nu}$
(296) $\quad+{}^{\star}U_{\beta}{}^{\mu}{}_{\delta}{}^{\nu}\,{}^{\star}J_{\gamma\mu\nu}+{}^{\star}U_{\beta}{}^{\mu}{}_{\gamma}{}^{\nu}\,{}^{\star}J_{\delta\mu\nu}+{}^{\star}U_{\gamma}{}^{\mu}{}_{\delta}{}^{\nu}\,{}^{\star}J_{\beta\mu\nu}.$
In particular, for any future timelike vector $T$ satisfying $g(T,T)\geq c_{0}>0$ in the region of interest, there exists a constant $C=C(c_{0})$ such that
(297) $\big|\mathsfbf D^{\alpha}Q_{\alpha\beta\gamma\delta}[U]\;T^{\beta}T^{\gamma}T^{\delta}\big|\leq C(c_{0})\,|U|\,|J|.$
Now we schematically obtain the commuted Bianchi system. For an integer $m\geq 1$, define the mm–th commuted Weyl field
$W^{(m)}:=\mathsfbf D^{m}W,$
where $\mathsfbf D$ denotes the spacetime Levi–Civita connection and $\mathsfbf D^{m}$ indicates iterated covariant differentiation. Since $\mathsfbf D$ preserves the metric and the Hodge operator, $W^{(m)}$ retains the Weyl algebraic symmetries. However, commuting the Bianchi system yields a nontrivial divergence:
(298) $\mathsfbf D^{\alpha}W^{(m)}_{\alpha\beta\gamma\delta}=J^{(m)}_{\beta\gamma\delta},\qquad\mathsfbf D^{\alpha}{}^{\star}W^{(m)}_{\alpha\beta\gamma\delta}={}^{\star}J^{(m)}_{\beta\gamma\delta}$.
Using the Ricci commutator and vacuum $\mathrm{Rm}=W$, one obtains the schematic structure
(299) $J^{(m)}=\sum_{i=0}^{m-1}\mathsfbf D^{\,i}W*\mathsfbf D^{\,m-1-i}W,\qquad{}^{\star}J^{(m)}=\sum_{i=0}^{m-1}\mathsfbf D^{\,i}W*\mathsfbf D^{\,m-1-i}W,$
where * denotes a universal finite sum of contractions depending only on $m$. Consequently,
(300) $|J^{(m)}|\leq C_{m}\sum_{i=0}^{m-1}|\mathsfbf D^{\,i}W|\,|\mathsfbf D^{\,m-1-i}W|,$
for a constant $C_{m}$ depending only on $m$.
Now we control the commuted Bel-Robinson current in the cone integral. Fix $p\in\mathcal{R}_{a}$ and $\tau\in(0,\tau_{*}]$ as in the zeroth–order argument, and set $D_{\tau}(p)$ to be the truncated causal past domain with boundary $\Sigma_{t(p)}(p)\cup\Sigma_{t(p)-\tau}(p)\cup\mathcal{N}^{-}_{\tau}(p)$. Let $T=\partial_{t}$ be the time vector field associated to the ADM foliation, and let $\pi_{T}$ denote its deformation tensor. Define the commuted Bel–Robinson current
$P^{(m)\alpha}:=Q^{\alpha}{}_{\beta\gamma\delta}[W^{(m)}]\;T^{\beta}T^{\gamma}T^{\delta}.$
Expanding $\mathsfbf D_{\alpha}P^{(m)\alpha}$ by the Leibniz rule and using symmetry of $Q$, one has the exact identity
(301) $\mathsfbf D_{\alpha}P^{(m)\alpha}$ $=\big(\mathsfbf D^{\alpha}Q_{\alpha\beta\gamma\delta}[W^{(m)}]\big)\,T^{\beta}T^{\gamma}T^{\delta}+\frac{3}{2}\,Q[W^{(m)}]_{\alpha\beta\gamma\delta}\,\pi_{T}^{\alpha\beta}T^{\gamma}T^{\delta}.$
Define the source term
$\mathcal{S}_{m}:=\big(\mathsfbf D^{\alpha}Q_{\alpha\beta\gamma\delta}[W^{(m)}]\big)\,T^{\beta}T^{\gamma}T^{\delta}.$
By (297) with $U=W^{(m)}$ and (298), we have the pointwise bound
(302) $|\mathcal{S}_{m}|\leq C(c_{0})\,|W^{(m)}|\,|J^{(m)}|.$
Define the mm–th order curvature energy and null flux by
$E_{m}(\tau;p):=\int_{\Sigma_{t(p)-\tau}(p)}Q[W^{(m)}](n,T,T,T)\,d\mu_{\Sigma_{t(p)-\tau}},\qquad F_{m}(\tau;p):=\int_{\mathcal{N}^{-}_{\tau}(p)}Q[W^{(m)}](L,T,T,T)\,d\mu_{\mathcal{N}}.$
Applying the divergence theorem to (301) on $D_{\tau}(p)$ yields the exact commuted energy–flux identity
(303) $E_{m}(0;p)+F_{m}(\tau;p)=E_{m}(\tau;p)+\frac{3}{2}\int_{\mathcal{D}_{\tau}(p)}Q[W^{(m)}]_{\alpha\beta\gamma\delta}\,\pi_{T}^{\alpha\beta}T^{\gamma}T^{\delta}\,dV+\int_{\mathcal{D}_{\tau}(p)}\mathcal{S}_{m}\,dV.$
As in the zeroth-order case, the dominant property of $Q$ implies the pointwise inequality
(304) $\big|Q[W^{(m)}]_{\alpha\beta\gamma\delta}\,\pi_{T}^{\alpha\beta}T^{\gamma}T^{\delta}\big|\leq C\,|\pi_{T}|\,Q[W^{(m)}](n,T,T,T),$
for a universal constant $C$. Using $dV=N\,dt\,d\mu_{\Sigma_{t}}$, we obtain
(305) $\left|\int_{\mathcal{D}_{\tau}(p)}Q[W^{(m)}]_{\alpha\beta\gamma\delta}\,\pi_{T}^{\alpha\beta}T^{\gamma}T^{\delta}\,dV\right|\leq C\int_{0}^{\tau}\|\pi_{T}\|_{L^{\infty}(\Sigma_{t(p)-s}(p))}\,E_{m}(s;p)\,ds.$
It remains to estimate the source contribution. By (302) and (300),
(306) $\int_{\mathcal{D}_{\tau}(p)}|\mathcal{S}_{m}|\,dV$ $\leq C_{m}\sum_{i=0}^{m-1}\int_{\mathcal{D}_{\tau}(p)}|W^{(m)}|\,|\mathsfbf D^{\,i}W|\,|\mathsfbf D^{\,m-1-i}W|\,dV.$
To close the estimate one uses the following standard splitting: two factors are placed in $L^{2}$ and the remaining factors are placed in $L^{\infty}$, supported by the bootstrap bounds and the Sobolev inequality bounds. It follows from Cauchy–Schwarz in $dV$ that
(307) $\int_{\mathcal{D}_{\tau}(p)}|\mathcal{S}_{m}|\,dV\leq C_{m}\,a^{-3/2+\delta}\int_{0}^{\tau}E_{m}(s;p)\,ds.$
(Here $C_{m}$ depends only on $m$, but is independent of $a$.) Combining (303), (305), and (307), and discarding the nonnegative flux term, yields
(308) $E_{m}(\tau;p)\leq E_{m}(0;p)+C\int_{0}^{\tau}\|\pi_{T}\|_{L^{\infty}(\Sigma_{t(p)-s}(p))}\,E_{m}(s;p)\,ds+C_{m}\,a^{-3/2+\delta}\int_{0}^{\tau}E_{m}(s;p)\,ds.$
Invoking (284) and $\tau\leq 1$, we obtain
(309) $E_{m}(\tau;p)\ \leq\ E_{m}(0;p)\,\exp\!\big(Ca^{-3/2+\delta}\big),$
and consequently, for instance at $\tau=\tfrac{3}{4}$,
$E_{m}\!\left(\tfrac{3}{4};p\right)\ \leq\ (1+o_{a\to\infty}(1))\,E_{m}(0;p).$
This completes the gauge–invariant $L^{2}$–control of $\mathsfbf D^{m}W$ on truncated null cones in the interior. We now turn to the evolution estimates for the metric coefficients $(g_{ij},k_{ij},N,X)$, which will be carried out in spacetime harmonic gauge.
Set
$\Omega:=J^{+}(\mathcal{M}^{1/a}_{int})\cap\{-a\leq t\leq-a+1\}.$
We argue in spacetime harmonic gauge on $\Omega$, i.e. we work with spacetime harmonic coordinates
$\Box_{g}x^{\mu}=0\qquad(\mu=0,1,2,3),$
equivalently $\Gamma^{\mu}:=g^{\alpha\beta}\Gamma^{\mu}_{\alpha\beta}=0$. In this gauge, the vacuum Einstein equations reduce to a quasilinear wave system
(310) $\Box_{g}g_{\mu\nu}=Q_{\mu\nu}(\partial g,\partial g),$
where $Q_{\mu\nu}$ is a universal quadratic form (no zeroth order terms). Moreover the constraints $\Gamma^{\mu}=0$ propagate (standard: $\Gamma^{\mu}$ satisfies a homogeneous linear wave equation with coefficients depending on $g$), hence (310) is equivalent to $\text{Ric}(g)=0$ on $\Omega$.
Because $\mathcal{M}^{1/a}_{int}$ has diameter $\simeq a$ in the chosen interior chart (recall that the $H$-radius of $\mathcal{M}^{1/a}_{int})$ is $O(a)$), global $L^{2}$-based homogeneous Sobolev norms on $\mathcal{M}^{1/a}_{int}$ can be $\mathcal{O}(a)$ even if $\|\partial g\|_{L^{\infty}}$ is $\mathcal{O}(a^{-1})$. To obtain pointwise control without losing powers of $a$, we work with uniformly local Sobolev norms.
Fix a smooth cutoff $\chi\in C_{c}^{\infty}(B_{2}(0))$ with $\chi\equiv 1$ on $B_{1}(0)$, and for each $y\in\mathbb{R}^{3}$ define $\chi_{y}(x):=\chi(x-y)$. For a sufficiently large integer $s$ set
$\|u(t)\|_{H^{s}_{\mathrm{ul}}(\Sigma_{t}\cap J^{+}(\widetilde{\mathcal{M}}_{1}))}:=\sup_{y\in\mathbb{R}^{3}}\ \|\chi_{y}\,u(t)\|_{H^{s}(\Sigma_{t}\cap J^{+}(\mathcal{M}^{1/a}_{int}))}$
(and similarly for $L^{2}_{\mathrm{ul}},\,L^{\infty}_{\mathrm{ul}}$). The crucial point is that the Sobolev embedding
(311) $\|u(t)\|_{L^{\infty}_{\mathrm{ul}}}\ \lesssim\ \|u(t)\|_{H^{2}_{\mathrm{ul}}}$
has a constant depending only on $\chi$ and the local geometry, and is therefore independent of $a$ (the estimate is just the standard $H^{2}(B_{2})\hookrightarrow L^{\infty}(B_{1})$ after localization).
Write $m$ for Minkowski in spacetime harmonic coordinates and h:=gmh:=g-m. On $\Sigma_{-a}\cap\mathcal{M}^{1/a}_{int}$, the assumptions (280)–(282) give
(312) $\|\partial g(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-1},\qquad\|k(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-3/2},\qquad\|N(-a)-1\|_{L^{\infty}}+\|\nabla N(-a)\|_{L^{\infty}}\ \lesssim\ a^{-3/2},$
and for he higher derivatives, we have
(313) $\|\partial^{k+1}g(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-1},~\|\partial^{k}k(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-3/2},$
[…]
To run a quasi-linear energy argument, we must control a finite number of derivatives in $L^{2}_{\rm ul}$. This is supplied by the smoothness of the data together with the curvature control coming from the characteristic development on $H_{-a-1/a}$ (as assumed in the statement). Concretely, we use the following standard input:
Because $\Gamma^{\mu}=0$, the deformation tensor of $\partial_{\alpha}$ is $\mathcal{O}(\partial g)$, and one has local (unit-scale) hyperbolic estimates relating $\partial^{2}g$ to $\text{Rm}(g)$ plus quadratic terms in $\partial g$. Therefore, the Bel-Robinson energy bounds available from the characteristic development (and their propagated Cauchy analogues on $\Sigma_{-a}$ restricted to $\mathcal{M}^{1/a}_{int}$) imply that for $s\leq 3$,
(314) $\|\partial g(-a)\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int})}+\|\partial_{t}g(-a)\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int})}\ \leq\ Ca^{-1},$
(315) $\|\partial^{k+1}g(-a)\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int})}+\|\partial^{k}\partial_{t}g(-a)\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int})}\ \leq\ Ca^{-k-1},~k\geq 2$
where $C$ is numerical and independent of $a$. (global $L^{2}$ norms may scale like $O(a)$, but the supremum over unit balls is $O(a^{-1})$)
For each $y\in\mathbb{R}^{3}$ define the localized energy on $\Sigma_{t}$:
$E_{s,y}(t):=\sum_{|\alpha|\leq s-1}\int_{\Sigma_{t}\cap J^{+}(\mathcal{M}^{1/a}_{int})}\Big(|\partial_{t}(\chi_{y}\partial^{\alpha}h)|^{2}+|\nabla(\chi_{y}\partial^{\alpha}h)|^{2}\Big)\,dx,$
and set the uniformly local energy
$E_{s}^{\mathrm{ul}}(t):=\sup_{y\in\mathcal{M}^{1/a}_{int}}E_{s,y}(t).$
We bootstrap, on $t\in[-a,-a+1]$, the smallness condition
(316) $\|\partial g(t)\|_{L^{\infty}_{\mathrm{ul}}(\Sigma_{t}\cap J^{+}(\mathcal{M}^{1/a}_{int}))}\ \leq\ C_{0}\,a^{-1+\delta},$
for a large numerical $C_{0}$ to be chosen.
Commuting (310) with $\partial^{\alpha}$ and multiplying by $\partial_{t}(\chi_{y}\partial^{\alpha}h)$, one obtains, after integration by parts on the (truncated) domain of dependence of $\operatorname{supp}\chi_{y}$ and using that the cutoff derivatives produce only lower-order terms supported in $B_{2}(y)\setminus B_{1}(y)$,
(317) $\frac{d}{dt}E_{s,y}(t)\ \leq\ C\Big(\|\partial g(t)\|_{L^{\infty}_{\mathrm{ul}}}\,+\,1\Big)\,E_{s,y}(t),$
where $C$ is numerical. Here the “+1+1” absorbs harmless commutators from $\chi_{y}$ (independent of $a$). Taking the supremum over $y$ yields
(318) $\frac{d}{dt}E_{s}^{\mathrm{ul}}(t)\ \leq\ C\Big(\|\partial g(t)\|_{L^{\infty}_{\mathrm{ul}}}\,+\,1\Big)\,E_{s}^{\mathrm{ul}}(t).$
Under the bootstrap (316) and for $a\gg 1$ (so that $a^{-1+\delta}\ll 1$), we obtain
$\frac{d}{dt}E_{s}^{\mathrm{ul}}(t)\ \leq\ C^{{}^{\prime}}(a^{-1+\delta}+1)\,E_{s}^{\mathrm{ul}}(t),$
with $C^{\prime}$ numerical. Hence, by Grönwall on an interval of length $1$,
(319) $E_{s}^{\mathrm{ul}}(t)\ \leq\ e^{C^{\prime}(a^{-1+\delta}+1)}\,E_{s}^{\mathrm{ul}}(-a)\ \qquad\text{for all }t\in[-a,-a+1],$
where $C^{\prime}$ is numerical and independent of $a$. Now using (314)) and uniformly local Sobolev (311) applied to $\partial g(t)$, and the definition of $E_{s}^{\rm ul}$,
(320) $\|\partial g(t)\|_{L^{\infty}_{\mathrm{ul}}}\ \lesssim\ \|\partial g(t)\|_{H^{2}_{\mathrm{ul}}}\ \lesssim\ \big(E_{s}^{\mathrm{ul}}(t)\big)^{1/2}\ \lesssim\ a^{-1}\lesssim a^{-1+\delta}a^{-\delta}\leq\frac{1}{2}a^{-1+\delta}$
and similarly
(321) $\|\partial^{k}g(t)\|_{L^{\infty}_{\mathrm{ul}}}\ \lesssim\ \|\partial^{k}g(t)\|_{H^{2}_{\mathrm{ul}}}\ \lesssim\ \big(E_{s}^{\mathrm{ul}}(t)\big)^{1/2}\ \lesssim\ a^{-k}\lesssim a^{-k+\delta}a^{-\delta}\leq\frac{1}{2}a^{-k+\delta}$
for sufficiently large $a\gg 1$.
Consequently, choosing $C_{0}$ large and $a\geq a_{0}(\delta)\gg 1$, we improve (316) and close the bootstrap:
(322) $\|\partial g\|_{L^{\infty}_{\mathrm{ul}}(\Omega)}\ \lesssim\ a^{-1+\delta},$
(323) $\|\partial^{k+1}g\|_{L^{\infty}_{\mathrm{ul}}(\Omega)}\ \lesssim\ a^{-k-1+\delta}$
In particular, the same bound holds for the usual $L^{\infty}$ norm on $\Omega$.
We now relate $(N,X)$ to the metric components. In ADM form with shift,
$g_{00}=-N^{2}+g_{ij}X^{i}X^{j},\qquad g_{0i}=g_{ij}X^{j}.$
Since $g_{ij}$ remains uniformly equivalent to $\delta_{ij}$ on $\Omega$ (by integrating (322) in time and using the initialization), we can solve $X^{j}=g^{ji}g_{0i}$ and obtain
(324) $\|X\|_{L^{\infty}(\Omega)}\ \lesssim\ \|g_{0i}\|_{L^{\infty}(\Omega)}.$
But $g_{0i}$ satisfies the same reduced wave equation (310) and has the same initial smallness (in the interior construction one has $g_{0i}(-a)=a^{-1}$ and $\partial g_{0i}(-a)=\mathcal{O}(a^{-1+\delta})$), hence (322) implies, after integrating $\partial_{t}g_{0i}$ over a unit time interval,
$\|g_{0i}\|_{L^{\infty}(\Omega)}\ \lesssim\ a^{-1+\delta}.$
Thus
(325) $\|X\|_{L^{\infty}(\Omega)}\ \lesssim\ a^{-1+\delta},$
which in particular yields the stated $\|X\|_{L^{\infty}}\lesssim a^{-1+\delta}$. But, we can obtain a better estimate given the initial estimate $X||_{L^{\infty}(\mathcal{M}^{1/a}_{int})}\lesssim a^{-3/2}$. This follows from the fact that $X=g_{0i}$ verifies
$\Box X_{i}=Q_{i}(\partial g,\partial g)$
and therefore the wave equation higher order energy estimate and Sobolev inequality implies
$X||_{L^{\infty}(\Omega)}\lesssim a^{-\frac{3}{2}}$
Similarly, from $g_{00}+1=-(N^{2}-1)+g_{ij}X^{i}X^{j}$, we obtain
$\|N-1\|_{L^{\infty}(\Omega)}\ \lesssim\ \|g_{00}+1\|_{L^{\infty}(\Omega)}+\|X\|_{L^{\infty}(\Omega)}^{2}\ \lesssim\ a^{-3/2},$
which is the middle estimate in the proposition.
Use the exact ADM identity
(326) $k_{ij}=-\frac{1}{2N}\big(\partial_{t}g_{ij}-\nabla_{i}X_{j}-\nabla_{j}X_{i}\big).$
one obtains
$k||_{L^{\infty}(\Omega)}\lesssim a^{-1+\delta}.$
But this is not optimal for our purpose. We can obtain a better estimate by directly integrating the evolution equation for $k$ and the evolution equation for $k$
$\partial_{t}k_{ij}=-\nabla_{i}\nabla_{j}N+N(\text{Ric}_{ij}+k_{ik}k^{k}~_{j}-\mbox{tr}_{g}kk_{ij})+L_{X}k_{ij}$
which yields the improved estimate
$k||_{L^{\infty}(\Omega)}\lesssim a^{-\frac{3}{2}+\delta}$
since $\text{Ric}||_{L^{\infty}(\Omega)}\lesssim a^{-2+\delta}$.
Combiningthe estimates yields, throughout $\Omega$,
$\|k\|_{L^{\infty}}\lesssim a^{-3/2+\delta},\qquad\|N-1\|_{L^{\infty}}\lesssim a^{-3/2+\delta},\qquad\|X\|_{L^{\infty}}\lesssim a^{-3/2+\delta},$
for any fixed $\delta\in(0,\tfrac{1}{2})$, provided $a\geq a_{0}(\delta)\gg 1$. All implicit constants are numerical (they depend only on the fixed cutoff $\chi$ and unit-scale Sobolev constants, hence do not depend on $a$). This completes the proof. ∎
Remark 9.
Notice that estimates $|k|\lesssim a^{-\frac{3}{2}+\delta}$ is only true in the time range [a,a+O(1)][-a,-a+O(1)] given initial data $|k|\lesssim a^{-\frac{3}{2}}$. For a larger time interval e.g., of length $O(a)$ would trivially ruin this due to large deformation.
Remark 10.
In principle, one could also perform the analysis on the spacetime domain
$\bigcup_{t\in[-a,-a-1/a+3/4]}\Phi_{t}\bigg(\Phi^{-1}_{t=-a-1/a+\frac{3}{4}}(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+3/4})\bigg)$ since one already posses the estimates (276-278) on the initial domain $\Phi^{-1}_{t=-a-1/a+\frac{3}{4}}(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+3/4})$ by the Characteristic development.
Corollary 7.1 (Quantitative increase of the Schoen–Yau radius).
Let $a\gg 1$ and let $\mathcal{M}^{1/a}_{\mathrm{int}}\subset\mathcal{M}_{-a}$ be the interior region constructed in Proposition 7.2. There exists $a_{0}\gg 1$ such that for all $a\geq a_{0}$,
$\mathrm{Rad}\!\left(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+3/4}\right)\;\geq\;\mathrm{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})\left(1+\frac{1}{10a}\right).$
Proof.
Let $T$ denote the future–directed unit normal to the foliation $\{\mathcal{M}_{t}\}$, and let $\Phi_{T}(t,\cdot)$ be its flow. For $t\in[0,3/4]$, set
$g(t):=(\Phi_{T}^{-1})^{*}g$
to be the induced metric on $\Phi_{T}(t,\mathcal{M}^{1/a}_{\mathrm{int}})\subset\mathcal{M}_{-a-1/a+t}$. Along the flow of $T$, the metric satisfies the exact evolution equation
(327) $\partial_{t}g_{ij}=-2Nk_{ij}+\mathscr{L}_{X}g_{ij}.$
By Proposition 7.2, on $\Phi_{T}(t,\mathcal{M}^{1/a}_{\mathrm{int}})$ we have the uniform bounds
$|k|\lesssim a^{-3/2+\delta},\qquad|N-1|+|X|\lesssim a^{-3/2+\delta},\qquad t\in[0,3/4].$
Integrating (327) in time yields the bilinear–form comparison
(328) $(1-Ca^{-3/2+\delta})\,g(0)\;\leq\;g(3/4)\;\leq\;(1+Ca^{-3/2+\delta})\,g(0),$
where the inequalities are understood in the sense of quadratic forms.
Since the Schoen–Yau radius is stable under $C^{0}$ quasi–isometries, it follows from (328) that
(329) $\mathrm{Rad}\!\left(\Phi_{T}(3/4,\mathcal{M}^{1/a}_{\mathrm{int}})\right)\;\geq\;\mathrm{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})\bigl(1-O(a^{-3/2+\delta})\bigr).$
Next, observe that $\Phi_{T}(3/4,\mathcal{M}^{1/a}_{\mathrm{int}})$ is a proper subset of
$J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+3/4},$
since $\Phi_{T}$ follows the timelike normal flow rather than null generators. To quantify the additional geometric thickness, we work near u=au=-a in double–null coordinates $(u,\underline{u},\theta^{A})$ and define the radial function
$r:=\underline{u}-u.$
From the null–gauge estimates along the initial hypersurface,
$|\Omega-1|\lesssim a^{-3/2},\qquad|b|\lesssim a^{-3/2},$
it follows that radial curves are uniformly close to unit–speed geodesics. Consequently,
(330) $\inf_{\theta^{A}\in\mathbb{S}^{2}}\int_{a}^{a+3/4}\sqrt{g_{ij}\!\left(\frac{dx^{i}}{dr},\frac{dx^{j}}{dr}\right)}\,dr=\frac{3}{4}+O(a^{-3/2}).$
Using the normalization of the interior region,
$\mathrm{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})=\frac{3\pi}{4}\,a+O(a^{-2}),$
the thickness (330) may be rewritten as $\frac{3}{4}+O(a^{-\frac{3}{2}})=\frac{4}{3\pi a}\,\mathrm{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})+O(a^{-3/2})$. We conclude that
$\mathrm{Rad}\!\left(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+3/4}\right)\;\geq\;\mathrm{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})\Bigl(1+\frac{4}{3\pi a}-O(a^{-3/2+\delta})\Bigr).$
For $a$ sufficiently large, the error term is dominated by the linear gain, yielding
$\mathrm{Rad}\!\left(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+3/4}\right)\;\geq\;\mathrm{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})\left(1+\frac{1}{10a}\right),$
which completes the proof. ∎
7.3. Completion of the proof of the main theorem
By the interior gluing/construction, the correction tensors are supported in a compact subset of $\Omega$ that is a positive distance away from $\partial\Omega$. In particular, the induced metric $\gamma$ and second fundamental form $k^{\top}$ on $\partial\Omega$ coincide exactly with those inherited from the double–null interface sphere, hence so do the null expansions. Therefore the generalized mean–curvature barrier quantity
$c_{*}\;:=\;\min_{\partial\Omega}\Big(H_{\partial\Omega}(\gamma)-\big|\mbox{tr}_{\partial\Omega}k\big|\Big)\;>\;0$
and the Schoen–Yau (HH–)radius bound
$\text{Rad}(\Omega)=R_{*}\;<\;\frac{3\pi}{2c_{*}}$
hold on $\partial\Omega$ with the same constants as in the interface geometry.
We now apply the Schoen–Yau barrier criterion (in the form used in Theorem 2.2), which asserts: if a compact initial domain $U\subset\mathcal{M}_{-a}$ has boundary satisfying
(331) $\min_{\partial U}\big(H_{\partial U}-|\mbox{tr}_{\partial U}k|\big)\;<\;\frac{3\pi}{2\,\text{Rad}(U)},$
then $U$ contains no marginally outer trapped surface. Since (331) holds for $U=\Omega$ by the preceding paragraph, it follows that there is no MOTS contained in $\Omega\subset\mathcal{M}_{-a}$.
Let $D^{\prime}\subset D_{a,1}$ denote the truncated characteristic development bounded by the outgoing null hypersurface $H_{-a-1/a}$. Define
$\mathcal{M}^{1/a}_{\mathrm{int}}\;:=\;\big(\mathcal{M}_{-a}\cap(D_{a,1}\setminus D^{\prime})\big)\;\cup\;\mathcal{M}_{1}.$
On $\mathcal{M}^{1/a}_{\mathrm{int}}$ we have $\mbox{tr}\chi>0$ by construction of the semi–global double–null region (and the choice of truncation), and the boundary $\partial\mathcal{M}^{1/a}_{\mathrm{int}}$ is a union of interface spheres along which the barrier geometry is inherited from the interface data up to the compactly supported interior corrections (which, again, do not touch the boundary). In particular, the initial slice contains no MOTS in $\mathcal{M}^{1/a}_{\mathrm{int}}$: indeed, a MOTS would have vanishing outward null expansion and hence cannot be contained in a region foliated by spheres with strictly positive outward expansion; equivalently, one may use the same Schoen–Yau barrier exclusion on the compact domains bounded by the relevant interface spheres (as above) to preclude the existence of a MOTS in $\mathcal{M}^{1/a}_{\mathrm{int}}$.
In the next step we study the quantitative radius gain under controlled evolution. By Corollary 7.1 (propagation of the Schoen–Yau radius under evolution in spacetime harmonic gauge, using the interior propagation estimates), for all sufficiently large $a$ one has
(332) $\text{Rad}\!\left(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+3/4}\right)\;\geq\;\text{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})\Big(1+\frac{1}{10a}\Big).$
Denote the evolved compact domain by
$\Omega_{3/4}\;:=\;J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+3/4}.$
Then (332) reads $\text{Rad}(\Omega_{3/4})\geq\text{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})(1+\frac{1}{10a})$.
Now we study the stability of the boundary barrier functional along the transported interface. We next quantify the change of the boundary barrier quantity $H-|\kappa|$ along the interface spheres transported from $S_{-a-1/a,\,1/a}$ to $S_{-a-1/a,\,3/4}$ through a time slab of length $3/4$. Here $\kappa:=\mbox{tr}_{\partial}k$ is the trace of $k$ tangential to the interface sphere.
On the semi-global double-null region, the null structure equations and the established point-wise bounds yield
$|\widehat{\underline{\chi}}|\lesssim a^{1/2}|u|^{-2},\qquad\mbox{tr}\underline{\chi}=-\frac{2}{|u|}+O(|u|^{-2}),$
along the relevant generators. Together with the curvature–flux control (already proved in Section 3), the generalized mean–curvature functional $\mathsfbf H$ associated to the transported interface admits the representation
$\mathsfbf H$ $=-\frac{|u_{\infty}|}{a}\,\mbox{tr}\underline{\chi}(u_{\infty},0)+\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{\prime}|\,|\widehat{\underline{\chi}}|^{2}(u^{\prime},0)\,du^{\prime}$
(333) $\quad+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{\prime}|^{2}}\int_{u_{\infty}}^{u^{\prime}}|u^{\prime\prime}|^{2}|\widehat{\underline{\chi}}|^{2}(u^{\prime\prime},0)\,du^{\prime\prime}\,du^{\prime}.$
Moreover, on the interior Cauchy development in spacetime harmonic gauge, we have the pointwise bounds (for any fixed $\delta\in(0,\frac{1}{2})$)
$|k|\leq Ca^{-3/2+\delta},\qquad|\partial g|\leq Ca^{-1+\delta}$
on the relevant slab, and the transport equation $\partial_{t}g_{ij}=-2Nk_{ij}+\mathscr{L}_{X}g_{ij}$ together with the corresponding evolution equation for $k$ implies that the induced metric and second fundamental form on the transported interface spheres vary by $O(a^{-3/2+\delta})$ in $C^{1}$ and $O(a^{-5/2+\delta})$ in the scalar barrier functional over unit time. In particular, for the two interface spheres $S_{-a-1/a,\,1/a}$ and $S_{-a-1/a,\,3/4}$ by direct integration and the estimates on the Ricci coefficients,
(334) $\Big|(H-|\kappa|)\big|_{S_{-a-1/a,\,3/4}}-(H-|\kappa|)\big|_{S_{-a-1/a,\,1/a}}\Big|\;\leq\;Ca^{-5/2},$
and the comparison between the geometric barrier quantity and $\mathsfbf H$ at the initial and transported interfaces takes the form
$(H-|\kappa|)\big|_{\partial\Omega_{3/4}}$ $=\mathsfbf H+O(a^{-5/2}),$
(335) $(H-|\kappa|)\big|_{\partial\mathcal{M}^{1/a}_{\mathrm{int}}}$ $=\mathsfbf H+O(a^{-5/2}).$
All error terms are uniform in $a$ (for $a$ large) and depend only on the bootstrap constants already fixed.
We now use the quantitative form of the criterion in Theorem 2.2(c): any compact domain $U$ whose boundary satisfies
(336) $(H-|\kappa|)\big|_{\partial U}\;>\;\frac{3\pi}{2\,\text{Rad}(U)}$
must contain a MOTS in its interior a lá [44].
We compare (336) for the initial domain $U=\mathcal{M}^{1/a}_{\mathrm{int}}$ and the evolved domain $U=\Omega_{3/4}$.
First we prove the (i) failure of (336) on the initial slice. From the explicit interface computation (using (333)) and the initial radius normalization
$\text{Rad}(\mathcal{M}_{1})=\frac{3\pi}{4}(a-1)+O(a^{-1}),$
one has
(337) $\frac{3\pi}{2\text{Rad}(\mathcal{M}_{1})}-\frac{2}{a}=\frac{2}{a^{2}}+O(a^{-3}),$
and also
(338) $\frac{3\pi}{2\text{Rad}(\mathcal{M}_{1})\sqrt{1+\frac{1}{10a}}}-\frac{2}{a}=\frac{8}{5a^{2}}+O(a^{-3}).$
Assume condition (c) of Theorem 2.2, namely
(339) $\frac{17}{10a^{2}}<\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{\prime}||\underline{\hat{\chi}}|^{2}(u^{\prime},\epsilon)\,du^{\prime}+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{\prime}|^{2}}\int_{u_{\infty}}^{u^{\prime}}|u^{\prime\prime}|^{2}|\underline{\hat{\chi}}|^{2}(u^{\prime\prime},\epsilon)\,du^{\prime\prime}\,du^{\prime}<\frac{19}{10a^{2}}.$
Combining (339) with (333) and the comparison (335) on the initial interface yields the quantitative defect
(340) $\Big(H-|\kappa|\Big)\big|_{S_{-a,0}}-\frac{3\pi}{2\text{Rad}(\mathcal{M}_{1})}\leq-\frac{1}{10a^{2}}+O(a^{-5/2}).$
Hence, for $a$ sufficiently large so that the $O(a^{-5/2})$ error is dominated by $\frac{1}{10a^{2}}$, we obtain
(341) $(H-|\kappa|)\big|_{\partial\mathcal{M}_{1}}<\frac{3\pi}{2\,\text{Rad}(\mathcal{M}_{1})}.$
In particular, the strict inequality (336) fails for $\mathcal{M}_{1}$ (and therefore for $\mathcal{M}^{1/a}_{\mathrm{int}}$), and by Steps 1–2 there is no MOTS in $\mathcal{M}^{1/a}_{\mathrm{int}}$.
Now we prove the (ii) validity of (336) on the evolved domain. By the radius gain (332), the right-hand side in (336) decreases by an amount of order $a^{-2}$ when passing from $\mathcal{M}^{1/a}_{\mathrm{int}}$ to $\Omega_{3/4}$. More precisely, using (338) together with (332) and the comparison (335) at time a1/a+3/4-a-1/a+3/4, we infer
$\big(H-|\kappa|\big)\big|_{\partial\Omega_{3/4}}-\frac{3\pi}{2\text{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})}\;\geq\;\frac{1}{10a^{2}}+O(a^{-5/2}).$
Therefore, for $a$ sufficiently large,
(342) $(H-|\kappa|)\big|_{\partial\Omega_{3/4}}>\frac{3\pi}{2\,\text{Rad}(\Omega_{3/4})}.$
Thus (336) holds strictly on $\Omega_{3/4}$.
the compact domain $\Omega_{3/4}$ and using (342), we conclude that $\Omega_{3/4}$ contains a marginally outer trapped surface in its interior. By Steps 1–2, no MOTS is present in $\mathcal{M}^{1/a}_{\mathrm{int}}\subset\mathcal{M}_{-a}$ on the initial slice. Therefore this MOTS is created strictly by the vacuum Einstein evolution from regular initial data, i.e. it is a dynamically formed MOTS.
This completes the proof of the main theorem.